A Complete Bibliography of Publications in the Journal of the Association for Information

Science and Technology

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

30 April 2024 Version 1.29

Title word cross-reference

#BlockSidewalk [MMFA20].

(N < 1 000) [LN17]. $110.00 [Bow19a, Ma22b, Ma22a]. $112.00 [vR17a]. $125.00 [Dor19, Kno24]. $126 [QH18]. $129.00 [All19, Gui17, Uns18]. $14.99 [Fou19]. $149.50 [Rou17]. $15.95 [How17, Saw19]. $17.95 [Hai16]. $175.00 [Hay18]. $18.00 [Soe18]. $21.53 [Saw19]. $22.95 [Gor20a]. $25.00 [FJ19]. $26.00 [BI17, Kos20, Zim17]. $27.00 [Meh22]. $28.00 [Afz22, Dut20, Hal19]. $29.95 [Jon17, See17]. $30 [Ols16]. $30.00 [Fri17, Sal20, Web20]. $34.00 [Cle18]. $34.95 [Huv22a]. $35.00 [Adl23, Ben18, Bow19b, Con20, Zha19]. $38.00 [Rhe17]. $39.95 [Whi17]. 45 [TW16b]. $49.95 [Far23]. 5 [YL18]. $50.00 [Kos21]. $51.96 [MS19]. $54.99 [Buc20, He18]. $59.95 [Bru18, Cho19]. $60.00 [Agu18]. $64.95 [Saw19].

$65.00 [Bat21, Haw18, Kea22]. $69.95 [O’H18]. $79.95 [Abb17]. $79.99 [Ama19, Jul20]. $80.00 [Erd18]. $82.95 [Col17b]. $85.00 [Bea17, Lev20]. $90.00 [Hjø17a, Sko23]. $95.00 [Aji17, Mat20]. $99.00 [Fre18]. $99.95 [Clo18]. > [The17]. ^{st} [Kon16]. _{14} [SS14c]. g [Bar15, Pra14b, Rou15]. h [Bar15, Cri15, Epi14, Mal15, Pra14b, Pra14c, Sch15, YL14]. N [DGR19].

-century [Kon16]. -gram-based [DGR19]. -index [Bar15, Cri15, Mal15, Pra14c, Rou15, Sch15, YL14]. -indices [Pra14b]. -plots [Epi14]. -year [YL18].

. [Sha21b].

1.0 [SKKM22]. 10-year [GAB^{+}18]. 10.1002 [Ano14]. 10.1002/asi.20981 [Ano14]. 10SENT [MDJ^{+}19]. 12 [DPN16, RAMP22]. 12-month [MPB^{+}20]. 12th [PMGB^{+}21]. 15th [PMGB^{+}21]. 19 [BK23, CCM23, Hod21, LBC^{+}22, NB23, RCM22, SBS23, SJT^{+}21, VFZ^{+}22, WZFZ22, ZGL^{+}22, Sha21a]. 1960s [Eng17b]. 1990s [Moo20].

2 [Adl23, Sko23]. 2.0 [DGCC17, TTH16]. 2013 [Bea17]. 2021 [TKA^{+}23]. 21st [dP20]. 2nd [SWKK16, Wei14b].

3rd [Haw18].

40-year [KH17]. 4th [Col17b].

60 [Ano14].

7 [Far23].

978 [Adl23, Far23, Sko23]. 978-0-08-102474-4 [Dor19]. 978-0-12-417112-1 [Abb17]. 978-0-19-871918-2 [Hai16]. 978-0-231-18620-9 [Ma22a]. 978-0-231-19552-2 [Sko23]. 978-0-231-19876-9 [Ma22b]. 978-0-262-02821-9 [Ols16]. 978-0-262-02881-3 [Whi17]. 978-0-262-03512-5 [BI17]. 978-0-262-03517-0 [Jon17]. 978-0-262-03624-5 [Cle18]. 978-0-262-03687-0 [Con20, Haw18]. 978-0-262-03709-9 [Bow19b]. 978-0-262-03744-0 [Ben18]. 978-0-262-03773-0 [Sal20]. 978-0-262-03785-3 [Hal19]. 978-0-262-03828-7 [Zha19]. 978-0-262-03859-1 [Bat21]. 978-0-262-03902-4 [Web20]. 978-0-262-04303-8 [Kos21]. 978-0-262-04350-2 [Adl23]. 978-0-262-52538-1 [Rhe17]. 978-0-262-52781-1 [Fri17]. 978-0-262-52820-7 [Zim17]. 978-0-262-52851-1 [How17]. 978-0-300-23401-5 [Kos20]. 978-0-367-18816-0 [Sto20]. 978-0-367-21944-4 [Swa22]. 978-0-367-90285-8 [Owe21]. 978-0-525-57451-4 [Afz22].

978-0-691-17954-4 [Kea22]. 978-0-7591-2333-5 [Bea17]. 978-0-8122-4923-1 [O’H18]. 978-0-8166-9043-5 [See17]. 978-0-8173-2060-7 [Far23]. 978-0-8389-1722-0 [Buc20]. 978-0-9817584-1-1 [Soe18]. 978-0-9817584-2-8 [Soe18]. 978-0-9817584-3-5 [Soe18]. 978-1-107-12216-1 [Bow19a]. 978-1-108-79539-5 [San22]. 978-1-119-09294-0 [Bru18]. 978-1-4422-4992-9 [Aji17]. 978-1-4739-1632-6 [Hay18]. 978-1-4773-0824-0 [Hjø17a]. 978-1-5095-1746-6 [Saw19]. 978-1-5095-3445-6 [Rey21]. 978-1-5095-3642-9 [Gor20a]. 978-1-5179-0025-0 [FJ19]. 978-1-5179-0154-7 [Dut20]. 978-1-5179-0715-0 [Meh22]. 978-1-5179-0946-8 [Huv22a]. 978-1-5381-1245-8 [Lev20]. 978-1-5381-7390-9 [Kno24]. 978-1-57387-499-1 [Rou17]. 978-1-62705-917-6 [Agu18]. 978-1-68173-082-0 [Fou19]. 978-1-68173-083-7 [Cho19]. 978-1-68173-300-5 [MS19]. 978-1-78330-275-8 [Huv20]. 978-1-78330-338-0 [Pan21]. 978-1-78560-968-8 [Col17b]. 978-1-78756-722-1 [Mat20]. 978-1-78973-364-8 [Fro21]. 978-1-84334-749-1 [Sav18b]. 978-1-84334-750-7 [Erd18]. 978-1-85604-850-7 [Col17a]. 978-1-876924-67-6 [Clo18]. 978-3-11-029803-1 [vR17a]. 978-3-11-037616-6 [Bud17]. 978-3-11-041816-3 [QH18]. 978-3-11-057525-5 [dP20]. 978-3-11-062954-5 [Han21]. 978-3-319-30148-8 [Gui17]. 978-3-319-56755-6 [Fre18]. 978-3-319-60776-4 [He18]. 978-3-319-71195-9 [Ama19]. 978-3-319-73514-6 [All19]. 978-3-319-92455-7 [Jul20]. 978-3-319-98487-2 [Fra21, Sha21a]. 978-3-658-14679-5 [Uns18]. 978-90-485-3741-9 [Sha21b]. 978-94-6298-619-0 [Sha21b].

Abel [Ma22b]. abeyance [Roh21]. Abingdon [Owe21, Swa22]. Abingdon-on-Thames [Swa22]. aboutness [OKB17]. above [Wor19]. abroad [PYH16]. absence [PN18]. abstract [ONZ23]. Abstracting [ZZL^{+}14]. abstracts [LJ22, PDS22]. Abuses [BI17]. academia [HRL21, JA16, Mah17]. Academia.edu [TK14]. Academic [CA20, Jor19, OGL19, PKB^{+}16, RJS^{+}17b, TM20, AT14, AMDFHVC24, BRS16, BDH17, CHH14, CCJ16, CKLP17, Cox23, GSG20, GJ22, GAS^{+}20, GB19b, HCKH15, HS22, Hoo17, Jam20, JHJ15, JDHL17, KPWS20, Kno24, KT16a, LL17, LOD^{+}19, LB22, LWM^{+}23, PB15, PBG^{+}17, RCM22, Rou17, Sil20, SL24, TD15, TM16, TK17c, VRF14, VCG14, YL14, ZTLV15, Hal14, OA14, TK14]. academics [Fas18, MSPP21]. acceptance [FvdWR^{+}22, MBVB17, TMBA20]. accepted [OR15]. access [BLWP14, Bjö14, Bor16c, BGS22, DQAMM^{+}24, FSPC16, IB16, JML22, Jam17, Jam20, JPF16, KN23, LMJ21a, LMJ21b, LWA^{+}17, LCWL18, MZ22, Moo20, MBZ^{+}22, NE17, ØC19, PSB^{+}14, PSB16, PSB17, QH18, RJS^{+}17b, She21, TD15, TLD17, VLBG^{+}16, WCP^{+}19, YL18, vLTW18, Bud17].

Accessibility [SR16, Aji17, WvdHS16, Zha14b]. Accessing [FHO23, GLR16]. Accidental [Erd18]. accommodation [LF21]. accomplishments [Lia17]. according [AKK^{+}17]. account [Cri15, FSK^{+}16, ZRG16]. Accounting [Hal19]. accounts [HBH^{+}16, KISS15]. accuracy [LJ17, TPR17, vdSvdB23]. Accurate [KB17]. accurately [SJL19]. achieved [WZL21]. achievements [RN16]. Achieving [KC20, CK17, Du14a]. acknowledgment [Ano14, SMAB09, THL17]. Acknowledgments [DFB14, GRH^{+}17, DPHP17]. acquiring [BWGD21]. Acquisition [MHJ19, LOM22]. acquisitions [ZZE17]. across [APRRC15, DFB14, RBJ22, Sin16, YSAM^{+}21, ZK14]. action [Fro21, KWR20, SJT^{+}21, XHM^{+}20, ZZY23]. actionable [KWR20]. actions [Zem17]. active [KDS^{+}17, SGV14, TGC^{+}21]. activists [VO23]. activities [FJF^{+}21, Gau16, Gor19, KL22, LSHH23, RC16]. activity [BGS22, FA19, GSG22]. actor [HL20]. actor-topic [HL20]. actors [TC17]. ad [DLA22, GBL^{+}18, LKRS18, MNNK15, WSWH20]. Adam [San22]. Adaptation [SKC18]. adapting [TG17]. Adaptive [DMMF17, WLZB14, HPLC19, Rog15]. Add [FJT23]. Add-Health [FJT23]. added [FC18]. Adding [AKAS14, Ma21, YZ17]. address [Bor16a]. addresses [REVR22, Sav15b, Sav16b]. Addressing [LBWE^{+}23, St.17, VO23]. adherence [ZYZ^{+}16]. Adjustable [PS14a]. adjustment [Hal17]. administration [RKHMRH21]. administrators [PH22]. adolescents [BM14d, Gau16, WOJBG17]. adoption [BVH15, Bur17, KT14a, KCL14, NGTKR15, SZZ16, SWSZ19, ZDZ21, ZLF20]. adult [Cho20]. adults [CPL22, QHH20, ZZY23]. advancement [WKJW16]. Advances [Alb15, BRS15, BR15a, BR15b, CCMT15, CS15a, EMK^{+}15, GSB^{+}16, Kon16, KS14b, Mat15, MOM^{+}15, OMM^{+}14, Sha14, ZS14a]. Advancing [KWLV22]. advantage [MPB^{+}20, MZ22, TW16a]. Adventures [DSBG20]. Adverse [GRSFV15a, HSL24, LSC18]. advertisements [SL21a]. advertising [SKO24]. Aesthetic [Mar15]. affect [ALC^{+}14, Che22, ZPL^{+}20, FJ19]. affecting [CJ16a, KS16, OY15, THC15]. affective [KTZ20, LI22]. Affelt [Fro21]. affiliations [HRL21, SC21]. Affordances [Kit19, Kan22]. Africa [KN23]. African [JWWV18, JWJ^{+}19, LMA^{+}22, SWKK16]. after [VO23]. aftermath [H17]. against [LML23]. Agarwal [Fou19]. Age [Kos20, NPG^{+}20, Ols16, SR14, TK17b, Uns18, BVd16, FC18, JRL20, KvOS^{+}17, LWLH18, Owe21, TLA^{+}16, XBL^{+}22, YR16, YZQT20, ZVW20, Erd18, Hai16, Han21, Web20, Wei14b, Far23, Lev20]. age-related [KvOS^{+}17]. agency [ALAB19, JJ14, Wes15]. agent [GRSFV15c]. agentic [Gre18, VO23]. Aggregated [LdMAdN16, LdN17]. aggregating [SM22]. aggregation [DP23, MTY14, OA15, TZK20]. Agnes [dP20]. agonistic [Eck18]. agriculture [ZZSX16]. AI [JV17]. aids [FT16, MVL^{+}14]. aimed [ALMP16]. AIS [She17, Sil18, SWLH17]. ALA [Buc20]. Alabama [Far23, MJ23]. Alan [San22]. alcoholism [CY14]. Alex [Mar15]. Alger [Lev15]. Algorithm [GB19a, KK19a, CDL18, Gia16, KS17, LXLL14].

Algorithmic [SAW21, BOR^{+}23, Kos21, Sha21b]. algorithmically [CL21]. algorithms [IABL16, OSvE16, San22]. Ali [Wei14a]. align [DSBG21]. Alistair [Baw14]. Allen [Wil14a]. Allocation [WLC16]. Alon [Aji17]. Alone [GKM^{+}21]. Alphabetic [Hen19]. also [XHM^{+}20]. alternative [GNVQdMA16, SBIT14, RA14]. Altmetric [WK19, CZW15, FDC22, HLM16, TKA^{+}23]. altmetrics [AZE^{+}19, Bor16b, BH16, CZW15, DBH18, LMP21, MT14, MTHL15, SWLH17, TW16b]. always [JL18, Rub15]. amazon.com [KT16b]. Amer [Sea20]. America [Lev20, Bur20, PH22, Lev20]. American [Ano14, GTAS22, HS14, JWWV18, LMA^{+}22, Ma23b]. American/Black [LMA^{+}22]. Americans [JWJ^{+}19]. among [BNG14, BBH15, Bor17, CPL22, CL17, DAT^{+}14, HHL21, HK21, JWWV18, JSW14, LS22, LBM19a, MVL^{+}14, PTGL17, QHH20, SZ20, Yan14]. amounts [BELR16]. Amplifying [TLD17]. Amsterdam [Abb17, Sha21b]. Amy [Fro21]. Anabel [Hay18]. Anable [FJ19]. analyses [Ano17n, Bor16b, CZ16, DBBJ16, SC17]. Analysis [CCM23, HS17, KB17, LCH^{+}22, MS19, SR16, TCP15, ZGBIL17, dSBG19, Bad14, BCK17, Bel16, BWL15, BB15, BM15b, Bor16a, Bor17, CGGS17, CL14, CK14, CHT^{+}15, CTM16, Daq20, DZC^{+}14, DB17b, GD18, GZBA22, GB19b, HPS^{+}14, HL20, HGCKDR16, HDZ^{+}15, HWW15, HWC17, JLC16, JZC20, JL23, KBD^{+}14, KTS^{+}15, KS16, KY17, KBS18, Kim19a, KT17c, LSM^{+}14, LKKS17, LMP21, LHR16, LCD^{+}14, LGD20, LJ22, Lin15, LR14, LCSL15, LK16, LB22, MM18, Mal15, MSB21, MC18, MG21, MJ23, MDJ^{+}19, MTHL15, MRKT15, Mor24, MWD17, NV17, OS15, OSD22, OA14, Ort22, PYH16, PGFG^{+}20, Pet19, RJGM16, RJS19, Rou17, RLAB^{+}24, RBJ22, SRG^{+}16, Sch18, SS17, SDUDL24, SBIT15, SGS21, SKKM22, SKZ^{+}14, SJK15, SKL17, SEZG22, THL17, TT17, TM15, TW16b, TCF^{+}24, TJV14, UM23, VG18, VN18, WW15]. analysis [WZK^{+}18, WSZC22, WLTY19, XCL^{+}21, XDSC16, Yan15, YCK16, YZ16, ZDW18, ZLCB15, ZY15, ZZLS16, ZS14c, ZS14b, ZYL18, CPL22]. Analytics [CHH^{+}17, CCJ16, ELB17, JRL20, JAG^{+}20, Mar23, MVA22, SMYL15, ZPPW24]. analyze [AST^{+}22]. Analyzing [AS14, LWM^{+}19, NH15, RSAS14, RTS^{+}17, RGJCTS16, TZS^{+}22, ÇÖ18, DC22, GOdR18, HSK15, LZSW17, RB14, Olt15]. Anatoly [Ma22a, Sko23]. Anatomy [BLWP14, FS18]. anchoring [WN15]. Andresen [Han21]. Andrew [Clo18, Fri17, Rog15]. Angewandte [MWD17]. angle [LZZ^{+}21, YB18]. angles [NSK17]. Anja [Kea22]. Anja-Silvia [Kea22]. Ann [Kea22]. Anne [Clo18]. Annette [Cle18]. anniversary [GAB^{+}18]. annotate [PHZF16]. annotated [AST^{+}22]. Annotating [HT14]. annotation [CML^{+}15, PMGB^{+}21, Sch18]. annotations [HT14]. annual [QSH14, Bea17]. Anonymity [CS15b, NB17]. Anonymous [See17, Tei16]. anonymously [Ior14]. answer [SMRR16, Wes15]. answerers [LS18]. answering [BK19, LCSL15, LJ17, ZLZG14, LS18]. answers [DMR^{+}17, FV21, ZCZ^{+}20]. ante [MBD15]. Antecedents [O’B17, RKL14, LHM17, NGTKR15]. Anthony [Hjø17b, Kea22, Wil14b]. anticipation [LJI18]. Antonio [Kos21]. Anxiety [JV17]. anxious [GKM^{+}21]. Anyone [Whi17]. Aparac [Cho19]. Aparac-Jelusić [Cho19]. API [LSX18]. app [Shm18, XDDG18]. applicants [VCG14]. Application [MKK17, BTMS16, KGS19, LKKS17, LBG^{+}21, PDS22, PGFG^{+}20, Sav15b, STF19, YCK16]. applications [FvdWR^{+}22, Hof17, MS16, MdSLdO^{+}23, TSRG16, WL14a, YHL^{+}21]. applied [Haj20, KLG14, WHvdZvdB17]. Applying [BR20, Cox23, Lai15, KT14a, XJBK17]. appraising [CCJ16]. Approach [KMC20, Rog15, ADC15, Ama22, AGM19, BM14c, Bor16a, BKM15, CCM23, CCLL14, CDGN15, Col15, CCJ16, CTM16, DP23, DMR^{+}17, DBBJ16, Esk20, GNVQdMA16, GKM^{+}21, GB19b, GS18, HPLR23, Hal17, HM18, HWW^{+}17, JKS^{+}16, JL23, KDS^{+}17, KD15, KCK16, Kue23, LL17, LS17, LYWF14, LML16, LK16, LI22, LM15b, LJLH17, MRAF16, MBML24, MNNK15, MRMCMVUL14, MRKT15, dSMdAGQ15, MBD15, PKK^{+}24, Riv15, RPK16, SAHZ21, SBIT15, SK23, SJK15, SH14, TJ20, TDOF^{+}16, VN18, VT16, WL20, WHvdZvdB17, WVZ20, WLZ^{+}20, WLG23, YH16, YR23, ZW23, ZvSvB14, Car14b, Cha14]. Approaches [Sav18a, Sav18b, YZ17, Ano18-27, CAS21, CV20, HL20, LML18, MDJ^{+}19, dLPMR16, PSB16, SC17, SAW21, SFL^{+}23, YY16]. appropriately [ØC19]. approximately [Rou16]. APPs [MS19, WZL21, WXLBO15, ZJHT23]. Arabic [KF16]. Arafat [Bat21]. Aral [Afz22]. Archaeology [Mak14, VT16]. archeology [DDB16]. archeology-specific [DDB16]. archetypes [LB20]. Architects [Far23]. architecture [AFB20, BOHK19, Bur14]. architectures [Swa22]. Archival [Clo18, DC15, Far23, FT16, FHO23, SS14b, Skö18]. Archived [Web20]. Archives [Han21, Ack21, BSG19, CC22, ESD22, HHL21, KPWS20, MWMB18]. Archiving [Cus15, DM14, LOD^{+}19]. area [JSY^{+}16, PGFG^{+}20, Poo17]. areas [WSZC22]. Argue [DS15]. argument [Bor14a, IQDS16]. argumentation [WSZC22]. Arp [Fri17]. arrangement [MMHA23, TFR15]. arrangements [GBS15]. Art [CDL18, Daq20, DM14, HSK15, KLG14]. Article [Ior14, BDH17, CK14, DTW16, JML22, KF16, PSB16, Tei17, TKA^{+}23]. article-level [JML22]. article-processing [PSB16]. articles [AGGTGL18, ALCD14, BAGL16, CHH14, CKR14, DBH18, GRSFV15d, GSG20, HCKH15, HLW18, Hen14, HR19, KT17b, KT17a, KISS15, KAL24, LR14, MT18a, McC15, MPB^{+}20, MTHL15, MRKT15, NBH16, ONZ23, OY15, OVE14, RPK16, SR15, SR16, TM15, TW16b, TK17b, TJV14, YCL20]. artifacts [MSM^{+}21]. Artificial [JLB^{+}23, LWM^{+}23, ØJW^{+}21, PPC19, WLS23, Cox23, LGZ14, PdlRR^{+}22, SKO24, SHKL23, SK23, SFG^{+}23]. Artists [Gor20b]. arts [Har14]. arts-informed [Har14]. artworks [CPP^{+}16]. ARWU [DD16, ZMJ14]. arXiv [Ior14, LSM^{+}14, MGK20]. asexual [KWLV22]. Ashley [O’H18]. Ashoori [Bat21]. asi.20981 [Ano14]. ASK [LJ17]. asker [ZPL^{+}20]. askers [Kan22]. asking [CS16b, RBJ18a].

Aspect [LZG^{+}23]. aspects [Vec14, XDDG18]. Assembling [Kha19]. assess [DS15, HB14, KT16b, KTA17, McC15]. Assessing [Hof17, IQDS16, JSW14, LCB^{+}18, MHMW17, RDPF16, ST22, WL20, WG19, DGCC17, Daq20, Gia18, KT16a, VCG14]. Assessment [BMM20, WW15, XLL16, AD15, AP20, BC16, BM14a, CS15a, Cho15, Cho20, GRC21, JGY^{+}19, JMV19, MMB17, MH15, RN16, SKC18, TR17, YZ17, ZGBIL17, Zub18]. assessments [ZBBW18]. assets [DAT^{+}14]. assigned [ST16c]. assignment [BRL20, RMDF23, XDSC16]. assignments [STAH14]. assist [GN17]. assisted [BST15, DP23]. Assisting [DC22, LCWC23]. associated [XBCH18, XBL^{+}21]. association [NF15]. Associations [CPL22, Par19, TM20]. associative [MBAG16]. assumptions [vdB16]. assurance [PMSA14]. asteroid [SS14c]. astronomy [CH16, DSBG20, DSBG21]. astrophysics [KH17]. asylum [LI22]. asymmetrical [Epi14]. asymmetry [RBJ22]. asynchronous [TZS^{+}22]. at-risk [QH20]. athletes [Mir14]. Atlas [Whi17]. attachment [SLM15, ZBDX18]. attempt [BH17b]. attention [AMDFHVC24, LZZ^{+}21, SMRR16, SR15, SZJ^{+}21]. attenuation [LF21]. Attitudes [CGGS17, CJ22, KCK16, RJS^{+}17b, WB21, HT14]. attract [TS21]. attribute [PMGB^{+}21, SLLL17, YZYR21]. attributes [GTZŽ14, SWKK16, wYKY17]. attribution [CMS23, HB14, HKBB18, JMV19, Sav16a, Sta18]. attributions [Daq20]. atypical [HLW18]. Aubrey [FJ19]. Audience [GS22, ØC19, PYH16, TK17b]. Audience-group-based [GS22]. audio [STCS17]. auditing [Rou17]. auditory [AG15a]. Audunson [Han21]. augment [LSX18]. augmentation [XCL^{+}21]. Augmenting [Hal20]. Austin [Hjø17a]. Australia [Clo18, HH18]. Austrian [MBD15]. Authentic [WS22, KKZS20]. authenticity [BCK17]. Author [GG16, Hen19, Hsi17, Kim19a, KMC20, LDK^{+}14, LC21, NE17, XDSC16, YCL20, AHJY18, Cri15, Don16, FVGL14, Gia16, KGS19, Kim19b, KK20, KKOS21, KW14, LK14, ML16, OA14, PGFG^{+}20, PS19, SGLF17, SWKK16, The17, ZYZ^{+}16, ZLY23, ZS14c, ZS14b, ZS16, ZSC18]. Author-based [Kim19a]. authorial [KS14a]. authoritative [CK16a, DMR^{+}17]. authoritativeness [Daq20]. Authority [RVGPS19, RVdAGP21, TLA^{+}16, KLC^{+}14]. Authors [BLB16, AKK^{+}17, GG16, Mil14, WCP^{+}19, YRH16, Ben18]. Authorship [Sav19, AMM^{+}19, CMS23, JMV19, KS17, Sav16a, Sta18, SWKK16, TCF^{+}24]. Autobiography [The17]. autocomplete [BS16]. Automated [AMM^{+}19, HL20, KF16, Ano18-27, HBH^{+}16, LML18, San22, SKO24]. Automatic [BDA17, RVGPS19, SBZ^{+}14, GSB^{+}16, KBD^{+}14, KT15, KT16a, LM15b, RVdAGP21, SRG^{+}16, Vec14, PdlRR^{+}22, WL14a]. Automatically [ALCD14, DMR^{+}17, Sil17]. Automating [SHKL23]. autonomy [RY18, San22]. avoidance [CV20]. awakened [LS16]. award [DD16]. aware [GGS22, HWZ22, LX17, LSC18, SVK17, SKKM22, SLCS18, YH16]. awareness [LS17, PKK^{+}24, SBS23, Xia14].

B [Lev20, Sko23]. B. [Mat20]. backbone [JL23]. backer [DLD^{+}21]. backpackers [HHPI21]. Backstage [Ano14, SMAB09]. backup [KMD16]. bacterial [MC18]. bad [MSK^{+}14]. badge [YHLG21]. Badia [Kos21]. bag [DLA14, DGGdST19]. bag-of-words [DLA14]. Baidu [VC15]. Balkans [YZQT20]. Baran [QH18]. Barcelona [MMFA20]. barriers [CKL16, KL22, WKJW16]. Barry [Fri17, Lin14b]. Barth [Ama19]. base [HL16, LWPG^{+}19, ZS14c, AGGTGL18]. Based [Cox14, GMB14, KMC20, MHJ19, MS19, NTNP19, ADD^{+}18, ACS18, AHJY18, AGMR18, Ama22, Ano17m, AGM19, ATKD^{+}18, Bar15, Bea16, BM15b, Bor16a, BH16, Bor17, BKM15, CWdR17, CCC14, CK14, Che16, CZ16, CS15b, CDGN15, CBAL15, Col15, CML^{+}15, DLA22, DZC^{+}14, DGGdST19, DGR19, DLD^{+}21, DS21, EBD15, ECJF21, EHdBvO17, FT17, FLJD23, FdGPC18, GS22, Göd16, GLR16, GLZ^{+}24, GBZ^{+}23, HB21, HRN^{+}23, HT17, JL23, KDS^{+}17, KD16, Kim19a, KKOS21, Lai15, LJK21, LBM19a, LBA16, LXLL14, LCD^{+}14, LYWF14, LLHF15, MS16, MRAF16, MDJ^{+}19, MŽA17, MKK17, MRMCMVUL14, MTY14, NSK17, Pal16, PHH^{+}20, dLPMR16, PMGB^{+}21, PS19, RN16, RPK16, RPK17, RH21, Sav18a, SB18, STAH14, TZK20, TWB^{+}18, TC17, THC15, VG18, VCG14, VT16, WW15, WSWH20, WMM^{+}23, WFRN14, Xia23, XDDG18, YY16, YC20, ZZLS16, ZZL^{+}14, ZPL^{+}20]. based [ZYS17, ZHGY18, dHSM17, BL14, Ano17n]. bases [DDB16, LG14]. Basic [Fri17, MAM19]. Basis [SV20, LdMAGB15]. Bates [Soe18]. batik [YXJ18]. battleground [ML22]. Bayesian [IABL16]. Bazaar [ALAB19]. BBC [KT17b]. be [AT14, BCK17, LdN17, Lue20, OR15, PN18, WZL21]. beam [BM14a]. beauties [LS16, YB18]. beauty [MA18b]. because [Bor17]. becomes [vdSvdBG^{+}14]. becoming [ZGE22, Dor19]. Bedford [Swa22]. before [VO23]. Behavior [CWdR17, Fou19, MHJ19, MCSB19, OGL19, AGBK19, AP17, ADL15, AGJ^{+}16, BR20, BCAL15, BS16, BM14d, BBC^{+}22, BV21, Che22, CLL15, CS16c, CPL22, CL17, FN17a, Fre15, GTAS22, GJ16, HEEB^{+}22, JWJ^{+}19, KvOS^{+}17, Kel21, KTS^{+}15, KTPE14, KSL19, KIMM17, LWA^{+}17, LOM22, LCLM18, lLTBH15, LMM^{+}22, Lue15, MVL^{+}14, NLP16, RJS19, RTS^{+}17, RJS17a, Rut22, STS14, Sav17, SL21a, SZZ16, SLS^{+}19, STF19, SH19, TT17, VVP^{+}21, VSJR19, WOJBG17, WJB22, WFRN14, XZ19, YHLG21, ZFH^{+}19, Zim17, Col17b]. Behavior-based [CWdR17]. behavioral [KT14a, MBVB17, Zha14c]. Behaviors [MS19, BJR19, CMB17, CWL^{+}23, FSPC16, GBS15, KS16, KY17, KCK16, LJT^{+}15, LCK16, LHXS22, OB19, RJS^{+}17b, SYF17, TJ20, WS22, WDTC20, ZLCB15, ZY15]. Behaviour [Col17a]. behind [DQAMM^{+}24, Mil14, MD15]. beholders [OMM^{+}14]. being [BKM15, EH15, EAB17, EA18, JAG^{+}20, LPL14, SH14, ZJHT23]. Belief [Whi14, Bi22, Kel21]. believing [OHP16]. Belmont [Xia23]. belonging [CY22]. benchmark [MSI^{+}19, ZRdR21]. Benchmarks [TCP15]. beneficial [ST16b]. benefit [MK15, VXN18, WTT^{+}14, ZPL^{+}20]. benefits [Kon16, RDS20, ZS14b]. Benford [MA18b]. Bergman [Jon17]. Berkeley [Soe18]. Berlin [Bud17, Fra21, Han21, QH18, vR17a]. Berners [MSdGS19].

Berners-Lee [MSdGS19]. Bernhard [Sha21b]. best [BdMA14, JdR17, Rou16, YL18, ZWW18]. Better [CBvEW18, Hil14a, KTS^{+}16, LYC^{+}22, Mah22, vdVZK21]. Between [RKHMRH21, ACS18, ALAB19, ATKD^{+}17, Bar15, BCZ14, CZZY15, Che16, Che22, CPL22, DBH18, DTW16, Eng17b, Epi14, EWX^{+}16, FJF^{+}21, FV21, FSPC16, GKM^{+}21, Hil14b, HCC15, HQH^{+}22, HLC15, JT17, JA16, JWJ^{+}19, KR15, LRB19, LRB21, MMHA23, MBML24, Men21, Mil14, MK15, Min16, POK15, RPK16, RPK17, Tab14, TM20, TS21, WMS^{+}15, WSvPK16, WXLBO15, XBL^{+}21, XZW^{+}19, YRWH18, Zem17, ZRG16]. Beyond [DLA14, DB17a, Hof17, TZK20, Zha14b, BOR^{+}23, GBL^{+}18, HHC^{+}14, MBVB17, PSB17, WBH18, Wor19, vdSvdB23, O’C16]. Bias [GRSFV15b, LML23, SWSZ19, JSY^{+}16, Kos21, LCH^{+}22, SCIF20, SDT22, VZRP15]. biased [BST15, YHS16]. Biases [NLFM19, PKK^{+}24, ZCvdH15, ZCYN23]. Biber [Bow19a]. Bibliogifts [Cab16]. bibliographic [DC22, LRB19, LLHF15, MŽA17, SAW21, ZS14c, ZYS17, ZQC^{+}23]. bibliographies [DC22]. biblioleaks [Cab16]. Bibliometric [Dor19, NF15, WVvLZ23, AGMR18, BWL15, BM15b, BH17b, FW19, HH18, HGCKDR16, HK14, IABL16, Kra17, LKKS17, LCRGTS14, MB14, MCDMM14, Pet19, Pra14a, Pra14d, Pra15a, Pra15b, RJGM16, SKZ^{+}14, TCF^{+}24, TSRG16, VCG14, Wan18, WHvdZvdB17, ZGCB15, vLTW18]. Bibliometric-enhanced [WVvLZ23]. Bibliometrics [O’C16, BTMS16, CJ16b, LMP21, LB16, Riv15, ZL16, Zit15, BI17]. bibliométrie [Zit15]. Bielby [Uns18]. Big [EMK^{+}15, Fri15, GBZ^{+}23, LWLH18, MF16, Söd23, Ten16]. bigger [EMK^{+}15]. Bigram [DLA14]. Bigram-enhanced [DLA14]. bilingual [ECJF21, LN19, MMFD15]. Bill [FM14]. bind [Cro15a]. Bingley [Col17b, Mat20]. Bio [CKL16]. Bio- [CKL16]. biochemical [ST16b]. biodiversity [LGD20]. biographies [OR15]. bioinformatics [SKZ^{+}14, SKL17, YRWH18]. biology [HB16]. Biomedical [TCP15, AH17, Ano18z, CSOD16, DCLH21, FT17, HPS^{+}14, Hod21, KBS18, LM15b, MWPF17, PYW18, WV15, ZTA20]. biomedicine [HPS^{+}14]. biotope [MC18]. birth [GOdR18, Kno24]. birtherism [DL22]. bisexual [KWLV22]. bit [TTH16]. Bitly [FCT^{+}21]. Black [KP18, TZH23, LMA^{+}22]. black-boxes [TZH23]. Blair [Kea22]. Blaise [O’C16, Rou17]. Blanchette [SR14]. blind [SDT22, XBCH18, XBL^{+}21]. blindness [Coc22]. Blocked [OHT^{+}16]. blog [SBIT14]. blogged [SBIT15]. blogs [SBIT14]. bodies [VP19]. Boellstorff [Hak14]. Bohannon [H17]. bon [Zit15]. Bonde [Con20]. Bonnie [Hak14, Lan15]. Book [Abb17, Adl23, Afz22, Agu18, Aji17, All19, Ama19, BI17, Bat15, Bat21, Baw14, Bea17, Ben18, Bor14a, Bow19a, Bow19b, Bru18, Bru14, Buc20, Buc15, Bud17, Car14b, Cha14, Cho19, Cle18, Clo18, Col17a, Col17b, Con20, Cox14, Cox15, Cus15, Dor19, Dow15, Dut20, Erd18, Far23, FM14, FJ19, Fou19, Fra21, Fre18, Fri17, Fro21, Gor20a, Gui17, Hai16, Hak14, Hal19, Han21, Haw18, Hay18, He18, Hjø17a, Hol15, How17, Huv20, Huv22a, Jon17, Jul20, Kea22, Kno24, Kos20, Kos21, Lan15, Lev15, Lev20, Lin14b, Ma22b, Ma22a, Mar15, Mat20, MPB^{+}20, Med15, Meh22, MS19, O’C16, O’H18, Ols16, OM15, Owe21, Pan21, QH18, Rey21, Rhe17, Rog15, Rou17, Sal20, San22, Sar16, Sav18b, Saw19, SK14a, See17, Sha21a, Sha21b, Sko23]. Book [Soe18, SR14, Sto20, Sve15, Swa22, Ten16, The17, Uns18, War15, Web20, Wei14a, Wei14b, Whi17, Zha19, Zim17, Zit15, vR17a, AP17, Bur20, Col20, OVE14, TK17a, WMX^{+}22, WLZ^{+}20, ZvSvB14, ZGCB15]. Booklover [CJ16a]. bookmark [GB16]. bookmarks [MTK16]. books [AT14, DPN16, KT16a, KT16b, KTA17, KT17b, KT17a, PN18, ZK14, ZBBW18, AT14, KT15]. Boolean [Hjø15c]. boost [HB15, ŠF17]. boosted [SMYL15]. Boosting [SLLL17]. Border [Ano17n, CZ16]. borders [BNG14]. Borgman [Ten16]. born/Korean [LS22]. born/US [LS22]. Börner [Whi17]. Bose [Jan15]. Boston [Bat21, dP20]. bot [HBH^{+}16]. BoTG [DGGdST19]. both [Jam17]. bottom [EBD15]. bottom-up [EBD15]. boundaries [EWX^{+}16, KK19b, Wor19]. Boundary [HAJ^{+}17]. bounded [McC15]. Bowker [Mat20]. box [LCWC23, WMS^{+}15]. boxes [TZH23]. Bradford [HHC^{+}14]. brain [HBS15, WBH18]. Brazil [FvdWR^{+}22, MdSLdO^{+}23, RA14, SHMC17]. Brazilian [MCDMM14]. Brea [Tho22]. Breach [JVF19, Jam20]. breakthrough [SC17]. breakthroughs [WMM^{+}23, WTvR16]. breast [Rub15]. BRICS [BWL15, VR17b]. bridge [PdlRR^{+}22]. Bridging [BCZ14, JA16, OR21, ZZY23]. Brief [Bar15, BM14a, BM14c, BM14b, BWL15, BM15a, BB15, CKR14, Cri15, DD16, Egg14b, Epi14, Fia14, Fur20, Hal14, HCKH15, HBH^{+}16, Hil14a, Hjø15b, Jan15, KISS15, LPL14, LHR16, Mal15, MD15, NF15, NLP16, OR15, OK16, Pra14b, Pra14a, Pra14c, RSAS14, RL15a, RD16, Sch15, SF15, SWKK16, Car14a]. Broadcasting [ZDZ21]. broken [RSAS14]. Brown [Bud17]. Browsing [Bat20, MCSB19, HT17, HS22, LJT^{+}15]. Bruce [Sko23]. Brügger [Web20]. brutal [Sha21a]. budget [WMS^{+}15]. Building [Fri17, TKB^{+}17, BBC^{+}22, GAS^{+}20, LPB19]. built [WKC14]. bullet [LZG^{+}23]. Bunz [Saw19]. burden [CK22]. Burdens [SR14]. Burgess [Buc20]. Burke [Lev20, Lev15]. bursting [KYFZ15]. Bushfires [CN14]. Business [JPF16, AGM19, Vec14, YP16]. businesses [Kno24]. button [GSG20]. Byström [Huv20].

C [Jan15, DC14, SB16]. C-sanitized [SB16]. C-word [DC14]. CA [Agu18, Cho19, Fou19, Hay18, MS19, Soe18]. Cabinet [Huv22a]. Cady [Bru18]. Cailin [Kos20]. Calculating [LCD^{+}14]. calculus [FvdWR^{+}22, LI14a, LJI18, MK15]. Call [Kra17, Fro21, St.17, XHM^{+}20]. Cambridge [Adl23, BI17, Ben18, Bow19a, Bow19b, Cle18, Con20, Dor19, Erd18, Fri17, Hal19, Haw18, How17, Jon17, Kos21, Ols16, Rey21, Rhe17, Sal20, San22, Sav18b, Sha21a, Web20, Whi17, Zha19, Zim17]. campuses [SLeH21]. Can [AT14, GKM^{+}21, KT16b, LdN17, MTK16, Sto20, WZL21, Whi17, ZGCB15, HB15, Kit19, ZMJ14]. Canada [AC18, QHH20]. cancer [EAB17, RHG23, Rub15]. candidate [BYSB16]. cannot [Kos21]. can’t [MT20]. Capabilities [Cha14]. capability [IQDS16, SDUDL24]. capacities [Hil14b]. capacity [ST16a]. capital [AMDFHVC24, CLSS14, CBH16, dSMdAGQ15, ZH17]. Capitalism [Meh22]. capture [RLAB^{+}24]. Capturing [GISW15]. Capurro [Uns18]. car [LO19]. Carbo [Car14a]. carbon [RLAB^{+}24]. Cards [Lev15]. career [BDX^{+}18, XBL^{+}22, ZZ21b]. careers [ML16]. careful [SM18]. Carnot [BR15b]. cascade [MCY^{+}21]. Case [Ama19, Col17b, OGL19, Tho22, BS16, CY22, ESD22, FJT23, Gia18, GAS^{+}20, HT17, Hjø15c, HR19, Ior14, JT17, Kan22, KIH16, LSHH23, LR14, MGIGW16, NBH16, PPLJ20, Riv15, Sav15a, ZYZ^{+}16, ZGB16, ZGEJ17, dHSM17]. case-based [HT17, dHSM17]. cases [LB16]. Cassidy [O’C16, Rou17, vR17a]. Castro [San22]. catalog [WCC^{+}17]. cataloging [CJ16a]. catalysts [LBC^{+}22]. Categorical [ZGBIL18]. categories [BOR^{+}23, FV21, LB16, MTHL15, TS21]. categorization [AFL^{+}14, HRN^{+}23, KLG14, Oh17]. categorizing [SSC17]. category [CHLRC14]. Catelijne [Rhe17]. cathedral [ALAB19]. causal [FJF^{+}21]. cause [CK21a, NILCP15]. cause-effect [CK21a]. caution [Kra17]. cave [HNX18]. caveats [LSX18, THL17]. Celia [Hak14]. Censorship [LZ21, Olt15, SYS17]. Center [CPL22]. Centered [HNX18, AFB20, KWLV22, LSC18, SAHZ21, She19, WXSL23]. central [SKZ^{+}14]. centrality [Fie16, HLM16]. centric [PSB17]. centuries [PMGB^{+}21]. century [Kon16, SWKK16, dP20]. cessation [ZY15]. chained [ACBC22]. chains [TC17]. challenge [HLB^{+}17, Sal20]. Challenges [Aba24, Bud17, Cho19, YC20, GVB17, SSC17, SS14a]. Cham [All19, Ama19, Fre18, Gui17, He18, Jul20]. Chandos [Dor19, Erd18, Sav18b]. Change [MBZ^{+}22, Bor16a, Cho23, Cox23, HM14, HK21, IESG20, RLAB^{+}24, SLE21, YCL20, ZGBIL17, Bor14a, HL16]. changes [OHP16, SRYL17, WLS23, Cox14]. Changing [Cle18, LG14, OHP16, WTvR16]. channel [Men21, SZ20]. channels [CHH^{+}17]. Chaos [BK23, SBS23]. chapters [OVE14]. Characteristics [CS16c, AKK^{+}17, AP17, HYD^{+}20, KCK16, LZ22, Par19, PSB^{+}14, TJ20, TPR17, WLC^{+}14]. characterization [LXLL14]. Characterize [MH18]. Characterizing [CWL^{+}23, SAB^{+}18]. Characters [ZZL20]. charges [PSB16]. charging [MBZ^{+}22]. Charles [Jul20]. Charting [ACBC22]. chat [KTZ20]. chatbots [WLS23]. ChatGPT [LWM^{+}23]. chats [MYK15, Zem17]. Chatting [TGV^{+}16]. checking [HE22]. Chemie [MWD17]. chemistry [Mil14, NE17]. Chen [Sko23]. Chicago [Buc20]. chicken [RYEP18]. child [MD16]. child-tablet [MD16]. Children [Bow19b, LYWF14, THWS14, WMX^{+}22]. chimeras [HR19]. China [Sko23, CW18, KHZ^{+}22, PYH16, TPZ15, WHZ23, ZZY23, ZJHT23]. Chinese [HS14, HY17, HCC15, HCS21, LZ21, Ma23b, SJL19, TSY15, YRH16, YRWH18, YZ16]. Chinese-speaking [HCC15]. Chirag [Fre18]. Choice [Cha14, BJSM18]. choices [SL21b, ZK14]. Christian [Zim17]. Christine [Ten16]. Christoph [Han21]. Christopher [Sko23]. chronic [BV21]. CI [SSA22]. Cicilie [Han21]. Cinderella [Fra21]. circles [Mil14]. circumstances [BJR19]. circumvention [SYS17]. Ciro [Mat20]. citable [Hen14]. Citation [Ano17m, Bel16, BH17b, FN17a, GMB14, HK14, KB17, LPW14, LBW19a, LR14, MM18, MCY^{+}21, dLPMR16, Rou16, ACS18, APRRC15, AKK^{+}17, BM14c, BM15a, Bor16a, Bor17, CCLL14, CDL22, Col15, CL17, CHLRC14, CWL^{+}23, DZC^{+}14, Don16, DTW16, Epi14, FJT23, FW17, FN17b, GRSFV15d, GNVQdMA16, HLW18, HPLC19, HR19, HHC^{+}14, HQH^{+}22, JdR17, JLC16, JZC20, JL23, KBS18, KT17c, LdMAdN16, LdN17, LCD^{+}14, MT16, MPB^{+}20, MZ22, MH16, MWD17, OSvE16, OY15, PYW18, PRRC16, Riv15, RGJCTS16, RPK16, RPK17, RD16, Rou17, SAA^{+}23, SC17, Sil18, SKA18, ŠF17, ST16b, TSY15, TW16a, TM20, The20, VG18, WL14a, WL14b, YY16, YCL20, YB18, ZDW18, ZS14b, ZS16]. Citation-Based [GMB14, Ano17m, dLPMR16, RPK16, RPK17]. citation-rank [BM14c]. Citations [LCHL14, MT18b, NE17, WC14, BL14, Bor16a, CZZY15, CL17, CZW15, CWL^{+}23, DBH18, Fan18, HPS^{+}14, JLC16, KT15, KT17b, KT17a, KAL24, KH17, LPL14, LB19a, LLHF15, MDL^{+}18, OMTK17, OA14, PM18, SBIT14, Sil17, SLM15, SLE21, TM16, TK17c, TS21, TKS^{+}23, WL14a, WVvLZ23, ZYL18, ZSC18, ZCYN23, ZTLV15]. cite [Sil17, XLKC22]. cited [AKK^{+}17, BLB16, BWL15, BB15, BM15b, BBH15, DT18, Fan18, HK14, HR19, LPL14, Mil14, PS14b, RN16, ZCW17, ZWW18]. CiteSpace [PHC17]. Cities [dP20, LHR16, MMFA20]. Citing [XLKC22, BM14b, Fan18, GG16, HBDL19, LCD^{+}14, Mil14, SRYL17, WW15]. Citizen [NTNP19, Haj20, LCNP17, WKJW16]. city [HPLR23, MVA22, SVB22]. Ciufolini [Ior14]. civil [Jam20]. claims [BL14]. clarification [TZS^{+}22]. class [Ko15, KSL19, Pra14a]. classes [BOR^{+}23]. Classical [Hjø15c]. Classification [Ort22, AFL^{+}14, ATKD^{+}17, BK19, Bi22, EBD15, GSB^{+}16, GNVQdMA16, GB19b, HCD17, KF16, Ko15, LRB19, LLB14, MR14, MRMCMVUL14, PRRC18, PdlRR^{+}22, SAA^{+}23, SRG^{+}16, ST16c, TZH23, TGV^{+}16, VAGR15, WKC14, ZSMF15]. classifications [SAW21]. classifier [OR15]. Classifying [GB16, KK19a, KMFD15, PDS22, wYKY17, AKAS14, ZRdR21]. classroom [Kon16]. Claudio [Gui17]. Claypool [Agu18, Cho19, Fou19, MS19]. Clayton [Clo18]. CLEF [TCP15]. CLEU [MSI^{+}19]. Click [STF19, GBL^{+}18, FCT^{+}21]. clicked [FCT^{+}21]. Climate [Cho23, RLAB^{+}24]. Climatic [HL16]. Clinical [KDS^{+}17, KZB17, SYG17, TM16, vdVZK21]. Clinton [San22]. Cloonan [Sal20]. Close [LBM^{+}19b, WLS23]. Closer [CBvEW18]. closing [GZBA22]. closure [KZ23]. cloud [LWA^{+}17, Ric18]. clouds [LZD16]. club [Ort22, RR21]. clubbing [TSY15]. clubs [BPB21]. cluster [GB19b, YZ16]. Clustering [DMMF17, CPP^{+}16, GSG22, GNVQdMA16, GLZ^{+}24, LLHF15, MMFD15, MGKA16, Sav15b, ZHGY18]. clustering-based [GLZ^{+}24, ZHGY18]. Clusterization [GRBCS15]. clusters [Che16, GK14]. CNN [FR23]. Co [DT18, Hoo17, LN17, LZA^{+}20, BSVZ17, ML16, MGKA16, ØJW^{+}21, PHH^{+}20, PHJ23, SJK15, TCF^{+}24, VN18, YLL20, Yu14, ZL16]. co-author [ML16]. co-cited [DT18]. co-comment [SJK15]. co-constitutive [ØJW^{+}21]. Co-contributorship [LZA^{+}20]. co-corresponding [TCF^{+}24]. co-creation [PHJ23]. co-embedded [MGKA16]. co-link [VN18]. Co-Occurrence [Hoo17, ZL16]. co-occurrence-based [PHH^{+}20]. Co-saved [DT18]. co-training [Yu14]. co-tweeted [DT18]. co-utility [BSVZ17]. Co-word [LN17]. coauthors [KD15]. Coauthorship [KD15, OVE14, Fie16, GK14, HWW15, KD16, LPW14, MCDMM14, RPK17, XOL^{+}18, ZQC^{+}23]. Cocitation [CBvEW18, DC14, MP23, ZS14c]. Cocluster [JJC19]. cocreated [FC18]. cocreation [PPLJ20]. Code [Mar15]. Codes [FHW17]. codification [LJI18]. Coding [Mar15, ZvSvB14, Cle18]. coefficients [ZL16]. cognition [Che22, LCWC23, MR17]. Cognitive [KvOS^{+}17, SP22, ZS20, KTZ20, KTPE14, KKZS20, LG14, ZS14a]. cognitively [AMOM16]. cohesion [DMdO16, GK14]. cold [ACL17, MBAG16, Joh17]. Cole [Jul20]. Colin [Lev15, Lev20]. Collaboration [LBW19a, RCM^{+}17a, Bor17, BDLM18, CH16, CYP15, CKL16, GK14, GK24, GISW15, HWZ22, HWW15, IQDS16, Kim19b, KMFD15, LBLL24, PTGL17, PS14b, RPK16, RD16, SHMC17, SC21, ST16b, TM20, TGC^{+}21, Tsi16, XBL^{+}22, YZQT20, ZBDX18, ZGEJ17, LGST15]. collaborations [HWZ20, LI21a, LZA^{+}20, ML16, She19]. Collaborative [BWGD21, HABT18, Sha14, ASTU16, ADL15, BSVZ17, CHH14, GEBP22, HSU16, Kha19, LK14, MR17, TT17, ZHGY18]. collaboratively [DGCC17]. collaborators [BDX^{+}18, TT17]. Collabportunities [GISW15]. Collecting [ZS19]. collection [TFR15]. collections [ABLS17, CPP^{+}16, CS16c, JKS^{+}16, LYL21, LPB19, LK14, MdSLdO^{+}23, SS14b, SF15, TK17d, WLW^{+}17]. Collective [PS20, GOdR18, Luy16, SJT^{+}21, SF20, TPZ15]. College [KP18, KCL14]. collide [RAW22]. Color [XZ19]. Columbia [Ma22b, Ma22a, Sko23]. combat [JSY^{+}16]. Combination [KMC20, FdGPC18, GNVQdMA16, MDJ^{+}19, PSB16]. combined [CLA^{+}16, LL17, SJK15]. Combining [FHO23, KKO15, LKY^{+}20, WVvLZ23, AKC^{+}18, Bi22, GHKB^{+}17, dLPMR16, QH20]. Command [See17]. Comment [She21, LMJ21b, SJK15]. Comments [PS20, LZG^{+}23, Ort22, TBP^{+}15]. commentsphere [SYS17]. commerce [XZ19]. commercial [LP16, WHX^{+}14]. commissioned [Hen14]. common [ADD^{+}18, CL17, PYW18]. commons [CJ22, EJ14, RH21]. commons-based [RH21]. communal [HDT23]. Communicating [SBS23, TK15]. Communication [BM14a, BM14c, BM14b, BWL15, Bud17, Car14b, CKR14, Egg14b, Epi14, Fia14, Hal14, HCKH15, Hil14a, Jan15, LPL14, LHR16, MD15, NF15, OR15, OGL19, Pra14b, Pra14a, Pra14c, RSAS14, RD16, CY14, CKL16, Esk20, FA17, H17, Hil14b, HHB17, KT14a, LI14a, PYH16, Pet19, TKB^{+}17, Tsi16, WG19, ZZY23, vR17a]. Communications [BB15, Dut20, Hjø15b, KISS15, RL15a, SF15, Bar15, BM15a, Cri15, DD16, HBH^{+}16, Mal15, MA18a, NLP16, OK16, Sch15, SWKK16]. communicative [DOM^{+}14, Sny14]. communities [CW18, CLL15, FHO23, HLC15, IESG20, JW15, KWLV22, LYC^{+}22, MT20, NT17a, ØC19, QLM^{+}22, RCM^{+}17a, She19, VO23, WHWZ21, WCO21, YRWH18, ZLZG14, Zha16, ZB19]. Community [CCC14, DJLS22, DM14, GB19a, GAS^{+}20, BK19, CH16, CY14, HYD^{+}20, KZ23, KLC^{+}14, KWLV22, LMA^{+}22, LCWL18, NBH16, PPLJ20, PH22, QZWC21, Rub15, SJK15, SHW^{+}15, Wor19, YC20, ZLZG14, ZLF20, ZWTX23, LS18, Mat20]. community-based [YC20]. community-inclusive [YC20]. Community-led [DJLS22]. companies [CAS21, Bow19b]. Companion [Kea22]. Comparative [LKKS17, OA14, VGC17, JMV19, LS22, Mal15, OS15, POK15, TT17, WSZC22, YRWH18]. compare [Hal14, PLPQB20]. compared [LdMAdN16]. Comparing [BMG^{+}17, DGR19, dLPMR16, ZYZ^{+}16, EHdBvO17, HM18, RB14]. Comparison [CK16a, Sav18a, ST16c, WC14, XBL^{+}21, Ano18-27, AFL^{+}14, AGJ^{+}16, Bar15, BM15a, BVH15, CK21a, CZW15, DC14, GIEVVGS19, HLC15, JDHL17, LN17, LML18, OSvE16, RN16, TTH16, TS21, VC15, XDSC16]. comparisons [LHR16]. compatibility [Ma23b]. competencies [Cox23]. competition [Hsi17, PTGL17]. competitive [JJ14]. complement [RHG23]. Complementary [LCSL15, AGMR18, Don16]. complete [Cho15]. completion [Gre16]. Complex [Bor15c, Gre18, WCTS15]. complexity [BR15b, HK17, LBW^{+}19b, OHP16, SJ17, SW18, TFR15, WOJBG17, WCTS15, vdSvdBG^{+}14]. compliance [Bor16c]. compliments [SDUDL24]. component [FS18]. composite [DBBJ16]. comprehensibility [vdSvdB23]. comprehension [KWR20]. comprehensive [AFL^{+}14, ZS14a]. compression [FDR16]. computation [PTGL17]. Computational [FHW17, SWKK16, Daq20, GHKB^{+}17, JSY^{+}16, RJGM16, RDPF16]. Computer [Cle18, CY14, FA17, HHB17, IQDS16, Kim19a, MD16, NBH16, OBD^{+}22, OHP16, SCIF20, SDT22, VS15]. computer-mediated [CY14, FA17, HHB17]. computer-supported [OHP16]. computers [BNG14, Kos21]. Computing [SK14a, RC16]. concentrated [BC16]. concentration [Bar15, Egg14b, KAL24, Pra15a, Rou15]. concept [BR20, CL17, Kue23, LS21, LBA16, Ma23b, SH14, VJH15]. concept-based [LBA16]. concept-symbols [CL17]. conceptions [DC15, Skö18]. Concepts [Agu18, Fou19, MS19, ADD^{+}18, Bar14, Kel21, MR14, McC15, MDC^{+}16]. conceptual [DJLS22, JSLC16, Sav17, TH16, YLL20]. Conceptualization [CS15a, SV20, Eck18, SSS20, VJH15]. concern [SM18]. concerning [CRGG17]. Concerns [LK20, VCG17, LS14, MK15, Min16, QHH20, TMBA20]. condensed [SHW^{+}15]. Conditions [SVC22, Hod21, MF16]. Conduct [Cox14]. conference [ACBC22, Kim19a, LB19a, SDT22]. Conferences [VS15]. confidential [Zim17]. configurations [Fin17, STD15, WS22]. Conflicting [HH18]. conflicts [Lev20, SFL^{+}23, Med15]. conflictual [BV21]. Congress [KBC17, LML16]. Congressional [GAB^{+}18]. Congresswomen [EOG16]. congruity [Che16]. Conjoint [LMP21]. conjunction [Cri15]. connect [LCM17]. Connecting [CC17]. connection [Che15]. connections [CL14, LYL21, Wor19]. Consciousnes [Col20, Jul20]. Consensus [AGMR18]. Consensus-based [AGMR18]. Consequences [PHJ23, HSL24, VCG17]. considerations [Bur17, WCO21]. Considering [Sta22, ADD^{+}18, KD15, TPR17]. consistency [GLZ^{+}24, JJC19, Pra14b, Pra14d]. consistent [PR16]. Consolidated [SSS20, Cho15]. Consortium [MSdGS19]. conspiracy [DL22]. constitutions [RFF^{+}18]. constitutive [ØJW^{+}21]. constraints [Kit19]. construal [TDOF^{+}16]. construct [Sil17]. Constructing [BOR^{+}23, MYK15, TH16]. Construction [Men21, PhD20, CBAL15, HL16, IQDS16, JDZ24, Kha19, Mad16, PB15, TL22, TK17d, ZGEJ17]. consultant [JS15]. Consumer [LWLH18, CK16a, DMR^{+}17, JL18, MS16, Par19, PEG^{+}19, RYEP18, SZZ16, Zha14b, ZT22]. consumers [JL18, ZSX15]. consumption [HSU16]. contact [FvdWR^{+}22]. containers [BBC^{+}22]. containing [XC17]. contains [Bor16a]. contemporary [KPC^{+}15, MB23, YZ16]. Content [Bea16, DZC^{+}14, dLPMR16, ADD^{+}18, AKK^{+}17, BCZ14, Bor16b, Che16, CY22, CYYG17, DGCC17, LJT^{+}15, LKKS17, LJ22, Mat22, MJ23, MOM^{+}15, MSTP^{+}22, NSK17, SLI^{+}15, SBIT15, YCK16, ZY15, ZYL18, TJV14]. Content- [dLPMR16]. Content-based [Bea16, DZC^{+}14, ADD^{+}18, Che16, NSK17]. Contentious [Med15, HS17]. Contents [Ano16n, Ano16o, Ano16p, Ano16q, Ano16r, Ano16s, Ano16t, Ano16u, Ano16v, Ano16w, Ano16x, Ano16y, Ano17o, Ano17p, Ano17q, Ano17r, Ano17s, Ano17t, Ano17u, Ano17v, Ano17w, Ano17x, Ano17y, Ano17z]. contests [BJSM18]. Context [Fou19, LK20, MHJ19, PKB^{+}16, dFB22a, Ano14, Che22, CDL22, CK22, DLA14, DLA16, DP23, DS21, FMNG18, GB17, GSB^{+}16, Hjø15a, JT17, JLB^{+}23, JL23, JRS16, LX17, MS16, RKHMRH21, SMAB09, SYF17, SVK17, SAHZ21, SM18, Tab14, Wu19]. context-aware [LX17, SVK17]. context-based [DS21, MS16]. context-dependent [DLA14, DLA16]. context-driven [SAHZ21]. context-free [Hjø15a]. Contexts [MS19, ZZL20, JPF16, ZK14]. Contextual [KL22, LHM17, PMGB^{+}21, SSA22, WVZ20]. contextualization [GJ22]. Contextualizing [Fre15, MSTP^{+}22]. contingency [SWSZ19, WvdHS16]. continued [AJWB16, CJ16a, WHWZ21]. continuity [HDT23]. contraction [BK23]. Contribution [See17, CLL15, Cri15, Fan18, LCNP17, ZFH^{+}19]. Contributions [Sav17, YRH16, LWPG^{+}19, VCJ22, WW16]. Contributor [Hen19, JW15]. contributorship [LZA^{+}20]. Control [Eng17a, BV21, FW17, Pet19, RDS20, RH21, SLS^{+}19, SK14b]. controlled [MPB^{+}20, EJ14]. controlling [PH22]. controversial [VZRP15]. conventional [Sto20]. convergence [BMG^{+}17, HBG19, Rub15]. convergent [Bor15a]. Conversation [ZZL20]. conversations [BH15, TZS^{+}22]. conversion [GMRS17, KTZ20]. Coopmans [Rhe17]. coordinated [RC16]. Copeland [Hai16]. coping [Che15, Zha19]. CoPs [DB15]. Copy [dFB22b]. Copyright [Ben18, FM14, Owe21]. Core [Vin17, IABL16, ZZL^{+}14]. corpora [AST^{+}22, CC22, Che16, KS14a, LN17, Ma23b, WLC^{+}14]. Corporate [MH18, JJ14]. corporations [AFB20]. corpus [Ano17m, MSI^{+}19, GMB14]. Correct [vdB16, Rou16]. correction [Pet19]. correlate [CZW15, SBIT14]. correlation [LC21]. Correlations [WMS^{+}15, ATKD^{+}17]. correspondence [KISS15]. corresponding [TCF^{+}24]. Cortada [Kno24]. Cosine} [ZL16]. Cost [WTT^{+}14, AWC18, MK15, PSB16, PSB17, VXN18, ZPL^{+}20, Jam17]. cost-benefit [VXN18]. Costly [ML16]. costs [JPF16]. could [MWMB18]. counting [ZS16]. Countries [KSB20, APRRC15, BWL15, BGS22, GRS16, LHR16, NILCP15, TM20]. country [GNVQdMA16, KT14a, KGP^{+}20, RLAB^{+}24, TM15, ZCYN23]. counts [Bor16b, MT16, MT18a, TS16]. coupling [JGY^{+}19, MBML24, ZS14c]. couplings [CdRM21]. coupon [GMRS17]. Course [Hoo17]. Course-Subject [Hoo17]. courses [ELB17]. Cover [Ano15a, Ano15b, Ano15c, Ano15d, Ano15e, Ano16a, Ano16b, Ano16c, Ano16d, Ano16e, Ano16f, Ano16g, Ano16h, Ano16i, Ano16j, Ano16k, Ano16l, Ano17a, Ano17b, Ano17c, Ano17d, Ano17e, Ano17f, Ano17g, Ano17h, Ano17i, Ano17j, Ano17k, Ano17l, Ano18a, Ano18b, Ano18c, Ano18d, Ano18e, Ano18f, Ano18g, Ano18h, Ano18i, Ano18j, Ano18k, Ano18l, Ano19a, Ano19b, Ano19c, Ano19d, Ano19e, Ano19f, Ano19g, Ano19h, Ano19i, Ano19j, Ano19k, Ano19l, Ano20a, Ano20b, Ano20c, Ano20d, Ano20e, Ano20f, Ano20g, Ano20h, Ano20i, Ano20j, Ano20k, Ano20l, Ano21a, Ano21b, Ano21c, Ano21d, Ano21e, Ano21f, Ano21g, Ano21h, Ano21i, Ano21j, Ano22a, Ano22b, Ano22c, Ano22d, Ano22e, Ano22f, Ano22g, Ano22h, Ano22i, Ano22j, Ano22k, Ano22l, Ano23a]. Cover [Ano23b, Ano23c, Ano23d, Ano23e, Ano23f, Ano23g, Ano23h, Ano23i, Ano23j, Ano24a, Ano24b, Ano24c, Ano24d, WMX^{+}22]. Coverage [GRH^{+}17, DS21]. COVID [BK23, CCM23, Hod21, LBC^{+}22, NB23, RCM22, SBS23, SJT^{+}21, VFZ^{+}22, WZFZ22, ZGL^{+}22, Sha21a]. COVID-19 [BK23, CCM23, Hod21, LBC^{+}22, NB23, RCM22, SBS23, SJT^{+}21, VFZ^{+}22, WZFZ22, ZGL^{+}22, Sha21a]. Cox [Mar15]. Craig [Huv22a]. crawls [MWMB18]. Create [She19]. created [DGCC17, LB19b]. Creating [BH15, Col15]. Creation [BK14, LMM^{+}22, Mah22, Mat22, PHJ23, RAW22, SC21, SH14]. creative [FMG16]. creativity [CCR24, LKKP16, LMM^{+}22, Owe21]. credibility [CS15a, Cho20, GRC21, UH17]. credit [XDSC16]. credit-assignment [XDSC16]. Crestani [He18]. crises [XHM^{+}20]. crisis [RAMP22, VO23]. crisis-informatics [RAMP22]. criteria [TG17, XLKC22, ZSX15]. Critical [Mor24, PRM22, Adl23, CK21b, EMK^{+}15, Mar16, MA18a, MJ23, RAMP22, SB18, YP16]. critical- [RAMP22]. criticism [CS16a]. criticize [Ior14]. Cronin [O’C16, Rou17]. Cross [LMMO15, LZZ^{+}21, AP20, GVB17, KKO15, KH17, ML22, MPB^{+}20, MSI^{+}19, PMSA14, RLAB^{+}24, WDTC20, ZFH^{+}19]. cross-country [RLAB^{+}24]. cross-device [WDTC20]. cross-disciplinary [ML22]. cross-domain [PMSA14]. Cross-language [LMMO15, KKO15, MSI^{+}19]. cross-level [ZFH^{+}19]. cross-lingual [AP20].

Cross-modal [LZZ^{+}21]. cross-publisher [MPB^{+}20]. cross-sector [GVB17]. cross-validation [KH17]. Crossing [BNG14]. Crossroads [Wei14b, SHKL23]. Crowd [HYD^{+}20, Vas15, CPP^{+}16, RR17, Xia23, Fre18]. crowd-sourced [CPP^{+}16]. crowdfunders [LB20]. crowdfunding [DLD^{+}21, LZ22]. crowdsourced [LSX18, lLTBH15, TKB^{+}17]. crowdsourcing [Cab16, NLP16, RDPF16, Xia14, ZGEJ17]. Cryptographic [SR14]. CSCO [Hoo17]. CT [Kos20]. cueing [SP22]. cues [UH17, ZGL^{+}22]. Cultivating [Erd18]. Cultural [Bea17, Cho19, HDT23, JW17, KSB20, KC20, TMBA20, CPP^{+}16, Coc22, HNX18, LJI18, LYL21, MdSLdO^{+}23, RFF^{+}18, SM22, TKB^{+}17, WZ18]. Culture [Cox15, SR14]. Cultures [Uns18, DSBG20, FSPC16, PYH16]. Curation [BMM20, CCP15, CA20, DSBG20, DSBG21, HPL^{+}22, May16, Poo17, SJT^{+}21]. curatorial [CCP15]. Current [Fia14, PSB^{+}14]. currently [Bor14b]. curriculum [UM23, WXSL23, ZWH^{+}23]. curtain [MD15]. customer [CLL15]. cutoff [Smo17]. CV [KT14b]. CWTS [OSvE16]. Cyber [FMD19, ADM^{+}20]. Cyber-human [FMD19]. Cyberbullying [Bow19b, BKM15]. cycle [BM23, CKL16, MT20, SJN14]. Cyclic [Sav18a, JZC20].

D [FM14, Fri17]. Daemons [Dut20]. Dalkir [Haw18]. Dan [MS19]. dangers [SKO24]. Danh [All19]. Danish [Hen19]. Dark [SKO24, SRPS16]. dashboards [MR19]. Data [All19, Bad14, Bru18, CCMT15, CS16a, FMNG18, GRH^{+}17, GGC^{+}19, JVF19, LGD20, Mar16, Olt15, SJT^{+}21, Ten16, VC15, WC14, War15, WKJW16, Yoo17, ZSAW23, ZWH^{+}23, ZYL18, AdUCDPH19, ACS18, AGM19, AKC^{+}18, Bi22, BSG19, BWGD21, BB15, BTMS16, Bor16a, BH16, Bor17, BM23, BVd16, BOHK19, BJSM18, CCM23, CCP15, CPL22, CKLP17, Daq20, DB17a, DSBG20, DAT^{+}14, EMK^{+}15, EJ14, ESD22, FJT23, FKY16, FW17, Fon21, FC18, Fri15, GHKB^{+}17, GBZ^{+}23, HPLR23, HRN^{+}23, HD22, HHL21, HPL^{+}22, HCS21, JdR17, JQMY24, JRL20, KGS19, KDS^{+}17, KS16, KY17, KKK24, Kra17, LdMAGB15, LJ22, LWLH18, MF16, MdSLdO^{+}23, MB14, May16, MGK20, MBD15, OSD22, Owe21, Pan21, PYW18, PG20, QH20, RSAS14, RKHMRH21, RGJCTS16, RL15a, Rou16, SSC17, SAHZ21, Sil17, Sil18, SM22, SFL^{+}23]. data [Sta22, SBA^{+}15, SHW^{+}15, TDOF^{+}16, TD15, Tho22, UM23, VGN20, WDL21, WMS^{+}15, WB21, WLZB14, WXSL23, XCL^{+}21, YC20, ZRdR21, ZT22, ZPPW24, ZF19, vLTW18, CCP15, DAT^{+}14, Gor20a, KBH^{+}22, MSdGS19, PB16, Ols16]. data-driven [BVd16]. Data-Frame [PB16]. data-mining [TDOF^{+}16]. data-sharing [KS16]. database [WMS^{+}15]. databases [DQAMM^{+}24, Hjø15c]. dataset [DSBG21, DCLH21, FJT23, RMDF23, ZLY23]. datasets [JQMY24, Zub18]. dataspace [MNNK15]. date [PBG^{+}17]. David [Bud17, OM15]. Day [Ols16, BNG14]. days [Bor14d]. DBpedia [RSAS14]. Dealing [DPPV14, MB14]. dearest [BR20]. death [LPL14, RCM22]. debate [NV17, WvdHS16]. Debons [Wil14b]. Deborah [Gor20a]. debt [MSK^{+}14]. debutante [Moo20]. decades [DPHP17, HM14]. Decay [LK16, OK16, REVR22]. deceive [CM23]. deceived [BCK17]. decentralization [RH21]. Decentralized [SP17]. Deception [HHB17, RL15b]. Decimal [Lev15]. decision [GN17, Hjø15c, MR19, MAP15, OHP16, PR16, San22, SKO24, SYG17, WCTS15, XZ19, YR16]. decision-making [Hjø15c, SKO24, WCTS15]. Decisions [LK20, BVH15, CJ16b, GHCW17, JL18, MBD15, vdVZK21]. declining [LLG14]. decomposition [AHJY18, Gia16]. decrease [KAL24]. deep [GLZ^{+}24, HWZ22, SEZG22, TZH23, WL20, ZW23, ZTA20]. deep-learning [WL20]. Deeply [HK21]. defections [JW15]. define [FJT23]. Defining [KMJR18, Mat22, SHKL23]. Definition [VCG14, Kue23, SSS20]. definitional [Hjø15a]. Definitions [Fur20]. delay [GZBA22]. delayed [HR19, YB18]. delays [SRYL17]. deletion [LZ21]. deliberation [XA14]. Delicious [OS15]. demand [CBJ19, GLZ^{+}24, STF19, ZW23]. Demarcating [Sil20]. democracy [Li14b]. Democratic [Han21]. demotivating [St.17]. Denise [Swa22]. denoising [ZHGY18]. department [ZZSX16]. departmental [WW16]. dependency [BK19, LRB21]. dependent [ATKD^{+}17, BJR19, DLA14, DLA16, Fie16]. depression [BK23]. Depth [MS19, CB15, WZK^{+}18]. depth-level [WZK^{+}18]. derived [CCP15]. Dérives [Zit15]. Describing [MSM^{+}21, CDL18, CK22]. Description [MMHA23, Mat22, Smi15]. descriptions [lLTBH15]. Descriptive [MGKA16, AG15b, LQ14, Los16]. descriptors [dC16]. deserts [LB19b]. Design [BMM20, FC22, HWW^{+}17, YLL20, AG15a, ADL15, BKM15, BJSM18, HPLR23, HEEB^{+}22, IESG20, LM15a, Li14b, LYWF14, Mak20, MF16, NLP16, PKK^{+}24, Pil14, RC16, Ric18, VOBB15, WXSL23, ZYS17, LCK16, Soe18, WB15]. design-based [LYWF14]. designation [MWD17]. Designer [MHJ19]. Designing [GBL^{+}18, IESG20, QLM^{+}22, YLL20]. despite [MK15]. destinies [MP23]. detailed [BK14, TCF^{+}24]. detect [Ama22, vdSvdB23]. Detecting [MBBL14, RCM^{+}17b, SGS21, MBML24, ZRdR21, MSK^{+}14]. Detection [GB19a, Ano17m, ALMP16, BDA17, CCC14, CTM16, DC14, Don16, GDO17, GIEVVGS19, HKBB18, Pal16, PDS22, PBN^{+}20, dLPMR16, RVGPS19, TND15, VG18, Yu14, GMB14]. determinants [FvdWR^{+}22, HWZ20, JW15, ZXPS18]. determination [Zha16]. determine [OR15]. Determining [KCL14, KW14]. Detwyler [Ma22a, Sko23]. developed [KT14a]. developers [ELB17]. Developing [EBD15, GB19b, HSSE20, LYWF14, SBA^{+}15, SWSZ19, BGS22, Mar16]. Development [Cha14, Cho19, Hjø17a, KC20, PDŠB16, CK17, GVF^{+}24, GS18, HL16, Kim15, LR14, MSM^{+}21, Olt15, PSB17, RBJ18b, SS14a, She19, TTH16, TH16, Yoo17]. Developments [CKLP17, Sav18a]. deviance [SLS^{+}19]. device [WDTC20]. devices [CBJ19, GLR16, KTS^{+}16, MS19]. Dewey [Lev15]. DH [ZSG^{+}22]. diabetes [St.17]. diachronic [MSdGS19]. diagnoses [LXW18]. diagnosis [ALMP16]. dialogue [ZRdR21]. Diane [Hjø17a, Pan21]. diet [ZGEJ17]. Differences [APRRC15, Sin16, WSvPK16, BM14d, CHLRC14, DS15, DBH18, DTW16, KvOS^{+}17, KMKL15, LBM19a, TS16]. Different [CBvEW18, Ano18-27, AGJ^{+}16, BIW20, BELR16, CY14, Hoo17, HCDÅ15, LB20, LCHL14, LLB16, LK14, LML18, PYH16, RBJ22, TJ20, TPR17, ZZ21b]. Differential [YLHW21]. differentiate [MYK15]. differentiation [CCLL14, LHM17]. Difficulty [LLH19, HK17, JMV19, KLH17, Tab14]. Diffusion [GTAS22, JGCR15, KBS18, LCM17, LK16, MDL^{+}18, She16, She17, SL24, TND15, TLD17, Yan16, ZDW18]. Digging [PG20]. Digital [Abb17, BMM20, BSG19, Cus15, Dut20, Erd18, FM14, FMG16, Han21, Jon17, LB22, ML22, PhD20, Poo17, RAMP22, Saw19, Uns18, WCdF^{+}22, Web20, CLSS14, Cho16, CJ22, CC17, Coc22, DJLS22, DJ20, DC15, FLJD23, Fin17, Fra21, Gia18, GW18, GARA^{+}16, JPS^{+}19, KLG14, KA22, KIMM17, LL17, MMB17, MBW^{+}14, Mak20, MM20, MdSLdO^{+}23, OSD22, Owe21, PMSA14, PG20, RC16, Rey21, SCIF20, SS14b, SEZG22, TLA^{+}16, TR17, XBCH18, XBL^{+}21, YLL20, ZSG^{+}22, ZVW20, Cho19, Dow15, Hol15, Wei14b, Zim17]. Digitally [Meh22, Zha19]. digitization [Mak14, Con20]. Digitized [LYL21, CS16c, KL22]. digitizing [LS14]. Dijk [Rey21]. DIKW [McD21]. dilemma [HHB17]. dilemmas [BM23, EMK^{+}15]. dimension [YZ17]. dimensional [Pra14a]. Dimensions [H17, ZS16, ACBC22, AGM19, KPC^{+}15, Sta22, WK16, YSAM^{+}21]. diplomacy [Sto20]. direct [KS14a]. directed [KD15, Sea20]. directory [ZZSX16]. Dirichlet [WLC16]. dirty [MA18b]. disabilities [GM19, WH15]. Disability [O’H18]. disabled [AMOM16, ALMP16]. disadvantage [SVC22]. disadvantaged [BJR19]. disaggregation [FDR16]. Disambiguation [DMMF17, KMC20, FVGL14, KD16, Kim19b, KK20, KKOS21, LDK^{+}14, LLHF15, SGLF17, ZLY23]. Disaster [SSR^{+}20, PBN^{+}20, RJS19, SLLW20, WWL16]. disaster-related [RJS19]. Disasters [PKA20]. Discapnet [ALMP16]. discern [BCK17]. disciplinarity [TDOF^{+}16]. Disciplinary [VCJ22, WW16, Yan16, DS15, FSPC16, ML22, TS16]. Discipline [Bat15, Sve15, BLGbH21, Hoo17, HWW15, LB22, TK17b, WW15, YY16]. discipline-level [YY16]. Disciplines [GGC^{+}19, DFB14, KMFD15, Yan14]. disclose [MK15]. Disclosing [HQH^{+}22, KCZ^{+}14]. disclosure [CBH16, CMB17, LKRS18, ZGB16]. discount [GMRS17]. Discourse [XC17, DS15, Mor24]. discourses [KWLV22]. discover [Her23, Abb17]. discoveries [WTvR16]. Discovering [CC22, LJK21, SHQ^{+}16, TC17, WLC^{+}14]. Discovery [Erd18, KLC^{+}14, LSC18, TJ20, WFRN14]. Discrepancies [CHH14]. Discrete [Smo17]. discriminant [MGKA16]. discursive [Mad16]. Discussion [Fan18, OGL19, dSBG19, BBM16, Gau16, GZBA22, GJ16, Sav15a, TMDZ21a]. discussions [HS17]. disease [LKKS17, SKA18, WHYJ17]. Diseases [CKR14]. disentangled [KPC^{+}15]. disincentives [SWL18]. disjoint [ADD^{+}18]. dislike [GSG20]. disobedience [Jam20]. disparities [CW18]. Disposed [Meh22]. Disposed-Racial [Meh22]. dispositions [GBS15]. disproportionately [TM15]. Disrupting [Adl23]. disruption [GB17, MB20, WMM^{+}23]. disrupts [LBLL24]. Disseminating [KT14b, TK15]. Dissemination [JQMY24, CHH^{+}17, LCL16, PYH16, Yan15]. Distance [dSBG19, DBBJ16, HHPI21, KNY^{+}14, LF21, YL17b]. distant [OB19]. Distill [LSX18]. distinct [ZBKG14]. distortions [AD15]. Distortive [KD16]. distraction [Gre16]. Distributed [BC16, BOHK19, KLC^{+}14, GBS15, TNL21]. Distribution [BBH15, BAGL16, MPB^{+}20]. distributional [Hal17]. Distributions [BM14a, Hof17, HBG19, MH16, PLPQB20, PRRC16, PSS22, ST16a, ŠF17, TTH16]. divergence [BMG^{+}17]. diverse [MMHA23, SLS20]. diversification [OA15, YL17a, YSAM^{+}21]. Diversity [ZRG16, BDX^{+}18, BLGbH21, JW17, RR17, Shm18, TL22]. divide [LWB16, MM20, RC16, RAMP22, Rey21]. division [LZA^{+}20]. Do [BS16, CL17, CZW15, DSBG21, EOG16, HR19, JL18, KTZ20, KISS15, LSX18, MT18b, RHG23, SBIT14, TD15, TS21, TKA^{+}23, vdVZK21, Bor16b, EH21, HPL^{+}22, Kim15, Kit19, MT18a, SVC22, SL21b, YL18, JJ14, Bor14a]. doctor [LHXS22]. document [AG15b, AHJY18, ABLS17, AFB20, Che16, Gia16, HRN^{+}23, KKZS20, LB15, LX17, LZ19, MGKA16, SB16, STCS17]. document-centered [AFB20]. Documentary [KK19a, Sea20]. Documentation [Ols16, ØC19, LSX18]. Documenting [GW18, TWB^{+}18, LGD20, MWMB18]. Documents [Mos18, SLE21, CSOD16, DQAMM^{+}24, DPPV14, DC15, FDR16, Hen14, KW14, Rob15, SJN14, SHQ^{+}16, Buc15, SR14]. Does [AD23, Hjø17b, LB19a, NILCP15, PTGL17, RD16, SJL19, SDT22, TW16a, AMDFHVC24, BM15a, Bor16a, Cj15, FLJD23, Hal14, LCM17, RDPN21, SR15, YHLG21]. DOI [Ano14]. Doing [Web20]. Domain [ACBC22, Smi15, Smi17a, FJT23, HRN^{+}23, Hoo17, KKK24, Lin14a, LZ19, LM15b, MGCTW15, MKK17, PMSA14, RVGPS19, SGLF17, SZZ14, THWS14, VOBB15, VT16, WKC14, ZLCB15, SK14a]. Domain-independent [Smi15, Smi17a]. domain-specific [HRN^{+}23, SGLF17, SZZ14]. Domain-topic [ACBC22]. domains [JL23, YL18]. dominant [XZ19]. Donald [Col17b, Cus15, CY22, Tei17]. Don’t [BCK17, MJ23, SSA20]. Dorothea [Cha14]. double [SDT22]. double-blind [SDT22]. Douglas [Bow19a]. Dousa [Car14b]. Down [MB20]. download [MH16]. downloads [CJ16b, KH17]. DPCVocab [CCP15]. draw [Har14]. draw-and-write [Har14]. Drive [Col20, Jul20]. driven [BVd16, Cab16, DMR^{+}17, Du14b, Hil14a, LML16, Meh23, SAHZ21, SBA^{+}15]. drivers [LL17]. drives [LMM^{+}22]. driving [DBH18]. dropping [FN17a]. drug [CK16a, CMB17, LSC18, SKA18, WHYJ17]. drug-disease [SKA18]. drug-symptom-disease [SKA18]. dual [CL14, LZZ^{+}21]. dual-map [CL14]. Duff [Baw14]. Duguid [Kea22]. During [PKA20, BK23, GHCW17, HCC^{+}21, Hjø15c, JDZ24, LCWC23, LXW18, RAMP22, Roh21, SJT^{+}21, Smi15, SS14c, TG15, VO23, ZJHT23, vdSvdB23]. dwell [LB15, WZK^{+}18]. dwelling [LS22]. Dyadic [HB14, FdGPC18]. Dynamic [CCT14, DC14, BK14, Fia14, YY16, ZXPS18, BL14]. Dynamics [Cox14, GHCW17, HLW18, HHB17, KR15, LdN17, Lin14a, Whi14, XBL^{+}22, Yan15]. dyslexia [BS16].

e-book [Fou19, MS19]. E-books [ZK14]. e-commerce [XZ19]. e-learning [PRM22]. E-mail [KR15, KISS15]. E-mails [KISS15]. E-prints [LSM^{+}14]. Early [CN14, PBN^{+}20, Bor14d, WK19]. Earth [SS14c]. East [QHH20]. Ebook [QH18, Ama19]. ecology [DOM^{+}14, JPS^{+}19]. Economic [Sil20, DCLH21, MSK^{+}14, NILCP15]. Economics [Hen19, GD18, Hal14, McC15, Sha21a]. economy [HSU16, MSB21, SSS20]. ecosystem [GJ22, Kue23]. ecosystems [YC20]. Ed [Zha19, Haw18]. Edge [KMC20, LXLL14, SLE21]. Edited [Bea17, Buc20, Huv20, Lan15, O’C16, O’H18, OM15, Pan21, Rou17, Zim17, Bat15, Buc15, Car14b, Cus15, Dow15, Sve15, War15]. edition [Col17b, Wei14b, MWD17]. Editor [Bor16b, vR17a, Man22, TSRG16, Bor14c, Bor14d, Bor14b, Bor15c, BBH15, Bor15b, Bur17, CB14, FN17b, Hjø17b, LWB16, Mah17, MMALCOM15, MB14, Mir14, Pra14d, Pra15a, Pra15b, Rou15, RR15, TD15, Tei16, VCG17, VR17b, ZMJ14, dC16, vdB16]. Editorial [Ano16m, Cro14, Cro15a, Cro15b, LHS24, MWPF17, Mos17, Mos18, SB14, ZVW20, GJ22, TMDZ21b]. Eds [Clo18, Han21, Hay18, Rhe17, Sko23, QH18]. Education [HM14, KC20, dSBG19, JRL20, JAG^{+}20, MR19, MAAL^{+}16, PRM22, PDŠB16, UM23]. educational [CLA^{+}16, GBZ^{+}23, JRL20, Kon16, SP22, TK17c, WWB21, YSAM^{+}21]. EEG [GHCW17]. Effect [KK20, WH15, AGBN16, BDH17, BGS22, CK21a, CHLRC14, EAB17, JL18, JQMY24, KCK16, KAL24, LPL14, LF21, LBM19a, Lia17, LYC^{+}22, MWD17, SLLW20, ŠF17, TSY15, WSWH20, YR16, ZGEJ17, ZXLC18]. effective [Ano18-27, GAS^{+}20, HB15, KZB17, LSC18, LML18, SZZ14, STCS17]. effectively [WLC^{+}14]. Effectiveness [AGGTGL18, CY22, HE22, Lew15, LK14, Pet19, RJK20, SY22, VLBG^{+}16, WKC14]. Effects [MGCTW15, POK15, SRM^{+}17, TPR17, WOJBG17, WN15, WV18, Xia14, CZZY15, CH16, DTW16, FW17, FN17b, GBS15, Gre16, HHSP22, IESG20, KD16, LS14, NB23, PTGL17, RKL14, RC16, RR17, RR21, SRG^{+}16, SP22, SP17, SWSZ19, Tsi16, UH17, WZFZ22, WK16, XLKC22, ZGL^{+}22]. Efficient [KBD^{+}14, BOHK19, KS17, YXAL23]. effort [FSK^{+}16, GRSFV15b]. Egbert [Bow19a]. egg [RYEP18]. Egghe [Dor19, Rou15]. Ego [Jor19, POK15]. Ego-Networks [Jor19]. Egocentric [MP23]. EHR [WLW^{+}17]. elaboration [Sav17, KCK16]. electric [RDS20]. Electronic [SR14, AF17, CCJ16, Jam20]. Elham [Bat21]. elicitations [BR16]. eliciting [CLA^{+}16]. elite [LLG14]. Elitist [BH15]. Ella [Sto20]. Ellis [Sav17].

Elsevier [Abb17, JML22]. elusive [AS24]. email [GDO17, REVR22]. emails [LC21]. Embedded [FC18, Du14a, MGKA16, OMTSLC15, TNL21]. Embedding [LSL23, ALCD14, LKY^{+}20, ZQC^{+}23]. embeddings [KDS^{+}17, LS21]. embodied [Lue20]. embodiment [KWLV22, Lue15, PPC19]. Emerald [Col17b, Fro21, Mat20]. emergence [Roh21]. emergency [RAMP22]. emergent [ACBC22]. Emerging [Oli21, CC22, GOdR18, MSTP^{+}22, RRHL17, Wan18]. Emily [Buc20]. Emme [Sav18b]. emotional [FJF^{+}21, GRC21, KZ23, WCO21]. emotions [LJK21, Lop14, SZWS22]. Empirical [CBH16, LCP15, MCSB19, ZZ20, AFL^{+}14, BH17a, BH17b, CGGS17, FLJD23, GLR16, HSSE20, LL17, LP16, LOM22, LKRS18, LM15b, PHZF16, PB16, QSH14, ZLF20, PMSA14]. empirically [CCP15]. empiricism [MF16]. employ [BB15, PHZF16]. Employee [SV17]. employees [DAT^{+}14]. empower [FHO23]. Empowering [ELB17]. empowerment [ZJHT23]. EMR [Bur17]. Emulation [Ack21]. enables [ADM^{+}20]. enabling [BOHK19]. encountered [PKB^{+}16]. encountering [MB20, SZWS22]. Encounters [LBM^{+}19b, RCM22]. encourage [NLP16]. end [CBAL15, KMD16]. ending [PSY^{+}21, SFL^{+}23]. enforcement [TGC^{+}21]. engage [GBS15]. engaged [CPL22]. Engagement [LBM^{+}19b, NTNP19, AGBK19, ALC^{+}14, GAS^{+}20, LZ22, LCLBY17, NT17a, O’B17, RC16, RJK20, SP22, TLS^{+}23, VOBB15, WHWZ21]. Engaging [EH15, HS17, WXLBO15]. engine [AS24, AKC^{+}18, LKRS18, LHS24, Mor24, PKK^{+}24, SKO24, UH17]. engineers [Fre15]. engines [KPWS20, Lew15, SL21a, SLH22, vdVZK21, Sha21b]. English [Ano14, BNG14, HCC15, JMV19, Kim15, Ma23b, MSI^{+}19, SMAB09, The20]. English-speaking [HCC15, The20]. English-written [Ano14, SMAB09]. Enhance [NTNP19, Hen14, SDUDL24, Sta18, WWB21]. Enhanced [Don16, DLA14, WVvLZ23, YXAL23]. enhancement [PHH^{+}20]. Enhancing [LBA16, ZZ21a]. enough [MT20]. enriched [LZSW17]. enriching [KBC17]. Ensemble [SKL17, WLC16]. entailment [KMJR18]. enterprise [CBM17, Mat22, WTT^{+}14]. enthusiast [LO19]. enticed [MK15]. entities [AG15b, BSZ^{+}14, DP23, FT17, GOdR18, KWR20, MC18, MMFD15, WMM^{+}23]. Entity [PS20, SKKM22, AG15b, CLA^{+}16, GDO17, KS14b, LMMO15, PMGB^{+}21]. Entity-aware [SKKM22]. entity-oriented [KS14b]. entrepreneur [LZ22]. entropy [BR15a, SLLW20]. environment [BM14d, Gre16, PSB16, RBJ18b, TTH16, XBL^{+}21, Pan21]. environmental [LBWE^{+}23]. environments [Ano18-27, LBG^{+}21, LK14, LML18, MBW^{+}14, Mak20, NB23, OB19, VP19, Wei14a]. episodes [WLC^{+}14]. Epistemic [ZFH^{+}19, DS15, Hod21, Kel21, LJ22, Oli21, PSY^{+}21, SE16]. epistemicide [PSY^{+}21]. epistemics [Kel21]. Epistemology [Kel21, WBH18, Fri15]. epistles [Sav19]. equal [ZTLV15]. equality [Jam17]. equally [WL14a]. equate [vdVZK21]. equation [TW16a]. Equipped [Eck18]. era [KPWS20, LL17]. Eric [Ma22a]. Erik [Han21]. ERP [KK19a]. ERP/MMR [KK19a]. Erratum [Ano14, Ano17m, Ano17n, Ano18m, Ano21k]. error [Don16]. errors [LCH^{+}22]. essay [VVP^{+}21]. estimate [JLC16]. estimates [Don16]. Estimating [Sav16a, VLBG^{+}16]. estimation [WL14a, WTT^{+}14]. estranged [Lue15]. Eszter [Zim17]. Ethical [HHB17, Jam20]. Ethics [FHW17, LWM^{+}23, Mar23, San22, SFG^{+}23, Tei17, Buc20]. Ethnicity [KKOS21]. Ethnicity-based [KKOS21]. Ethnography [Hak14]. Euclidean [Hal20]. EUR [Fra21, Han21, Sha21a, Sha21b]. Europe [KHZ^{+}22, RN16]. European [Cho19, AMM^{+}19, GRS16, KGP^{+}20]. evacuation [LBG^{+}21]. evaluate [GISW15, Hoo17]. Evaluating [ABLS17, Lew15, STS15, CCLL14, GSB^{+}16, HCD17, TTH16, ZHGY18]. Evaluation [Agu18, BI17, MŽA17, OSvE16, SVK17, SAW21, TCP15, WC14, WKC14, ZBKG14, Zit15, AG15a, AGMR18, ALCD14, Bor15c, BTMS16, CHH^{+}17, FSK^{+}16, GGS22, GLR16, HWW^{+}17, HLB^{+}17, HNX18, LL17, LCP15, LKKS17, LCWC23, MT14, MBD15, PLPQB20, PRRC18, RJS17a, SVC22, SB18, WB15, WSvPK16, ZSX15, ZSC18, ZYS17, ZF19, dRR15, Zit15]. evaluations [SVB22]. evaluative [LB16]. event [BDA17, KK19b, Pal16, WLC^{+}14, wYKY17, ZS19]. event-related [KK19b, ZS19]. events [FW17, HCC^{+}21, LGD20, PBN^{+}20, PHC17]. Everyday [Cox14, Hal19, CS16a, CPL22, Gre18, HSSE20, HB15, IB16, ST22, Sin16, WV15]. Everyone [TNL21]. everything [KR15]. everywhere [SYF17, TNL21]. Evidence [GRS16, HYD^{+}20, JVF19, JSY^{+}16, LBC^{+}22, MZ22, OMTK17, SR14, ZXPS18, KKO15, KT17b, KT17a, LML16, TM16, BC16, NILCP15, SDT22, TKA^{+}23, ZGE22]. evidentiary [TFR15]. Evolution [TJV14, UM23, HDZ^{+}15, KMFD15, LJI18, ML21, MT18a, MCY^{+}21, NF15, RFF^{+}18, SMW14, SHQ^{+}16, TL18, WWL16, ZZE17, GRBCS15, YZQT20]. evolutionary [AJWB16, JL23, ZLZG14, ZZZL17]. evolve [BVd16]. evolved [AJWB16]. ex [MBD15]. ex-ante [MBD15]. ex-post [MBD15]. examination [CJ16a, GK14, ML22]. examine [YY16]. Examining [FvdWR^{+}22, GKM^{+}21, HBH^{+}16, IESG20, LBW^{+}19b, MMHA23, RYEP18, SY20, HBG19, LB20, LB15, RHG23]. example [PR16, QH20, Sav16b]. Excellence [TKA^{+}23, BC16, BWL15, PS14b, VR17b]. excellent [BdMA14]. Excesses [Zit15]. exchange [JDHL17, ZY15, Kan22]. exchanged [PN18]. exchanges [CY14, PHJ23]. exciting [BR15a]. excluded [DD16]. exclusion [DQAMM^{+}24]. execution [SSC17]. exercise [BC16]. exist [HR19]. exodus [ZXPS18]. expand [BSZ^{+}14]. expanded [ST22, Skö18]. Expanded^{T M} [HK14]. expansion [BST15, BK23, CDGN15, JKS^{+}16, PMD14, SBZ^{+}14, TL18, XLL16]. expectation [MBVB17]. expectations [MS16, Wes15]. expected [Bor16a]. experience [AvS15, AvS16, Gor19, GBL^{+}18, HLB^{+}17, LP16, LSC18, MS16, QLM^{+}22, RR17, Rut21, TJ20, Zha14c, Bat21]. experience-aware [LSC18]. Experiences [RR20, HSL24, KWLV22, WCP^{+}19]. experiment [LCRGTS14, MGK20, RSAS14, ZGCB15]. experimental [Che16, LK14, SLI^{+}15, WHvdZvdB17]. experimentation [Her23]. experiments [MSI^{+}19, Smi17b]. expert [HRN^{+}23, SB18, WTT^{+}14]. expert-in-the-loop [HRN^{+}23]. expertise [Smi15, Smi17a]. experts [AD23, CK14, EHdBvO17, RAW22, ZGEJ17]. explainable [ZW23]. Explaining [AJWB16, SSdMCM20]. explanation [Egg14a]. Explication [WK16]. Explicit [OA15, YSAM^{+}21]. Exploiting [BSVZ17, JSY^{+}16, MMFD15, WLC^{+}14, SGLF17]. exploration [HPLC19, LP16, LRB21, Mad16, VRF14]. explorative [SC17]. Exploratory [CBM17, OGL19, AS14, AGJ^{+}16, BTMS16, Cho20, DDB16, Gau16, HABT18, HS22, LCWC23, MJ23, OSD22, PB15, RJK20, SHGI16, Wat14]. explore [GHKB^{+}17, GS18]. Exploring [ACS18, AKK^{+}17, Eng17b, GBL^{+}18, JRS16, Kim15, Lia17, LCWL18, PKA20, SH19, Yu14, vLTW18, ABLS17, LCM17, Fou19]. explosion [Hil14a]. exponential [Smo17]. Exposure [NLFM19]. express [MT20]. expressed [RBJ18b]. expression [Smi17b, Mar15]. expressions [FJF^{+}21, KTZ20]. extended [Cho20, FLJD23, FvdWR^{+}22, SWSZ19, WZL21]. Extending [YP16]. extensive [DMdO16, WDL21, CZW15]. extent [AMDFHVC24]. external [OMTSLC15]. extra [LI21a, NLP16]. extra-role [NLP16]. extra-scientific [LI21a]. extracted [FT17]. Extracting [JL23, LS17, MRAF16, STCS17, ZLZG14, KT15]. Extraction [BDA^{+}18, HWC17, ZZL20, Bar14, CLA^{+}16, KDS^{+}17, LBWE^{+}23, Vec14, WHYJ17, YHL^{+}21, ZZ21a, VT16]. Extreme [BMG^{+}17]. extremism [WWB21]. Eye [KTS^{+}15, GHCW17, KTS^{+}16, WLTY19, vdSvdB23]. Eye-tracking [KTS^{+}15, GHCW17, vdSvdB23]. eyes [OMM^{+}14].

F [Buc20, Pan21]. F1000 [DTW16, WC14]. F1000Prime [Bor15a, Bor17]. Fabio [He18]. Facebook [QH18, FR23, OS15, PHJ23, Roh21, VXN18, WXLBO15, ZGB16]. faced [dC16]. Facet [Col17a, Huv20, Pan21]. faceted [NH15, RJK20]. facets [LRB21, QH18]. Facial [SSA20]. facilitate [Mah22]. Facilitating [HPLC19, WFRN14]. facilities [VP19]. facility [Söd23]. Facing [Bud17, FDC22]. fact [HE22]. Factor [KHZ^{+}22, DD16, SRYL17, XZ19]. factorization [SLLL17]. Factors [GEBP22, LJLH17, MAAL^{+}16, OY15, St.17, THC15, ZPL^{+}20, Che16, CJ16a, DBH18, KCL14, KS16, LWA^{+}17, Li14b, TMBA20, XBCH18, XBL^{+}21, YP16]. Factory [Dow15]. facts [Kno24]. faculty [FMG16]. Fagerlid [Han21]. fail [dC16]. failed [Sto20]. failures [AD15]. FAIR [GGS22]. Fairness [CJ22, GGS22, BOR^{+}23, MBD15]. Fairness-aware [GGS22]. fairy [Fra21]. Fake [Fro21, Kos21, SGS21]. fall [SR15]. False [Kos20, LCRGTS14, Tei17]. fame [Fas18, VRF14, WZ18]. familiarity [AZE^{+}19, NPG^{+}20]. families [BV21, MŽA17, ZBBW18]. family [DLA22, FLJD23]. far [PdlRR^{+}22]. fast [LLHF15]. fatigue [RKL14]. favorites [GB16]. feasibility [ACL17, ZS14b]. feature [AFL^{+}14, ATKD^{+}18, MC18]. feature-based [ATKD^{+}18]. features [AKAS14, ATKD^{+}17, CMS23, FdGPC18, LDS^{+}19, MDC^{+}16, SJ17, SMRR16, TDOF^{+}16, WSWH20, WLC^{+}14, ZW23]. federal [Olt15]. Feedback [vdSvdB23, AG15a, CDGN15, JBD^{+}19, LYC^{+}22, PHH^{+}20, TBP^{+}15, THC15, YH16]. feedback-based [CDGN15]. feeling [SFL^{+}23]. Feelings [FJ19]. Feist [Owe21]. Female [KP18, SCIF20, The20, TF21]. feminist [FC22]. Fenwick [Dut20]. Festschrift [Uns18]. few [BR15a, SAB^{+}18]. fibromyalgia [Che22]. fiction [AP17, BF18, EHdBvO17]. fictional [BAK^{+}17]. Fidelia [Car14b]. fiduciaries [JRL20]. Field [Bru18, BH17b, BGS22, Chi19, HM14, Ior14, Meh23, RJGM16, TMDZ21a, TMDZ21b, ZDP^{+}23, ZGCB15]. field-normalized [BH17b]. field-specific [Chi19]. fields [GD18, LB16, MBBL14, PYW18, SAW21, TW16b, TM20, TKS^{+}23, TH16, ZRG16]. fieldwork [TWB^{+}18]. Fietkiewicz [Ama19]. Fifth [LBM^{+}19b]. Fight [FM14]. figures [ZTA20]. file [BIW20, DJ20]. file-sharing [BIW20]. files [BWF14, TD15]. Filing [Huv22a, KR15]. Filipino [AC18]. Film [Sea20]. films [WMS^{+}15]. Filtering [YL17a, BSVZ17, DS21, ESWC18, HABT18, LJT^{+}15, ZCW17, ZHGY18]. finance [Hal14]. financial [ESD22, GN17, LLB14, RDPN21, SKKM22]. find [ZB19]. Finding [DO14, Yan14, AGGTGL18, DMR^{+}17, FT16, HB16, SVK17, WSvPK16, WV15]. fine [LS17, LWL^{+}18]. fine-grained [LS17, LWL^{+}18]. Finland [AST^{+}22]. Finnish [DBH18]. Fiorella [Cox15]. firm [QSH14]. Firms [JVF19]. first [CCM23, LB19a, RBJ18a, RBJ18b]. first-time [RBJ18a, RBJ18b]. fit [Fro21, YHLG21]. Five [DPHP17, VGN20]. Fix [GKM^{+}21]. flagging [CY22]. Flanders [DDB16]. flavors [WVvLZ23]. flexible [Oh17, YHL^{+}21]. Flick [KBC17]. Flickr [KBC17, OS15]. flipping [MZ22]. Floridi [BR20]. flow [HWZ20, Lai15, Olt15, WWL16]. flows [BELR16, DDB16, MT14]. flyby [SS14c]. focus [HL16]. focused [LCLBY17]. Follow [SLE21, JL18]. followees [TG17]. followers [GKM^{+}21]. following [MRAF16]. footprint [PM18]. Ford [Col17a]. forecasting [ZW23, SHGI16]. forename [KK20]. Forensically [Hod21]. forever [LMJ21a, LMJ21b, She21]. Form [BD16]. Form-ing [BD16]. Formal [Fri17, PYW18, KMKL15]. formality [DB15]. Formalization [RH21]. Format [DB17b]. formation [YR23]. formative [AG15a]. formats [BBC^{+}22, CY14, QLM^{+}22]. formula [Pra14c]. formulation [KBD^{+}14, Smi17a]. forum [Gau16, SMRR16]. Forums [dSBG19, CYYG17, CMB17, LSC18]. Foscarini [Cox15]. foundations [Bat21, Buc20]. Four [ESD22, HBS15]. Fourth [SK14a]. Fox [FR23]. fragmentation [VCJ22]. Frame [PB16]. Framework [TKA^{+}23, BBC^{+}22, CCP15, CK22, DGCC17, DJLS22, GSB^{+}16, Gor24, GEBP22, GLZ^{+}24, HPLR23, Hoo17, HCD17, HLB^{+}17, LWL^{+}18, MJ23, Meh23, MMHA23, PSY^{+}21, PEG^{+}19, RB14, SSA22, SMYL15, Sil17, Smi17b, TLS^{+}23, TC17, WCTS15, ZF19]. Frameworks [BMM20, MMB17]. Framing [TMDZ21a, WFRN14, WLG23]. France [GD18]. Francois [SR14]. Franklin [Meh22]. fraud [ML16]. FRBR [MŽA17]. FRBR-based [MŽA17]. Frederick [Ma22b]. Free [BGS22, Jam17, BLGbH21, DQAMM^{+}24, Hjø15a, Kim19b]. free-access [DQAMM^{+}24]. Free-to-publish [Jam17]. free-to-read [Jam17]. frequency [CO16, iCV18, KLH17, Sin16, ZS16]. frequency-weighted [ZS16]. frequently [BWL15]. freshmen [GBZ^{+}23]. Friedman [Aji17]. Friend [LK20]. friends [GS22, WXLBO15, GS22]. front [ZS14c]. Full [DS21, DAT^{+}14, McC15, MDC^{+}16, MH16, ONZ23, ZYL18]. full-text [McC15, MDC^{+}16, ONZ23, ZYL18]. full-time [DAT^{+}14]. function [Bor14c, CB15, Zha14a, ZSC18]. functional [XDDG18]. functionality [lLTBH15, XDDG18, GTZŽ14]. functionality-based [XDDG18]. functions [BS16, Bi22]. fund [MBD15]. funded [PG20, WYZL21]. Funding [GRH^{+}17, LBW19a, THL17, GRS16, HPL^{+}22, JGCR15, MBD15, PSB16, RHG23, Vas15]. Fundulaki [All19]. Further [Bur17]. fusion [WLZB14]. Future [CA20, HQH^{+}22, LB19a, Lia17, PSB^{+}14, SBIT14, SK23, ZZ21b]. future-oriented [SK23]. futurities [FC22]. Fuzzy [CK16b, Don16, LL17, MTY14, Oh17, VJH15, ZHGY18]. fuzzy-SEM [LL17].

G [Sar16]. G. [Ama19]. gained [Smi15]. Gaining [Smi17a]. galleries [CC22]. Game [MHJ19, CDL18, CBAL15, EHdBvO17, MSM^{+}21, RC16, SMW14]. Games [FJ19, JSLC16, LCP15, PTGL17, Mir14]. Gans [Sha21a]. gap [JA16, Sha21a]. gaps [OR21]. gatekeeping [PT21]. gather [ZBKG14]. Gaussian [IABL16]. Gay [MJ23, KWLV22]. gaze [ALC^{+}14]. GBTM [BL14]. Gender [BM14d, DP23, IS22, RJS17a, WYZL21, ZCYN23, HK21, Sin16, The17, TF21, SCIF20]. gender-polarized [TF21]. Gendered [SSA20]. Gene [MR14]. general [DGCC17, DGGdST19, WKC14]. generalism [SLS20]. generate [CCR24]. generated [CL21, Par19, WLG23]. Generates [KB17]. Generating [AG15b, DB17a]. generation [BST15, DZC^{+}14, LZD16, LGZ14, Rob15, RVdAGP21, YHL^{+}21]. Generic [KK16, ZGPB17]. Genetic [GB19a, IABL16]. genius [LS16]. geNov [AH17]. Genre [BBC^{+}22, MG21, SA15, The17]. genre-theoretical [SA15]. Genres [Fin17]. geo [AG15b]. geo-located [AG15b]. geobiology [TWB^{+}18]. geodata [DDB16]. Geoff [Mar15]. geographic [AS14, RDPF16]. geographical [BCZ14]. geography [SHMC17]. geolocation [ESWC18, LYCX14]. geometric [WLTY19]. geospatial [Li14b]. Germany [Uns18]. get [JWWV18, MT18a, SLE21]. Getting [CLSS14, MM20]. Giedion [See17]. Gillian [Cox15]. Gilliland [Clo18]. Gingras [BI17, Zit15]. girls [YB18]. GIS [SE16]. Gitelman [Buc15, War15]. Given [Col17b]. Giving [OSD22]. GKC [SSA22]. GKC-CI [SSA22]. Global [KSB20, Söd23, XHM^{+}20, YZQT20, BWGD21, BK14, Hil14a, LCM17, Mat20, NSK17, ST16a, WCC^{+}17]. Glushko [Bat15, Sve15]. GmbH [Han21]. Gnoli [Gui17]. goal [WLG23]. goal-framing [WLG23]. goals [CK17]. Goeing [Kea22]. Going [MBVB17]. Gold [PSB17, JPF16, ZLY23, vLTW18]. Gold-centric [PSB17]. Good [MSK^{+}14, Egg14b]. Goodreads [KTA17, TK17a, Wor19]. goods [Hof17]. Google [AGGTGL18, AT14, BS16, BTMS16, GD18, Hal14, JRS16, KT15, KT17c, LCRGTS14, LN19, OA14, RLAB^{+}24, VRF14, VC15]. googling [HSL24]. gorilla [WMX^{+}22]. Gould [Bor14a]. governance [FMNG18, SSA22]. Government [LBW19a, DAT^{+}14, SFL^{+}23, TMBA20, ZZSX16, Bor16c]. governments [HPLR23, Kno24]. grade [CBAL15]. graduate [WXSL23]. Grafton [Kea22]. Graham [Saw19]. grained [LS17, LWL^{+}18]. gram [DGR19]. grammatical [ÇÖ18]. grand [HLB^{+}17]. grants [WYZL21]. Graph [KMC20, MKK17, DGGdST19, JGY^{+}19, LSL23, MRAF16, ZYS17]. Graph-based [MKK17, DGGdST19, MRAF16, ZYS17]. graphic [LM15a]. graphics [SR16]. graphs [CHH14, DGGdST19, Egg14a, Fie16, HCKH15]. gratitude [DPHP17, MT20]. GRAW [JGY^{+}19]. Great [Sea20, SK14a]. greater [HHC^{+}14]. greatly [Poo17]. Greek [JMV19]. green [BLWP14, JPF16, TPZ15]. Gregory [Sav18b]. Gremm [Ama19]. ground [CS16a, Roh21]. grounded [BMG^{+}17, Bur14]. grounds [PHJ23]. Group [Col17b, QZWC21, GS22, HHB17, JHJ15, Lai15, Roh21, Sav15a, STAH14, BL14]. Group-Based [BL14]. Groups [BDA^{+}18, FW17, HRB14, Lai15]. Growth [BM15b, MBZ^{+}22, NILCP15, PSB^{+}14, Sav17, SL24]. Gruyter [Bud17, Han21, QH18, dP20, vR17a]. guanxi [OD16]. Guest [SB14]. guidance [MKK17, ZZLS16]. Guide [Agu18, Dor19, Mat15]. guided [Smi15, Smi17a]. Guideline [TM16]. Guidelines [VGN20]. Guns [Dor19].

H [Bor14a, Hjø17a, CZZY15]. H-index [CZZY15]. Håkon [Han21]. Haase [Hay18]. habits [Lin15]. Hack [Sea20]. Handbook [Hak14, Hay18, Bru18, Sar16]. handling [SAB^{+}18]. Hands [RR20]. Hans [Han21]. Hans-Christoph [Han21]. happens [Her23]. happy [TR16]. hardback [Aji17, Bat21, Owe21]. Hardcover [Bow19b, Bru18, Bud17, Fre18, Gui17, Hal19, Hay18, Hjø17a, Jon17, QH18, Abb17, Afz22, All19, BI17, Bea17, Ben18, Bow19a, Cle18, Col17b, Con20, Far23, Han21, Haw18, Jul20, Kea22, Kno24, Kos20, Kos21, Lev20, Ma22b, Ma22a, Mat20, O’H18, Ols16, Rou17, Saw19, Sha21b, Swa22, Web20, Whi17, Zha19, vR17a]. Hargittai [Zim17]. Harmonizing [KBH^{+}22]. Harnessing [O’C16, TPZ15]. hashtag [SS17]. hashtag-oriented [SS17]. Hashtags [PT21, ÇÖ18, KYFZ15]. Hastings [Bea17]. Haunss [Med15]. Haven [Kos20]. Hawkins [Cus15]. Hayot [Ma22a]. hazardous [SRPS16]. hbk [Sko23]. HDP [DC14]. Health [PEG^{+}19, SRPS16, TCP15, WYZL21, CW18, CLSS14, CK16a, CHJ20, Cho20, CS15b, CYYG17, CV20, DMR^{+}17, Fox20, GB17, GJ16, GLZ^{+}24, HH14, IESG20, JWJ^{+}19, KCZ^{+}14, LS22, LS14, LMA^{+}22, LSC18, LHXS22, LJLH17, LCWL18, NGTKR15, Par19, RJS17a, RBJ18b, SZ20, TBP^{+}15, TM16, VFZ^{+}22, WHWZ21, XHM^{+}20, Zha14b, ZBKG14, Zha14c, ZSX15, ZZLS16, Zha16, ZCZ^{+}20, ZLF20, ZT22, ZPPW24, ZGEJ17, ZB19, FJT23]. health-related [LMA^{+}22, Zha14c]. healthcare [PPC19, TBP^{+}15]. Heinström [Huv20]. helix [CYP15, LPU15]. help [AJWB16, KT16b, XBCH18, ZB19]. help-seeking [XBCH18]. helpfulness [CB15, TZK20]. hemodialysis [VP19]. Henningsen [Han21]. Henrik [Han21]. Herbert [McC15]. Herbjørn [Han21]. Herd [SZZ16]. here [SYF17]. Heritage [Bea17, Cho19, KC20, LI22, HNX18, KGS19, MdSLdO^{+}23, SM22, Cus15]. Herman [FM14]. heterogeneity [QZWC21]. heterogeneous [JSY^{+}16, KBH^{+}22, LDS^{+}19, PBN^{+}20, QH20, CdRM21]. heuristics [Bor15c, Du14a, SGLF17]. hidden [AHJY18, LMP21, OHT^{+}16, ZXLC18]. Hierarchical [SZJ^{+}21, MBML24, MŽA17, Sil17, SAW21, SHQ^{+}16]. hierarchy [GK14]. high [BH15, Cri15, Kim15, KSL19, LWPG^{+}19]. high-quality [BH15]. high-tech [LWPG^{+}19]. Higher [JRL20, JAG^{+}20, MR19, MAAL^{+}16, PRM22, PDŠB16, RD16, SBIT14]. Highly [BLB16, AKK^{+}17, BB15, BBH15, BK14, ESWC18, HK14, JKS^{+}16, RN16, RD16, ZCW17, PS14b]. hijacked [Aba24]. Hispanic [BNG14]. Hiss [Lev15]. Historian [SS14b]. Historical [Kea22, KKZS20, Daq20, FA17, Hjø15b, JKS^{+}16, KS14a, KL22, MBBL14, MGK20, OBD^{+}22, SS14b, SE16]. histories [CL17]. historiographic [HGCKDR16]. History [Buc15, Huv22a, Web20, CBAL15, KL22, MKK17, SA15, SE16, ZGCB15, See17, Sko23]. HIV [KP18, MVL^{+}14]. HIV/AIDS [MVL^{+}14]. Hobohm [Han21]. Hoboken [Bru18]. hoc [DLA22, MNNK15]. holding [WMX^{+}22]. hole [MB20]. holes [GK14]. Holistic [PG21, HLB^{+}17]. homelessness [SFL^{+}23]. Homophily [ZBDX18]. Honor [Uns18]. hot [LdN17]. hotel [ZW23]. hour [Sea20]. House [Afz22]. household [BVH15]. Howes [Rog15]. Hub [BGS22]. Huertas [Gui17]. Human [FHW17, HPLR23, Mat15, O’H18, Rog15, SAHZ21, DL22, FMD19, GHKB^{+}17, Gor20a, Hjø15c, LJK21, LGZ14, LMM^{+}22, MOM^{+}15, MVA22, MAM19, NBH16, PS14a, PTGL17, SV17, SDUDL24, She19, ST16c, SH14, WvdHS16, WXSL23, YR23, ZZ21a]. human-assigned [ST16c]. Human-centered [SAHZ21, She19, WXSL23]. human-computer [NBH16]. human-information [PS14a]. human-oriented [MVA22]. human-robot [YR23]. Human-supervised [HPLR23]. humanitarian [VGN20]. humanities [BTMS16, CS16c, CC17, Coc22, DM14, FW19, GW18, HH18, KGP^{+}20, LB22, ML22, MT14, OBD^{+}22, OVE14, Poo17, PG20, SEZG22, TK17d, WCdF^{+}22, ZSG^{+}22]. humility [Gor22]. Humphreys [Hal19]. hurdles [LBWE^{+}23]. Hurricane [VO23, LBG^{+}21]. Hybrid [PSB17, JML22, NB23, NE17, PSB16]. hybridization [Cj15]. Hype [Afz22, CS16a]. Hyperlinks [OMTSLC15, KT14b, Vau16, YCK16]. hypertension [JWWV18]. Hypertext [Fin17]. Hypervisibility [KP18]. hypothesis [FN17a, JJC19].

I-distance [DBBJ16]. I3 [YL14]. iAggregator [MTY14]. Ian [Huv20]. Ibekwe [Car14b]. Ibekwe-SanJuan [Car14b]. Ibero [GTAS22]. icons [HBS15]. ICT [SH19]. ICT4D [FHO23, HDT23, Meh23, PHJ23, TLS^{+}23, ZDP^{+}23]. ICTs [Cha14, NT17a]. idea [LYC^{+}22, MSPP21]. Ideas [Kos20, CCR24]. Ideation [MHJ19, LM15a, Xia14]. Identification [CL17, MGIGW16, XBCH18, IS22, YB18]. identifiers [BD16]. identify [LL17, ZCW17]. Identifying [AGM19, BBM16, KGS19, LS18, LDS^{+}19, MC18, MRKT15, SC17, SJK15, XDDG18, ZZZL17, CCLL14, PSY^{+}21, St.17, Wan18]. identities [DS15, Mad16, Fou19]. identity [FLJD23, JDZ24, Kit19, LYL21, MO20, SWL18, Wu19]. ideology [Mar16]. idiomatic [MD16]. If [MWMB18, Bor16a, DD16, KW14, vdVZK21]. iField [ZSAW23, WXSL23, ZWH^{+}23]. iFQ [OSvE16]. Ignorance [Gre19]. II [Lev20, Soe18]. III [Soe18]. illicit [CMB17]. Illness [Zha19, BV21, Che15]. illustration [FW17]. ILT [PDŠB16]. I’m [SSA20]. Image [Ano15a, Ano15b, Ano15c, Ano15d, Ano15e, Ano16a, Ano16b, Ano16c, Ano16d, Ano16e, Ano16f, Ano16g, Ano16h, Ano16i, Ano16j, Ano16k, Ano16l, Ano17a, Ano17b, Ano17c, Ano17d, Ano17e, Ano17f, Ano17g, Ano17h, Ano17i, Ano17j, Ano17k, Ano17l, Ano18a, Ano18b, Ano18c, Ano18d, Ano18e, Ano18f, Ano18g, Ano18h, Ano18i, Ano18j, Ano18k, Ano18l, Ano19a, Ano19b, Ano19c, Ano19d, Ano19e, Ano19f, Ano19g, Ano19h, Ano19i, Ano19j, Ano19k, Ano19l, Ano20a, Ano20b, Ano20c, Ano20d, Ano20e, Ano20f, Ano20g, Ano20h, Ano20i, Ano20j, Ano20k, Ano20l, Ano21a, Ano21b, Ano21c, Ano21d, Ano21e, Ano21f, Ano21g, Ano21h, Ano21i, Ano21j, Ano22a, Ano22b, Ano22c, Ano22d, Ano22e, Ano22f, Ano22g, Ano22h, Ano22i, Ano22j, Ano22k, Ano22l, Ano23a]. Image [Ano23b, Ano23c, Ano23d, Ano23e, Ano23f, Ano23g, Ano23h, Ano23i, Ano23j, Ano24a, Ano24b, Ano24c, Ano24d, ATKD^{+}17, ATKD^{+}18, Bea16, Chr17, CML^{+}15, LXLL14, LWL^{+}18, lLTBH15, NSK17, TR17, THC15]. images [HD22, HSK15, KLG14, LYCX14, SCIF20, TGV^{+}16]. imaging [MAS^{+}15]. Imagining [AC18]. imbalance [WYZL21]. Imitating [DB15]. imitation [MDL^{+}18]. immersive [Rob15]. immigrant [FLJD23, SH19]. Immigration [Ano17n, CZ16]. Impact [BM20, GVB17, GVF^{+}24, Hal20, JVF19, KHZ^{+}22, LBW19a, LYL21, SRM^{+}17, AT14, ACS18, APRRC15, BS16, BM15a, Bor17, Cri15, Egg14b, GD18, GRSFV15d, GRC21, GRS16, HH19, HBH^{+}16, HGCKDR16, HLM16, HQH^{+}22, Kel21, KT16a, KT17a, KA22, LBWE^{+}23, LGST15, lLTBH15, LHXS22, Mat22, MHMW17, MDC^{+}16, Meh23, MH15, ML16, MWD17, OMTK17, PRRC18, Pra14b, Pra15a, RCM22, RDPN21, RHG23, SRYL17, ST16b, SZWS22, TLD17, TW16a, TK17b, The20, WL14b, WZ18, WVvLZ23, WTvR16, WK19, XBL^{+}22, Yan15, YZ17, YL14, ZCvdH15, ZZ21b]. impact-driven [Meh23]. Impacts [EA18, KT16b, KTA17, YLHW21, YZYR21]. impaired [STS14, STS15]. Impairments [BM20]. impeding [St.17]. Imperfect [ZCvdH15]. implement [Meh23]. implementation [ALCD14, LYC^{+}22, PSB17, ZYS17]. implementing [JPF16, YP16]. implication [TGC^{+}21]. Implications [ZSC18, HM14, HBH^{+}16, HHL21, LCK16, Mak20, YZ17, YR23]. Implicit [HWC17, GIEVVGS19]. Importance [Tei16, JLC16, WCO21]. important [KT17a, MWD17, WL14a, YRWH18]. Imposing [DB15]. impression [LB19a]. impressionistic [UM23]. improve [BSVZ17, Cho16, ZYS17]. improved [Gia16, Mat20]. improvement [STF19, SR16]. Improving [AFRH18, BK19, CDGN15, CML^{+}15, PMD14, PS19, SY22, SK14b, ZGPB17, LYOW15, PKK^{+}24, PS14a]. imputation [MBD15]. In-Depth [MS19]. in-links [OMTSLC15]. in-situ [HWW^{+}17]. In-text [ZS14b, ZSC18, SAA^{+}23, ZS16]. inbox [KR15]. incessantly [Hil14b]. incident [ADM^{+}20]. inclusive [YC20]. incognizance [St.17]. inconsequential [Pra15b]. incorporating [LLD^{+}17, Du14b]. incorporation [McC15]. Increase [CKR14, DMdO16, SMRR16]. increases [Bor17, TSY15]. Increasing [LCNP17]. Incremental [SGLF17, BDA17, ECJF21]. independent [AMM^{+}19, RPK16, Smi15, Smi17a]. Index [Hal20, HK14, Lev15, Bar15, BH17a, CZZY15, Cri15, Fas18, Fia14, LCRGTS14, Mal15, NSK17, Pra14c, RGJCTS16, Rou15, Sch15, YL14, GVF^{+}24, LPW14, VC15, WL14b]. indexed [MRKT15]. indexer [RMDF23]. Indexing [CSOD16, WLC16, Ano18-27, BRL20, CK14, Chr17, Göd16, GSB^{+}16, LM15b, LML18, SP17, YCK16, Ols16]. indicator [BH17b, DBBJ16, HLM16, LB15, Pra14a, WMM^{+}23, WK19, YL14, ZRG16]. Indicators [Agu18, HH19, O’C16, CZW15, DB17a, HH18, HBH^{+}16, Hen14, LCRGTS14, Pra14d, TKS^{+}23, VCG14, Vin17, YZ17, ZSX15, ZvSvB14]. indices [IABL16, Pra14b]. Indigenous [Coc22, HDT23, LYL21]. indirect [KS14a]. individual [BBM16, GBS15, JDZ24, KS16, LZA^{+}20, POK15, RY18, WL14b, ZS14a]. individualistic [Ma21]. individuals [CY22, GM19, TJ20]. Indo [AMM^{+}19]. Indo-European [AMM^{+}19]. indoor [RTS^{+}17]. inequalities [CLSS14]. inequality [CHLRC14, CK22, Hil14b, YZ16]. Infectious [CKR14]. Inflated [Hen14]. inflectional [JKS^{+}16]. Influence [TND15, BGS22, CLL15, DL22, FLJD23, Jan15, KCL14, Kim15, KS14a, Li14b, LTD^{+}15, LLD^{+}17, LWM^{+}19, LJLH17, LCWL18, MPT15, MAAL^{+}16, OHP16, PGK24, PB15, PB16, RR17, Rou16, SL24, SKZ^{+}14, SLCS18, YLHW21, ZTLV15]. influenced [SRYL17]. influencer [RJS17a]. influencers [RVGPS19]. Influences [LBW19a, SL21a, WS22, XA14, ZWTX23]. influential [HR19]. inform [ADL15, CJ16b]. Informal [PYW18]. informality [DB15]. Informatics [Bea17, Gre19, Meh22, Mos17, SV20, Ano17n, CZ16, Mar16, MSW^{+}19, PRM22, PPC19, PPLJ20, RAMP22, WCO21]. Information [Agu18, Alb15, AGBK19, Ano14, Ano16n, Ano16o, Ano16p, Ano16q, Ano16r, Ano16s, Ano16t, Ano16u, Ano16v, Ano16w, Ano16x, Ano16y, Ano17o, Ano17p, Ano17q, Ano17r, Ano17s, Ano17t, Ano17u, Ano17v, Ano17w, Ano17x, Ano17y, Ano17z, Ano18n, Ano18o, Ano18p, Ano18q, Ano18r, Ano18s, Ano18t, Ano18u, Ano18v, Ano18w, Ano18x, Ano18y, Ano19m, Ano19n, Ano19o, Ano19p, Ano19q, Ano19r, Ano19s, Ano19t, Ano19u, Ano19v, Ano19w, Ano19x, Ano20m, Ano20n, Ano20o, Ano20p, Ano20q, Ano20r, Ano20s, Ano20t, Ano20u, Ano20v, Ano20w, Ano20x, Ano21l, Ano21m, Ano21n, Ano21o, Ano21p, Ano21q, Ano21r, Ano21s, Ano21t, Ano21u, Ano21v, Ano21w, Ano22m, Ano22n, Ano22o, Ano22p, Ano22q, Ano22r, Ano22s, Ano22t, Ano22u, Ano22v, Ano22w]. Information [Ano22x, Ano23k, Ano23l, Ano23m, Ano23n, Ano23o, Ano23p, Ano23q, Ano23r, Ano23s, Ano23t, Ano24e, Ano24f, Ano24g, Ano24h, BLB16, Baw14, BRS15, BR15a, BM20, BV21, BJR19, Buc20, CCMT15, Car14b, Che15, CK21a, CM23, CK17, Col17a, Col20, CK21b, Cox15, DDB16, EMK^{+}15, Erd18, Far23, Fou19, Fre18, GW18, Gor19, Gor20b, Gre18, Hai16, He18, Her23, Hjø17a, Huv20, HEEB^{+}22, JKS^{+}16, JDHL17, JWWV18, Jul20, KSB20, KS14b, KC20, KP18, LKKP16, LOM22, Lin15, LCL16, Lue15, Ma23a, MHJ19, Mak20, MdSLdO^{+}23, Mat15, MVL^{+}14, MOM^{+}15, MAP15, MS19, NTNP19, NLFM19, O’H18, OMM^{+}14, Ols16, OM15, PKA20, PH22, PWK19, Rou17, RBJ22, RR20, Rut24b, Rut24a, Sar16, Sav18a, SS14a, Sha14, Soe18, TK17d, VCJ22, VO23, Wei14a, Wor19, YZ16, ZS14a]. Information [ZZ20, ZPPW24, ZVW20, AJWB16, AG15a, ADD^{+}18, Adl23, Aji17, AFB20, AH17, Ano18z, ADL15, AGJ^{+}16, Bad14, BK23, BR15b, BR20, BCAL15, BS16, BBC^{+}22, Bur14, CAS21, CO16, CPP^{+}16, CLSS14, CBH16, Che22, CLL15, CK16a, CHJ20, CS15a, Cho20, CPL22, CN14, Cho23, Cho16, CKL16, CS15b, CC17, CBM17, Col17b, CMB17, CV20, CK22, DO14, DL22, Du14b, DS21, EH21, Eng17a, Eng17b, Esk20, FA19, FVGL14, FC22, FC18, Fre15, FHO23, Fro21, GGS22, Gau16, GM19, GLR16, GSB^{+}16, GISW15, GTAS22, Gor22, GEBP22, GJ22, GJ16, GLZ^{+}24, GMW19, HCC^{+}21, HB21, HWW^{+}17, Har14, HRB14, HSSE20, HABT18, Hil14a, Hil14b, Hjø15b, HB15, Hof17, HLB^{+}17, HK17, HBS15, HWW15, HK21, HAJ^{+}17, Huv22b, IB16, JBD^{+}19, JT17, JW17, JDZ24, JJ14, JRS16, Joh17]. information [JWJ^{+}19, JRL20, KT14a, KvOS^{+}17, KZ23, Kel21, KDS^{+}17, KKO15, KWLV22, KMD16, Kno24, KSL19, Kon16, KKK24, Kos21, KCZ^{+}14, Kue23, KL22, LCH^{+}22, LM15a, LCK16, LB19b, LBG^{+}21, LS22, LSHH23, Lev20, LHS24, LB16, LBA16, LS14, Li21b, LWLH18, LMA^{+}22, LB15, LLB16, LJ17, LLB20, LO19, Lop14, LMM^{+}22, LPB19, LI22, LJLH17, ML21, Ma21, Mad16, MBW^{+}14, MB20, McD21, MAS^{+}15, MR17, MJ23, Meh23, Men21, MŽA17, MK15, Min16, MNNK15, MP19, MWPF17, NGTKR15, NB23, New23, NF15, OR21, OS15, OHT^{+}16, Oh17, OB19, ONZ23, Oli21, Olt15, PLPQB20, PS14a, PSY^{+}21, PPLJ20, PGFG^{+}20, Pil14, PMGB^{+}21, PEG^{+}19, PDŠB16, PG21, PB15, PKB^{+}16, PBG^{+}17, PT21, PHJ23, RJS19, RDS20, RAW22, Roh21, RLAB^{+}24, RJS17a, Rub15, RR21, RBJ18a, RBJ18b, Rut22, SVK17, SRPS16, Sav17, ST22]. information [SZ20, Sha21a, She16, SSA22, Sin16, SA15, Sko23, Sny14, SLLW20, St.17, Sta18, SWL18, SLeH21, SH19, SEZG22, SWSZ19, SZWS22, SKC18, Tab14, TG15, TBP^{+}15, TMDZ21a, TMDZ21b, TT17, TZS^{+}22, TLA^{+}16, TGC^{+}21, TR16, THWS14, TJV14, VG18, Vau16, VP19, VFZ^{+}22, VSJR19, WW16, WHX^{+}14, WL20, WG19, Wat14, WWL16, WVvLZ23, WB15, WH15, WJB22, WV15, WWB21, WFRN14, WCO21, WvdHS16, WLZB14, WLW^{+}17, WVZ20, XHM^{+}20, YR16, YZQT20, Zha14b, ZBKG14, Zha14a, Zha14c, ZLCB15, ZSX15, ZZY23, ZS14c, ZDP^{+}23, ZGL^{+}22, ZGB16, ZYS17, ZB19, ZH17, BR15a, Bur20, Mak20, VT16, Huv22a, Huv20, Kea22, Lev15, Ma22b, Ma22a, Uns18]. Information-Based [MHJ19]. information-searching [BS16]. Information-Seeking [KP18, Sav18a, CHJ20, CPL22, Fre15, Gau16, RJS19, RJS17a, Sav17, Sin16, TT17, TZS^{+}22, WFRN14]. information-sharing [CKL16]. Informational [CY14, Sav15a]. informed [EH15, Har14]. Informetric [Egg14a]. Informetrics [vR17a, Hjø17b]. informing [HEEB^{+}22, Mak20, PHJ23]. infosphere [MB23]. infrastructure [ESD22, FA17, MdSLdO^{+}23, She19]. infrastructures [BSG19, DDB16, MHMW17]. infrequent [RN16]. INFUSE [TND15]. ing [BD16]. initial [KD16, RCM22, YLHW21]. initial-based [KD16]. Initiated [LBM^{+}19b]. initiative [HWW^{+}17]. injustice [JW15, Oli21, PSY^{+}21]. innovating [WLS23]. Innovation [MDL^{+}18, ZDZ21, LPU15, LWPG^{+}19, LYC^{+}22, Poo17, She17, ZDW18]. innovations [FMG16, LA14, SL24]. inputs [Hoo17]. Inquiry [HKBB18, CBAL15]. inquiry-based [CBAL15]. inscriptions [ML22]. insider [HKBB18]. Insight [GHKB^{+}17]. Insights [KZ23, DTDT21, TPZ15]. inspire [She19]. Instagram [FR23]. instant [OD16, YB18]. instead [BM14a]. institute [BTMS16, Bor16a]. Institutes [WYZL21]. institution [BdMA14]. Institutional [KS16, PSB16, BD16, BM15a, Bor16a, HRL21, JPF16, KN23, KISS15, LSHH23, May16, RKHMRH21, STD15, Sil20, SC21]. Institutions [BdMA14, BLB16, BB15, CC22, GG16, JRL20, MR19, PS14b]. instruction [Smi15]. instrument [SKC18]. integral [Esk20, LMM^{+}22]. Integrated [OBD^{+}22, BOHK19, GK14, YL14]. Integrating [JBD^{+}19, VG18, MKK17, MBVB17]. integration [ADM^{+}20, LI22, OR21]. Integrative [Tho22]. Integrity [Mos17, Jam17]. Intellectual [MHJ19, Med15, CL17, Gor22, Gor24]. Intelligence [LWM^{+}23, Cox23, JLB^{+}23, JJ14, LCH^{+}22, LGZ14, NPG^{+}20, ØJW^{+}21, PPC19, PdlRR^{+}22, RRHL17, SKO24, SK23, SFG^{+}23, WLS23, YP16, SHKL23]. intelligence- [SHKL23]. Intelligence-written [LWM^{+}23]. intelligent [She19]. intensive [LWPG^{+}19, SDUDL24]. intent [JBD^{+}19, O’C16]. intention [HLC15, KT14a, MBVB17, POK15, XLKC22, ZPL^{+}20]. intentions [KBC17, MVL^{+}14]. intents [KWR20]. inter [MYK15]. inter-post [MYK15]. Interacting [FT16, GBS15, XBCH18]. Interaction [JDZ24, Lan15, LBM^{+}19b, Rog15, CV20, JL18, KKK24, KL22, MF16, MD16, NH15, NBH16, PS14a, QZWC21, RJK20, TG15, ZF19]. interactionist [WV15]. Interactions [Che22, HCC15, CCC14, CLL15, CdRM21, DL22, GKM^{+}21, JRS16, SKA18]. Interactive [RJK20, GIEVVGS19, GN17, JSLC16, LCP15, LdMAGB15, LdMAdN16, LQ14, PHC17, PB15, WN15, YXAL23]. intercultural [WG19]. Interdisciplinarity [LG14, HL16, LdMAGB15, LI21a, Mad16, ZRG16]. interdisciplinary [Eng17b, GVB17, LI22, MD16, OBD^{+}22, RD16, SVC22, ZGE22, Gui17]. interest [ASTU16, ALC^{+}14, SS14c, vdSvdB23]. interesting [vdSvdBG^{+}14]. interests [DS21, SB18]. interface [Li14b, NH15, STS15, SK14b, ZBKG14]. Interfaces [GLR16]. interhuman [LI14a]. intermediaries [BJR19]. intermediary [BR16]. intermediates [WV18]. International [All19, Ama19, Fre18, Fur20, Jul20, LPW14, LBW19a, MWD17, OB19, CH16, CYP15, GK24, OMTSLC15, PS14b, PG20, RD16, ST16b, TLA^{+}16, YRWH18]. internationalization [LPW14]. Internationally [Ama22]. Internet [Dow15, Lev20, Saw19, Wei14b, BM14d, HRB14, Lin14a, LXW18, NT17a, ST16a, WMS^{+}15, WB15, Dut20]. Interoperability [CA20, VT16]. interpersonal [TKB^{+}17, ZH17]. interplay [ALAB19, JT17, Min16]. interpretable [LWL^{+}18, TZH23]. interpretation [BdMA14, LB15]. interpretations [AFRH18]. interprofessional [WCTS15]. Interrater [Bor15a]. Interrupting [PSY^{+}21]. interruptions [HHSP22]. intersections [ZDP^{+}23]. intersex [KWLV22]. intervene [dRR15]. intervention [CS15b, LMA^{+}22, ZY15]. interventions [NLP16]. Intimate [Wes15]. Intrainstitutional [WLW^{+}17]. Intrigue [Lev15]. intrinsic [ADD^{+}18]. introduce [GSG20]. Introduction [Cox14, GMW19, ZVW20, Col17a]. Intuitive [MD16]. invariant [BAGL16]. inventors [YL17a]. inverse [BS17]. Investigating [AZE^{+}19, MHJ19, MB20, MPT15, RCM22, STS14, Smi17b, SKA18, SZWS22, WLTY19, GARA^{+}16, LHXS22]. Investigation [Haj20, JVF19, SRM^{+}17, AT14, DBH18, Fox20, HHPI21, LKRS18, LTD^{+}15, QLM^{+}22, SL21b, TS16, Vau16, WS22, ZJHT23]. Investigative [CAS21]. Investment [YR16, HYD^{+}20, dSMdAGQ15]. invisible [Mor24]. involvement [EWX^{+}16, POK15, XJBK17]. Iorio [Ior14]. IR [FSK^{+}16, FS18, GIEVVGS19]. IRBs [Xia23]. Irini [All19]. Iris [Abb17]. Irrelevant [KK19a]. IS&R [Hjø15b]. ISBN [Abb17, Adl23, Afz22, Agu18, Aji17, All19, Ama19, BI17, Bat21, Bea17, Ben18, Bow19a, Bow19b, Bru18, Buc20, Bud17, Cho19, Cle18, Clo18, Col17a, Col17b, Con20, Dor19, Dut20, Erd18, Far23, FJ19, Fou19, Fra21, Fre18, Fri17, Fro21, Gor20a, Gui17, Hai16, Hal19, Han21, Haw18, Hay18, He18, Hjø17a, How17, Huv20, Huv22a, Jon17, Jul20, Kea22, Kno24, Kos20, Kos21, Lev20, Ma22b, Ma22a, Mat20, Meh22, MS19, O’H18, Ols16, Owe21, Pan21, QH18, Rey21, Rhe17, Rou17, Sal20, Sav18b, Saw19, See17, Sha21a, Sha21b, Sko23, Soe18, Sto20, Swa22, Uns18, Web20, Whi17, Zha19, Zim17, dP20, vR17a]. ISBN-13 [Agu18, Saw19]. iSchool [SAHZ21, WCdF^{+}22]. iSchools [CK17, KC20, UM23, ZZE17]. ISI [MRKT15]. ISI-indexed [MRKT15]. Islamic [Sto20]. Isolated [RBJ18a]. ISP [BCAL15, CBAL15]. Israel [ZGB16]. Israeli [BM14d]. Issue [Ano16n, Ano16o, Ano16p, Ano16q, Ano16r, Ano16s, Ano16t, Ano16u, Ano16v, Ano16w, Ano16x, Ano16y, Ano17o, Ano17p, Ano17q, Ano17r, Ano17s, Ano17t, Ano17u, Ano17v, Ano17w, Ano17x, Ano17y, Ano17z, Ano18n, Ano18o, Ano18p, Ano18q, Ano18r, Ano18s, Ano18t, Ano18u, Ano18v, Ano18w, Ano18x, Ano18y, Ano19m, Ano19n, Ano19o, Ano19p, Ano19q, Ano19r, Ano19s, Ano19t, Ano19u, Ano19v, Ano19w, Ano19x, Ano20m, Ano20n, Ano20o, Ano20p, Ano20q, Ano20r, Ano20s, Ano20t, Ano20u, Ano20v, Ano20w, Ano20x, Ano21l, Ano21m, Ano21n, Ano21o, Ano21p, Ano21q, Ano21r, Ano21s, Ano21t, Ano21u, Ano21v, Ano21w, Ano22m, Ano22n, Ano22o, Ano22p, Ano22q, Ano22r, Ano22s, Ano22t, Ano22u, Ano22v, Ano22w, Ano22x, Ano23k, Ano23l, Ano23m]. Issue [Ano23n, Ano23o, Ano23p, Ano23q, Ano23r, Ano23s, Ano23t, Ano24e, Ano24f, Ano24g, Ano24h, LHS24, AGBK19, Ano18z, GMW19, MWPF17, TMDZ21b, WJB22, ZSG^{+}22, ZSAW23, ZDP^{+}23]. issues [HH14, May16, TS21]. Italian [Riv15]. Italy [AD15]. item [BSVZ17, GARA^{+}16]. iterative [PB15]. Ivan [He18].

J [Ama19, Bat15, Buc20, Bud17, Clo18, Rog15, Sav18b, Soe18, Sve15, Tei17, Wei14b]. J. [Jan15]. J.Berkeley [Soe18]. Jack [Hai16, Sko23]. Jairo [Mat20]. James [Ben18, Kno24, Kos20]. Jan [Rey21]. Janet [Rhe17]. Jannica [Huv20]. Jannis [Lan15]. Japan [KIMM17]. Jared [Uns18]. JASIS&T [WJB22]. JASIST [Ano18z, LHS24, MWPF17, ZWW18, ZSAW23, ZDP^{+}23, ZSG^{+}22]. Jean [SR14]. Jean-Francois [SR14]. Jeffrey [How17]. Jehane [Sea20]. Jelusić [Cho19]. Jerks [BH15]. Jesse [Bow19a]. Jochumsen [Han21]. John [Bru18, Buc20, H17]. Johnson [Far23]. Join [RR21]. Joint [CHT^{+}15, ZZL20, QH20]. Jonathan [Ma22b, O’H18, Wei14b]. Jorge [Adl23]. Joshua [Sha21a]. Journal [Ano14, Hal20, KHZ^{+}22, LdMAGB15, LHR16, AGMR18, BELR16, CGGS17, CJ16b, GNVQdMA16, Hal17, HM18, HL16, Hsi17, Kim19a, LSM^{+}14, LG14, LdMAdN16, LdN17, ML22, McC15, MBZ^{+}22, NF15, Ort22, OVE14, PR16, PSB16, SB18, SRYL17, SKL17, SBA^{+}15, TJV14, WCP^{+}19, WW15, WHZ23, NF15]. journalism [Tei17, YR16]. journalistic [SHKL23]. journals [Aba24, BLGbH21, CCMT15, CCLL14, Epi14, GG16, Hal14, HM14, KPWS20, LMJ21a, LMJ21b, LLG14, MT16, MZ22, MH16, NE17, PSB17, RA14, RPK16, RD16, RJS^{+}17b, She21, SRYL17, SMYL15, ŠF17, VS15, WHZ23, XHC^{+}15, YL18, ZRG16, ZAE^{+}19]. journey [Du14b, Owe21, PBG^{+}17]. judgment [BM14d, ZGBIL18]. judgments [Du14b, LPB19, PSS22, WHX^{+}14, Wat14]. Julia [Ama19]. Julian [Owe21]. Jumping [Tab14]. just [Vas15]. justice [Adl23, FC22, JW15, Meh23].

K-pop [SJK15]. Kaixin001 [ZXPS18]. Kaja [Ama19]. Kallinikos [Lan15]. Karim [Sea20]. Kathrin [QH18]. Katriina [Huv20]. Katsiaryna [QH18]. Katy [Whi17]. Keeping [PBG^{+}17, DO14, SBH21]. Kelly [Uns18]. Kennerly [Ma22b]. Kent [Wil14a]. kernel [PHH^{+}20]. Ketchikan [Soe18]. key [DTDT21, LL17, WMX^{+}22]. Keyphrase [ZZL20, ZZ21a]. keyphrases [MRAF16]. keyword [CK14, CML^{+}15]. keywords [BRS16, HS22, LK14, PGFG^{+}20, ZYZ^{+}16, PDS22, Ma22b]. Kimiz [Haw18]. Kind [LBM^{+}19b, Bar14]. Kindle [Saw19]. kiss [LPL14]. Kleine [Cha14]. Knautz [QH18]. knot [LO19]. Know [ØC19]. Knowledge [Ama19, BDA^{+}18, Buc15, Car14b, CHH^{+}17, CBAL15, Gui17, HLC15, KB17, Med15, QZWC21, SLS^{+}19, SC21, SH14, Whi17, Ama22, AGBN16, Bad14, Bar14, BL14, BSG19, BH15, CRGG17, FJT23, GTZŽ14, HDT23, HL16, Hjø15c, IQDS16, JS15, JT17, JPS^{+}19, JGCR15, KLC^{+}14, Kha19, Kim15, KN23, Lai15, LG14, LWPG^{+}19, LSL23, LYL21, LGZ14, LB15, LK16, LLB16, LJ17, LZ19, LXW18, Mah22, Mat22, MOM^{+}15, MSW^{+}19, MKK17, MT14, MRMCMVUL14, ØC19, PGK24, Par19, PPC19, PB15, QLM^{+}22, RLAB^{+}24, SL21a, SS14a, SDUDL24, She17, SW18, SK23, Smi15, Smi17a, Sta22, TKB^{+}17, TLS^{+}23, VT16, WMM^{+}23, Yan14, Yan16, YZ17, ZLCB15, ZFH^{+}19, ZWTX23, ZS14c, Adl23, Haw18, Swa22]. knowledge-based [MRMCMVUL14, VT16]. knowledge-diffusion [JGCR15]. Knowledge-dissemination [CHH^{+}17]. knowledge-intensive [LWPG^{+}19, SDUDL24]. Knowledge-sharing [HLC15, JS15]. Known [BM20, WMX^{+}22]. Knox [Buc20]. Korea [POK15]. Korean [LS22, SH19]. Korean-born [LS22]. Korean-born/Korean-dwelling [LS22]. Korean-born/US-dwelling [LS22]. Korean-dwelling [LS22]. Krystyna [Abb17]. Kuhlthau [BCAL15, CBAL15].

L [Hak14, Ior14, Sav18b, Ten16]. Label [PS20]. labeling [EHdBvO17, MC18, MSTP^{+}22, SAW21, ZTA20]. labelling [HWC17]. labels [GN17]. Labor [Dow15, Hod21, LZA^{+}20, dFB22a]. laborers [BNG14]. lag [HDZ^{+}15]. LAGOS [ZLY23]. LAGOS-AND [ZLY23]. land [OB19]. landing [GBL^{+}18]. landscape [LB22]. landscapes [LB19b]. Language [MSPP21, AMM^{+}19, BNG14, DGR19, DLD^{+}21, Eng17a, HB21, HB15, KKO15, Kim15, LMMO15, LBA16, LWM^{+}23, MSI^{+}19, MGIGW16, RL15a, SP17, SL21b, TPR17]. language-independent [AMM^{+}19]. languages [AMM^{+}19, JKS^{+}16, LMMO15, RBJ22]. Lanham [Aji17, Bea17, Lev20]. Large [WHYJ17, ACS18, Ano17m, BC16, Bur14, FN17a, GARA^{+}16, HABT18, KTS^{+}15, KD16, LWM^{+}23, OSD22, RMDF23, WCP^{+}19, ZLY23, GMB14]. Large-scale [WHYJ17, ACS18, Ano17m, BC16, FN17a, HABT18, KD16, RMDF23, WCP^{+}19, GMB14]. Larsen [Han21]. latent [FdGPC18, SLLL17, WLC16]. Latin [CMS23]. Lau [Clo18]. Law [Ben18, SR14, Wei14b, BS17, iCV18, HHC^{+}14, MA18b, RPK16, Smo17, TGC^{+}21]. laws [TGC^{+}21]. Laypeople [CHJ20]. Lazar [O’H18]. LDA [LX17]. Lea [Ma22a]. lead [HDZ^{+}15]. lead-lag [HDZ^{+}15]. leader [SLE21, WZFZ22]. leaders [GKM^{+}21]. leadership [HWZ20, HWZ22, LCB^{+}18, RY18, TPZ15]. Leading [VCJ22, LPW14, PS14b, SLE21]. learn [BCK17, VZRP15]. Learned [TCP15, SBH21, ZS16]. learners [Kim15]. Learning [BSZ^{+}14, BH15, KWLV22, PSS22, Sil17, SYG17, TG17, ADM^{+}20, BDA17, CLL15, CTM16, CML^{+}15, GKM^{+}21, GLZ^{+}24, IB16, JRL20, JAG^{+}20, KF16, KDS^{+}17, KKOS21, KSL19, LKY^{+}20, MGKA16, O’B17, PRM22, QSH14, SGV14, ST16c, TZH23, WL20, WH15, XLL16, YH16, ZW23, ZvSvB14, Kim15]. learning-based [CML^{+}15]. least [KT14a, OHP16]. lecture [KK16]. Lectures [Agu18, Fou19, MS19]. led [DJLS22, Moo20]. Lee [Hal19, Lin14b, MSdGS19]. legacy [BM23]. legal [CJ22, LCHL14, WVvLZ23]. legitimacy [Sil20, SL24]. Leiden [BdMA14, BM15a, Egg14a, LBM19a, MD15, PRRC18]. Lem [Bru14]. lens [Coc22, Ma21]. Leo [Dor19]. Leonardi [Lan15]. lesbian [KWLV22]. less [TF21]. Lessons [SBH21, TCP15, ZS16]. Let [LN19]. Letter [Bor14c, Bor14d, Bor14b, Bor16b, Bur17, CB14, Mah17, MB14, Mir14, Pra14d, Rou15, Tei16, TSRG16, VCG17, VR17b, ZMJ14]. Letters [Bor15c, BBH15, Bor15b, FN17b, Hjø17b, JVF19, LWB16, Man22, MMALCOM15, Pra15a, Pra15b, RR15, TD15, dC16, vdB16]. Leung [Adl23]. level [Bor16a, CHT^{+}15, CHLRC14, DPN16, DTW16, JML22, KBD^{+}14, LG14, LBM19a, LTD^{+}15, LLD^{+}17, RL15b, She17, WZK^{+}18, WLW^{+}17, Yan15, YY16, ZFH^{+}19, vLTW18]. Levels [CBvEW18, LPU15, LWPG^{+}19, LCHL14]. Leveraging [BRS16, Fre18, WL14b]. Lexical [XCL^{+}21, Ama22, KKO15, MRKT15, VAGR15]. lexical-based [Ama22]. lexicons [Eng17a]. LGBTQ [Kit19, MJ23]. LGBTQIA [KWLV22, JDZ24]. Liang [MS19]. Libcitations [WZ18]. Liberating [Mad16]. LibGen [Cab16]. librarians [Fro21, HCC15, OBD^{+}22, SCIF20, YLL20, Zem17]. librarianship [YLL20]. librarie [DSBG21]. Libraries [Abb17, CA20, Cho19, Ack21, CC22, Cho16, CJ16b, CCJ16, CKLP17, GAS^{+}20, HCS21, MM20, MJ23, OSD22, RCM22, XBCH18, Han21]. Library [CK21b, DSBG20, MCSB19, MH18, OM15, RMDF23, ZZ20, Adl23, Aji17, Cox23, Eck18, GARA^{+}16, GJ22, LL17, LB16, Li21b, LYWF14, ML21, PGFG^{+}20, She16, SSdMCM20, TR17, TJV14, WW16, XBL^{+}21, Zem17, BLB16, KBC17, VCJ22]. LibraryThing [Wor19]. license [MMFA20]. Life [Cox14, BM23, CPL22, CKL16, Cro14, HCC^{+}21, HK21, JDZ24, MB23, SJN14, Sin16, Hal19]. lifecycle [DB17b, Roh21, SF20]. lifecycles [Sta22]. lightweight [YXAL23]. Like [GB16, Kit19]. Likelihood [KCK16]. Limitations [GRH^{+}17, ZS14b]. limited [AKC^{+}18, She16]. line [GBZ^{+}23]. lingual [AP20]. linguistic [BCK17, LBW^{+}19b, SBZ^{+}14, TDOF^{+}16]. linguistics [Eng17b, JSY^{+}16, RJGM16]. link [Lue15, STF19, VN18, ZZL^{+}14]. linkage [BCZ14, Che16, LCD^{+}14]. linkage-based [Che16]. linkages [MBML24, XZW^{+}19]. linked [Daq20, HS22, Pan21, KBH^{+}22, All19]. LinkedIn [ZGB16]. Linking [CPP^{+}16, Zha14a, CLA^{+}16, DB17a, GDO17, LMMO15]. links [ALCD14, FCT^{+}21, HPLC19, OMTSLC15, RSAS14]. LIS [BLB16, MT16, Mat15]. Lisa [Buc15, Col17b, War15]. list [SL21b]. lists [BD16, PN18, WHZ23]. Literacy [Cle18, BR20, DJLS22, Kue23, LI22, Mat20, PDŠB16, SKC18]. literary [FMG16, Sko23]. literature [BL14, Bel16, BGS22, CL21, Cox23, HPS^{+}14, Hof17, HB16, JSY^{+}16, KBS18, LKKP16, MMB17, SM22, SWLH17, WW16, WL14a, WHYJ17, WFRN14]. literature-based [WFRN14]. LiTeWi [CLA^{+}16]. Lithuanian [MM20]. Little [Ten16]. Littlefield [Aji17, Bea17, Kno24, Lev20]. live [KTZ20]. loan [GZBA22]. Local [NSK17, BPB21, LB19b, LX17, OB19]. located [AG15b]. Locating [SFG^{+}23]. Location [SLCS18, AKK^{+}17, GS22, GLR16, LS17]. Location-aware [SLCS18]. location-based [GLR16]. lockdown [BK23]. logarithmic [MA18b]. logic [CHT^{+}15]. logics [SHKL23]. logs [LZSW17]. Lois [Kra14]. London [Col17a, Fro21, Huv20, Pan21, Sto20, Mir14]. long [CL17, ESD22, HHPI21, SYG17]. long-distance [HHPI21]. long-term [CL17, ESD22]. longer [Cro15a]. longitudinal [HT17, KH17, SMW14, YL18, Zub18]. look [BH17a, WLS23]. Looking [GB17, Col17b]. looks [Kit19]. lookup [AGJ^{+}16, TR17]. loop [HRN^{+}23]. Loosen [RH21]. López [Gui17, Adl23]. López-McKnight [Adl23]. losing [RH21]. Lost [GBZ^{+}23]. Lotka [BS17, Smo17]. Louise [Pan21]. loyalty [SSdMCM20, WZL21]. luck [MBW^{+}14]. Lui [Sko23]. Luke [Hay18]. Lunin [Kra14]. Lupton [Gor20a]. lurkers [HLC15]. Lynch [Rhe17]. Lynne [Mat20].

M [Abb17, Buc20, Car14b, Col17b, Erd18, Lan15, Sko23]. MA [BI17, Bat21, Ben18, Bow19b, Cle18, Con20, Dor19, Erd18, Hal19, Haw18, How17, Jon17, Ols16, Rhe17, Sal20, Sav18b, Sha21a, Web20, Whi17, Zha19, Zim17]. machine [AKC^{+}18, BDA17, Cj15, CTM16, ECJF21, GKM^{+}21, KF16, KKOS21, Mat20, SAA^{+}23, ZvSvB14, ALYS16, Afz22, Mat20]. machine-learning [CTM16, ZvSvB14]. machines [Mos18, SLH22]. Macro [She17]. Macro-level [She17]. mail [KR15, KISS15]. mails [KISS15]. Main [JZC20, JL23, LCHL14, LK16]. Mainka [dP20]. maintaining [EH15, MSM^{+}21]. maintenance [ESD22, Hod21, LSHH23]. major [ACBC22, GJ16, HCC^{+}21, SWKK16]. make [TT17]. maker [KSL19]. makers [EH21]. makerspace [EH21]. makerspaces [Li21b, MMHA23]. makes [BdMA14, KZB17, MSPP21]. Making [Huv22b, LMM^{+}22, Mor24, Pil14, TZH23, AC18, BWGD21, GB17, GN17, Hjø15c, HK21, JDZ24, LPB19, MAP15, OHP16, SKO24, WCTS15, MBW^{+}14]. Makri [Erd18]. maladaptation [GBZ^{+}23]. Malaysia [AT14]. Male [TF21, SCIF20]. malevolent [ZRdR21]. Management [Cox15, Haw18, Huv20, ADM^{+}20, BM23, BOHK19, CCLL14, CKLP17, CK22, DO14, DJ20, DTDT21, FA19, GD18, HHL21, HCS21, KR15, LWA^{+}17, LSHH23, Mat22, McC15, MVA22, NB23, PG20, RKHMRH21, Ric18, RVdAGP21, RPK16, SDUDL24, SWL18, SLeH21, TLS^{+}23, TK17d]. Managing [DAT^{+}14, EJ14, Jon17, KP18, BV21, GS22]. Mandarin [Ma23b]. mandate [Bor16c, VLBG^{+}16]. mandates [NE17]. Mandatory [JVF19]. manifold [Kos21]. manipulate [LCRGTS14]. manufacturing [LWPG^{+}19]. manuscript [KBH^{+}22]. manuscripts [SMYL15]. many [Mir14, PHC17]. Map [ST16c, BK14, CL14, Hoo17, Whi17]. mapped [LdN17]. Mapping [BPB21, Epi14, KKZS20, WHvdZvdB17, GRBCS15, KNY^{+}14, MMB17, MdSLdO^{+}23, RRHL17, YL17b]. Maps [LHR16, BH16, LdMAGB15, LdMAdN16, LN17, TH16, YL17a]. Marcia [Soe18]. Marcin [All19]. marginality [KWLV22]. marginalization [GM19]. marginalized [CK21b, FHO23, NT17a, VO23]. María [Gui17, VO23]. markers [MSPP21]. market [CZZY15, LWM^{+}19]. Marketakis [Fra21]. marketing [Du14b]. markets [RDPN21]. Markov [AHJY18, CHT^{+}15]. Masking [Sta18]. Mass [Con20, Mos17, RDPN21]. massively [ELB17, SMW14]. mastering [HSSE20]. mastery [CPL22]. match [SAB^{+}18]. Matched [FW17, LdMAdN16]. Matching [RL15a, Don16, GLZ^{+}24, MWD17, OSvE16, YHL^{+}21, YXAL23]. material [Pil14]. Materiality [VP19, Lan15]. mathematical [YL14]. mathematics [BGS22]. matrices [ZL16]. matrix [SLLL17, YL14, DMdO16]. Matter [MT18b, AD23, JRL20, KTZ20, KPWS20, KLG14, LB19a, PTGL17, SHW^{+}15, Vas15, ZW23]. matters [LGST15, MS16, PKB^{+}16]. Matthew [Uns18]. maturity [CKLP17, TTH16]. Matusiak [Abb17]. Maximization [Rog15, SLCS18]. maximum [HM18]. maximum-normalized [HM18]. may [MPT15, dC16]. McKelvey [Dut20]. McKemmish [Clo18]. McKnight [Adl23]. McLean [Mar15]. MD [Aji17, Bea17, Lev20]. Me [SSA20, Kit19, LHXS22, Rub15]. mean [Bor16b, TZK20]. Meaning [Col20, HK21, AC18, iCV18, JDZ24, LI14a, Jul20]. meaning-frequency [iCV18]. Meaning-making [HK21, JDZ24]. meaningful [Gor19, HK21, LI22]. Means [NTNP19]. measure [ADC15, ADD^{+}18, Egg14b, FSK^{+}16, HSSE20, LSL23, MYK15, Rou15, Sch15]. measured [AT14]. measurement [BLGbH21, LKY^{+}20, LdMAGB15, LI21a]. measurements [KD16]. Measures [Pra14b, Pra15a, Fie16, FMP16, GNVQdMA16, HH18, HBS15, MGCTW15]. Measuring [Bor16c, CDL22, KLH17, KS14a, KH17, LPU15, PS14b, YL17b, YCL20, ZDW18, ZTLV15, dSBG19, AH17, TK15, WB21, CO16, DS15]. mechanism [HB14]. mechanisms [AJWB16, ZS14a, ZS20]. Mechanization [See17]. mechanology [Sha21b]. Mechtild [Sar16]. Medford [Rou17]. Media [Bow19b, BVd16, Buc15, Hal19, Hay18, Mos17, Saw19, Ano17n, AMDFHVC24, Bi22, BDH17, BKM15, CY14, CZ16, CdRM21, Du14a, FJF^{+}21, Fin17, GBL^{+}18, HCC^{+}21, HHL21, JSLC16, KKK24, KA22, LCP15, LBG^{+}21, LJK21, LZ22, LWM^{+}19, LZ21, MVA22, MSTP^{+}22, MGIGW16, OS15, Par19, PBN^{+}20, PT21, RDPN21, SBS23, SP22, Sin16, SCIF20, Sto20, SWLH17, Vau16, WZFZ22, ZCZ^{+}20, ZAE^{+}19]. media-embedded [Du14a]. mediated [Bat21, CY14, FA17, HHB17, QZWC21]. Mediating [BSG19]. mediation [IQDS16, SDUDL24, WTT^{+}14]. mediations [SE16]. medical [AF17, Ano14, ATKD^{+}18, HCDÅ15, LS21, LK14, SMAB09, TPR17, TW16b, WHYJ17, dHSM17]. medically [ATKD^{+}17]. Medicine [CKR14, RMDF23]. Medieval [PMGB^{+}21]. medium [LN17]. medium-sized [LN17]. Medline [RL15a]. Medline/PubMed [RL15a]. MedlinePlus [Zha14c]. Meese [Ben18]. mega [WCP^{+}19]. mega-journal [WCP^{+}19]. megacitation [ZvSvB14]. megajournals [SLS20]. Meikle [Saw19]. MELIBEA [VLBG^{+}16]. Members [LML16]. Memoir [The17]. Memoriam [Car14a, Kra14, Wil14b, Wil14a]. memories [GOdR18, HS14]. Memory [WSWH20, BELR16, Far23, Luy16]. men [BBH15, MVL^{+}14]. Mendeley [BH16, JHJ15, MT16, MT18a, MT14, MTHL15, MTK16, ST16b, TM15, TW16b, TS16, ZCW17]. Mental [dC16, CK14, CYYG17]. mention [WL14b]. mentioned [ESWC18, HB16]. mentions [DCLH21, FDC20, ZYL18]. menu [WH15]. Mercedes [Saw19]. Merging [Jor19]. message [ZGL^{+}22]. messages [BBM16, SMRR16, VAGR15]. messaging [OD16]. Meta [Cro14, DPHP17, LCWC23]. Meta-life [Cro14]. meta-synthesis [DPHP17]. Metadata [AKC^{+}18, FA17, Fur20, PMSA14, BRS16, HD22, Hod21, HNX18, JML22, KBH^{+}22, LBWE^{+}23, LCP15, LSL23, LQ14, LGZ14, May16, MA18a, YHL^{+}21]. Metadata. [How17]. metaphor [HHK^{+}17]. meter [RDS20]. method [Bel16, CL14, Col15, CLA^{+}16, EBD15, GISW15, HRN^{+}23, JGY^{+}19, KT15, KT16a, LLHF15, MDJ^{+}19, SGV14, SWKK16, Vec14, XLKC22, XZW^{+}19, YHL^{+}21, YR23, YXJ18]. Methodological [AD15, SV20, ALYS16, She17]. Methodology [MH18, Cho15, VCJ22]. Methods [Cox14, FC22, AFL^{+}14, Bea16, BAG17, BS17, BIW20, DC14, Fox20, GHKB^{+}17, KS14b, Lai15, LS22, LPB19, ML21, OA15, She16, She17, Skö18, WCC^{+}17, WLZB14, XLL16, ZJHT23, Hak14, Hay18, Hol15]. Metric [Dor19, RR15, ZBBW18, AH17, HM18]. Metric-Wise [Dor19]. Metric-wiseness [RR15]. Metrics [NT17b, OM15, AZE^{+}19, CdRM21, DTW16, FDC20, FCT^{+}21, HH19, Hal17, HM18, KH17, LMP21, Man22, Rou17, SBIT14]. Mexican [HGCKDR16]. MF [ATKD^{+}18]. MF-Re-Rank [ATKD^{+}18]. mHealth [ZJHT23]. Michael [Agu18, O’H18, Rhe17]. Michele [Ma22b, Sal20]. Microblog [ZZL20]. microblogging [LTD^{+}15, LLD^{+}17, LCL16]. microblogs [AFRH18, BDA17, ZZ21a]. microscope [Rou17]. Microsoft [Hal14, OA14]. middle [BCAL15]. might [Cox23]. migrants [AC18, VGN20]. migration [SBH21]. milestone [BDX^{+}18]. millennials [Kit19]. million [SS17]. Milosevic [Bow19b]. mind [OHP16, VGN20]. mine [ALYS16]. Mining [ATKD^{+}17, BDA^{+}18, KWR20, LJT^{+}15, McC15, ALMP16, GZBA22, JRL20, KCZ^{+}14, LCB^{+}18, Lai15, LSC18, LZG^{+}23, TDOF^{+}16, VGC17, VC15, Sav18b]. Minneapolis [Dut20, FJ19, Huv22a, Meh22, See17]. Minnesota [Dut20, FJ19, Huv22a, Meh22, See17]. Minty [Sto20]. minutes [Sea20]. MIREX [HLB^{+}17]. mirrored [SS14c]. misbehavior [CM23]. misconduct [CB14, GJ22, Ior14]. Misinformation [Kos20, DL22, FV21, WZFZ22]. misleading [Tei17, ZMJ14]. mismeasure [MM18]. missing [DQAMM^{+}24, Lue15, MBD15]. misspelled [SF15]. mix [ZJHT23]. mix-methods [ZJHT23]. mixed [Fox20, GEBP22, HWW^{+}17, LS22, PEG^{+}19, WCC^{+}17, XLKC22, YR23]. mixed-initiative [HWW^{+}17]. mixed-method [XLKC22, YR23]. mixed-methods [Fox20, WCC^{+}17]. mixture [ST16a]. Mizzaro [He18]. MMR [KK19a]. MN [Dut20, FJ19, Huv22a, See17]. Mobile [STD15, Ama22, BNG14, CBJ19, FvdWR^{+}22, GLR16, GBL^{+}18, GMRS17, HH14, HHSP22, JT17, KTS^{+}16, MS16, RDPF16, TPZ15, TG15, WHX^{+}14, WZL21, XDDG18, He18, MS19]. modal [LZZ^{+}21]. modality [ATKD^{+}18]. Model [BCAL15, CBAL15, AWC18, AHJY18, AvS16, AGM19, ATKD^{+}18, BF18, BK14, BVH15, CHT^{+}15, CBH16, DLA16, DLA22, DGGdST19, FdGPC18, Göd16, HB21, HB14, JSLC16, KT14a, LTD^{+}15, LZSW17, LQ14, LX17, Lop14, MG21, Men21, MP19, QH20, RDPF16, SJN14, SB16, Sav17, ST22, ST16a, SBA^{+}15, SWSZ19, SZJ^{+}21, TND15, VCG14, Wan18, WSWH20, WZL21, WWL16, WCTS15, WFRN14, WXSL23, YLL20, ZS14a, ZFH^{+}19, ZHGY18, KCK16, WLC16]. Modelers [FHW17]. Modeling [AvS15, BL14, CK21a, CJ16b, HLW18, HSK15, KTPE14, LRB19, SZ20, ZZL20, AFRH18, AMOM16, BMG^{+}17, FW17, GLZ^{+}24, JBD^{+}19, KvOS^{+}17, LN17, LBA16, LX17, LJLH17, MAM19, PS19, ST16c, TW16a]. Modelling [XOL^{+}18]. models [ACBC22, AG15b, ATKD^{+}17, BR20, Bor16a, CK14, CS15a, DGR19, GIEVVGS19, GTAS22, GJ16, LWM^{+}23, Mal15, QSH14, RA14, SM22, TZH23, TKB^{+}17, TWB^{+}18, WZK^{+}18, dC16]. moderated [SDUDL24]. moderates [MS16]. moderating [AGBN16, LZ22]. moderation [MSTP^{+}22, ZFH^{+}19]. modern [AFL^{+}14, Bor14d, BM15b, Kno24]. modification [SY20, SY22]. monadic [FdGPC18]. Monash [Clo18]. monetized [Jam20]. money [NE17]. monitor [vLTW18]. Monodisciplinary [LBLL24]. month [MPB^{+}20]. months [PHC17]. monumental [Sal20]. MOOC [ELB17]. mood [HCD17, HY17, Lop14]. moods [LWM^{+}19]. moral [BM14d]. More-than-human [Gor20a]. Mores [SH19]. Morgan [Agu18, Cho19, Fou19, MS19]. mortgage [GZBA22]. Most [KB17, BWL15, BB15]. mothers [LS22, PH22, RBJ18a, RBJ18b]. motivates [NBH16]. motivating [EH15, St.17]. Motivation [AGBN16, CBH16, LCD^{+}14, ZFH^{+}19]. Motivational [Kan22]. Motivations [KBC17, LOD^{+}19, OS15, WCP^{+}19, CS16b, CKL16, MTK16, NE17]. mouth [HBG19]. Mouton [vR17a]. movement [Moo20]. movements [BCZ14, CL14, KTS^{+}16, WLTY19]. moves [LJ22]. movie [TL22, WMS^{+}15, WN15]. movies [EHdBvO17, MG21]. Moving [BOR^{+}23, Hil14b]. much [Bor16a, Cj15, Hil14a]. multi [AG15b, AHJY18, AKC^{+}18, DP23, Gia16, LZZ^{+}21, LB15, STS14, TLS^{+}23, YR23]. multi-angle [LZZ^{+}21]. multi-author [AHJY18, Gia16]. multi-context [DP23]. multi-document [AG15b]. multi-engine [AKC^{+}18]. multi-session [LB15, STS14]. multi-stakeholder [TLS^{+}23]. multi-study [YR23]. multichannel [LCWL18]. Multidimensional [CWL^{+}23, MH15, O’C16, SRM^{+}17, CCLL14, LZSW17, ZGPB17, MTY14]. Multidisciplinary [Car14b, CGGS17, CZW15, DYB14, LBLL24]. multidiscipline [JDHL17]. multilevel [KS16, KY17, LJLH17, MGKA16, MBD15]. Multilingual [KGP^{+}20, ECJF21, SL21b]. multimodal [GIEVVGS19, HCD17, SGS21, WSWH20]. multinational [WG19]. Multiobjective [GB19a]. multipanel [ZTA20]. multiplayer [SMW14]. Multiple [SMRR16, STF19, Che16, HRL21, LLHF15, MB23, MBD15, YSAM^{+}21, ZLCB15, ZCvdH15]. multiplexity [LBG^{+}21]. Multisensory [Rob15]. multisided [SLH22]. Multiverse [Clo18]. multiview [DGCC17]. multiviewpoint [ZGEJ17]. multiword [Vec14]. municipal [ZF19]. municipality [SBH21]. Murals [HNX18]. museum [MKK17]. Museums [Han21, Ack21, CC22, KLG14]. Music [TJ20, HCD17, HY17, HLB^{+}17, HK17, LP16, LCK16, LWA^{+}17, LRB19, LRB21]. musical [LKKP16].

Mutharaju [All19]. Mutual [LI14a]. my [Fox20, GS22, LHXS22, MBW^{+}14, MT20, WV15]. myth [Mad16].

Name [DMMF17, KMC20, Don16, FVGL14, FN17a, KD16, Kim19b, KK20, KKOS21, LDK^{+}14, LLHF15, SGLF17, ZLY23]. name-dropping [FN17a]. Named [PS20, DP23, MMFD15, PMGB^{+}21]. names [Lin14a]. naming [PSY^{+}21]. Nanna [Con20]. nanoscientist [CKL16]. nanotechnology [JGCR15]. narcissistic [GKM^{+}21]. Nardi [Hak14, Lan15]. Naresh [Fou19]. Narrative [KN23, BKM15]. narrative-based [BKM15]. narratives [Kha19]. Nathan [Far23]. National [TL18, TWB^{+}18, Tho22, WYZL21, AD15, CDL22, JPF16, KHZ^{+}22, LPU15, LWPG^{+}19, PS14b, RFF^{+}18, WW16, vLTW18, RMDF23]. nations [LBM19a, The20]. native [GBL^{+}18]. natural [Eng17a, Gre16, RJS19, RPK17]. naturalistic [LOM22, VSJR19]. Nature [Bat20, Kim15, lLTBH15, Bor15b, BH17a, MMALCOM15]. navigability [WK16]. navigation [ALMP16, NPG^{+}20, WKC14, ZZLS16]. Neal [Buc20]. Neal-Schuman [Buc20]. Need [Col20, Jul20, Bor14a, Cj15, Hjø17b, MP19]. needs [Col17b, DM14, HRB14, LCK16, MNNK15, RBJ18b]. negation [CTM16]. negative [Ko15, LSX18]. negotiate [SFL^{+}23]. neighbor [CML^{+}15]. Neither [ALAB19]. neonatal [AF17]. Netflix [Sea20]. Network [dSBG19, Ano17n, AMOM16, BLB16, BELR16, CSOD16, CCC14, CZ16, CdRM21, EAB17, EA18, GRSFV15d, HCC^{+}21, JLC16, JL23, Kan22, KD15, KBS18, LCM17, LBG^{+}21, LA14, LCSL15, LX17, LZA^{+}20, MSB21, MBML24, MK15, Min16, MAM19, QH20, RJGM16, RKL14, TK14, TK17a, YXAL23, ZY15, ZZLS16, ZZY23, ZQC^{+}23, ZTA20, ZZ21b, HLM16, LPW14, TK14]. Networked [Ten16, Far23, Fin17, MA18a, Lin14b]. Networking [Jor19, OGL19, DOM^{+}14, HCC15, JHJ15, JDHL17, KT14a, LOD^{+}19, LCL16, Shm18, Wu19, ZXPS18]. Networks [GB19a, Jor19, BH16, DT18, FW17, FN17b, GK14, HWZ22, HL20, IABL16, JGCR15, JSY^{+}16, JZC20, KD15, KD16, Kim19b, LKKS17, LdMAdN16, LDS^{+}19, MCDMM14, MP23, OD16, RVGPS19, SYS17, SLM15, SK14b, SLCS18, SEZG22, TG17, Tsi16, VS14, XOL^{+}18, YY16, ZZL^{+}14, ZGB16, ZB19]. Neural [SAA^{+}23, DGR19, ECJF21, HWZ22, JWJ^{+}19, QH20, SEZG22, ZTA20]. neural- [DGR19]. neural-based [ECJF21]. neuro [GMW19]. neuro-information [GMW19]. Neuroimaging [HBS15]. neurophysiologic [JBD^{+}19]. Neuropsychological [MP19]. newcomer [OB19]. newcomers [Lin15]. News [KT17b, SYS17, AvS15, AvS16, ACL17, Fro21, HS14, KF16, LCLBY17, LMM^{+}22, MMFD15, O’B17, RDPN21, SLI^{+}15, SGS21, SKKM22, TC17, WLC^{+}14, ALC^{+}14, FR23, Kos21]. newspaper [JKS^{+}16]. newspapers [KL22, OBD^{+}22]. newswork [SHKL23]. newsworthy [ALCD14, LMM^{+}22]. next [DZC^{+}14]. Ngarrindjeri [HDT23]. niches [SLS20]. Niels [Web20]. Nigel [Col17a]. NIL [GDO17]. Nivedita [Jan15]. NJ [Bru18, Rou17]. No [Ten16, Cro15a, SAB^{+}18]. Nobel [FW17, Fie16, LR14, RN16, YCL20]. Noble [FN17b]. noise [LCH^{+}22]. nomadism [JPS^{+}19]. non [AT14, Ma21, ZGEJ17]. non-experts [ZGEJ17]. non-individualistic [Ma21]. non-Western [AT14]. noncomputational [TWB^{+}18]. noninstitutional [LC21]. nonliteral [MGIGW16]. nonmovable [HNX18]. nonprofit [SFL^{+}23]. nonsensical [CL21]. nonserial [Chi19]. nonspeech [AG15a]. nonwork [CK22]. nor [ALAB19]. Norm [Hen19]. normal [GB17]. normalization [Col15, Hal17, ZL16]. normalized [BTMS16, Bor17, BH17b, Egg14b, HM18]. Normative [Baw14, Riv15]. norms [HD22, JS15, POK15, SSA22]. Norwegian [SBH21]. nostalgia [Sta22]. note [CRGG17, Hjø15b, OK16, SM18, ZXLC18]. note-taking [ZXLC18]. Notification [JVF19]. notion [Skö18]. Noujaim [Sea20]. novel [Col15, GS18, LCWC23, MRAF16, WMM^{+}23, XZW^{+}19]. novelty [AH17, LML23, LBC^{+}22, TL22]. novices [EHdBvO17, WV18]. nudge [WZFZ22]. nudges [GS22]. Nuechterlein [Wei14b]. Nugent [Sko23]. number [AKAS14, BM15b, Bor16a, KD15, SBIT14, WL14b]. numbers [BM14a, CKR14, Fie16]. nutrition [GN17]. NY [Afz22, Gor20a, Sto20].

O [Col17b]. Oaks [Hay18]. object [LJ22]. objectionable [LJT^{+}15]. objective [CK21a]. Objects [VP19, HAJ^{+}17, LGD20, Pil14, Skö18]. obliteration [McC15]. Observational [GGC^{+}19]. Observations [LMA^{+}22]. observe [DS15]. observer [Men21]. obstetric [AF17]. obtain [EH21]. occupational [SCIF20]. Occurrence [Hoo17, PHH^{+}20, ZL16]. Ochiai} [ZL16]. O’Connor [Kos20]. odds [BM15a, Ko15]. OECD [NILCP15]. Ofer [Jon17]. off [MDJ^{+}19, WXLBO15]. off-the-shelf [MDJ^{+}19]. offensive [MYK15]. office [WMS^{+}15]. Older [Cho20, PKA20, CPL22, QHH20, ZZY23]. Oliver [Cox15]. Olympic [Mir14]. omission [PRM22]. On-demand [CBJ19]. on-site [MKK17]. oncology [ACBC22, CKR14]. One [YHLG21, BdMA14, HBDL19, TTH16, TGV^{+}16, YRH16]. one-bit [TTH16]. ongoing [BH15]. Online [BDA^{+}18, Bow19b, CMB17, Jor19, LHM17, NLFM19, QHH20, RYEP18, dSBG19, ALC^{+}14, BCK17, BBM16, BPB21, CW18, CBH16, CLL15, CHJ20, CS16c, CS16b, Cho20, CYYG17, DMR^{+}17, Du14a, ELB17, EWX^{+}16, FMG16, Gor24, HCC^{+}21, HS17, HYD^{+}20, HBG19, IQDS16, IESG20, JDZ24, JW15, KZ23, KLC^{+}14, Kit19, KT16a, LCLBY17, LCLM18, LSC18, LDS^{+}19, LHXS22, LZG^{+}23, Lop14, LCWL18, MT20, NV17, O’B17, ØC19, PPLJ20, PN18, PEG^{+}19, PH22, QLM^{+}22, QZWC21, RC16, RVdAGP21, Roh21, RJS17a, Rub15, RBJ18a, RBJ18b, Sav15a, SZ20, SYS17, SMW14, SZZ16, SLS^{+}19, SK14b, Tsi16, WOJBG17, WHWZ21, Wes15, WSvPK16, WCO21, Wu19, Wu23, Xia14, XA14, YR16, YZYR21, Zem17, ZY15, Zha16, ZLF20, ZW23, Bow19a, Zim17]. Ontological [MR14, AFB20]. Ontologies [AF17, Fri17, BOHK19, CLA^{+}16, SS14a]. Ontology [TKB^{+}17, ADD^{+}18, AF17, BAK^{+}17, Göd16, KMKL15, ZGPB17, ZGEJ17, MR14, Fri17]. ontology-based [Göd16]. Onward [Ano16m]. Open [Bjö14, HD22, KBH^{+}22, LMJ21a, PSB^{+}14, QH18, AdUCDPH19, Aji17, BLWP14, Bor16c, BJSM18, DB17a, ELB17, FSPC16, GVF^{+}24, JML22, Jam20, JHJ15, JPF16, KN23, LMJ21b, LYC^{+}22, MZ22, Moo20, MBZ^{+}22, NE17, NB15, PSB16, PSB17, RJS^{+}17b, She21, TD15, TLD17, VLBG^{+}16, WCP^{+}19, YL18, YLHW21, YC20, ZF19, vLTW18, GVF^{+}24, LMJ21b, She21]. Open-access [PSB^{+}14, PSB16, WCP^{+}19, YL18]. OpenAIRE [AGGTGL18]. openness [Mah22, NT17b, QLM^{+}22, TL22]. operate [MMFA20]. Operating [Lin14b]. operationalization [CS15a, HHK^{+}17, LB16]. operationalize [Meh23]. Operationalizing [CdRM21]. operator [MTY14]. Opinion [HBG19, JA16, Lue15, Rob15, SRPS16, Vas15, CLSS14, Hen14, HWC17, LTD^{+}15, VGC17, Yu14]. opinions [LCHL14]. Opportunities [GVB17, GISW15, YC20]. opportunity [AWC18, JWWV18]. Optimization [ZZSX16, GMRS17, PDS22]. option [Hal14]. Order [Hen19, BD16, KD15, QZWC21, Sha21b]. Organization [Cox14, RAW22, GTZŽ14, HWW^{+}17, Hjø15c, MdSLdO^{+}23, Gui17]. Organizational [MMB17, ADM^{+}20, JS15, LCB^{+}18, SV17, WKC14]. Organizations [ZZ20, Bur14, CM23, HB14, LHR16, YL17a, YC20]. Organizations-Related [ZZ20]. Organizing [Lan15, MSM^{+}21, Bat15, Sve15]. Orientation [XBL^{+}21, WH15]. orientations [MSK^{+}14]. oriented [AWC18, BKM15, CML^{+}15, KS14b, MVA22, SS17, SK23, WV18]. orienting [LHS24]. origins [iCV18]. Orthogonality [LRB21]. other [FW19, GG16, Ior14, MSPP21, VGN20]. Our [Cus15, BR20, Coc22, HS14, Jon17]. outcome [Tsi16, ZZE17]. Outcomes [Cho19, CO16, GEBP22, LHM17, MHMW17, O’B17, PEG^{+}19, SDUDL24, Sin16]. outpace [AD23]. outpatient [VP19]. output [GD18, LQ14, NILCP15]. outputs [AKC^{+}18, GRS16, WYZL21]. outrage [FV21]. Outsiders [CK21b]. Outsiders-within-Library [CK21b]. outstanding [Lia17]. Overcoming [VZRP15, WKJW16]. overextension [TL18]. Overflow [GSG22, YHLG21]. Overlay [BH16, KNY^{+}14, RRHL17]. overlays [CL14, LdMAGB15, LdMAdN16]. Overview [JdR17, Bjö14, TCF^{+}24]. overwhelmed [RBJ18a]. Owen [Kos20]. own [LB22, MBW^{+}14, SLeH21, TM15]. ownership [Ric18]. Oxford [Hai16]. Oxymoron [War15].

P [Bor14a, SK14a, KF16]. P-Stemmer [KF16]. P100’ [BM14c, BM14c]. PA [O’H18]. PAC [Rou16]. page [GBL^{+}18, WH15]. PageRank [CCLL14]. pages [MA18b]. Pairs [CBvEW18, CDGN15, WHYJ17]. pandemic [BK23, RAMP22, SBS23, Sha21a]. Pandemics [LBC^{+}22, ZJHT23]. Pao [Ma22a]. paper [Fan18, LJ22, MWD17, Buc15]. Paperback [Agu18, Dut20, Fri17, Soe18, Uns18, Zim17, Buc20, Cho19, Clo18, Col17a, Dor19, Erd18, FJ19, Fra21, Fro21, Gor20a, Hai16, How17, Huv20, Huv22a, Meh22, Pan21, Rey21, Rhe17, San22, Saw19, Sha21a, Sto20]. PaperPoles [HPLC19]. papers [Ano14, BLB16, BRS16, BdMA14, BWL15, Bor17, CL21, CCM23, CK21a, JQMY24, LB19a, LGD20, LCRGTS14, LWM^{+}23, MP23, PDS22, PS14b, RN16, SMAB09, WSZC22, ZWW18]. Paradigm [Hal20, TMDZ21b, PG21, TMDZ21a, TNL21]. paradox [Fox20, Wu19]. Parallel [Hen14, AKC^{+}18]. Parameterized [PS20]. parameters [Los16]. parametric [Hal17, HM18]. Paranormal [The17]. Paranormal-Romance [The17]. paratext [Ano14, SMAB09]. parent [HM18]. Parental [EWX^{+}16]. parents [GM19]. Park [TWB^{+}18, Tho22]. Parkinson [LKKS17]. parliamentarian [AST^{+}22]. parse [KLH17]. parse-tree [KLH17]. parsed [AST^{+}22]. partial [SY20, SY22]. participate [HSU16]. Participation [dSBG19, JHJ15, KZ23, KPC^{+}15, YLHW21]. participatory [BKM15, PKK^{+}24, YLL20]. partitioning [KKOS21]. partner [Wes15]. partnerships [YLL20]. Passive [TGC^{+}21]. past [LYC^{+}22, Sal20, PSB^{+}14]. Patent [JdR17, KNY^{+}14, KT17c, LCD^{+}14, CZZY15, GS18, HHC^{+}14, LSL23, PM18, YL17a, YL17b]. patenting [VCG17]. patento [dSMdAGQ15]. patento-scientometric [dSMdAGQ15]. patents [OMTK17, XZW^{+}19]. path [JZC20, JL23, LK16, YY16]. path-based [YY16]. paths [GS18, LCHL14, QZWC21, YR16, Yan14, YL17a]. pathways [ZZZL17]. Patient [HCDÅ15, LSC18, MAS^{+}15, KLC^{+}14, LHXS22, LXW18, WCTS15, WLW^{+}17, ZYZ^{+}16]. Patient-centered [LSC18]. patient-level [WLW^{+}17]. patients [BK23, EAB17]. patrons [Zem17]. pattern [NSK17, PDS22]. Patterns [CL14, KSB20, Chi19, CMS23, DPPV14, DFB14, GB19b, HDZ^{+}15, LOM22, LCLM18, NH15, PS14b, RCM^{+}17b, RPK17, SY20, Söd23, TL18, YXJ18, Zha14c, ZY15]. Paul [Kea22, Lan15]. Pauline [Sav19]. pay [ZPL^{+}20]. payment [ZPL^{+}20]. payment-based [ZPL^{+}20]. Payne [Rog15]. paywalls [Jam20]. PDF [Sha21b, Hod21]. peak [MF16, Roh21]. Pearce [Hak14]. Peer [Tei16, Bor14a, Bor15a, GRSFV15b, GRSFV15c, H17, KH17, LCNP17, MBD15, NB15, NB17, RH21, SDT22, ZCvdH15, RR21]. peers [BM15a, JDZ24]. Penguin [Afz22]. Pennington [Pan21]. Pennsylvania [O’H18]. People [Cox15, OR21, AGBK19, AMOM16, ALMP16, CCR24, HSU16, HK21, Ior14, KT14a, LS18, Li21b, MK15, NBH16, St.17, VXN18, WH15]. Perceived [MO20, JSW14, LCB^{+}18, SSdMCM20, YCL20]. perceiving [SZWS22]. Percentiles [BM14a]. perception [HSL24, RR21]. Perceptions [BYSB16, JW15, SYF17, AGBN16, HCDÅ15, LWA^{+}17, LBG^{+}21, SHW^{+}15, WWL16, Kon16]. percolating [GK14]. perform [BM15a]. Performance [MC18, CBM17, KTS^{+}15, LZ22, LJLH17, PRRC18, Pra14a, QSH14, RPK16, RPK17, SZZ14, Sch15, SHGI16, SJL19, TL18, YZQT20, YL14]. performances [FS18]. periodical [Chi19]. perish [Hal14]. persistence [Zub18]. persistent [BDLM18]. Person [DMMF17, GDO17, LMMO15]. person-entity [LMMO15]. Personal [CK22, DO14, GMRS17, Jor19, Min16, Tsi16, CBJ19, Esk20, FA19, JPS^{+}19, Kel21, KMD16, LS14, LWLH18, MK15, Oh17, SB18, Sta22, TK17d, WXLBO15, Cus15]. Personal-discount [GMRS17]. personalities [BYSB16]. Personalization [JPS^{+}19, LLB20, CWdR17, KCK16]. Personalized [BDA^{+}18, LZD16, SZJ^{+}21]. Personalizing [LB15, SLI^{+}15]. personally [Gor19]. Persons [PKA20]. perspective [BWF14, CLL15, CZW15, Fan18, FA17, FvdWR^{+}22, HBDL19, KTZ20, Kim19b, LOM22, LML23, LXW18, LBW^{+}19b, MVA22, SLS^{+}19, SHKL23, SA15, SBZ^{+}14, TL22, WV15, WvdHS16, Wu19, YZYR21, Zha16, ZT22, ZPL^{+}20, ZZ21b]. Perspectives [OBD^{+}22, CdRM21, JAG^{+}20, LA14, PRM22, RAMP22, SLLW20, XBL^{+}22, MAS^{+}15, Gor20a, Cho19]. persuasion [FN17a]. pervasive [Rob15]. Pew [CPL22]. Pham [San22]. phenomena [XOL^{+}18]. phenomenological [SH14]. phenomenon [HHC^{+}14, Men21]. Philadelphia [O’H18]. Philip [Wei14b]. phones [BNG14]. photo [LYCX14]. photos [KBC17]. phrase [Hjø15b, KBD^{+}14]. phrase-level [KBD^{+}14]. Phuoc [All19]. physical [Ior14, Men21, SC21]. Physician [IB16, LXW18, ZLF20]. Physics [PDS22, SHW^{+}15]. physiological [WLTY19]. picture [TR17, WMX^{+}22]. pictures [TGV^{+}16]. piece [Lue15]. piling [KR15]. PIM [DO14]. Pioneer [Hai16]. piracy [CJ22]. Pirate [Jam20]. Pirates [Ben18]. Pitfalls [HCKH15]. Pits [Tho22]. place [WV15]. Places [PWK19, ASTU16, OR21]. plagiarism [Ano17m, dLPMR16, VG18, GMB14]. plain [LZ21]. planning [MVA22, Sav15a]. platform [Cab16, ELB17, GAB^{+}18, LHXS22, MPB^{+}20]. platformized [Ma23a]. platforms [DB17a, HH14, KA22, SCIF20, SK23, ZPL^{+}20]. play [CCLL14, PTGL17]. players [PTGL17]. Playground [Dow15]. Playing [FJ19]. plea [Bor16b]. PLOS [YRH16]. plot [MG21]. plots [BM14a, Epi14]. plural [MB23]. plus [ZYZ^{+}16, PDS22]. poetry [FMG16, PdlRR^{+}22]. polarity [HWC17, MRMCMVUL14, VAGR15]. polarization [HBG19]. polarized [TF21]. Policies [Bow19b, DAT^{+}14, JS15, NB17, SBA^{+}15]. Policy [Wei14b, Ano17n, AD23, CZ16, HCS21, Olt15, PSB17]. politeness [NB15]. Political [Hen19, Mar15, NT17a, MSB21, WZFZ22]. politicians [RAW22]. Politics [Con20, FM14, Med15, Joh17, Lev20, Olt15]. Polity [Gor20a, Rey21, Saw19]. Polity/Digital [Saw19]. Pomerantz [How17]. pop [HY17, SJK15]. popular [ACL17, PHZF16]. popularity [Du14a, MSPP21]. populations [MMHA23, VGN20]. portal [MAS^{+}15, ZZSX16, ZZLS16]. portals [ZF19]. portfolio [CL14, LHR16]. posed [Aba24]. position [HB21, LB22]. position-based [HB21]. positions [HCDÅ15, LA14, VCG14]. positive [Ko15, Tsi16]. Possessed [Dut20]. possessions [BR20]. possibilistic [CSOD16]. possibilities [BRS15, PSB^{+}14, vLTW18]. possible [AP20, Hjø15a]. post [HWW15, MYK15, MBD15]. post-web [HWW15]. posters [HLC15]. Postpublication [Tei16]. posts [LJ17, ZCZ^{+}20]. Potential [WC14, KK19b, LS18, LCLBY17, RD16, ST22, SC17, SBIT14]. Potentialities [BRS15]. £49.95 [Col17a]. poverty [GM19, HRB14]. Power [MSB21, LXW18, RPK16, Smo17, XBL^{+}22]. power-law [RPK16]. Powering [Wei14a]. pp [Abb17, Adl23, Afz22, Agu18, Aji17, All19, Ama19, BI17, Bat21, Bea17, Ben18, Bow19a, Bow19b, Bru18, Buc20, Bud17, Cho19, Clo18, Col17a, Col17b, Con20, Dor19, Dut20, Erd18, Far23, FJ19, Fou19, Fra21, Fre18, Fri17, Fro21, Gor20a, Gui17, Hai16, Hal19, Han21, Haw18, Hay18, He18, Hjø17a, How17, Huv20, Huv22a, Jon17, Jul20, Kea22, Kno24, Kos20, Kos21, Lev20, Ma22b, Ma22a, Mat20, Meh22, MS19, O’H18, Ols16, Owe21, Pan21, QH18, Rey21, Rhe17, Rou17, Sal20, San22, Sav18b, Saw19, See17, Sha21a, Sha21b, Sko23, Soe18, Sto20, Swa22, Uns18, Web20, Whi17, Zha19, Zim17, dP20, vR17a]. Practicability [Ano17m, GMB14]. Practical [Agu18, PSY^{+}21]. Practice [Cox14, Rhe17, Bur14, CC17, FC22, Gre18, HCS21, LI22, Mat15, New23, RCM^{+}17a, STD15, Sil18, Smi17a, Sny14, YR23, ZSC18, Soe18, Abb17, Haw18]. Practice-Based [Cox14]. Practices [CCP15, FMG16, GGC^{+}19, HD22, KC20, KP18, Ack21, ACS18, AGBK19, CK17, CKL16, Coc22, CHLRC14, DO14, DM14, DAT^{+}14, FA19, GG16, GW18, HDT23, HHL21, HEEB^{+}22, JdR17, JS15, JT17, KWLV22, Li21b, Lin15, LO19, Mat22, MR19, PH22, RKHMRH21, RGJCTS16, ST22, SDUDL24, SE16, Tab14, TGC^{+}21, VO23, WJB22, ZPPW24]. practitioners [JJ14, VJH15]. pragmatic [KMKL15]. pre [HWW15]. pre-web [HWW15]. precarity [VO23]. Predatory [Mar23, Sil20, XHC^{+}15]. predict [CJ16b, QSH14, WZL21]. predictable [GZBA22]. Predicting [DLD^{+}21, GBZ^{+}23, LLB16, MDC^{+}16, SSC17, VVP^{+}21, WHWZ21, ZLCB15, ZQC^{+}23, ACL17, LB19a, SAB^{+}18, ZZ21b]. prediction [DPN16, FdGPC18, GMRS17, KYFZ15, LYCX14, QH20, SZZ14, WLTY19]. predictions [Mar23]. Predictive [WZK^{+}18, GZBA22, IABL16, Mar23, MBD15]. predictor [VRF14]. prediscussion [ZXLC18]. predominantly [LMA^{+}22]. Preference [TL22, OHP16, TJ20, WN15]. preferences [DLD^{+}21, Hsi17, KIMM17, LMP21, MAS^{+}15, PGK24, ZK14]. preferential [SLM15, ZBDX18]. prehistory [Moo20]. prehypertension [JWJ^{+}19]. premodern [KBH^{+}22]. preparation [WDTC20]. preprint [KPWS20]. preprints [HDZ^{+}15]. prepublication [LML23]. presence [BPB21, LHXS22, ZAE^{+}19, DFB14]. Presentation [OGL19, GBL^{+}18, KTS^{+}16, Min16, QLM^{+}22]. presentations [TK17c]. Presenting [MŽA17]. Preservation [PhD20, Ack21, KMD16, KA22, MMB17, Min16, Sal20, SBH21, Fra21]. Preserving [CA20, Cus15, FHO23, VS14]. presidential [KA22, BYSB16]. Press [Adl23, BI17, Bat21, Ben18, Bow19a, Bow19b, Cle18, Con20, Dut20, Far23, FJ19, Fri17, Gor20a, Hai16, Hal19, Haw18, Hjø17a, How17, Huv22a, Jon17, Kea22, Kos20, Ma22b, Ma22a, Meh22, O’H18, Ols16, Rhe17, San22, Saw19, See17, Sha21a, Sha21b, Sko23, Soe18, Web20, Whi17, Zha19, Zim17, Kos21, Sal20]. Prevalence [CL21]. preventing [St.17, WWB21]. Prevention [KP18]. price [Lin14a]. primary [CRGG17, Hof17]. primer [Aji17, AD23]. priming [Smi17b]. Princeton [Kea22]. principal [GRSFV15c]. principal-agent [GRSFV15c]. principle [DLA22]. principles [Xia23]. Print [KIMM17, Fro21, ZK14]. prints [LSM^{+}14]. priorities [LSHH23, Skö18]. prioritization [dSMdAGQ15, STF19]. priority [RVdAGP21]. Privacy [HH14, LK20, VJH15, BR20, CS15b, Esk20, Fox20, FvdWR^{+}22, GS22, JAG^{+}20, KCK16, LS14, Ma23b, MS16, MGK20, MK15, PRM22, QHH20, RDS20, RB14, SB16, SSR^{+}20, SK14b, TMBA20, VGN20, VOBB15, VS14, VXN18, WXLBO15, Wu19, WVZ20, ZVW20]. privacy-preserving [VS14]. privacy/privacy [SSR^{+}20]. private [KA22, LWB16]. prize [FN17b, LR14, FW17]. Prizes [RN16]. Probabilistic [TZK20, LWL^{+}18]. probabilities [Ko15]. probability [DLA22, Sav16a]. probably [Rou16]. problem [ALMP16, GRSFV15c, Sin16, Cox15]. problem-solving [Sin16]. problematic [Sin16]. Problematizing [H17]. Problems [Cox14, HB16, KYFZ15, MB14, She16, SLeH21]. procedural [Smi15]. procedures [PDS22]. process [AdUCDPH19, CHJ20, JPF16, Kim15, Lop14, MB20, PMSA14, SZ20, SHGI16, TWB^{+}18, WDL21, WN15, ZS14a, ZXPS18]. processed [HBS15]. Processes [Jor19, Sav18a, CCP15, KSL19, LGD20, PB15, TK17d]. processing [JWJ^{+}19, LI14a, PSB16, VAGR15]. Produced [Sea20]. product [LZ22, LXLL14]. production [RH21, Yan16]. Productivity [SKZ^{+}14, BS17, WYZL21]. products [GN17, LLB14]. Professional [FHW17, Jor19, ZZ20, ZGB16, Bea16, CJ22, HLC15, TBP^{+}15, TK17c, Vin17]. professionals [Du14b, Fro21, OM15]. professions [Cox23, PSY^{+}21, Soe18]. Professor [Hjø17b]. proficiency [Kim15]. profile [LSHH23, SMYL15, ZXLC18]. profile-boosted [SMYL15]. Profiles [OGL19, OA14]. Profiling [VS14]. program [Mar16, ZY15]. programmatic [HL16]. programmers [SCIF20]. Programming [Cle18]. programs [SP17, WWB21]. project [CBAL15, DLD^{+}21, SVC22, SHW^{+}15]. projects [EH21, SSC17, WYZL21, YLHW21]. proliferation [WHZ23]. prominent [CY22]. promises [HCKH15]. promote [HDT23]. promoted [MPB^{+}20]. promotion [BDH17]. Proof [SR14]. Propagation [PS20, CML^{+}15, JGY^{+}19]. propensity [MWD17]. Proper [Zit15, Rou15]. properties [Bar14, HPL^{+}22, PS14a]. Property [Med15]. proportion [BdMA14]. proportions [CKR14]. Proposal [Fia14]. proposals [SVC22, SR16]. Proposed [MH18, WFRN14]. protect [Fox20, VXN18]. protected [BOR^{+}23]. Protecting [Bow19b]. protection [CBH16, QHH20, WLZ^{+}20]. protocol [DMR^{+}17]. protocol-driven [DMR^{+}17]. prototype [MŽA17, WB15]. provenance [MWMB18, TWB^{+}18]. proverb [ZGPB17]. provide [KISS15, VS14, ZMJ14]. provided [JML22]. provider [KCK16, TBP^{+}15]. Providing [Sav15a]. provincial [LPU15]. provision [Min16]. Proximity [HWZ22, BCZ14, HWZ20, LSL23, SC21]. Proximity-aware [HWZ22]. proxy [OMTSLC15, Rou16, XBL^{+}22]. Pruning [KMC20]. Pruning-Based [KMC20]. pseudo [PHH^{+}20, YH16]. pseudo-relevance [PHH^{+}20, YH16]. pseudocollaboration [GISW15]. pseudonyms [Ior14]. pseudorelevance [THC15]. PSI [LWL^{+}18]. psychological [EAB17, EA18, MYK15]. psychology [McC15]. Public [ZCZ^{+}20, Ano17n, CZ16, CS15b, DTDT21, DAT^{+}14, Eck18, GJ16, JGCR15, KISS15, KA22, KCZ^{+}14, LWB16, LML16, MM20, MJ23, PG20, Sto20, VFZ^{+}22, ZZLS16, ZPPW24]. public-funded [PG20]. public/private [LWB16]. Publication [Ano14, CBvEW18, GRH^{+}17, MBBL14, SRM^{+}17, ŠF17, AdUCDPH19, Bor16a, CL14, GB19b, Hsi17, HQH^{+}22, JLC16, Kim19a, Man22, PSB16, PSB17, Pra15b, SRYL17, SHMC17, YHL^{+}21, SMAB09]. publications [Bor14d, BM15b, BTMS16, CC16, Chi19, Cri15, DQAMM^{+}24, DFB14, DCLH21, FDC20, FCT^{+}21, Fas18, HPLC19, HDZ^{+}15, ML22, MSdGS19, MT18a, Mil14, RHG23, SC17, SAW21, XZW^{+}19, ZCW17, ZYL18, ZCYN23, ZGE22]. publish [Hal14, Jam17, NE17]. Publisher [Sav18b, JML22, MPB^{+}20]. publisher-provided [JML22]. Publishers [Agu18, Aji17, ZGCB15]. publishes [XHC^{+}15]. Publishing [All19, Ama19, Clo18, Col17a, Col17b, Dor19, Erd18, Fre18, Hay18, Huv20, Jul20, Mat20, Pan21, BGS22, FMG16, Jam17, Jam20, KBH^{+}22, KIH16, KGP^{+}20, LWM^{+}23, MO20, Moo20, MBZ^{+}22, RA14, RD16, RJS^{+}17b, Sil20, SL24, vLTW18]. PubMed [LDK^{+}14, RL15a, Sch18, SKZ^{+}14]. PubPeer [Ort22]. Puerto [VO23]. purchase [CJ16b, JL18, KTZ20, Shm18]. pursued [WZL21].

QA [Kan22, CS16b, JDHL17, Sav15a, SWSZ19, WLG23, ZWTX23, ZPL^{+}20]. QA [LCSL15]. Qatar [Ama19]. QS [DBBJ16]. Qualified [Hal19, Mir14]. qualitative [BWGD21, Fas18, RLAB^{+}24, ZS20]. qualities [BL14]. Quality [HH19, ZSX15, Bor17, BH15, Cho15, DGCC17, JW17, LL17, MO20, PMSA14, PS14a, Pet19, Pra14d, Rou16, SSdMCM20, SHW^{+}15, TKA^{+}23, TKS^{+}23, Tsi16, VVP^{+}21, WL20, WHWZ21, WvdHS16, YCL20, YH16, Zha14b, ZT22, ZvSvB14]. quality-aware [YH16]. Quan [Hay18]. Quan-Haase [Hay18]. quantification [HH19]. quantify [WL14b]. Quantifying [NLFM19, WMM^{+}23]. Quantitative [MH18, SRG^{+}16]. Quantity [Pra14d, CMS23, WHWZ21]. quantum [BRS15]. queer [KWLV22]. querier [KZB17]. Queries [TCP15, THL17, BSZ^{+}14, HT17, RCM^{+}17b, SAB^{+}18, SYG17, VC15]. Query [BST15, SZZ14, THWS14, AS14, ATKD^{+}17, CDGN15, JKS^{+}16, KBD^{+}14, KWR20, KZB17, Lew15, LZSW17, LJLH17, PMD14, RJK20, SY20, SY22, Smi17a, Smi17b, SL21b, SBZ^{+}14, TPR17, XLL16]. Query-biased [BST15]. Query-performance [SZZ14]. Querying [All19]. Question [BR16, LS18, BK19, Coc22, Fon21, LCSL15, LJ17, Lue20, SMRR16, Wes15, ZLZG14, ZWTX23, dRR15]. question-answering [LCSL15]. questionnaire [TLA^{+}16]. questions [BVd16, CS16b, DMR^{+}17, WLG23]. quickly [MT18a].

R [Adl23, Aji17, Far23, O’C16, Rou17, vR17a, RL15a]. R01 [WYZL21]. Raan [Hjø17b]. rabbit [MB20]. Race [Erd18, Adl23, Man22]. Racial [Meh22]. radiology [wYKY17]. Raf [Dor19]. Rafael [Agu18, Fou19, MS19, Uns18]. Raghava [All19]. Ragnar [Han21]. Rain [SHGI16]. Rainie [Lin14b]. Rains [Zha19]. Random [Afz22, SLM15]. randomized [MPB^{+}20]. rank [AWC18, BM14c, GNVQdMA16, OA15, PSS22, SGV14, TW16a, XLL16, YH16, ZMJ14, ZGCB15, ATKD^{+}18]. rank-oriented [AWC18]. Ranking [ATKD^{+}18, Hal14, KS14b, AD15, CML^{+}15, DLA22, DD16, DBBJ16, Hal17, HM18, JSY^{+}16, JJC19, LWL^{+}18, SKL17, BdMA14, Egg14a, MD15, PRRC18]. ranking-oriented [CML^{+}15]. rankings [AGMR18, BELR16, Bor14c, DB17a, KHZ^{+}22, LBM19a, SB18, ZS16, BM15a]. Raphael [Cho19]. Rasmussen [Pan21]. rates [BM15b, OY15, TM20]. rating [CB15, Fas18, Fia14, MKK17, TZK20]. ratings [TZK20]. Ratio [BH17b]. rationale [Xia14, XC17]. rationale-containing [XC17]. rationality [McC15]. ratios [BM15a]. Raw [War15]. Rawls [Hof17]. RCR [BH17b]. Re [ATKD^{+}18, DO14, GM19, LHS24, SHKL23, MMALCOM15, SVK17]. Re-Finding [DO14, SVK17]. Re-orienting [LHS24]. Re-Ranking [ATKD^{+}18]. re-revisited [MMALCOM15]. Re-situating [GM19]. reach [Söd23]. react [HH18]. reaction [LSC18]. read [Jam17, MT18a, TM15]. Readability [JVF19, AP20, DPN16, JGY^{+}19]. Reader [AP17, DS21, MT18a, Ma22a]. readers [OMM^{+}14, TM15, TK17a, ZK14]. readership [BH16, Du14a, MT16, MT14, MTK16, OMM^{+}14, ST16b, TW16b, TS16, ZCW17]. reading [BF18, GHCW17, LCLBY17, ZZ21a, vdSvdB23, KIMM17]. readjustment [HCC^{+}21]. reads [MTHL15]. ready [MM20]. Real [ZSMF15, BCZ14, DC22, FJF^{+}21, KK19b]. Real-time [ZSMF15, DC22, KK19b]. realism [MR14]. realities [Oli21]. reality [LWM^{+}23]. realization [MP19]. really [AD23, ZW23, vdVZK21]. reasoning [HT17, KKZS20, All19]. Reasons [LLH19, DQAMM^{+}24]. rebuttal [Col20]. received [WHWZ21]. reception [Bor14d]. Recherche [Zit15]. Recognition [PS20, SSA20, HY17, HR19, Mah17, PMGB^{+}21, SLLL17, Vin17, YB18]. Recognizing [LBM^{+}19b]. Recolecta [AGGTGL18]. recommend [BRS16, SMYL15]. recommendation [AWC18, BAG17, DYB14, HWZ22, Lai15, LF21, LLD^{+}17, MBAG16, MKK17, SVK17, SZJ^{+}21, TL22, THWS14, XDDG18]. Recommendations [MGK20, WC14, BSVZ17, DTW16, WN15]. recommender [SVB22, ZHGY18]. Recommending [PN18, TG17]. reconstructing [Hod21]. record [LSM^{+}14, SBH21]. record-keeping [SBH21]. Records [DL22, Ric18, AF17, AKC^{+}18, DTDT21, HCDÅ15, KBC17, KA22, LS14, Cox15]. Recovering [JLC16]. recycling [GRBCS15]. redaction [SB16]. redistribution [Hil14b]. reduce [BS16, MM20, SDT22]. Reducing [RC16, ZCvdH15, CW18]. reduction [SLLW20]. redundancies [LI14a]. redundancy [LJI18]. referees [CGGS17, ZCvdH15]. reference [Chi19, Col15, RCM^{+}17a, RCM22, YHL^{+}21, Zem17, MBBL14]. references [BAGL16, BM15b, CDL22, Mil14, TM16, ZRG16]. referencing [ACS18]. referential [LQ14, dC16]. referral [LMA^{+}22]. refined [SC17]. Refining [CCLL14]. Reflect [Hal20, MTK16, TKA^{+}23]. reflected [AS24]. Reflections [BWGD21]. reflexivity [ZFH^{+}19]. reformulate [SYG17]. reformulation [AS14]. reformulations [LJLH17]. refugee [OR21]. refugees [LI22]. regional [LPU15, LWPG^{+}19, Söd23]. register [EBD15, TDOF^{+}16, Bow19a]. registries [GTZŽ14]. regression [Bor16a, ZLCB15]. Rejoinder [FN17b, Hjø17b]. relate [HPL^{+}22]. Related [ZZ20, Bor17, CCT14, Esk20, KvOS^{+}17, KK19b, LMA^{+}22, RJS19, SC17, Sta18, Zha14c, ZCZ^{+}20, ZS19]. Relatedness [Ma23b, LKY^{+}20, MGCTW15]. Relating [Gor22]. relation [ACS18, DM14, FJF^{+}21, JAG^{+}20]. relational [HLM16, Riv15]. Relations [MT18b, BK19, CK21a, Hjø15a, KK16, LPW14, LdMAdN16, LdN17, SKA18, Sto20, XC17, LA14]. relationship [BCZ14, CZZY15, FSPC16, HQH^{+}22, LRB19, Lue15, MMHA23, ØJW^{+}21, RPK16, RPK17]. Relationships [RBJ18b, Eng17b, Epi14, FV21, JWJ^{+}19, JSW14, LSM^{+}14, LRB21, RYEP18, ZZZL17, Mil14]. Relative [LBW19a, Cri15, BH17b]. Relevance [WHX^{+}14, CDGN15, DLA16, GIEVVGS19, GHCW17, JBD^{+}19, JJC19, LZSW17, LPB19, MCY^{+}21, MTY14, PHH^{+}20, PSS22, RDPF16, Rut21, SLH22, VR17b, Wat14, WVvLZ23, YH16, ZGBIL18, dFB22a]. relevancy [AH17, LCHL14]. Relevant [KK19a]. reliability [Bor15a, Wat14]. remote [NB23, RAMP22]. repertoires [CLSS14]. Replicability [LWB16, VR17b]. replies [ZLF20]. report [Wu23]. reports [AS24, QSH14, wYKY17]. repositories [AGGTGL18, Bjö14, HD22, KN23, PMSA14, PSB^{+}14, SZZ14]. repository [FSPC16]. represent [SJL19, TD15]. Representation [HY17, AF17, DGGdST19, Li14b, LQ14, MdSLdO^{+}23, Par19, Sav16b, Sny14, WSZC22, Rhe17]. representations [ABLS17, Che15, KDS^{+}17, PS14a]. representative [Lew15]. Representing [BAK^{+}17, FW19, Vec14]. Reprioritizing [CK21b]. republication [Pet19]. reputation [AMDFHVC24, CB15, PGK24, RVdAGP21]. Request [LK20]. Requests [Zem17, HT17]. requirements [Eng17a]. reranking [FT17]. Research [Agu18, BI17, Bud17, CA20, CPL22, GD18, HCKH15, Hay18, HWZ20, HCS21, May16, RN16, She16, She17, SV20, SBA^{+}15, SHW^{+}15, TKA^{+}23, TWB^{+}18, WC14, Yan15, Zit15, AD15, AGBK19, Ano14, BC16, BWGD21, Bor15c, BTMS16, BM23, BVd16, CCM23, CH16, CK21a, CKL16, Col17b, CJ16b, CKLP17, DPHP17, DJ20, DM14, DTW16, DCLH21, FDC22, FMP16, FSPC16, GRBCS15, GW18, GRS16, GK24, GTAS22, GVB17, HD22, HWZ22, HGCKDR16, HWW15, HEEB^{+}22, JdR17, Kel21, Kim15, Kno24, KT14b, KT17b, KKZS20, KL22, LSHH23, LHS24, LBM19a, Lia17, LYWF14, Lue15, LWM^{+}23, ML21, Mak20, MBBL14, Mat15, Mat20, MHMW17, MOM^{+}15, MH15, MTHL15, Mor24, MBD15, NGTKR15, NILCP15, OBD^{+}22, OK16, OMM^{+}14, OY15, PRRC18, PG21, PG20, PHJ23, RMDF23, RKHMRH21, RHG23, SMAB09]. research [SVC22, SB18, SBIT15, SMYL15, SS14b, SR16, SWL18, SLeH21, ST16b, TSY15, TW16a, TM16, TS21, TKS^{+}23, TH16, VCJ22, VOBB15, VCG17, Wan18, WSZC22, WG19, WHvdZvdB17, WTvR16, WK19, WVZ20, Wu23, Xia23, YRWH18, ZYZ^{+}16, ZZY23, ZS14c, ZSC18, Zim17, ZZE17, ZZ21b, dRR15, MT14, SBIT14, Clo18]. Researcher [SWL18, LSHH23, LMP21, Pil14, PBG^{+}17, XLKC22]. researchers [BM14a, BB15, DC22, ELB17, Fia14, HHL21, KISS15, Dor19]. ResearchGate [HLM16, JDHL17, LOD^{+}19, OGL19, TK15, TK17b]. researching [CAS21, CBAL15]. resistance [Eck18, MJ23]. Resonance [Rut21]. resource [ALYS16, LMA^{+}22, RC16, Ric18, SDUDL24, YCK16]. resources [BH15, CH16, CCJ16, DMR^{+}17, LMA^{+}22, PHZF16, TPZ15].

Response [Bur20, LMJ21b, ADM^{+}20]. responses [KK19b, PKB^{+}16, ZRdR21]. Responsibilities [FHW17]. responsible [SAHZ21]. restorers [LO19]. restricted [Olt15]. result [CK21a, OA15, SL21b]. Results [TLA^{+}16, WCP^{+}19, BdMA14, FT17, KTS^{+}16, LZ19, PMSA14, PGK24, RLAB^{+}24, SAB^{+}18, UH17, YSAM^{+}21, YZ16, ZSX15, ZGBIL17]. resumption [WDTC20]. retail [RTS^{+}17]. Rethinking [Chr17, MZ22, SW18]. Retraction [SRM^{+}17, LC21, Sch18]. retrenchment [TL18]. Retrieval [Agu18, Fou19, GGC^{+}19, MS19, AH17, Ano18z, ADL15, ATKD^{+}17, ATKD^{+}18, BK19, Bea16, BWF14, BBM16, CK16b, CCT14, DLA22, DP23, Eng17a, Eng17b, GGS22, Göd16, GSB^{+}16, HB21, HT17, Hjø15b, Hjø15c, HB15, HLB^{+}17, HK17, JBD^{+}19, JKS^{+}16, KKO15, LCK16, Lew15, LBA16, LZZ^{+}21, LCSL15, LB15, LLB20, LPB19, LK14, LJLH17, MWPF17, NSK17, PLPQB20, THC15, WVvLZ23, WLZB14, WLW^{+}17, ZLCB15, ZGPB17, ZYS17, dHSM17, He18]. retrieval-assisted [DP23]. Retrieving [LS18, WH15, YXJ18]. retweeting [SLLW20]. retweets [AKAS14]. reuse [AdUCDPH19, BSG19, CK16b, FKY16, HPL^{+}22, KY17, MSI^{+}19, PYW18, Tho22, WDL21, WB21]. reuser [Yoo17]. reveal [Hoo17]. Reverse [TR17, MZ22]. reverse-flipping [MZ22]. Revew [Sea20]. Review [Adl23, Afz22, Agu18, Aji17, All19, Ama19, Bat15, Bat21, Bea17, Ben18, Bor14a, Bow19a, Bow19b, Bru18, Bru14, Buc20, Bud17, Car14b, Cho19, Cle18, Clo18, Col17a, Con20, Dor19, Dow15, Dut20, Erd18, Far23, FM14, FJ19, Fou19, Fra21, Fre18, Fri17, Fro21, Gor20a, GGC^{+}19, Hal19, Han21, Haw18, Hay18, He18, Hjø17a, Huv20, Huv22a, Jul20, Kea22, Kno24, Kos20, Kos21, Lev20, Lin14b, Ma22b, Ma22a, Mat20, MPB^{+}20, Med15, Meh22, MS19, O’H18, OM15, Owe21, Pan21, QH18, Rey21, Rhe17, Rog15, Rou17, Sal20, San22, Sar16, Sav18b, Saw19, SK14a, See17, Sha21a, She17, Sil18, Sko23, Soe18, Sto20, SWLH17, Swa22, Tei16, Ten16, Uns18, Web20, Wei14a, Whi17, Zha19, Zim17, BCK17, Bor15a, Bur20, CK16a, CB15, Col20]. review [CKR14, DJ20, EMK^{+}15, EA18, GRSFV15b, GRSFV15c, GEBP22, H17, JL18, Kel21, KN23, KH17, LPL14, LKKP16, LHXS22, MMB17, MOM^{+}15, NB15, NBH16, NB17, OMM^{+}14, PEG^{+}19, SM22, Skö18, SDT22, VGC17, WDL21, YZYR21, ZSX15, ZW23, ZCvdH15, Fri17]. reviewer [CB15, RMDF23]. reviewers [GRSFV15a]. Reviews [Abb17, BI17, Baw14, Buc15, Cha14, Col17b, Cox14, Cox15, Cus15, Gui17, Hai16, Hak14, How17, Jon17, Lev15, Mar15, O’C16, Ols16, Sha21b, SR14, Sve15, War15, Wei14b, Zit15, vR17a, Bel16, HK14, Hol15, JL18, KT16b, KTA17, Lan15, RYEP18, SZZ16, Vec14, Wu23, XDDG18, ZvSvB14]. revisited [BR15a, Bor15b, GAB^{+}18, MMALCOM15, Sav19, WvdHS16, Rhe17]. Revisiting [Coc22, FA19, Moo20, REVR22, TJ20, FC18]. Revitalizing [CPP^{+}16]. revolution [Bor14b, Kno24]. RFD [WZL21]. Rhetoric [Joh17, Far23, PG21]. rhetorical [LJ22, RL15b]. rhythmic [CMS23]. Rican [VO23]. rich [KSL19]. Richard [Hol15]. richness [YZYR21]. Rick [Gui17].

Rieder [Sha21b]. right [LB22]. Rights [FM14, O’H18, Esk20, Mat15]. Rigid [Oh17]. rise [HRL21, Wu23]. Risk [PhD20, HSL24, MVL^{+}14, QH20, SV17, TD15]. Risks [RDS20, FMNG18, GZBA22]. risky [HSL24]. Robert [Bat15, Fri17, Sve15]. Robertson [Huv22a]. robins [EOG16]. robot [LYWF14, YR23]. robots [PPC19]. Robustness [BELR16]. Rocchio [THC15]. Rocchio-based [THC15]. Rockin [EOG16]. Rogers [Hol15]. Role [CK17, Wei14a, AGBN16, ALAB19, BBM16, BJR19, CW18, Du14a, GJ16, HWZ20, LM15a, LZ22, LZ19, LMM^{+}22, LB22, MA18a, MR17, NLP16, PM18, SB18, SHMC17, Smi17b, SC21, TMBA20, WLTY19, ZB19, ZH17, dRR15]. roles [CCLL14, HS17, JGCR15, KZ23, LSM^{+}14, LB15, LDS^{+}19]. Romance [The17]. Ronald [Dor19, Ols16]. roost [EOG16]. roots [MBBL14]. Rosenbloom [SK14a]. Rousseau [Dor19]. routine [RL15a]. routing [SZZ14]. Routledge [Owe21, Sto20, Swa22]. Rowman [Aji17, Bea17, Kno24, Lev20]. RPYS [MBBL14]. Rubel [San22]. rule [EOG16]. rural [CW18, PH22]. Rusk [Sko23]. Russia [PS14b]. Russian [KGS19, LPU15]. Ruthven [Huv20].

S [Baw14, QH18, SK14a, TMDZ21a]. Sachi [Bat21]. SAGE [Hay18]. Sakr [All19]. sales [RYEP18, YZYR21]. salient [LXLL14]. Samantha [Bea17]. same [KW14]. sample [Lew15]. sampling [TJ20]. Samuel [Ma22b]. San [Agu18, Cho19, Fou19, MS19]. sanctuary [TGC^{+}21]. Sandvig [Zim17]. sanitization [SB16]. sanitized [SB16]. Sanitizing [Mos17]. SanJuan [Car14b]. sarcasm [MGIGW16]. satisfaction [CBM17, FKY16, LCL16, RY18, WLTY19]. satisficing [McC15]. satisfied [RY18]. satisfying [MNNK15]. saved [DT18]. Saving [HHL21]. Say [MJ23]. scaffolding [BD16]. Scagnetto [He18]. scalability [BIW20]. Scale [Kim19b, ACS18, Ano17m, BC16, BWGD21, FN17a, HABT18, KD16, Los16, RMDF23, RPK16, WCP^{+}19, WHYJ17, GMB14]. Scale-free [Kim19b]. scale-independent [RPK16]. Scales [TL22]. scaling [RPK17, SS14a]. scarce [FVGL14]. scatter [ZBKG14]. scatter/gather [ZBKG14]. scattering [HHC^{+}14]. Scenarios [MM20, FVGL14]. schema [HNX18]. schemas [XDSC16]. Scheme [LLH19, BLGbH21, GB19b, Ko15, WKC14]. scheme-free [BLGbH21]. Scholar [LGZ14, Moo20, CWL^{+}23, GW18, GB19b, KKK24, AGGTGL18, AT14, BTMS16, GD18, Hal14, LCRGTS14, OA14]. Scholar-led [Moo20]. Scholarly [CA20, O’C16, SRM^{+}17, SHMC17, SWLH17, TK17d, vR17a, BL14, CWL^{+}23, DQAMM^{+}24, EJ14, GRSFV15d, H17, HT14, JLC16, KT17a, KIH16, LSHH23, LQ14, LWM^{+}23, Mat20, MOM^{+}15, MH15, MH16, MO20, OMM^{+}14, Pet19, RJS^{+}17b, TLA^{+}16, TM15, ZLY23, ZAE^{+}19, ZCYN23, ZGCB15, Rou17]. scholars [HH18, Lia17, OBD^{+}22, Xia23]. Scholarship [Mos17, Ten16, FW19, SS14b, Wu23, TK15]. Scholz [Dow15]. school [BCAL15, Kim15, KSL19, RC16]. Schools [SV20, WXSL23]. Schuman [Buc20]. Sci [BGS22]. Sci-Hub [BGS22]. Science [Alb15, Ano14, BLB16, BRS15, BR15a, Bru18, CCMT15, CK21b, EMK^{+}15, FW19, Hen19, Jon17, KS14b, Mat15, MOM^{+}15, NTNP19, OMM^{+}14, PYH16, PWK19, Sha14, ZS14a, ZZ20, ZSAW23, Bad14, Bat21, BR15b, Bor14d, BM15b, BK14, CS16a, CS15a, CYP15, CdRM21, DFB14, DM14, ESD22, FCT^{+}21, Fon21, FHO23, GRBCS15, GSB^{+}16, GJ22, GB19b, GMW19, HPLR23, H17, Haj20, HS17, Hof17, HAJ^{+}17, JdR17, Jam17, Joh17, Kim19a, Kon16, LCNP17, LHS24, LG14, LB16, LCD^{+}14, LR14, LBLL24, ML21, MM18, MdSLdO^{+}23, MBML24, MSTP^{+}22, MBD15, NF15, SSC17, SAHZ21, She16, SFL^{+}23, Söd23, SEZG22, SDT22, ST16c, TLD17, TJV14, UM23, Vas15, VS15, WW16, WHvdZvdB17, WKJW16, WTvR16, WWB21, WCO21, WXSL23, Yan16, ZWH^{+}23, ZS14c, ACS18, CHLRC14, HK14, LdMAdN16, Mak20, OSvE16, RL15a]. Science [Sch18, SJL19, ŠF17, VCJ22, Sar16]. Science-Based [NTNP19]. science-society [CdRM21]. Sciences [LPW14, ALYS16, BTMS16, Ior14, KGP^{+}20, LdN17, MT14, OVE14, PGFG^{+}20, RPK17, YB18, ZGE22, Hjø17a]. Scientific [Bud17, FMP16, GMB14, KPWS20, KB17, MT18b, Rhe17, RA14, SK14a, ZZZL17, ADC15, Ano17m, AMDFHVC24, BS17, BAGL16, Bor14b, BWL15, Bor16a, BDLM18, BGS22, CL21, CCM23, CCP15, CC16, CKL16, CB14, DFB14, DAT^{+}14, FDC20, FCT^{+}21, Fas18, GRS16, HPLC19, HGCKDR16, HLM16, HB16, JSY^{+}16, JL23, JQMY24, KISS15, KMFD15, LGST15, LRB19, LI21a, LGD20, LML23, LBC^{+}22, LBW^{+}19b, LZA^{+}20, MDC^{+}16, Mil14, MP23, MH16, ML16, ONZ23, Olt15, REVR22, Rou16, SJL19, TL18, VG18, VR17b, WL14b, WSZC22, WMM^{+}23, WTvR16, XOL^{+}18, XZW^{+}19, XBL^{+}22, Yan14, YRWH18, YZQT20, ZBDX18]. scientific-scholarly [MH16]. scientifically [Tho22]. Scientist [KY17, FKY16, KS16, WB21]. scientists [Ama22, BM14b, Bor14d, BBH15, Bor17, HH19, OBD^{+}22, RD16, WKJW16]. scientometric [Hen14, dSMdAGQ15, VCG17, RRHL17]. scientometricians [Vin17]. scientometrics [dRR15, Bor14b]. SCImago [BdMA14, GNVQdMA16]. scope [ALC^{+}14]. scoping [GJ16]. Scopus [Aba24, GD18, LdMAGB15, LdMAdN16, MT16, TS21]. score [MWD17, OA15, PLPQB20, VLBG^{+}16]. scores [Bor17, Gre16, TKA^{+}23, WK19]. scrape [Wu23]. scrape-and-report [Wu23]. screen [LZG^{+}23]. screens [KTS^{+}15]. sculpting [Rut24a]. SDGs [CK17]. Search [Col20, Jul20, LLH19, MS19, SJ17, Wei14a, ASTU16, AS24, AP17, AGJ^{+}16, Bel16, CWdR17, CCR24, CML^{+}15, EH15, FT17, GISW15, Gor24, GVF^{+}24, Hal14, HE22, HSL24, HABT18, HHSP22, HS22, JJC19, KBD^{+}14, KvOS^{+}17, KTS^{+}15, KTS^{+}16, KPWS20, KTPE14, KS14b, Lew15, LKRS18, LHS24, LA14, LXLL14, LCWC23, LLB16, LZ19, Lop14, LJLH17, LYOW15, LN19, Mor24, NH15, OA14, OA15, PKK^{+}24, PGK24, RDPF16, RLAB^{+}24, RJK20, SY20, SY22, STS14, STS15, SAB^{+}18, SL21a, SKO24, SHGI16, Smi15, Smi17a, STCS17, SL21b, SLH22, TC17, UH17, VVP^{+}21, VRF14, VC15, VS14, VR17b, VSJR19, WOJBG17, WS22, WMX^{+}22, Whi14, WTT^{+}14, WV18, WLTY19, WLZ^{+}20, WDTC20, XJBK17, YSAM^{+}21, ZBKG14, ZLCB15, ZGBIL17, ZGPB17, vdVZK21, Bat21]. search-assisting [LCWC23]. searcher [LLB16]. searchers [STS15]. searches [Hjø15c, WSvPK16]. Searching [BM20, CCR24, GGC^{+}19, Soe18, Zha14c, BS16, CBM17, KWR20, NPG^{+}20, SF15, Zha14b]. Seattle [BNG14]. Seb [Meh22]. Sebastian [Med15]. second [AD15, CLSS14, KA22]. secondary [SLS^{+}19, Wat14]. secrets [Zim17]. section [ONZ23]. sector [GVB17]. secure [Min16]. Securing [SV17]. security [ADM^{+}20, HH14, SV17, VGN20, Ano17n, CZ16]. Seeing [OHP16, HB16]. Seeker [Fou19]. seekers [LI22]. Seeking [BM20, Fre18, KP18, Sav18a, Alb15, CHJ20, CPL22, CN14, Cho23, Col17b, Du14b, Fre15, Gau16, Gor22, GEBP22, Her23, JDZ24, LKKP16, LS22, LJ17, MR17, MAP15, PB15, RJS19, RJS17a, Sav17, SZ20, Sha14, Sin16, St.17, TG15, TT17, TZS^{+}22, WFRN14, XBCH18]. segmentation [DGR19, ZTA20]. Segmenting [ÇÖ18]. segments [XC17]. Selected [GGC^{+}19, Soe18]. selecting [ZH17]. Selection [LWA^{+}17, AFL^{+}14, CHJ20, GRSFV15a, LZ19, SZ20, SL21a, SK23, UH17, WvdHS16, Zha14b]. Self [FVGL14, Hal19, OGL19, CDL22, Don16, KCZ^{+}14, LOD^{+}19, LZZ^{+}21, MO20, Wu19, Zha16, ZT22, ZSC18, dC16]. self-archiving [LOD^{+}19]. self-attention [LZZ^{+}21]. self-citation [Don16]. self-citations [ZSC18]. self-determination [Zha16]. self-disclosing [KCZ^{+}14]. self-identity [MO20, Wu19]. Self-Presentation [OGL19]. self-references [CDL22]. self-referential [dC16]. self-tracking [ZT22]. Self-training [FVGL14]. Selves [Jor19, Gor20a]. SEM [LL17]. semantic [ADD^{+}18, HL20, JL23, KK16, KS14b, LKY^{+}20, LBA16, LWL^{+}18, MGCTW15, MSK^{+}14, MRAF16, Par19, Smi17b, SLE21, VG18, VT16, YXAL23, BOHK19, MSdGS19]. semantic-enhanced [YXAL23]. semantics [JSW14, LML16, Swa22]. SemGraph [MRAF16]. semicontrolled [SP17]. sensations [LJK21]. Sense [GB17, EH15, TT17, TZH23]. sensemaking [GBS15, TT17, ZS14a, ZS20, PB16]. sensitive [WLZ^{+}20]. sensitivity [DD16, DBBJ16, GMRS17]. Sentence [FDR16]. SEntFiN [SKKM22]. Sentiment [Ano17n, Bi22, CZ16, Pal16, ALC^{+}14, CHT^{+}15, CTM16, LZG^{+}23, MDJ^{+}19, RDPN21, SKKM22, XCL^{+}21, YCL20]. Sentiment-based [Pal16]. Separating [Jor19]. separation [HCC^{+}21]. sequence [KDS^{+}17, MC18]. Sequential [AHJY18]. serendipitous [SZWS22, WFRN14]. Serendipity [Erd18, SVB22, LHM17, MPT15, QLM^{+}22, MBW^{+}14]. series [LZG^{+}23]. service [BK19, CKLP17, JHJ15, KCZ^{+}14, LL17, LYWF14, SSdMCM20, WLZ^{+}20]. Services [Fou19, MS19, Aji17, CS16b, LP16, LWA^{+}17, LWPG^{+}19, MJ23, RCM22, Cho16, Agu18]. session [AWC18, LB15, RJK20, STS14]. session-specific [AWC18]. set [IABL16, ZRdR21, ZYL18]. sets [Col15]. setting [Gre16]. settings [SV17, SK14b, WS22]. settling [SH19]. Seven [KPC^{+}15, BVH15, KGP^{+}20, The20]. seven-country [KGP^{+}20]. seven-model [BVH15]. Severity [JVF19, CYYG17]. shadow [MR19].

Shah [Fre18]. Shangri [dHSM17]. Shangri-La [dHSM17]. Shaobo [MS19]. Shape [Jor19, OSD22, YXJ18]. Shaping [OD16, RLAB^{+}24, Rut24b]. share [AGBN16]. Shared [BWF14, Fou19, RCM^{+}17a, She16, CCM23, IQDS16, SBS23, SC21, RY18]. Sharing [LMA^{+}22, TR16, AGM19, BIW20, BSG19, Cab16, CKL16, CJ22, DAT^{+}14, Du14b, EJ14, HCC^{+}21, HSU16, HD22, HT14, HLC15, JS15, JGCR15, JWWV18, KS16, KN23, OS15, PYW18, Pil14, SSS20, SW18, SBA^{+}15, SWL18, WG19, WKJW16, ZGL^{+}22]. shelf [MDJ^{+}19]. Sherif [All19]. shift [ML21, PR16, TMDZ21a, TMDZ21b]. shine [SHGI16]. Shiri [Wei14a]. shopping [CLL15]. short [FCT^{+}21, KMD16, LZG^{+}23, YXAL23]. short-term [KMD16]. short-text [YXAL23]. shot [KK19b]. Shots [KK19a]. Should [GSG20, WXSL23]. Sichuan [ZZY23]. side [SRPS16]. sides [SKO24]. Sigfried [See17]. sighted [XBL^{+}21]. Signals [Mos17, GRC21, RVGPS19, RVdAGP21, WLTY19]. Signed [GB19a, NB15]. significance [LBM19a, PLPQB20]. significant [FW19, HK21, Tho22]. Signification [LJI18]. Silvia [Kea22]. similar [Bar14, GARA^{+}16]. similarities [KMKL15, MGKA16, ZL16]. Similarity [CBvEW18, ADD^{+}18, Col15, GARA^{+}16, LKY^{+}20, MYK15, WW15, ZRG16]. similarity-based [Col15]. Simon [McC15]. Simple [Hal20, Bor15c, Hal17, KS17, PHH^{+}20]. simplification [FDR16]. Simplified [LLH19]. simulated [HHSP22]. simulation [MGK20]. Sinan [Afz22]. since [Hof17, LGST15]. single [BM14a]. Sins [PRM22]. Sister [Jan15]. site [EAB17, JDHL17, KT14a, LOD^{+}19, MKK17, Sav15a, TK17a, TK17c]. Sites [Jor19, OGL19, CK16a, CJ16a, EA18, HCC15, Kan22, LCM17, MK15, Min16, Tho22, ZXPS18]. situ [HWW^{+}17]. situating [GM19]. Situation [Fou19]. situational [SBS23]. situations [HRB14, XBCH18]. Six [DTDT21, GRS16, SBH21, The20]. size [LGST15, Pra15a, RR17, TK17b, YHLG21]. sized [LN17]. sizes [LBM19a]. Sleeping [LS16, MA18b, YB18]. SlideShare [TK17c]. Sloan [Hay18]. Slovenia [KMFD15]. Small [OB19, BPB21, HRN^{+}23, KDS^{+}17, KTS^{+}15, LN17]. smaller [GRS16]. smart [LCLM18, MMFA20, MVA22, RDS20, YB18, dP20]. Smartphone [CS15b]. Smartphone-based [CS15b]. smartphones [FLJD23, KCL14]. Smith [Fri17]. smoking [ZY15]. smoothing [LBA16]. Social [Bow19b, CZ16, Cox14, FKY16, GRSFV15d, GB19a, Gre19, HCC^{+}21, Hal19, Hay18, Jor19, Lan15, LZ22, Lin14b, MT18b, NTNP19, OGL19, PPLJ20, PhD20, SBS23, SLS^{+}19, Sin16, SV20, ZZLS16, ZAE^{+}19, dSBG19, AMOM16, ALYS16, BCZ14, Bi22, BTMS16, BDH17, BKM15, BV21, CPP^{+}16, CCC14, CBH16, CLL15, CJ16a, CY14, CdRM21, DM14, EAB17, EA18, ESD22, FJF^{+}21, HH19, HHL21, HCC15, JdR17, JS15, JHJ15, JDHL17, KT14a, Kan22, KA22, KGP^{+}20, LCM17, LOD^{+}19, LJK21, LLD^{+}17, LWM^{+}19, Lia17, LYL21, LML16, LYCX14, LCSL15, LCL16, LJ17, LDS^{+}19, LZ21, LCWL18, MMFA20, Mar16, Meh23, MSW^{+}19, MVA22, Mil14, MK15, Min16, MT14, MSTP^{+}22, MAM19, MGIGW16, OS15, OVE14, PHZF16, POK15, Par19, PPC19, PBN^{+}20, PT21, RCM22, RKL14, RDPN21, RC16, RVGPS19, Rub15, SMRR16, SYS17, SMW14, SP22]. social [Shm18, SK14b, SLCS18, SWLH17, SWSZ19, TG15, TK17a, TG17, Vau16, VS14, WZFZ22, WCO21, Wu19, WLG23, XZ19, YLHW21, ZY15, ZCZ^{+}20, ZZY23, ZWTX23, ZXPS18, ZPL^{+}20, ZGB16, ZGE22, ZH17, ZZ21b, Ano17n, CLSS14, DOM^{+}14, LPW14, Sto20, TK14, Fre18, Pan21]. Social-media-based [CZ16, Ano17n]. socially [SAHZ21]. Societal [FHW17]. Society [Ama19, Ano14, Med15, Saw19, CdRM21, YZ16, Baw14]. socio [DS15, HL20, KZ23, WCO21]. socio-emotional [KZ23, WCO21]. socio-epistemic [DS15]. socio-semantic [HL20]. sociocultural [CK21b]. sociolinguistic [LJ22]. sociological [BF18]. sociology [CK21a, GD18, Riv15]. sociotechnical [SSC17]. Sofia [Adl23]. Softcite [DCLH21]. softcover [He18, See17]. Software [CA20, HB16, Ack21, Aji17, CK16b, DCLH21, Fre15, YRWH18, YLHW21]. solidarity [Ano14, SMAB09]. Solo [CHH14]. solutions [Bor15c]. solve [Kos21]. solving [KKK24, Sin16]. Some [Mal15, Egg14a, LR14]. Someone [SSA20]. something [Her23, Kit19]. Sonnenwald [Hjø17a]. Sons [Bru18]. sophisticated [FDR16]. Sorting [WHZ23]. sound [Gia18]. Source [KSB20, WC14, Aji17, CHJ20, CRGG17, PGK24, SMRR16, SBIT14, STAH14, WvdHS16, YZ17, YLHW21, Zha14b]. source-based [STAH14]. sourced [CPP^{+}16]. Sources [RR20, QH20, RHG23, Sil20, ZB19, ZH17]. South [KN23, POK15]. sovereignty [MMFA20]. space [MGKA16, WV15]. Spaces [Han21, KKZS20, RTS^{+}17, SLS^{+}19, VP19]. Spain [GD18, RA14]. Spamming [KIH16, SS17]. Spanish [Ano14, AGGTGL18, Bor16c, CCM23, DFB14, PMGB^{+}21, PdlRR^{+}22, SMAB09]. Spanish- [Ano14, SMAB09]. Spatial [Söd23, SE16, LF21, MMHA23]. spatiotemporal [STD15]. speakers [MSPP21]. speaking [HCC15, The20, Mar15]. Spear [Fri17]. Special [LHS24, AGBK19, Ano18z, GMW19, MWPF17, TMDZ21b, WJB22, ZSG^{+}22, ZSAW23, ZDP^{+}23]. Specialism [SLS20]. specialty [LG14, NF15]. specific [AWC18, AKAS14, Bor16a, Chi19, DDB16, HRN^{+}23, SGLF17, SZZ14, Zha14c]. Specifying [BdMA14]. Spectroscopy [MBBL14]. spectrum [CV20]. spectrums [Sil20]. speculation [CTM16]. speculative [FC22]. speech [KK16]. sphere [Esk20]. spheres [RAW22]. spillover [LF21]. Spiteri [Pan21]. spoken [STCS17]. sporting [BPB21]. spots [LdN17]. spread [WZFZ22, Kos20]. spreading [MSPP21]. Springer [All19, Ama19, Fra21, Fre18, Gui17, He18, Jul20, Uns18]. SQA [RCM^{+}17a]. square [BS17]. stabilities [WN15]. stability [FDC20, LS21]. stable [MDJ^{+}19]. Stack [GSG22, Kan22, YHLG21]. Stage [Sav18a, LS17, LB15, SGV14]. Stage-based [Sav18a]. stages [HWW15, LLB16, MYK15, RJS19, ZZ21b]. stakeholder [SBA^{+}15, TLS^{+}23, WWL16]. stakeholder-driven [SBA^{+}15]. Stakeholders [DSBG21, SFL^{+}23]. stamped [SHQ^{+}16]. standard [Mal15, ZLY23]. Standards [Fur20]. Stanislaw [Bru14]. stanzas [PdlRR^{+}22]. star [TZK20]. stars [DSBG21]. start [ACL17, MBAG16]. starting [GBZ^{+}23]. state [DM14, LWPG^{+}19, Sto20, TGC^{+}21, ZZE17, Sav15b, Sav16b]. state-of-the-art [DM14]. statements [CK21a, LBWE^{+}23, LML16]. states [RB14, POK15, RN16, SH19]. Statistical [LBM19a, KKO15, PLPQB20, RBJ22]. statistics [BS17, TW16a, Aji17]. stature [CHH^{+}17]. status [SL24, YCL20]. Stefano [He18]. Stein [O’H18]. STEM [SR16]. Stemmer [KF16]. step [JLC16]. Stephann [Erd18]. Stephen [Rog15, Zha19]. Steps [LI14a]. steroid [HGCKDR16]. Steve [Jon17, Rhe17]. stick [Fra21]. sticky [BL14]. still [Bor14a, KPWS20]. stimuli [HBS15]. Sting [H17]. STM [Bud17]. Stochastic [FT17]. Stock [Ama19, Sar16, LWM^{+}19, Sar16]. stone [HNX18]. stop [LPB19]. stopping [LPB19]. storage [Hjø15b, Mat22, RLAB^{+}24]. stories [HDT23, KT17b, SGS21]. Storing [All19]. Story [LCLBY17, McD21, Lev20, LMM^{+}22, TC17]. Story-focused [LCLBY17]. Storytelling [McD21]. Strategic [RRHL17]. strategies [DSBG21, KMD16, MBW^{+}14, NB15, QHH20, Sav16b, STAH14, WMX^{+}22, Yu14]. strategy [GAS^{+}20, Mat22, TPZ15]. stream [ZS19]. streams [GOdR18]. strength [Riv15, Shm18, Söd23, WL14a]. strengths [ADC15, ZZL^{+}14]. Striking [SLeH21]. string [KK20]. Structural [Shm18, DPPV14, GK14, LBWE^{+}23, LQ14, SBZ^{+}14, TW16a, VG18]. Structure [Jor19, ALAB19, ÇÖ18, CL17, Hoo17, HQH^{+}22, KMKL15, LBG^{+}21, LSHH23, Los16, RL15b, YY16]. Structured [ONZ23]. Structures [Swa22, BLB16, CK21a, KMFD15, MR14]. Structuring [LYOW15]. Struggling [KL22, WV15]. Stuart [OM15]. Student [JAG^{+}20, STAH14, BCAL15, KCL14, Kim15, OB19, QH20, RC16, SMRR16, Wat14]. Students [KP18, CBAL15, ELB17, SZ20]. studies [Adl23, BVd16, CO16, Coc22, ESD22, GEBP22, Haj20, LB16, Mad16, PEG^{+}19, RDS20, Tho22]. Study [Cox14, KSB20, OGL19, ZZ20, ACS18, AS14, BTMS16, BH17b, Cab16, CK14, Che16, Cho20, CBM17, CdRM21, DDB16, FJT23, FDC20, FLJD23, Gau16, Gia18, GJ16, GAS^{+}20, Har14, HSL24, HT17, HK14, HS22, HWW15, JML22, KMD16, KIH16, KMFD15, KGP^{+}20, LL17, LMJ21a, LMJ21b, LM15a, LOD^{+}19, LS22, LOM22, lLTBH15, LR14, LPB19, LK14, LM15b, MSdGS19, MD16, MGIGW16, PHZF16, POK15, PPC19, PPLJ20, PHC17, PB15, PHJ23, QSH14, RDPN21, SYS17, SLI^{+}15, She21, SMW14, SP22, SLeH21, TG15, UM23, WCC^{+}17, WS22, Wat14, WYZL21, XA14, XLKC22, YL18, YRWH18, YHLG21, YR23, ZYZ^{+}16, ZS20, ZCZ^{+}20, ZLF20, Ama19]. Studying [HE22, MWMB18, BBC^{+}22, Zim17]. Stuff [Jon17]. style [GBL^{+}18, TND15]. styles [CDL18, KTPE14, LCB^{+}18, LBW^{+}19b, Sin16]. stylistic [AKAS14]. subfield [GS18]. subfields [PDS22]. Subgroup [YR23]. Subject [KLG14, Ols16, Ano18-27, Bjö14, CHLRC14, HCDÅ15, LB16, LM15b, LML18, SP17, ST16c, TS21, WV18, WLZ^{+}20, ZRG16, ZZSX16, Hoo17]. subject-oriented [WV18]. Subjectivity [Ben18]. submit [PR16]. subnet [ZZL^{+}14]. subscriptions [PSB16]. subsentence [CHT^{+}15]. subsentence-level [CHT^{+}15]. subsequent [LPL14]. substantially [TM20]. substitute [RHG23]. substitution [YZYR21]. subsumers [ADD^{+}18]. succeed [Sto20]. Success [WSvPK16, AP17, BDX^{+}18, DLD^{+}21, YP16]. successes [LYC^{+}22]. successful [TL18]. SUDMAD [AHJY18]. Sue [Clo18]. suffering [HR19]. suggestion [RJK20]. suggestions [TPR17]. Sugimoto [Rou17, vR17a, O’C16]. sum [MOM^{+}15]. Summa [Bru14]. summaries [AG15b, MG21, STCS17]. summarization [CBJ19, FDR16, KK16, KK19b, ONZ23, YHS16]. summary [BST15, RVdAGP21]. superiority [The20]. Supervised [QSH14, AFL^{+}14, HPLR23, KKOS21]. supply [GLZ^{+}24]. Support [BDA^{+}18, CY14, FLJD23, HS22, IESG20, KLC^{+}14, MBW^{+}14, MR19, MJ23, OS15, Rub15, RBJ18a, Sav15a, SYG17, STCS17, WHWZ21, XJBK17, ZY15, ZB19]. supported [IQDS16, OHP16]. supporting [LM15a, LCWC23, NB17, ZBKG14, ZPPW24]. surname [KGS19]. surnames [KGS19]. Surroundings [Fou19]. Surveillance [New23, SSA20, TMBA20, ZPPW24]. Survey [Fur20, AS24, BAG17, CCMT15, CPL22, Col17b, JHJ15, LLB20, MTK16, SB18, WCP^{+}19, WB21]. survey-based [SB18]. Survival [PGFG^{+}20, Kan22]. sustainability [Cho16, CK17, TPZ15]. Sustainable [KC20, CK17, IESG20]. sustained [YLHW21, Zha16]. Sweden [HH18]. Switzerland [All19, Ama19, Fre18, Gui17, He18, Jul20]. syllabi [KT16a]. Syllabic [CMS23]. symbols [CL17]. symptom [SKA18]. synergies [She19]. synergistic [SWSZ19]. Synergy [LWPG^{+}19, LPU15, LI21a]. syntactic [VAGR15]. syntagmatization [Eng17a]. syntax [JSW14, VG18]. syntax-semantic-based [VG18]. Synthesis [Agu18, Fou19, MS19, DPHP17, GEBP22, NGTKR15, PEG^{+}19, She16, She17, YL14]. synthetic [WS22]. System [LBM^{+}19b, Soe18, AJWB16, AG15a, ALAB19, Fas18, FS18, Fia14, HT17, LSHH23, MR19, MGK20, NSK17, PMGB^{+}21, Ric18, XJBK17, ZZLS16, ZYS17, Lin14b]. System-Initiated [LBM^{+}19b]. Systematic [MdSLdO^{+}23, Bel16, EA18, LKKP16, MOM^{+}15, OMM^{+}14, PEG^{+}19, SM22, ZSX15]. Systematically [GHKB^{+}17]. systemic [SSS20]. systems [CS15b, FMD19, GTZŽ14, HBS15, HWW15, HEEB^{+}22, LCH^{+}22, LCK16, Lev20, LI14a, LPU15, LWPG^{+}19, NGTKR15, PRM22, PRRC18, SY22, SVK17, SVB22, SBH21, SWL18, YP16, ZHGY18, San22]. Szostak [Gui17].

T [Bor14a, Buc20, Hak14]. Table [Ano16n, Ano16o, Ano16p, Ano16q, Ano16r, Ano16s, Ano16t, Ano16u, Ano16v, Ano16w, Ano16x, Ano16y, Ano17o, Ano17p, Ano17q, Ano17r, Ano17s, Ano17t, Ano17u, Ano17v, Ano17w, Ano17x, Ano17y, Ano17z]. tables [CHH14, HCKH15, MA18b]. tablet [MD16]. Tackling [Cox15]. tactics [FV21, XJBK17, XBL^{+}21]. tag [BAG17, JSW14, MBAG16, QLM^{+}22, SZJ^{+}21, TTH16]. taggers [CK14]. tagging [Cho15, CS16c, DP23, lLTBH15, PHZF16, Pan21, WXLBO15]. tags [CPP^{+}16, EHdBvO17, KK16, KLG14, lLTBH15, LK14, PHZF16, XZ19]. tailored [KF16]. take [JS15, JWWV18]. take-up [JS15]. taken [LYCX14].

Takes [See17]. Taking [FSK^{+}16, Huv22b, ZXLC18, ZRG16]. tale [Fra21]. talent [ZZE17]. Talk [MCSB19, AST^{+}22, MWMB18, Xia23]. talks [MSPP21]. Tammera [Erd18]. Tar [Tho22]. target [Hil14b]. targeted [NLP16, SLCS18]. Task [HK17, LLH19, CBM17, Du14b, Gre16, Kim15, KKK24, Lai15, LB15, OHP16, SJ17, WOJBG17, WS22, WDTC20, ZFH^{+}19]. task-based [Lai15]. task-driven [Du14b]. task-solving [KKK24]. tasks [AGJ^{+}16, Bor15c, DC22, EH15, HHSP22, KKZS20, LB15, SJ17, STS14, SHGI16, SHW^{+}15, TT17, VSJR19, WSvPK16, Xia14, ZBKG14, ZXLC18]. Tatjana [Cho19]. taxonomic [MJ23]. Taxonomy [KB17, CDL18, HCC^{+}21, KF16, LJ17, NGTKR15, WKC14, ZRdR21]. Taylor [Hak14]. teach [WXSL23]. Teacher [DO14]. teachers [ELB17]. Teaching [Kon16, KT16a, MAAL^{+}16, RAMP22]. Team [LGST15, XBL^{+}22, DYB14, MR17, OHP16, SDUDL24, ZFH^{+}19]. teams [RY18, WG19, YR23]. teamwork [WTvR16]. tech [LWPG^{+}19]. Technical [KB17, AD23]. technique [Har14]. Techniques [Sav18b, Sha21b]. Technologiae [Bru14]. Technological [Hil14b, Lan15, MMFA20, KNY^{+}14, OMTK17, SDUDL24, YL17b]. technologies [Coc22, GS18, JS15, RRHL17, RLAB^{+}24, SRPS16, STD15, SK23, Cha14]. technologists [SHKL23]. Technology [Ano14, BVH15, SSA20, AGBK19, AGBN16, AD23, Bat21, CAS21, DB17b, FA19, GW18, Hil14a, Hof17, HHPI21, KT14a, KSL19, LL17, LB16, MBVB17, MBML24, NF15, PM18, WG19, ZZY23, FM14, O’H18, Rog15]. technology-mediated [Bat21]. technology-rich [KSL19]. technoscience [FC22]. TED [MSPP21, PYH16]. Teen [CBH16, EWX^{+}16]. Telecommunications [Wei14b]. Telerehabilitation [NTNP19]. television [SP17]. template [YHL^{+}21]. temples [HNX18]. Temporal [GSG22, GHCW17, LCLM18, MAM19, Sav18a, LS17, RJS19, RCM^{+}17b, SRG^{+}16, TS16, WLC^{+}14]. temporalities [MB23]. Temporality [BCAL15]. Tencent [LTD^{+}15]. Tensions [Kha19]. Term [HB21, BRL20, CL17, CLA^{+}16, DLA14, ESD22, KMD16, Ko15, SAW21]. term-weighting [Ko15]. terminologies [SM18]. Terminology [GTZŽ14, TPR17]. terms [LdMAdN16, RN16]. territories [ADC15]. test [FN17a, Gre16, LPB19, PDŠB16, PB16]. Testing [AvS16, MBD15, Riv15, SLM15, WG19, CBH16, MVL^{+}14, PDŠB16, SWSZ19]. tests [PLPQB20]. Texas [Hjø17a]. Text [BDA^{+}18, KCZ^{+}14, Sav15b, Sav16b, Sav18b, SEZG22, AFL^{+}14, Cab16, DFB14, DGGdST19, GZBA22, GOdR18, KF16, KLH17, Ko15, LCB^{+}18, LLB20, LZ21, McC15, MDC^{+}16, MH16, MRKT15, MSI^{+}19, ONZ23, SAA^{+}23, SRG^{+}16, TZH23, VG18, XC17, XZ19, YXAL23, ZS14b, ZS16, ZYL18, ZSC18, DMdO16]. text-segments [XC17]. text-sharing [Cab16]. textbooks [CLA^{+}16]. Texts [MT18b, Zem17, DMdO16, MSK^{+}14]. textual [CC22, DGGdST19, GHKB^{+}17, KMJR18, LML16]. Thames [Swa22]. Thanks [GB16, MT20]. their [Bor17, ÇÖ18, Che15, EH21, HHL21, HPL^{+}22, HCDÅ15, JAG^{+}20, KK16, KTPE14, LC21, MRKT15, SL21a, SWL18, TM15, Tsi16, VXN18, YCL20, YCK16]. Thelwall [Agu18]. them [MBW^{+}14]. Thematic [SV20]. theoretical [BBC^{+}22, GTAS22, KWLV22, Kue23, LOM22, Mal15, ST22, SA15, TKB^{+}17]. theories [DL22, Fon21, GTAS22, Haj20, SLM15, Car14b, vR17a]. Theorizing [JS15, Haj20]. Theory [Abb17, Baw14, Cox14, Haw18, PB16, Sil18, Soe18, WTvR16, AFB20, BMG^{+}17, Bur14, CLL15, CC17, Cox23, FLJD23, Hjø17b, LB19b, LBG^{+}21, LA14, LML16, MBVB17, Riv15, Rut22, SLLW20, STF19, WB15, WJB22, YR23, ZvSvB14, dFB22b, Adl23, Hjø17a]. Theory-changing [WTvR16]. theory-driven [LML16]. there [Bor14b, Rub15, SYF17, TSY15, WMX^{+}22]. Thesauri [Wei14a, Eng17a, Hjø15a, WKC14]. Thesaurus [KMKL15, Los16]. these [DQAMM^{+}24, GARA^{+}16, MWMB18]. Things [MCSB19, Saw19]. Thinking [LCWC23]. thirty [NGTKR15]. Thomas [Car14b]. Thought [SV20, FMD19]. Thousand [Hay18, TR17]. thread [BBM16]. Threads [OGL19, GZBA22]. threat [HKBB18]. Three [SC17, Pra14a, VP19, WK16]. three-class [Pra14a]. three-dimensional [Pra14a]. Threshold [CZZY15, DMMF17]. thresholds [BdMA14]. Thylstrup [Con20]. tie [Shm18]. ties [Cro15a, SMW14, ZB19]. Tijana [Bow19b]. Tim [MSdGS19]. Time [EHdBvO17, Fie16, GSG22, DC22, DAT^{+}14, KK19b, LB15, RBJ18a, RBJ18b, SHQ^{+}16, TSRG16, Vas15, WZK^{+}18, ZZ21a, ZGBIL17, ZSMF15, SR15]. Time-based [EHdBvO17]. Time-dependent [Fie16]. time-stamped [SHQ^{+}16]. Times [MB23, HCC^{+}21, JAG^{+}20, SBS23]. titled [Col20]. Titles [PDS22]. Today [Rou17, Eng17b]. together [TT17]. TOIM [LTD^{+}15]. Tom [Hak14]. tomographic [LKKS17]. Toni [Car14a]. Tonje [Han21]. Tonkin [Sav18b]. too [PdlRR^{+}22]. tool [AGMR18, DC22, DMdO16, DPN16, LCWC23, Mor24, ZCW17, dHSM17]. Tools [Sav18b, DM14, IQDS16, TK17d, WG19, WV18, ZSX15]. top [Bor15b, Hal14, MMALCOM15, SDT22, ZMJ14]. Topic [BLGbH21, KBS18, KK19a, LTD^{+}15, ACBC22, ABLS17, AFRH18, BMG^{+}17, CCT14, DC14, GLZ^{+}24, HL20, HDZ^{+}15, LN17, LB15, LLB16, LX17, Mat15, MG21, MCY^{+}21, PS19, SHQ^{+}16, Sta18, ST16c, XZW^{+}19, Yan15, ZZLS16]. topic-based [ZZLS16]. Topic-Irrelevant [KK19a]. Topic-level [LTD^{+}15, Yan15]. topic-related [CCT14, Sta18]. Topic-Relevant [KK19a]. topical [SKL17]. topicality [HSK15]. topics [CC22, ML21, SLLL17, VCJ22, VZRP15, Wan18, WZFZ22, ZZZL17, HRB14]. topology [SJK15]. Total [PSB17, OMTSLC15, PSB16]. tourist [ADL15]. Tourte [Sav18b]. trace [TH16, YL14]. traces [LML16, MA18a]. Tracing [MA18a, FvdWR^{+}22, MHMW17, Tab14, LCHL14]. tracked [JAG^{+}20]. Tracking [BCAL15, DC14, FA19, GHCW17, KTS^{+}15, ZT22, vdSvdB23]. trade [WXLBO15]. trade-off [WXLBO15]. trading [YZ17]. traditional [AZE^{+}19]. training [DJLS22, ECJF21, FVGL14, HRN^{+}23, Yu14]. trajectories [SL24]. Trajectory [BL14, SS14c, TH16]. transcribed [KK16]. transfer [Ama22, JGCR15, PMSA14]. transformed [Kno24]. Transforming [DTDT21, NF15]. transgender [HK21, KWLV22]. transient [BL14]. transition [HCC^{+}21, KA22, XOL^{+}18]. Transitioning [Ama19]. transitions [JDZ24, Rut22]. transitivity [ZBDX18]. translate [AMDFHVC24]. translation [AKC^{+}18, Cj15, ECJF21, KKO15, Mat20, SAA^{+}23, SKC18, Wor19]. translocal [AC18]. transmedia [BAK^{+}17]. transmission [LYL21]. transmitters [Ama22]. Transparency [BJSM18, JML22, Mah22]. transportation [BNG14]. travel [Sav15a]. Trebor [Dow15]. TREC [TCP15]. tree [KLH17]. trends [LWM^{+}19, MBZ^{+}22, TK17c, VRF14, ZSMF15, VC15]. triage [HABT18]. Triaging [CYYG17]. trial [LMA^{+}22, MPB^{+}20]. triangulation [Gre18, WHvdZvdB17, WB15]. tribute [Car14a]. trilateral [ZPL^{+}20]. triple [CYP15, LPU15]. triple-helix [LPU15]. triples [SKA18]. triplet [Bi22]. Trolling [FR23, SYF17, FV21, SF20]. Tropical [CKR14]. true [Hjø15a, Min16]. Trump [CY22, Tei17]. Trust [VFZ^{+}22, DL22, H17, JRL20, LXW18, RY18, Yoo17]. Trustworthiness [HKBB18, TLA^{+}16, DC15, HB14]. Truth [RL15b]. Trying [Her23]. tumor [wYKY17]. Turing [Hai16]. Tuscaloosa [Far23]. TV [LCLM18, LZG^{+}23]. Tweet [YHS16, CCT14]. Tweet-biased [YHS16]. tweeted [DT18, TGV^{+}16]. Tweeting [HPS^{+}14]. Tweets [ZZ20, AKAS14, CBJ19, FDC22, HPS^{+}14, LZD16, LS17, SS17, SLLW20, ZS19, HBH^{+}16, LYOW15]. twist [FSK^{+}16]. Twitter [AKAS14, BYSB16, CY22, ESWC18, EH15, FDC20, FCT^{+}21, FR23, Gia18, GAB^{+}18, GB16, GKM^{+}21, HBH^{+}16, HL20, IS22, Jor19, KYFZ15, LCB^{+}18, LYOW15, MRMCMVUL14, OS15, OMTSLC15, Pal16, SS14c, Vau16, VAGR15, YCK16, ZS19, ZGL^{+}22, ZSMF15, Zub18]. Twitter-specific [AKAS14]. Twittersphere [ZZ20, EOG16]. Two [RDS20, CMB17, FV21, JLC16, JGY^{+}19, KW14, LS17, QZWC21, Smi17b, WSZC22, HM14, SGV14]. two-stage [LS17, SGV14]. two-step [JLC16]. two-view [JGY^{+}19]. TX [Hjø17a]. Type [KB17, LBM^{+}19b, AG15b, AKAS14, DTW16, Ior14, LZ22, LB15, PGK24, St.17, ZWTX23]. Types [KKK24, Oh17, BR16, CCLL14, GARA^{+}16, HBS15, SJ17]. Typology [MCSB19, Cho20]. Tzitzikas [Fra21].

U.S [Ano17n]. U.S. [CZ16, JVF19, LWPG^{+}19, LML16, MSB21, ZZ20, ZF19, ZZE17]. ubiquitous [DJ20]. UK [Bow19a, Col17b, Fro21, Kos21, Mat20, Pan21, Rey21, PSB17, TGV^{+}16, TKA^{+}23]. unavailable [FDC20]. uncaptured [JLC16]. uncertainties [ZS16]. Uncertainty [DBBJ16, LN19, DD16, GN17, SLLW20, ZHGY18]. uncited [KAL24]. uncover [IABL16]. Uncovering [LML16, MR19, Vau16, DQAMM^{+}24]. under-cited [HR19]. underlying [LJK21, TSY15]. understand [DL22, WZL21, ZY15]. Understanding [ASTU16, ADL15, BDLM18, BDX^{+}18, CB15, CKL16, Gor20b, HHL21, KT14a, KTS^{+}16, LF21, LWA^{+}17, LS21, LZSW17, LK14, Ma21, NB17, Par19, PB15, PB16, RKHMRH21, Rog15, Skö18, SA15, SLLW20, TCF^{+}24, VOBB15, VSJR19, WDL21, WZFZ22, WDTC20, WLG23, Zha16, ZBDX18, ZXPS18, ZGL^{+}22, ZZ21b, BV21, CKLP17, CWL^{+}23, FMD19, FSPC16, KMD16, SE16, YP16, YLL20, Che15, EH15]. understandings [WCP^{+}19]. undocumented [VGN20]. unexpected [AJWB16]. unexplored [Poo17]. unfamiliar [OB19]. unfolds [MPT15]. Unified [ZTA20, MBVB17]. unifying [SSA22]. unimodal [GIEVVGS19]. unintended [MR19]. union [WCC^{+}17, Sav15b, Sav16b]. United [POK15, RN16, SH19]. units [Vec14]. Universe [MH18]. universities [AMDFHVC24, DB17a, LBM19a, PRRC18]. University [Bow19a, Clo18, Dut20, Far23, FJ19, Hai16, Hjø17a, Huv22a, Kea22, Kos20, Ma22b, Ma22a, Meh22, O’H18, San22, See17, Sha21b, Sko23, Bor14c, BM15a, CJ16b, DB17a, DBBJ16, GBZ^{+}23, HCS21, OMTSLC15, SZ20, SLeH21, PRRC16]. unlikely [BMG^{+}17]. unsupervised [AHJY18, DP23, Gia16, ST16c, Bar14]. Untangling [LO19, WCTS15]. untold [Lev20]. Updating [GNVQdMA16]. uploading [POK15]. uptake [BDH17]. urban [CW18, Lin15, SVB22, YZ16]. urban-rural [CW18]. Urdu [MSI^{+}19]. URL [OK16]. US-born [LS22]. US-born/US-dwelling [LS22]. US-dwelling [LS22]. USA [TGV^{+}16]. usability [TNL21]. usable [DC22]. usage [AZE^{+}19, CK14, CJ16a, GBL^{+}18, LMP21, LXW18, MRKT15, MAM19, SH19, WVvLZ23, Zit15]. Use [BMM20, KSB20, MHJ19, NB15, QH18, Zit15, AJWB16, AS24, BBM16, BM15a, Bor15c, BH16, CHH14, CO16, Che15, CM23, CDGN15, Cri15, Du14b, EAB17, EA18, FLJD23, Fon21, FvdWR^{+}22, GAB^{+}18, GTZŽ14, Gor22, HH19, HDT23, HH18, HHPI21, KT14a, Kra17, MBVB17, MAAL^{+}16, MAP15, NT17a, PHC17, RC16, Sin16, SS14b, St.17, SWLH17, SM18, WV15, Zha16, ZYS17, vLTW18]. used [HBS15, Ior14, LJ22, Ort22]. useful [JL18, MT16, VS14]. usefulness [JL18, JSW14, LB15, VAGR15]. User [ALC^{+}14, AMOM16, CS16b, DC15, HNX18, JHJ15, LP16, LCK16, LLD^{+}17, LCL16, PS20, SVB22, XJBK17, ZLF20, ZJHT23, AKAS14, AvS15, AvS16, BBM16, CBM17, ESWC18, EBD15, EA18, FSK^{+}16, GSG22, GAS^{+}20, HT17, HLB^{+}17, KTS^{+}15, KTPE14, LCLBY17, LZSW17, LWM^{+}19, lLTBH15, LYC^{+}22, MMHA23, MTHL15, MTK16, NGTKR15, PKK^{+}24, PPLJ20, PHC17, QH20, RCM^{+}17b, SY20, SL21a, SSdMCM20, SS14c, TBP^{+}15, TPR17, TG17, Vec14, VSJR19, WZL21, WHWZ21, WMS^{+}15, WLZ^{+}20, WLG23, XBCH18, XDDG18, XZ19, Zha14c, ZLCB15, ZY15, ZS20, ZXPS18, ZGBIL17, ZF19]. user-based [EBD15]. User-Centered [HNX18]. user-experience [AvS16, HLB^{+}17]. user-generated [WLG23]. User-level [LLD^{+}17]. Users [Ben18, Soe18, WCC^{+}17, Bea16, ESWC18, FJF^{+}21, HCC15, JHJ15, KBC17, KMD16, LWA^{+}17, MR19, PHZF16, STS14, SL21b, Tab14, TK17c, XBL^{+}21, QH18]. Uses [BI17, GRH^{+}17, KHZ^{+}22, SWL18, TK17c, WCC^{+}17]. Using [AST^{+}22, ALYS16, BCK17, DMMF17, DL22, FJT23, GB19a, GN17, HRB14, Hoo17, KK16, NLP16, PLPQB20, WWB21, YY16, ZY15, dSBG19, vdSvdB23, Aji17, AG15b, AGBN16, BK19, BDA17, BM15a, BTMS16, BSZ^{+}14, CPP^{+}16, Che16, Cho20, CY14, DP23, DBBJ16, EHdBvO17, GD18, GZBA22, GNVQdMA16, GISW15, GKM^{+}21, Hal14, Har14, HB16, KBD^{+}14, KGS19, KLH17, KDS^{+}17, KK19b, KKOS21, Ko15, LCNP17, Lew15, LdN17, LN17, LTD^{+}15, LWM^{+}19, LR14, LDS^{+}19, MC18, MG21, MDC^{+}16, MGK20, MŽA17, MBD15, ONZ23, OA15, PMD14, PDS22, PBN^{+}20, RJS19, RGJCTS16, SC17, SP17, SKZ^{+}14, SEZG22, TND15, TZH23, TDOF^{+}16, UM23, WLTY19, XBL^{+}22, YB18, ZLCB15, ZZ21a, ZL16, ZQC^{+}23]. UTAUT [KT14a]. Utility [Rog15, BSVZ17]. utilize [JJ14].

vaccination [NV17]. valence [JL18, PR16]. valence-consistent [PR16]. Valerie [Sal20]. Validation [WB21, KH17, PDŠB16, RDPF16, SKC18, TTH16]. validity [Bor15a, MBD15, Sch18]. Valuable [Lin14a]. valuation [LWLH18]. Value [FMNG18, RR20, AGM19, CZZY15, FC18, PPLJ20, PHJ23, RR21, SFL^{+}23, TBP^{+}15, TM16, TFR15, Meh22]. value-based [AGM19]. Values [FHW17, BKM15, HH18, MAM19, RCM^{+}17a]. values-oriented [BKM15]. vanished [LMJ21a, LMJ21b, She21]. variability [CS15a]. variant [Sch15]. Variation [Bow19a]. variations [Mal15, TKB^{+}17]. Varieties [SL24]. various [PHZF16]. vary [TM20]. varying [SJ17]. vector [NSK17]. Vee [Cle18]. Veni [Wu23]. venture [dSMdAGQ15]. verb [Eck18]. verification [AMM^{+}19, CV20, KS17, LMM^{+}22, PS19, SWKK16]. verifying [Roh21]. versa [NILCP15]. versus [CHH14, CK16a, DBH18, Hal14, HBG19, Kim19a, LCL16, MT16, RVGPS19, VS15, WS22, WvdHS16, YB18, ZDZ21, ZK14]. Vertesi [Rhe17]. Vertical [Huv22a]. very [MWD17, SAB^{+}18]. via [BCAL15, CCLL14, CML^{+}15, HWZ22, MGKA16, PS20, RDPF16, SGS21, SKA18]. VIC [Clo18]. vice [NILCP15]. vici [Wu23]. Victorian [CN14]. Video [FJ19, KK19b, CDL18, EHdBvO17, JSLC16, LCP15, LCLM18, MSM^{+}21, PGK24, SJK15]. videogames [Skö18]. Videos [KK19a, KK16, PYH16, SP22, TF21]. vidi [Wu23]. Vietnam [Luy16]. view [JGY^{+}19, LF21, VXN18]. viewed [TF21]. viewers [LCLM18]. viewing [LCLM18, SL21a]. Viewpoints [BRL20, SMRR16]. Views [SSA20]. village [YZQT20]. violence [Roh21, Wes15]. violent [WWB21]. VIP [MWD17]. virality [ZDZ21]. Virtual [Hak14, BCZ14, HB14, HLC15, LCNP17, RCM^{+}17a, RCM22, RY18, SW18, YZQT20, ZB19]. virtue [Gor24]. Virtuous [Gor24]. virus [ZCZ^{+}20]. virus-related [ZCZ^{+}20]. Visibility [Hal20, OHP16]. visible [Mor24, Pil14, RD16]. vision [RCM^{+}17a]. visitors [MKK17]. Visual [Alb15, Gor20b, KK19a, Sny14, CDL18, ELB17, HPLC19, HHK^{+}17, HBS15, LM15a, MNNK15, PS14a, SP22]. visualization [IQDS16, MŽA17, OHT^{+}16, WV18, YL17a, Aji17]. visualizations [MŽA17]. Visualizing [CC16, ZZZL17, KNY^{+}14]. visually [HS22, STS14, STS15]. Vocabulary [CCP15, Eng17a]. Voice [Tei16, SY20, SY22]. volatile [Sal20]. volatility [FDC22]. Vold [Han21]. Volume [Soe18, VRF14]. vote [GSG22]. voting [WMS^{+}15]. VQR [AD15]. VS [Uns18]. vulnerability [SV17]. vulnerable [PHJ23, VGN20].

W [Kno24, Sko23]. Waiting [BR15b]. wall [Cro15b]. Walter [Bud17, Han21, dP20]. wandering [Lin15]. Want [SSA20]. War [Lev20, Joh17, Luy16]. Ward [GNVQdMA16]. Warner [Owe21]. warning [CN14]. wars [Lev20, Bur20]. wary [CPL22]. was [DD16]. Washington [BNG14, TGC^{+}21]. waste [GRBCS15]. Watch [SSA20]. watching [KK19b]. water [WHvdZvdB17]. way [WLS23]. Wayback [ALYS16]. ways [PHC17]. WCs [LB16]. weak [ZB19]. Weatherall [Kos20]. Web [DMMF17, GVF^{+}24, MSdGS19, Web20, BCZ14, Bur14, CWdR17, Cho20, CCJ16, DM14, EBD15, HSL24, HWW15, JJC19, KTS^{+}15, KTPE14, KS14b, KT14b, Lew15, Li14b, MWMB18, Ort22, RSAS14, SAB^{+}18, SR15, SS14b, SM22, STF19, VC15, VN18, VS14, WSWH20, Whi14, WH15, YCK16, ZSX15, ZZSX16, ACS18, ALMP16, BOHK19, CS15a, CHLRC14, DGCC17, DOM^{+}14, HE22, LdMAdN16, MSdGS19, NPG^{+}20, OSvE16, OMTK17, RTS^{+}17, RL15a, Sch18, SJL19, ŠF17, SLH22, TTH16, VRF14, Agu18, OM15]. webometric [NV17]. webpage [WZK^{+}18]. website [ALMP16, MJ23, PGK24, WK16]. websites [AvS15, AvS16, ALYS16, CK16a, LCSL15, OMTSLC15, Wes15, XZ19]. week [TGV^{+}16]. weighted [Ano18-27, CDGN15, KBS18, LM15b, LML18, ZZL^{+}14, ZS16]. weighting [Ko15, SAW21]. weights [DLA14]. Weiser [Wei14b]. well [BM15a, BKM15, EAB17, EA18, ZJHT23]. well-being [BKM15, EAB17, EA18, ZJHT23]. Wellman [Lin14b]. wellness [ZT22]. We’re [JAG^{+}20, LS16]. Western [AT14]. wheel [SK14b]. whenever [JWWV18]. Where [LYCX14, Sto20, ZB19]. Whether [Fon21]. Which [BB15, KB17, YL18, OR15, TKS^{+}23]. While [SSA20]. Whither [HHPI21]. Whittaker [Jon17]. Who [MTHL15, XHC^{+}15]. Whole [RJK20]. Whole-session [RJK20]. Whose [SLH22]. Wide [MSdGS19]. wider [KT16b, KTA17]. Wiesbaden [Uns18]. wiki [AGBN16]. wikification [FdGPC18]. Wikipedia [ALAB19, AMDFHVC24, Bar14, JA16, Kon16, KT17b, KT17a, Luy16, Mah17, MAAL^{+}16, MOM^{+}15, OR15, OMM^{+}14, POK15, RBJ22, TLD17, Tsi16, WL20, XA14, ZCYN23]. Wikipedias [JW17]. wikis [ZFH^{+}19]. wilderness [HHPI21]. Wiley [Bru18]. will [HB15, OR15]. willingness [GBS15]. winners [LR14]. Winnipeg [AC18]. Wisdom [Fre18, BM14b, HYD^{+}20, McD21]. Wise [Dor19]. wiseness [RR15]. wish [PN18]. Withdrawal [Ior14]. Withdrawn [Ior14]. withholding [SLS^{+}19, ZWTX23]. within [CK21b, MA18a, RH21, SLS20]. without [RH21]. WL [WL14b]. WL-index [WL14b]. Wolfgang [Ama19, Sar16]. women [BBH15, JWWV18, PHJ23]. Woolgar [Rhe17]. word [CDGN15, DGR19, HBG19, JGY^{+}19, KDS^{+}17, LKY^{+}20, LZD16, LN17, DC14]. word-of-mouth [HBG19]. WordNet [LKY^{+}20, PMD14]. Words [ZZL20, BR15a, DLA14, Lin14a, OKB17, TR17]. Work [Cox14, Gor20b, Huv20, CS16a, CC17, CK21b, Cox23, DOM^{+}14, Du14b, ESD22, FMNG18, GBS15, IB16, JLB^{+}23, Kit19, LS16, NB23, ØJW^{+}21, OD16, Pra15b, SJ17, SW18, SFG^{+}23, STAH14, St.17, SWKK16, TFR15, VGN20, WLS23, Xia23, YCL20]. work-based [Xia23]. worker [Mat22]. workers [JT17]. Workflow [SM22, GHKB^{+}17]. workflows [OBD^{+}22, TWB^{+}18]. Working [SSA20, Sav18b]. Workplace [SSA20, Huv20]. Works [Soe18, Ior14, ZBBW18]. workspace [DOM^{+}14]. World [Lev20, dP20, BB15, CC16, CJ16a, EWX^{+}16, FJF^{+}21, HT14, LCM17, ØJW^{+}21, MSdGS19, Lan15, Ten16, Sal20]. WorldCat [WZ18]. WorldCat.org [WCC^{+}17]. Worlds [Hak14, PKA20, BK23, BAK^{+}17, Hen14, Hjø15a, OB19]. worldwide [BdMA14, LBM19a, PSB^{+}14]. worried [RBJ18a]. worth [MSPP21, TR17]. WOS [ZYZ^{+}16, LB16]. write [Har14, ØC19]. writing [CHH14, Cro15b, FMG16, LBW^{+}19b, STAH14, VVP^{+}21, Cle18]. writings [Jan15]. written [Ano14, KW14, LWM^{+}23, SMAB09]. wrong [SSdMCM20]. Wu [MS19]. Wylot [All19].

Xiao [Sko23]. xiaomi [PPLJ20]. Xie [Abb17]. XML [DPPV14, MNNK15, SJN14]. XV [Sto20]. xxxx [Adl23].

Yahoo [FV21, ZCZ^{+}20]. Yale [Kos20]. Year [MBBL14, GAB^{+}18, KH17, OK16, YL18]. years [BWGD21, NGTKR15]. Yellowstone [TWB^{+}18, Tho22]. Yiannis [Fra21]. yield [RD16]. York [Afz22, Gor20a, Ma22b, Ma22a, QHH20, Saw19, Sko23]. Young [Li21b, RBJ18a, WMX^{+}22, KT14a, MVL^{+}14, RBJ18b]. youth [YLL20]. YouTube [OS15, SJK15, TF21]. Yves [BI17, Zit15].

zen [DYB14]. zero [HBDL19]. zigzagged [TC17]. Zika [ZCZ^{+}20]. Zipf [iCV18]. zynergy [Pra14c]. zynergy-index [Pra14c].

References

Abalkina:2024:CPH

[Aba24] Anna Abalkina. Challenges posed by hijacked journals in Scopus. Journal of the Association for Information Science and Technology, 75(4):395–422, April 2024. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Abbas:2017:BRD

[Abb17] June Abbas. Book reviews: Discover Digital Libraries: Theory and Practice. Iris Xie and Krystyna M. Matusiak. Amsterdam: Elsevier, 2016. 388 pp. $79.95 (hardcover). (ISBN 978-0-12-417112-1). Journal of the Association for Information Science and Technology, 68(12):2841–2843, December 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Aletras:2017:ETR

[ABLS17] Nikolaos Aletras, Timothy Baldwin, Jey Han Lau, and Mark Stevenson. Evaluating topic representations for exploring document collections. Journal of the Association for Information Science and Technology, 68(1):154–167, January 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Allard:2018:IWT

[AC18] Danielle Allard and Nadia Caidi. Imagining Winnipeg: The translocal meaning making of Filipino migrants to Canada. Journal of the Association for Information Science and Technology, 69(10):1193–1204, October 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Abdo:2022:DTM

[ACBC22] Alexandre Hannud Abdo, Jean-Philippe Cointet, Pascale Bourret, and Alberto Cambrosio. Domain-topic models with chained dimensions: Charting an emergent domain of a major oncology conference. Journal of the Association for Information Science and Technology, 73(7):992–1011, July 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Acker:2021:EPS

[Ack21] Amelia Acker. Emulation practices for software preservation in libraries, archives, and museums. Journal of the Association for Information Science and Technology, 72(9): 1148–1160, September 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Arapakis:2017:FPP

[ACL17] Ioannis Arapakis, Berkant Barla Cambazoglu, and Mounia Lalmas. On the feasibility of predicting popular news at cold start. Journal of the Association for Information Science and Technology, 68(5):1149–1164, May 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Ahlgren:2018:ERB

[ACS18] Per Ahlgren, Cristian Colliander, and Peter Sjögårde. Exploring the relation between referencing practices and citation impact: a large-scale study based on Web of Science data. Journal of the Association for Information Science and Technology, 69(5):728–743, May 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Abramo:2015:VIS

[AD15] Giovanni Abramo and Ciriaco Andrea D’Angelo. The VQR, Italy’s second national research assessment: Methodological failures and ranking distortions. Journal of the Association for Information Science and Technology, 66(11):2202–2214, November 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Aspray:2023:DTR

[AD23] William Aspray and Philip Doty. Does technology really outpace policy, and does it matter? A primer for technical experts and others. Journal of the Association for Information Science and Technology, 74(8):885–904, August 2023. CO-DEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Abramo:2015:NAM

[ADC15] Giovanni Abramo, Ciriaco Andrea D’Angelo, and Flavia Di Costa. A new approach to measure the scientific strengths of territories. Journal of the Association for Information Science and Technology, 66(6):1167–1177, June 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Adhikari:2018:IIC

[ADD^{+}18] Abhijit Adhikari, Biswanath Dutta, Animesh Dutta, Deepjyoti Mondal, and Shivang Singh. An intrinsic information content-based semantic similarity measure considering the disjoint common subsumers of concepts of an ontology. Journal of the Association for Information Science and Technology, 69(8):1023–1034, August 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Arif:2015:UTC

[ADL15] Abu Shamim Mohammad Arif, Jia Tina Du, and Ivan Lee. Understanding tourists’ collaborative information retrieval behavior to inform design. Journal of the Association for Information Science and Technology, 66(11):2285–2303, November 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Adler:2023:BRK

[Adl23] Melissa Adler. Book review: Knowledge justice: Disrupting library and information studies through critical race theory. By Sofia Y. Leung, Jorge R. López-McKnight, Cambridge: The MIT Press. 2021. pp. 358. $35.00 (xxxx). ISBN: 978-0-262-04350-2. Journal of the Association for Information Science and Technology, 74(5):594–598, May 2023. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Ahmad:2020:HIC

[ADM^{+}20] Atif Ahmad, Kevin C. Desouza, Sean B. Maynard, Humza Naseer, and Richard L. Baskerville. How integration of cyber security management and incident response enables organizational learning. Journal of the Association for Information Science and Technology, 71(8):939–953, August 2020. CO-DEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Abella:2019:POD

[AdUCDPH19] Alberto Abella, Marta Ortiz de Urbina-Criado, and Carmen De-Pablos-Heredero. The process of open data publication and reuse. Journal of the Association for Information Science and Technology, 70(3):296–300, March 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Almeida:2017:ORE

[AF17] Mauricio Barcellos Almeida and Fernanda Farinelli. Ontologies for the representation of electronic medical records: the obstetric and neonatal ontology. Journal of the Association for Information Science and Technology, 68(11):2529–2542, November 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Almeida:2020:TDC

[AFB20] Mauricio B. Almeida, Eduardo R. Felipe, and Renata Barcelos. Toward a document-centered ontological theory for information architecture in corporations. Journal of the Association for Information Science and Technology, 71(11):

1308–1326, November 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Aphinyanaphongs:2014:CEC

[AFL^{+}14] Yindalon Aphinyanaphongs, Lawrence D. Fu, Zhiguo Li, Eric R. Peskin, Efstratios Efstathiadis, Constantin F. Aliferis, and Alexander Statnikov. A comprehensive empirical comparison of modern supervised classification and feature selection methods for text categorization. Journal of the Association for Information Science and Technology, 65(10): 1964–1987, October 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Alkhodair:2018:IIT

[AFRH18] Sarah A. Alkhodair, Benjamin C. M. Fung, Osmud Rahman, and Patrick C. K. Hung. Improving interpretations of topic modeling in microblogs. Journal of the Association for Information Science and Technology, 69(4):528–540, April 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Afzal:2022:BRH

[Afz22] Waseem Afzal. Book review: The Hype Machine. Aral, Sinan. New York, NY: Penguin Random House, 2020. 416 pp. $28.00 (hardcover). (ISBN 978-0-525-57451-4). Journal of the Association for Information Science and Technology, 73(5):752–754, May 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Absar:2015:DFE

[AG15a] Rafa Absar and Catherine Guastavino. The design and formative evaluation of nonspeech auditory feedback for an information system. Journal of the Association for Information Science and Technology, 66(8):1696–1708, August 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Aker:2015:GDM

[AG15b] Ahmet Aker and Robert Gaizauskas. Generating descriptive multi-document summaries of geo-located entities using entity type models. Journal of the Association for Information Science and Technology, 66(4):721–738, April 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Allen:2019:IBI

[AGBK19] David Allen, Lisa M. Given, Gary Burnett, and Stan Karanasios. Information behavior and information practices: a special issue for research on people’s engagement with technology. Journal of the Association for Information Science and Technology, 70(12):1299–1301, December 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Arazy:2016:MSK

[AGBN16] Ofer Arazy, Ian Gellatly, Esther Brainin, and Oded Nov. Motivation to share knowledge using wiki technology and the moderating effect of role perceptions. Journal of the Association for Information Science and Technology, 67(10): 2362–2378, October 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Abad-Garcia:2018:EOB

[AGGTGL18] Maria-Francisca Abad-García, Aurora González-Teruel, and Javier González-Llinares. Effectiveness of OpenAIRE, BASE, Recolecta, and Google Scholar at finding Spanish articles in repositories. Journal of the Association for Information Science and Technology, 69(4):619–622, April 2018. CO-DEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Athukorala:2016:ESD

[AGJ^{+}16] Kumaripaba Athukorala, Dorota Głowacka, Giulio Jacucci, Antti Oulasvirta, and Jilles Vreeken. Is exploratory search different? A comparison of information search behavior for exploratory and lookup tasks. Journal of the Association for Information Science and Technology, 67(11):2635–2651, November 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Arenas:2019:IBM

[AGM19] Alvaro E. Arenas, Jie Mein Goh, and Brian Matthews. Identifying the business model dimensions of data sharing: a value-based approach. Journal of the Association for Information Science and Technology, 70(10):1047–1059, October 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Aledo:2018:CBJ

[AGMR18] Juan A. Aledo, Jose A. Gámez, David Molina, and Alejandro Rosete. Consensus-based journal rankings: a complementary tool for bibliometric evaluation. Journal of the Association for Information Science and Technology, 69(7):936–948, July 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Aguillo:2018:BRW

[Agu18] Isidro F. Aguillo. Book review: Web Indicators for Research Evaluation: A Practical Guide. Synthesis Lectures on Information Concepts, Retrieval, and Services. Michael Thelwall. San Rafael, CA: Morgan & Claypool Publishers, 2016. 170 pp. $60.00 (Paperback). (ISBN-13: 978-1-62705-917-6). Journal of the Association for Information Science and Technology, 69(3):498–499, March 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

An:2017:GNM

[AH17] Xiangdong An and Jimmy Xiangji Huang. geNov: a new metric for measuring novelty and relevancy in biomedical information retrieval. Journal of the Association for Information Science and Technology, 68(11):2620–2635, November 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Aldebei:2018:SSU

[AHJY18] Khaled Aldebei, Xiangjian He, Wenjing Jia, and Weichang Yeh. SUDMAD: Sequential and unsupervised decomposition of a multi-author document based on a hidden Markov model. Journal of the Association for Information Science and Technology, 69(2):201–214, February 2018. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Ajiferuke:2017:BRS

[Aji17] Isola Ajiferuke. Book review: Statistics for library and information services: a primer for using open source R software for accessibility and visualization. Alon Friedman. Lanham, MD: Rowman & Littlefield Publishers, 2015. 376 pp. $95.00 (hardback). (ISBN 978-1-4422-4992-9). Journal of the Association for Information Science and Technology, 68(6):1594– 1595, June 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Abraham:2016:EUC

[AJWB16] Chon Abraham, Iris Junglas, Richard T. Watson, and Marie-Claude Boudreau. Explaining the unexpected and continued use of an information system with the help of evolved evolutionary mechanisms. Journal of the Association for Information Science and Technology, 67(1):212–231, January 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Arakawa:2014:ATS

[AKAS14] Yui Arakawa, Akihiro Kameda, Akiko Aizawa, and Takafumi Suzuki. Adding Twitter-specific features to stylistic features for classifying tweets by user type and number of retweets. Journal of the Association for Information Science and Technology, 65(7):1416–1423, July 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Ayala:2018:MRM

[AKC^{+}18] Brenda Reyes Ayala, Ryan Knudson, Jiangping Chen, Gaohui Cao, and Xinyue Wang. Metadata records machine translation combining multi-engine outputs with limited parallel data. Journal of the Association for Information Science and Technology, 69(1):47–59, January 2018. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

An:2017:ECH

[AKK^{+}17] Juyoung An, Namhee Kim, Min-Yen Kan, Muthu Kumar Chandrasekaran, and Min Song. Exploring characteristics of highly cited authors according to citation location and content. Journal of the Association for Information Science and Technology, 68(8):1975–1988, August 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Arazy:2019:NBN

[ALAB19] Ofer Arazy, Hila Lifshitz-Assaf, and Adam Balila. Neither a bazaar nor a cathedral: The interplay between structure and agency in Wikipedia’s role system. Journal of the Association for Information Science and Technology, 70(1):3–15, January 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Albertson:2015:AIS

[Alb15] Dan Albertson. Advances in information science: Visual information seeking. Journal of the Association for Information Science and Technology, 66(6):1091–1105, June 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Arapakis:2014:UEO

[ALC^{+}14] Ioannis Arapakis, Mounia Lalmas, B. Barla Cambazoglu, Mari-Carmen Marcos, and Joemon M. Jose. User engagement in online News: Under the scope of sentiment, interest, affect, and gaze. Journal of the Association for Information Science and Technology, 65(10):1988–2005, October 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Arapakis:2014:AEN

[ALCD14] Ioannis Arapakis, Mounia Lalmas, Hakan Ceylan, and Pinar Donmez. Automatically embedding newsworthy links to articles: From implementation to evaluation. Journal of the Association for Information Science and Technology, 65(1): 129–145, January 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Allemang:2019:BRL

[All19] Dean Allemang. Book review: Linked Data: Storing, Querying, and Reasoning. Sakr, Sherif, Wylot, Marcin, Mutharaju, Raghava, Le Phuoc, Danh, and Fundulaki, Irini. Cham, Switzerland: Springer International Publishing, 2018. 223 pp. $129.00 (hardcover). (ISBN: 978-3-319-73514-6). Journal of the Association for Information Science and Technology, 70(8):905–907, August 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Arbelaitz:2016:WMN

[ALMP16] Olatz Arbelaitz, Aizea Lojo, Javier Muguerza, and Iñigo Perona. Web mining for navigation problem detection and diagnosis in Discapnet: a website aimed at disabled people. Journal of the Association for Information Science and Technology, 67(8):1916–1927, August 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Arora:2016:UWM

[ALYS16] Sanjay K. Arora, Yin Li, Jan Youtie, and Philip Shapira. Using the Wayback Machine to mine websites in the social sciences: a methodological resource. Journal of the Association for Information Science and Technology, 67(8):1904– 1915, August 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Aman:2019:BRT

[Ama19] Mohammed M. Aman. Book review: Transitioning Towards a Knowledge Society: Qatar as a Case Study. Gremm, Julia, Barth, Julia, Fietkiewicz, Kaja J., and Stock, Wolfgang G.. Cham, Switzerland: Springer International Publishing, 2018. 244 pp. $79.99 (ebook). (ISBN 978-3-319-71195-9). Journal of the Association for Information Science and Technology, 70(7):769–771, July 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Aman:2022:IMS

[Ama22] Valeria Aman. Internationally mobile scientists as knowledge transmitters: a lexical-based approach to detect knowledge transfer. Journal of the Association for Information Science and Technology, 73(10):1418–1431, October 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Arroyo-Machado:2024:AMC

[AMDFHVC24] Wenceslao Arroyo-Machado, Adrián A. Díaz-Faes, Enrique Herrera-Viedma, and Rodrigo Costas. From academic to media capital: To what extent does the scientific reputation of universities translate into Wikipedia attention? Journal of the Association for Information Science and Technology, 75(4):423–437, April 2024. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Adamovic:2019:ALI

[AMM^{+}19] Sasa Adamovic, Vladislav Miskovic, Milan Milosavljevic, Marko Sarac, and Mladen Veinovic. Automated languageindependent authorship verification (for Indo-European languages). Journal of the Association for Information Science and Technology, 70(8):858–871, August 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Arbelaitz:2016:UMS

[AMOM16] Olatz Arbelaitz, José María Martínez-Otzeta, and Javier Muguerza. User modeling in a social network for cognitively disabled people. Journal of the Association for Information Science and Technology, 67(2):305–317, February 2016. CO-DEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Anonymous:2014:EBS

[Ano14] Anonymous. Erratum: “Backstage” solidarity in Spanishand English-written medical research papers: Publication context and the acknowledgment paratext. Journal of the American Society for Information Science and Technology 60(2) 2009, 307–317. DOI: 10.1002/asi.20981. Journal of the Association for Information Science and Technology, 65 (10):2164, October 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic). See [SMAB09].

Anonymous:2015:CIa

[Ano15a] Anonymous. Cover image. Journal of the Association for Information Science and Technology, 66(8):C1, August 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Anonymous:2015:CIb

[Ano15b] Anonymous. Cover image. Journal of the Association for Information Science and Technology, 66(9):C1, September 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Anonymous:2015:CIc

[Ano15c] Anonymous. Cover image. Journal of the Association for Information Science and Technology, 66(10):C1, October 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Anonymous:2015:CId

[Ano15d] Anonymous. Cover image. Journal of the Association for Information Science and Technology, 66(11):C1, November 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Anonymous:2015:CIe

[Ano15e] Anonymous. Cover image. Journal of the Association for Information Science and Technology, 66(12):C1, December 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Anonymous:2016:CIa

[Ano16a] Anonymous. Cover image. Journal of the Association for Information Science and Technology, 67(1):C1, January 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Anonymous:2016:CIb

[Ano16b] Anonymous. Cover image. Journal of the Association for Information Science and Technology, 67(2):C1, February 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Anonymous:2016:CIc

[Ano16c] Anonymous. Cover image. Journal of the Association for Information Science and Technology, 67(3):C1, March 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Anonymous:2016:CId

[Ano16d] Anonymous. Cover image. Journal of the Association for Information Science and Technology, 67(4):C1, April 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Anonymous:2016:CIe

[Ano16e] Anonymous. Cover image. Journal of the Association for Information Science and Technology, 67(5):C1, May 2016. CO-DEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Anonymous:2016:CIf

[Ano16f] Anonymous. Cover image. Journal of the Association for Information Science and Technology, 67(6):C1, June 2016. CO-DEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Anonymous:2016:CIg

[Ano16g] Anonymous. Cover image. Journal of the Association for Information Science and Technology, 67(7):C1, July 2016. CO-DEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Anonymous:2016:CIh

[Ano16h] Anonymous. Cover image. Journal of the Association for Information Science and Technology, 67(8):C1, August 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Anonymous:2016:CIi

[Ano16i] Anonymous. Cover image. Journal of the Association for Information Science and Technology, 67(9):c1, September 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Anonymous:2016:CIj

[Ano16j] Anonymous. Cover image. Journal of the Association for Information Science and Technology, 67(10):C1, October 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Anonymous:2016:CIk

[Ano16k] Anonymous. Cover image. Journal of the Association for Information Science and Technology, 67(11):C1, November 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Anonymous:2016:CIl

[Ano16l] Anonymous. Cover image. Journal of the Association for Information Science and Technology, 67(12):C1, December 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Anonymous:2016:EO

[Ano16m] Anonymous. Editorial: Onward. Journal of the Association for Information Science and Technology, 67(2):247–248, February 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Anonymous:2016:IITa

[Ano16n] Anonymous. Issue information — table of contents. Journal of the Association for Information Science and Technology, 67(1):1–2, January 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Anonymous:2016:IITb

[Ano16o] Anonymous. Issue information — table of contents. Journal of the Association for Information Science and Technology, 67(2):245–246, February 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Anonymous:2016:IITc

[Ano16p] Anonymous. Issue information — table of contents. Journal of the Association for Information Science and Technology, 67(3):491–492, March 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Anonymous:2016:IITd

[Ano16q] Anonymous. Issue information — table of contents. Journal of the Association for Information Science and Technology, 67(4):755–756, April 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Anonymous:2016:IITe

[Ano16r] Anonymous. Issue information — table of contents. Journal of the Association for Information Science and Technology, 67(5), May 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Anonymous:2016:IITf

[Ano16s] Anonymous. Issue information — table of contents. Journal of the Association for Information Science and Technology, 67(6), June 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Anonymous:2016:IITg

[Ano16t] Anonymous. Issue information — table of contents. Journal of the Association for Information Science and Technology, 67(7):1521–1522, July 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Anonymous:2016:IITh

[Ano16u] Anonymous. Issue information – table of contents. Journal of the Association for Information Science and Technology, 67(8):1791–1792, August 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Anonymous:2016:IITi

[Ano16v] Anonymous. Issue information – table of contents. Journal of the Association for Information Science and Technology, 67(9):2045–2046, September 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Anonymous:2016:IITj

[Ano16w] Anonymous. Issue information – table of contents. Journal of the Association for Information Science and Technology, 67(10):2307–2308, October 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Anonymous:2016:IITk

[Ano16x] Anonymous. Issue information – table of contents. Journal of the Association for Information Science and Technology, 67(11):2573–2574, November 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Anonymous:2016:IITl

[Ano16y] Anonymous. Issue information – table of contents. Journal of the Association for Information Science and Technology, 67 (12):2829–2830, December 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Anonymous:2017:CIa

[Ano17a] Anonymous. Cover image. Journal of the Association for Information Science and Technology, 68(1):C1, January 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Anonymous:2017:CIb

[Ano17b] Anonymous. Cover image. Journal of the Association for Information Science and Technology, 68(2):C1, February 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Anonymous:2017:CIc

[Ano17c] Anonymous. Cover image. Journal of the Association for Information Science and Technology, 68(3):C1, March 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Anonymous:2017:CId

[Ano17d] Anonymous. Cover image. Journal of the Association for Information Science and Technology, 68(4):C1, April 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Anonymous:2017:CIe

[Ano17e] Anonymous. Cover image. Journal of the Association for Information Science and Technology, 68(5):C1, May 2017. CO-DEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Anonymous:2017:CIf

[Ano17f] Anonymous. Cover image. Journal of the Association for Information Science and Technology, 68(6):C1, June 2017. CO-DEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Anonymous:2017:CIg

[Ano17g] Anonymous. Cover image. Journal of the Association for Information Science and Technology, 68(7):C1, July 2017. CO-DEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Anonymous:2017:CIh

[Ano17h] Anonymous. Cover image. Journal of the Association for Information Science and Technology, 68(8):C1, August 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Anonymous:2017:CIi

[Ano17i] Anonymous. Cover image. Journal of the Association for Information Science and Technology, 68(9):C1, September 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Anonymous:2017:CIj

[Ano17j] Anonymous. Cover image. Journal of the Association for Information Science and Technology, 68(10):C1, October 2017.

CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Anonymous:2017:CIk

[Ano17k] Anonymous. Cover image. Journal of the Association for Information Science and Technology, 68(11):C1, November 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Anonymous:2017:CIl

[Ano17l] Anonymous. Cover image. Journal of the Association for Information Science and Technology, 68(12):C1, December 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Anonymous:2017:ECB

[Ano17m] Anonymous. Erratum: Citation-based plagiarism detection: Practicability on a large-scale scientific corpus. Journal of the Association for Information Science and Technology, 68 (1):270, January 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Anonymous:2017:ESM

[Ano17n] Anonymous. Erratum: Social-media-based public policy informatics: Sentiment and network analyses of U.S. Immigration and Border Security. Journal of the Association for Information Science and Technology, 68(12):2847, December 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic). See [CZ16].

Anonymous:2017:IITa

[Ano17o] Anonymous. Issue information — table of contents. Journal of the Association for Information Science and Technology, 68(1):1–2, January 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Anonymous:2017:IITb

[Ano17p] Anonymous. Issue information — table of contents. Journal of the Association for Information Science and Technology, 68(2):271–272, February 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Anonymous:2017:IITc

[Ano17q] Anonymous. Issue information — table of contents. Journal of the Association for Information Science and Technology, 68(3):541–542, March 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Anonymous:2017:IITd

[Ano17r] Anonymous. Issue information — table of contents. Journal of the Association for Information Science and Technology, 68(4):809–810, April 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Anonymous:2017:IITe

[Ano17s] Anonymous. Issue information — table of contents. Journal of the Association for Information Science and Technology, 68(5):1071–1072, May 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Anonymous:2017:IITf

[Ano17t] Anonymous. Issue information — table of contents. Journal of the Association for Information Science and Technology, 68(6):1339–1340, June 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Anonymous:2017:IITg

[Ano17u] Anonymous. Issue information — table of contents. Journal of the Association for Information Science and Technology, 68(7):1597–1598, July 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Anonymous:2017:IITh

[Ano17v] Anonymous. Issue information — table of contents. Journal of the Association for Information Science and Technology, 68(8):1805–1806, August 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Anonymous:2017:IITi

[Ano17w] Anonymous. Issue information — table of contents. Journal of the Association for Information Science and Technology, 68(9):2035–2036, September 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Anonymous:2017:IITj

[Ano17x] Anonymous. Issue information — table of contents. Journal of the Association for Information Science and Technology, 68(10):2297–2298, October 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Anonymous:2017:IITk

[Ano17y] Anonymous. Issue information — table of contents. Journal of the Association for Information Science and Technology, 68(11):2523–2524, November 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Anonymous:2017:IITl

[Ano17z] Anonymous. Issue information — table of contents. Journal of the Association for Information Science and Technology, 68(12):2691–2692, December 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Anonymous:2018:CIa

[Ano18a] Anonymous. Cover image. Journal of the Association for Information Science and Technology, 69(1):C1, January 2018. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Anonymous:2018:CIb

[Ano18b] Anonymous. Cover image. Journal of the Association for Information Science and Technology, 69(2):C1, February 2018. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Anonymous:2018:CIc

[Ano18c] Anonymous. Cover image. Journal of the Association for Information Science and Technology, 69(3):C1, March 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2018:CId

[Ano18d] Anonymous. Cover image. Journal of the Association for Information Science and Technology, 69(4):C1, April 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2018:CIe

[Ano18e] Anonymous. Cover image. Journal of the Association for Information Science and Technology, 69(5):C1, May 2018. CO-DEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2018:CIf

[Ano18f] Anonymous. Cover image. Journal of the Association for Information Science and Technology, 69(6):C1, June 2018. CO-DEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2018:CIg

[Ano18g] Anonymous. Cover image. Journal of the Association for Information Science and Technology, 69(7):C1, July 2018. CO-DEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2018:CIh

[Ano18h] Anonymous. Cover image. Journal of the Association for Information Science and Technology, 69(8):C1, August 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2018:CIi

[Ano18i] Anonymous. Cover image. Journal of the Association for Information Science and Technology, 69(9):C1, September 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2018:CIj

[Ano18j] Anonymous. Cover image. Journal of the Association for Information Science and Technology, 69(10):C1, October 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2018:CIk

[Ano18k] Anonymous. Cover image. Journal of the Association for Information Science and Technology, 69(11):C1, November 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2018:CIl

[Ano18l] Anonymous. Cover image. Journal of the Association for Information Science and Technology, 69(12):C1, December 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2018:E

[Ano18m] Anonymous. Erratum. Journal of the Association for Information Science and Technology, 69(9):1177, September 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2018:IIa

[Ano18n] Anonymous. Issue information. Journal of the Association for Information Science and Technology, 69(1):1–2, January 2018. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Anonymous:2018:IIb

[Ano18o] Anonymous. Issue information. Journal of the Association for Information Science and Technology, 69(2):185–186, February 2018. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Anonymous:2018:IIc

[Ano18p] Anonymous. Issue information. Journal of the Association for Information Science and Technology, 69(3):345–346, March 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2018:IId

[Ano18q] Anonymous. Issue information. Journal of the Association for Information Science and Technology, 69(4):501–502, April 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2018:IIe

[Ano18r] Anonymous. Issue information. Journal of the Association for Information Science and Technology, 69(5):631–632, May 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2018:IIf

[Ano18s] Anonymous. Issue information. Journal of the Association for Information Science and Technology, 69(6):761–762, June 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2018:IIg

[Ano18t] Anonymous. Issue information. Journal of the Association for Information Science and Technology, 69(7):867–868, July 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2018:IIh

[Ano18u] Anonymous. Issue information. Journal of the Association for Information Science and Technology, 69(8):957–958, August 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2018:IIi

[Ano18v] Anonymous. Issue information. Journal of the Association for Information Science and Technology, 69(9):1071–1072, September 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2018:IIj

[Ano18w] Anonymous. Issue information. Journal of the Association for Information Science and Technology, 69(10):1179–1180, October 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2018:IIk

[Ano18x] Anonymous. Issue information. Journal of the Association for Information Science and Technology, 69(11):1289–1290, November 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2018:IIl

[Ano18y] Anonymous. Issue information. Journal of the Association for Information Science and Technology, 69(12):1403–1404, December 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2018:JSI

[Ano18z] Anonymous. JASIST special issue on biomedical information retrieval. Journal of the Association for Information Science and Technology, 69(3):500, March 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2018:TEA

[Ano18-27] Anonymous. Toward effective automated weighted subject indexing: a comparison of different approaches in different environments. Journal of the Association for Information Science and Technology, 69(7):956, July 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2019:CIa

[Ano19a] Anonymous. Cover image. Journal of the Association for Information Science and Technology, 70(1):C1, January 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2019:CIb

[Ano19b] Anonymous. Cover image. Journal of the Association for Information Science and Technology, 70(2):C1, February 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2019:CIc

[Ano19c] Anonymous. Cover image. Journal of the Association for Information Science and Technology, 70(3):C1, March 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2019:CId

[Ano19d] Anonymous. Cover image. Journal of the Association for Information Science and Technology, 70(4):C1, April 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2019:CIe

[Ano19e] Anonymous. Cover image. Journal of the Association for Information Science and Technology, 70(5):C1, May 2019. CO-DEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2019:CIf

[Ano19f] Anonymous. Cover image. Journal of the Association for Information Science and Technology, 70(6):C1, June 2019. CO-DEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2019:CIg

[Ano19g] Anonymous. Cover image. Journal of the Association for Information Science and Technology, 70(7):C1, July 2019. CO-DEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2019:CIh

[Ano19h] Anonymous. Cover image. Journal of the Association for Information Science and Technology, 70(8):C1, August 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2019:CIi

[Ano19i] Anonymous. Cover image. Journal of the Association for Information Science and Technology, 70(9):C1, September 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2019:CIj

[Ano19j] Anonymous. Cover image. Journal of the Association for Information Science and Technology, 70(10):C1, October 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2019:CIk

[Ano19k] Anonymous. Cover image. Journal of the Association for Information Science and Technology, 70(11):C1, November 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2019:CIl

[Ano19l] Anonymous. Cover image. Journal of the Association for Information Science and Technology, 70(12):C1, December 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2019:IIa

[Ano19m] Anonymous. Issue information. Journal of the Association for Information Science and Technology, 70(1):1–2, January 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2019:IIb

[Ano19n] Anonymous. Issue information. Journal of the Association for Information Science and Technology, 70(2):99–100, February 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2019:IIc

[Ano19o] Anonymous. Issue information. Journal of the Association for Information Science and Technology, 70(3):205–206, March 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2019:IId

[Ano19p] Anonymous. Issue information. Journal of the Association for Information Science and Technology, 70(4):305–306, April 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2019:IIe

[Ano19q] Anonymous. Issue information. Journal of the Association for Information Science and Technology, 70(5):417–418, May 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2019:IIf

[Ano19r] Anonymous. Issue information. Journal of the Association for Information Science and Technology, 70(6):533–534, June 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2019:IIg

[Ano19s] Anonymous. Issue information. Journal of the Association for Information Science and Technology, 70(7):641–642, July 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2019:IIh

[Ano19t] Anonymous. Issue information. Journal of the Association for Information Science and Technology, 70(8):773–774, August 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2019:IIi

[Ano19u] Anonymous. Issue information. Journal of the Association for Information Science and Technology, 70(9):909–910, September 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2019:IIj

[Ano19v] Anonymous. Issue information. Journal of the Association for Information Science and Technology, 70(10):1045–1046, October 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2019:IIk

[Ano19w] Anonymous. Issue information. Journal of the Association for Information Science and Technology, 70(11):1171–1172, November 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2019:IIl

[Ano19x] Anonymous. Issue information. Journal of the Association for Information Science and Technology, 70(12):1297–1298, December 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2020:CIa

[Ano20a] Anonymous. Cover image. Journal of the Association for Information Science and Technology, 71(1):C1, January 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2020:CIb

[Ano20b] Anonymous. Cover image. Journal of the Association for Information Science and Technology, 71(2):C1, February 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2020:CIc

[Ano20c] Anonymous. Cover image. Journal of the Association for Information Science and Technology, 71(3):C1, March 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2020:CId

[Ano20d] Anonymous. Cover image. Journal of the Association for Information Science and Technology, 71(4):C1, April 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2020:CIe

[Ano20e] Anonymous. Cover image. Journal of the Association for Information Science and Technology, 71(5):C1, May 2020. CO-DEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2020:CIf

[Ano20f] Anonymous. Cover image. Journal of the Association for Information Science and Technology, 71(6):C1, June 2020. CO-DEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2020:CIg

[Ano20g] Anonymous. Cover image. Journal of the Association for Information Science and Technology, 71(7):C1, July 2020. CO-DEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2020:CIh

[Ano20h] Anonymous. Cover image. Journal of the Association for Information Science and Technology, 71(8):C1, August 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2020:CIi

[Ano20i] Anonymous. Cover image. Journal of the Association for Information Science and Technology, 71(9):C1, September 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2020:CIj

[Ano20j] Anonymous. Cover image. Journal of the Association for Information Science and Technology, 71(10):C1, October 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2020:CIk

[Ano20k] Anonymous. Cover image. Journal of the Association for Information Science and Technology, 71(11):C1, November 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2020:CIl

[Ano20l] Anonymous. Cover image. Journal of the Association for Information Science and Technology, 71(12):C1, December 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2020:IIa

[Ano20m] Anonymous. Issue information. Journal of the Association for Information Science and Technology, 71(1):1–2, January 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2020:IIb

[Ano20n] Anonymous. Issue information. Journal of the Association for Information Science and Technology, 71(2):125–126, February 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2020:IIc

[Ano20o] Anonymous. Issue information. Journal of the Association for Information Science and Technology, 71(3):249–250, March 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2020:IId

[Ano20p] Anonymous. Issue information. Journal of the Association for Information Science and Technology, 71(4):377–378, April 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2020:IIe

[Ano20q] Anonymous. Issue information. Journal of the Association for Information Science and Technology, 71(5):501–502, May 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2020:IIf

[Ano20r] Anonymous. Issue information. Journal of the Association for Information Science and Technology, 71(6):617–618, June 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2020:IIg

[Ano20s] Anonymous. Issue information. Journal of the Association for Information Science and Technology, 71(7):727–728, July 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2020:IIh

[Ano20t] Anonymous. Issue information. Journal of the Association for Information Science and Technology, 71(8):869–870, August 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2020:IIi

[Ano20u] Anonymous. Issue information. Journal of the Association for Information Science and Technology, 71(9):995–996, September 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2020:IIj

[Ano20v] Anonymous. Issue information. Journal of the Association for Information Science and Technology, 71(10):1143–1144, October 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2020:IIk

[Ano20w] Anonymous. Issue information. Journal of the Association for Information Science and Technology, 71(11):1279–1280, November 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2020:IIl

[Ano20x] Anonymous. Issue information. Journal of the Association for Information Science and Technology, 71(12):1417–1418, December 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2021:CIa

[Ano21a] Anonymous. Cover image. Journal of the Association for Information Science and Technology, 72(3):C1, March 2021.

CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2021:CIb

[Ano21b] Anonymous. Cover image. Journal of the Association for Information Science and Technology, 72(4):C1, April 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2021:CIc

[Ano21c] Anonymous. Cover image. Journal of the Association for Information Science and Technology, 72(5):C1, May 2021. CO-DEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2021:CId

[Ano21d] Anonymous. Cover image. Journal of the Association for Information Science and Technology, 72(6):C1, June 2021. CO-DEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2021:CIe

[Ano21e] Anonymous. Cover image. Journal of the Association for Information Science and Technology, 72(7):C1, July 2021. CO-DEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2021:CIf

[Ano21f] Anonymous. Cover image. Journal of the Association for Information Science and Technology, 72(8):949–950, August 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2021:CIg

[Ano21g] Anonymous. Cover image. Journal of the Association for Information Science and Technology, 72(9):C1, September 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2021:CIh

[Ano21h] Anonymous. Cover image. Journal of the Association for Information Science and Technology, 72(10):C1, October 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2021:CIi

[Ano21i] Anonymous. Cover image. Journal of the Association for Information Science and Technology, 72(11):C1, November 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2021:CIj

[Ano21j] Anonymous. Cover image. Journal of the Association for Information Science and Technology, 72(12):C1, December 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2021:E

[Ano21k] Anonymous. Erratum. Journal of the Association for Information Science and Technology, 72(10):1319, October 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2021:IIa

[Ano21l] Anonymous. Issue information. Journal of the Association for Information Science and Technology, 72(1):1–2, January 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2021:IIb

[Ano21m] Anonymous. Issue information. Journal of the Association for Information Science and Technology, 72(2):139–140, February 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2021:IIc

[Ano21n] Anonymous. Issue information. Journal of the Association for Information Science and Technology, 72(3):267–268, March 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2021:IId

[Ano21o] Anonymous. Issue information. Journal of the Association for Information Science and Technology, 72(4):385–386, April 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2021:IIe

[Ano21p] Anonymous. Issue information. Journal of the Association for Information Science and Technology, 72(5):521–522, May 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2021:IIf

[Ano21q] Anonymous. Issue information. Journal of the Association for Information Science and Technology, 72(6):653–654, June 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2021:IIg

[Ano21r] Anonymous. Issue information. Journal of the Association for Information Science and Technology, 72(7):797–798, July 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2021:IIh

[Ano21s] Anonymous. Issue information. Journal of the Association for Information Science and Technology, 72(8):949–950, August 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2021:IIi

[Ano21t] Anonymous. Issue information. Journal of the Association for Information Science and Technology, 72(9):1097–1098, September 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2021:IIj

[Ano21u] Anonymous. Issue information. Journal of the Association for Information Science and Technology, 72(10):1215–1216, October 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2021:IIk

[Ano21v] Anonymous. Issue information. Journal of the Association for Information Science and Technology, 72(11):1321–1322, November 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2021:IIl

[Ano21w] Anonymous. Issue information. Journal of the Association for Information Science and Technology, 72(12):1459–1460, December 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2022:CIa

[Ano22a] Anonymous. Cover image. Journal of the Association for Information Science and Technology, 73(1):C1, January 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2022:CIb

[Ano22b] Anonymous. Cover image. Journal of the Association for Information Science and Technology, 73(2):C1, February 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2022:CIc

[Ano22c] Anonymous. Cover image. Journal of the Association for Information Science and Technology, 73(3):C1, March 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2022:CId

[Ano22d] Anonymous. Cover image. Journal of the Association for Information Science and Technology, 73(4):C1, April 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2022:CIe

[Ano22e] Anonymous. Cover image. Journal of the Association for Information Science and Technology, 73(5):C1, May 2022. CO-DEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2022:CIf

[Ano22f] Anonymous. Cover image. Journal of the Association for Information Science and Technology, 73(6):C1, June 2022. CO-DEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2022:CIg

[Ano22g] Anonymous. Cover image. Journal of the Association for Information Science and Technology, 73(7):C1, July 2022. CO-DEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2022:CIh

[Ano22h] Anonymous. Cover image. Journal of the Association for Information Science and Technology, 73(8):C1, August 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2022:CIi

[Ano22i] Anonymous. Cover image. Journal of the Association for Information Science and Technology, 73(9):C1, September 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2022:CIj

[Ano22j] Anonymous. Cover image. Journal of the Association for Information Science and Technology, 73(10):C1, October 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2022:CIk

[Ano22k] Anonymous. Cover image. Journal of the Association for Information Science and Technology, 73(11):C1, November 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2022:CIl

[Ano22l] Anonymous. Cover image. Journal of the Association for Information Science and Technology, 73(12):C1, December 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2022:IIa

[Ano22m] Anonymous. Issue information. Journal of the Association for Information Science and Technology, 73(1):1–2, January 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2022:IIb

[Ano22n] Anonymous. Issue information. Journal of the Association for Information Science and Technology, 73(2):141–142, February 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2022:IIc

[Ano22o] Anonymous. Issue information. Journal of the Association for Information Science and Technology, 73(3):345–346, March 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2022:IId

[Ano22p] Anonymous. Issue information. Journal of the Association for Information Science and Technology, 73(4):489–490, April 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2022:IIe

[Ano22q] Anonymous. Issue information. Journal of the Association for Information Science and Technology, 73(5):635–636, May 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2022:IIf

[Ano22r] Anonymous. Issue information. Journal of the Association for Information Science and Technology, 73(6):755–756, June 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2022:IIg

[Ano22s] Anonymous. Issue information. Journal of the Association for Information Science and Technology, 73(7):897–898, July 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2022:IIh

[Ano22t] Anonymous. Issue information. Journal of the Association for Information Science and Technology, 73(8):1063–1064, August 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2022:IIi

[Ano22u] Anonymous. Issue information. Journal of the Association for Information Science and Technology, 73(9):1199–1200, September 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2022:IIj

[Ano22v] Anonymous. Issue information. Journal of the Association for Information Science and Technology, 73(10):1363–1364, October 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2022:IIk

[Ano22w] Anonymous. Issue information. Journal of the Association for Information Science and Technology, 73(11):1511–1512, November 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2022:IIl

[Ano22x] Anonymous. Issue information. Journal of the Association for Information Science and Technology, 73(12):1663–1664, December 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2023:CIa

[Ano23a] Anonymous. Cover image. Journal of the Association for Information Science and Technology, 74(1):C1, January 2023. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2023:CIb

[Ano23b] Anonymous. Cover image. Journal of the Association for Information Science and Technology, 74(2):C1, February 2023. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2023:CIc

[Ano23c] Anonymous. Cover image. Journal of the Association for Information Science and Technology, 74(3):C1, March 2023. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2023:CId

[Ano23d] Anonymous. Cover image. Journal of the Association for Information Science and Technology, 74(4):C1, April 2023. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2023:CIe

[Ano23e] Anonymous. Cover image. Journal of the Association for Information Science and Technology, 74(5):C1, May 2023. CO-DEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2023:CIf

[Ano23f] Anonymous. Cover image. Journal of the Association for Information Science and Technology, 74(6):C1, June 2023. CO-DEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2023:CIg

[Ano23g] Anonymous. Cover image. Journal of the Association for Information Science and Technology, 74(8):C1, August 2023. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2023:CIh

[Ano23h] Anonymous. Cover image. Journal of the Association for Information Science and Technology, 74(9):C1, September 2023. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2023:CIi

[Ano23i] Anonymous. Cover image. Journal of the Association for Information Science and Technology, 74(10):C1, October 2023. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2023:CIj

[Ano23j] Anonymous. Cover image. Journal of the Association for Information Science and Technology, 74(12):C1, December 2023. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2023:IIa

[Ano23k] Anonymous. Issue information. Journal of the Association for Information Science and Technology, 74(1):1–2, January 2023. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2023:IIb

[Ano23l] Anonymous. Issue information. Journal of the Association for Information Science and Technology, 74(2):143–144, February 2023. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2023:IIc

[Ano23m] Anonymous. Issue information. Journal of the Association for Information Science and Technology, 74(3):301, March 2023. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2023:IId

[Ano23n] Anonymous. Issue information. Journal of the Association for Information Science and Technology, 74(4):381, April 2023. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2023:IIe

[Ano23o] Anonymous. Issue information. Journal of the Association for Information Science and Technology, 74(5):491, May 2023. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2023:IIf

[Ano23p] Anonymous. Issue information. Journal of the Association for Information Science and Technology, 74(6):599–600, June 2023. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2023:IIg

[Ano23q] Anonymous. Issue information. Journal of the Association for Information Science and Technology, 74(8):883–884, August 2023. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2023:IIh

[Ano23r] Anonymous. Issue information. Journal of the Association for Information Science and Technology, 74(9):1047–1048, September 2023. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2023:IIi

[Ano23s] Anonymous. Issue information. Journal of the Association for Information Science and Technology, 74(10):1157–1158, October 2023. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2023:IIj

[Ano23t] Anonymous. Issue information. Journal of the Association for Information Science and Technology, 74(12):1343, December 2023. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2024:CIa

[Ano24a] Anonymous. Cover image. Journal of the Association for Information Science and Technology, 75(1):C1, January 2024. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2024:CIb

[Ano24b] Anonymous. Cover image. Journal of the Association for Information Science and Technology, 75(2):C1, February 2024. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2024:CIc

[Ano24c] Anonymous. Cover image. Journal of the Association for Information Science and Technology, 75(4):C1, April 2024. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2024:CId

[Ano24d] Anonymous. Cover image. Journal of the Association for Information Science and Technology, 75(5):C1, May 2024. CO-DEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2024:IIa

[Ano24e] Anonymous. Issue information. Journal of the Association for Information Science and Technology, 75(1):1, January 2024. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2024:IIb

[Ano24f] Anonymous. Issue information. Journal of the Association for Information Science and Technology, 75(2):95, February 2024. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2024:IIc

[Ano24g] Anonymous. Issue information. Journal of the Association for Information Science and Technology, 75(4):375, April 2024. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anonymous:2024:IId

[Ano24h] Anonymous. Issue information. Journal of the Association for Information Science and Technology, 75(5):501–502, May 2024. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Anna:2017:RCB

[AP17] Mikkonen Anna and Vakkari Pertti. Reader characteristics, behavior, and success in fiction book search. Journal of the Association for Information Science and Technology, 68(9): 2154–2165, September 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Azpiazu:2020:CLR

[AP20] Ion Madrazo Azpiazu and Maria Soledad Pera. Is crosslingual readability assessment possible? Journal of the Association for Information Science and Technology, 71(6):644– 656, June 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Albarran:2015:DCI

[APRRC15] Pedro Albarrán, Antonio Perianes-Rodríguez, and Javier Ruiz-Castillo. Differences in citation impact across countries. Journal of the Association for Information Science and Technology, 66(3):512–525, March 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Aloteibi:2014:AGQ

[AS14] Saad Aloteibi and Mark Sanderson. Analyzing geographic query reformulation: an exploratory study. Journal of the Association for Information Science and Technology, 65(1): 13–24, January 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Andersson:2024:ESE

[AS24] Cecilia Andersson and Olof Sundin. The elusive search engine: How search engine use is reflected in survey reports. Journal of the Association for Information Science and Technology, 75(5):613–624, May 2024. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Andrushchenko:2022:UPA

[AST^{+}22] Mykola Andrushchenko, Kirsi Sandberg, Risto Turunen, Jani Marjanen, Mari Hatavara, Jussi Kurunmäki, Timo Nummenmaa, Matti Hyvärinen, Kari Teräs, Jaakko Peltonen, and Jyrki Nummenmaa. Using parsed and annotated corpora to analyze parliamentarians’ talk in Finland. Journal of the Association for Information Science and Technology, 73(2):288–302, February 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Aldosari:2016:UCS

[ASTU16] Misfer Aldosari, Mark Sanderson, Audrey Tam, and Alexandra L. Uitdenbogerd. Understanding collaborative search for places of interest. Journal of the Association for Information Science and Technology, 67(6):1331–1344, June 2016. CO-DEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Abrizah:2014:CIN

[AT14] A. Abrizah and Mike Thelwall. Can the impact of non-Western academic books be measured? An investigation of Google Books and Google Scholar for Malaysia. Journal of the Association for Information Science and Technology, 65 (12):2498–2508, December 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Ayadi:2017:MCB

[ATKD^{+}17] Hajer Ayadi, Mouna Torjmen-Khemakhem, Mariam Daoud, Jimmy Xiangji Huang, and Maher Ben Jemaa. Mining correlations between medically dependent features and image retrieval models for query classification. Journal of the Association for Information Science and Technology, 68(5):1323– 1334, May 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Ayadi:2018:MRR

[ATKD^{+}18] Hajer Ayadi, Mouna Torjmen-Khemakhem, Mariam Daoud, Jimmy Xiangji Huang, and Maher Ben Jemaa. MF-Re-Rank: a modality feature-based re-ranking model for medical image retrieval. Journal of the Association for Information Science and Technology, 69(9):1095–1108, September 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Aranyi:2015:MUE

[AvS15] Gabor Aranyi and Paul van Schaik. Modeling user experience with news websites. Journal of the Association for Information Science and Technology, 66(12):2471–2493, December 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Aranyi:2016:TMU

[AvS16] Gabor Aranyi and Paul van Schaik. Testing a model of userexperience with news websites. Journal of the Association for Information Science and Technology, 67(7):1555–1575, July 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Ackerman:2018:SSO

[AWC18] Brian Ackerman, Chong Wang, and Yi Chen. A sessionspecific opportunity cost model for rank-oriented recommendation. Journal of the Association for Information Science and Technology, 69(10):1259–1270, October 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Aung:2019:IFU

[AZE^{+}19] Htet Htet Aung, Han Zheng, Mojisola Erdt, Ashley Sara Aw, Sei-Ching Joanna Sin, and Yin-Leng Theng. Investigating familiarity and usage of traditional metrics and altmetrics. Journal of the Association for Information Science and Technology, 70(8):872–887, August 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Badia:2014:DIK

[Bad14] Antonio Badia. Data, information, knowledge: an information science analysis. Journal of the Association for Information Science and Technology, 65(6):1279–1287, June 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Belem:2017:STR

[BAG17] Fabiano M. Belém, Jussara M. Almeida, and Marcos A. Gonçalves. A survey on tag recommendation methods. Journal of the Association for Information Science and Technology, 68(4):830–844, April 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Bertin:2016:IDR

[BAGL16] Marc Bertin, Iana Atanassova, Yves Gingras, and Vincent Larivière. The invariant distribution of references in scientific articles. Journal of the Association for Information Science and Technology, 67(1):164–177, January 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Branch:2017:RTF

[BAK^{+}17] Frank Branch, Theresa Arias, Jolene Kennah, Rebekah Phillips, Travis Windleharth, and Jin Ha Lee. Representing transmedia fictional worlds through ontology. Journal of the Association for Information Science and Technology, 68 (12):2771–2782, December 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Barbu:2014:WKK

[Bar14] Eduard Barbu. What kind of knowledge is in Wikipedia? Unsupervised extraction of properties for similar concepts. Journal of the Association for Information Science and Technology, 65(12):2489–2497, December 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Bartolucci:2015:BCC

[Bar15] Francesco Bartolucci. Brief communications: A comparison between the g-index and the h-index based on concentration. Journal of the Association for Information Science and Technology, 66(12):2708–2710, December 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Bates:2015:BRD

[Bat15] Marcia J. Bates. Book review: The Discipline of Organizing — edited by Robert J. Glushko. Journal of the Association for Information Science and Technology, 66(2):432–436, February 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Bates:2020:NB

[Bat20] Marcia J. Bates. The nature of browsing. Journal of the Association for Information Science and Technology, 71(5): 616, May 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Bates:2021:BRS

[Bat21] Marcia J. Bates. Book review: Search foundations: Toward a science of technology-mediated experience. Arafat, Sachi and Ashoori, Elham. Boston, MA: MIT Press, 2019. 448, pp. $65.00 (hardback). (ISBN 978-0-262-03859-1). Journal of the Association for Information Science and Technology, 72(3):377–383, March 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Bawden:2014:BRN

[Baw14] David Bawden. Book reviews: A Normative Theory of the Information Society — by Alistair S. Duff. Journal of the Association for Information Science and Technology, 65(11): 2376–2377, November 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Bornmann:2015:BCW

[BB15] Lutz Bornmann and Johann Bauer. Brief communications: Which of the world’s institutions employ the most highly cited researchers? An analysis of the data from highlycited.com. Journal of the Association for Information Science and Technology, 66(10):2146–2148, October 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Brannon:2022:GCB

[BBC^{+}22] Brittany Brannon, Amy G. Buhler, Tara Tobin Cataldo, Ixchel M. Faniel, Lynn Silipigni Connaway, Joyce Kasman Valenza, and Christopher Cyr. Genre containers: building a theoretical framework for studying formats in information behavior. Journal of the Association for Information Science and Technology, 73(4):609–624, April 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Bornmann:2015:LED

[BBH15] Lutz Bornmann, Johann Bauer, and Robin Haunschild. Letters to the Editor: Distribution of women and men among highly cited scientists. Journal of the Association for Information Science and Technology, 66(12):2715–2716, December 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Bhatia:2016:IRI

[BBM16] Sumit Bhatia, Prakhar Biyani, and Prasenjit Mitra. Identifying the role of individual user messages in an online discussion and its use in thread retrieval. Journal of the Association for Information Science and Technology, 67(2):276–288, February 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Bonaccorsi:2016:DCR

[BC16] Andrea Bonaccorsi and Tindaro Cicero. Distributed or concentrated research excellence? Evidence from a large-scale research assessment exercise. Journal of the Association for Information Science and Technology, 67(12):2976–2992, December 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Beheshti:2015:TMS

[BCAL15] Jamshid Beheshti, Charles Cole, Dhary Abuhimed, and Isabelle Lamoureux. Tracking middle school students’ information behavior via Kuhlthau’s ISP Model:Temporality. Journal of the Association for Information Science and Technology, 66(5):943–960, May 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Banerjee:2017:DDU

[BCK17] Snehasish Banerjee, Alton Y. K. Chua, and Jung-Jae Kim. Don’t be deceived: Using linguistic analysis to learn how to discern online review authenticity. Journal of the Association for Information Science and Technology, 68(6):1525–1538, June 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Benjamin:2014:BVR

[BCZ14] Victor Benjamin, Hsinchun Chen, and David Zimbra. Bridging the virtual and real: the relationship between web content, linkage, and geographical proximity of social movements. Journal of the Association for Information Science and Technology, 65(11):2210–2222, November 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Beynon-Davies:2016:FII

[BD16] Paul Beynon-Davies. Form-ing institutional order: the scaffolding of lists and identifiers. Journal of the Association for Information Science and Technology, 67(11):2738–2753, November 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Bandaragoda:2017:AED

[BDA17] Tharindu Rukshan Bandaragoda, Daswin De Silva, and Damminda Alahakoon. Automatic event detection in microblogs using incremental machine learning. Journal of the Association for Information Science and Technology, 68(10): 2394–2411, October 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Bandaragoda:2018:TMP

[BDA^{+}18] Tharindu Rukshan Bandaragoda, Daswin De Silva, Damminda Alahakoon, Weranja Ranasinghe, and Damien Bolton. Text mining for personalized knowledge extraction from online support groups. Journal of the Association for Information Science and Technology, 69(12):1446–1459, December 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Botting:2017:ESM

[BDH17] Nicola Botting, Lucy Dipper, and Katerina Hilari. The effect of social media promotion on academic article uptake. Journal of the Association for Information Science and Technology, 68(3):795–800, March 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Bu:2018:UPS

[BDLM18] Yi Bu, Ying Ding, Xingkun Liang, and Dakota S. Murray. Understanding persistent scientific collaboration. Journal of the Association for Information Science and Technology, 69 (3):438–448, March 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Bornmann:2014:WPE

[BdMA14] Lutz Bornmann and Felix de Moya Anegón. What proportion of excellent papers makes an institution one of the best worldwide? Specifying thresholds for the interpretation of the results of the SCImago Institutions Ranking and the Leiden Ranking. Journal of the Association for Information Science and Technology, 65(4):732–736, April 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Bu:2018:UST

[BDX^{+}18] Yi Bu, Ying Ding, Jian Xu, Xingkun Liang, Gege Gao, and Yiming Zhao. Understanding success through the diversity of collaborators and the milestone of career. Journal of the Association for Information Science and Technology, 69(1): 87–97, January 2018. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Beaudoin:2016:CBI

[Bea16] Joan E. Beaudoin. Content-based image retrieval methods and professional image users. Journal of the Association for Information Science and Technology, 67(2):350–365, February 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Beaudoin:2017:BRA

[Bea17] Joan E. Beaudoin. Book review: Annual Review of Cultural Heritage Informatics: 2012–2013. Edited by Samantha K. Hastings. Lanham, MD: Rowman & Littlefield, 2014. 290 pp. $85.00 (hardcover). (ISBN 978-0-7591-2333-5). Journal of the Association for Information Science and Technology, 68(5):1335–1337, May 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Belter:2016:CAL

[Bel16] Christopher W. Belter. Citation analysis as a literature search method for systematic reviews. Journal of the Association for Information Science and Technology, 67(11): 2766–2777, November 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Bohlin:2016:RJR

[BELR16] Ludvig Bohlin, Alcides Viamontes Esquivel, Andrea Lancichinetti, and Martin Rosvall. Robustness of journal rankings by network flows with different amounts of memory. Journal of the Association for Information Science and Technology, 67(10):2527–2535, October 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Benson:2018:BRA

[Ben18] Sara R. Benson. Book review: Authors, Users, and Pirates: Copyright Law and Subjectivity. James Meese. Cambridge, MA: The MIT Press, 2018. 240 pp. $35.00 (hardcover) (ISBN 978-0-262-03744-0). Journal of the Association for Information Science and Technology, 69(11):1400–1402, November 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Birdi:2018:TNS

[BF18] Briony Birdi and Nigel Ford. Towards a new sociological model of fiction reading. Journal of the Association for Information Science and Technology, 69(11):1291–1303, November 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Buehling:2022:FAS

[BGS22] Kilian Buehling, Matthias Geissler, and Dorothea Strecker. Free access to scientific literature and its influence on the publishing activity in developing countries: The effect of Sci-Hub in the field of mathematics. Journal of the Association for Information Science and Technology, 73(9):1336–1355, September 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Bullard:2015:LEJ

[BH15] Julia Bullard and James Howison. Learning from Elitist Jerks: Creating high-quality knowledge resources from ongoing conversations. Journal of the Association for Information Science and Technology, 66(11):2267–2276, November 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Bornmann:2016:OMB

[BH16] Lutz Bornmann and Robin Haunschild. Overlay maps based on Mendeley data: the use of altmetrics for readership networks. Journal of the Association for Information Science and Technology, 67(12):3064–3072, December 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Bornmann:2017:ELN

[BH17a] Lutz Bornmann and Robin Haunschild. An empirical look at the Nature index. Journal of the Association for Information Science and Technology, 68(3):653–659, March 2017. CO-DEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Bornmann:2017:RCR

[BH17b] Lutz Bornmann and Robin Haunschild. Relative Citation Ratio (RCR): an empirical attempt to study a new fieldnormalized bibliometric indicator. Journal of the Association for Information Science and Technology, 68(4):1064–1067, April 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Bar-Ilan:2017:BRB

[BI17] Judit Bar-Ilan. Book reviews: Bibliometrics and Research Evaluation: Uses and Abuses. Yves Gingras. Cambridge, MA: MIT Press, 2016. 136 pp. $26.00 (hardcover). (ISBN: 978-0-262-03512-5). Journal of the Association for Information Science and Technology, 68(9):2290–2292, September 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Bi:2022:SCS

[Bi22] Yaxin Bi. Sentiment classification in social media data by combining triplet belief functions. Journal of the Association for Information Science and Technology, 73(7):968–991, July 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Bergman:2020:SDF

[BIW20] Ofer Bergman, Tamar Israeli, and Steve Whittaker. The scalability of different file-sharing methods. Journal of the Association for Information Science and Technology, 71(12): 1424–1438, December 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Bjork:2014:OAS

[Bjö14] Bo-Christer Björk. Open access subject repositories: an overview. Journal of the Association for Information Science and Technology, 65(4):698–706, April 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Buchanan:2019:IBD

[BJR19] Steven Buchanan, Cara Jardine, and Ian Ruthven. Information behaviors in disadvantaged and dependent circumstances and the role of information intermediaries. Journal of the Association for Information Science and Technology, 70(2):117–129, February 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Brunswicker:2018:TDC

[BJSM18] Sabine Brunswicker, Bjørn Jensen, Zhounan Song, and Ann Majchrzak. Transparency as design choice of open data contests. Journal of the Association for Information Science and Technology, 69(10):1205–1222, October 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Boyack:2014:CHD

[BK14] Kevin W. Boyack and Richard Klavans. Creation of a highly detailed, dynamic, global model and map of science. Journal of the Association for Information Science and Technology, 65(4):670–685, April 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Bae:2019:IQR

[BK19] Kyoungman Bae and Youngjoong Ko. Improving question retrieval in community question answering service using dependency relations and question classification. Journal of the Association for Information Science and Technology, 70(11): 1194–1209, November 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Bao:2023:CEC

[BK23] Xin Bao and Ping Ke. Chaos, expansion, and contraction: The information worlds of depression patients during the COVID-19 pandemic lockdown. Journal of the Association for Information Science and Technology, 74(8):971–989, August 2023. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Bowler:2015:CWB

[BKM15] Leanne Bowler, Cory Knobel, and Eleanor Mattern. From cyberbullying to well-being: a narrative-based participatory approach to values-oriented design for social media. Journal of the Association for Information Science and Technology, 66(6):1274–1293, June 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Baumgartner:2014:GBT

[BL14] Susanne E. Baumgartner and Loet Leydesdorff. Group-Based Trajectory Modeling (GBTM) of citations in scholarly literature: Dynamic qualities of “transient” and “sticky knowledge claims”. Journal of the Association for Information Science and Technology, 65(4):797–811, April 2014. CO-DEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Bauer:2016:HCP

[BLB16] Johann Bauer, Loet Leydesdorff, and Lutz Bornmann. Highly cited papers in Library and Information Science (LIS): Authors, institutions, and network structures. Journal of the Association for Information Science and Technology, 67(12):3095–3100, December 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Bu:2021:TDD

[BLGbH21] Yi Bu, Mengyang Li, Weiye Gu, and Win bin Huang. Topic diversity: a discipline scheme-free diversity measurement for journals. Journal of the Association for Information Science and Technology, 72(5):523–539, May 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Bjork:2014:AGO

[BLWP14] Bo-Christer Björk, Mikael Laakso, Patrik Welling, and Patrik Paetau. Anatomy of green open access. Journal of the Association for Information Science and Technology, 65(2): 237–250, February 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Bornmann:2014:BCD

[BM14a] Lutz Bornmann and Werner Marx. Brief communication: Distributions instead of single numbers: Percentiles and beam plots for the assessment of single researchers. Journal of the Association for Information Science and Technology, 65(1):206–208, January 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Bornmann:2014:BCW

[BM14b] Lutz Bornmann and Werner Marx. Brief communication: The wisdom of citing scientists. Journal of the Association for Information Science and Technology, 65(6):1288–1292, June 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Bornmann:2014:BCP

[BM14c] Lutz Bornmann and Rüdiger Mutz. Brief communication: From P100 to P100’: a new citation-rank approach. Journal of the Association for Information Science and Technology, 65(9):1939–1943, September 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Bouhnik:2014:GDM

[BM14d] Dan Bouhnik and Deshen Mor. Gender differences in the moral judgment and behavior of Israeli adolescents in the Internet environment. Journal of the Association for Information Science and Technology, 65(3):551–559, March 2014.

CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Bornmann:2015:BCH

[BM15a] Lutz Bornmann and Rüdiger Mutz. Brief communications: How well does a university perform in comparison with its peers? The use of odds, and odds ratios, for the comparison of institutional citation impact using the Leiden Rankings. Journal of the Association for Information Science and Technology, 66(12):2711–2713, December 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Bornmann:2015:GRM

[BM15b] Lutz Bornmann and Rüdiger Mutz. Growth rates of modern science: a bibliometric analysis based on the number of publications and cited references. Journal of the Association for Information Science and Technology, 66(11):2215–2222, November 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Berget:2020:WKA

[BM20] Gerd Berget and Andrew MacFarlane. What is known about the impact of impairments on information seeking and searching? Journal of the Association for Information Science and Technology, 71(5):596–611, May 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Bossaller:2023:RDL

[BM23] Jenny Bossaller and Anthony J. Million. The research data life cycle, legacy data, and dilemmas in research data management. Journal of the Association for Information Science and Technology, 74(6):701–706, June 2023. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Baumer:2017:CGT

[BMG^{+}17] Eric P. S. Baumer, David Mimno, Shion Guha, Emily Quan, and Geri K. Gay. Comparing grounded theory and topic modeling: Extreme divergence or unlikely convergence? Journal of the Association for Information Science and Technology, 68(6):1397–1410, June 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Becker:2020:DUA

[BMM20] Christoph Becker, Emily Maemura, and Nathan Moles. The design and use of assessment frameworks in digital curation. Journal of the Association for Information Science and Technology, 71(1):55–68, January 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Baron:2014:CNB

[BNG14] Luis Fernando Baron, Moriah Neils, and Ricardo Gomez. Crossing new borders: computers, mobile phones, transportation, and English language among Hispanic day laborers in Seattle, Washington. Journal of the Association for Information Science and Technology, 65(1):98–108, January 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Browne:2019:DDO

[BOHK19] Oliver Browne, Philip O’Reilly, Mark Hutchinson, and Nenad B. Krdzavac. Distributed data and ontologies: an integrated Semantic Web architecture enabling more efficient data management. Journal of the Association for Information Science and Technology, 70(6):575–586, June 2019. CO-DEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Bornmann:2014:BRD

[Bor14a] Lutz Bornmann. Book review: Do we still need peer review? An argument for change — by T. H. P. Gould. Journal of the Association for Information Science and Technology, 65 (1):209–213, January 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Bornmann:2014:LET

[Bor14b] Lutz Bornmann. Letter to the Editor: Is there currently a scientific revolution in Scientometrics? Journal of the Association for Information Science and Technology, 65(3): 647–648, March 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Bornmann:2014:LEF

[Bor14c] Lutz Bornmann. Letter to the Editor: On the function of university rankings. Journal of the Association for Information Science and Technology, 65(2):428–429, February 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Bornmann:2014:LER

[Bor14d] Lutz Bornmann. Letter to the Editor: The reception of publications by scientists in the early days of modern science. Journal of the Association for Information Science and Technology, 65(10):2160–2161, October 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Bornmann:2015:IRC

[Bor15a] Lutz Bornmann. Interrater reliability and convergent validity of F1000Prime peer review. Journal of the Association for Information Science and Technology, 66(12):2415–2426, December 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Bornmann:2015:LEN

[Bor15b] Lutz Bornmann. Letters to the Editor: Nature’s top 100 revisited. Journal of the Association for Information Science and Technology, 66(10):2166, October 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Bornmann:2015:LEC

[Bor15c] Lutz Bornmann. Letters to the Editor: Complex tasks and simple solutions: the use of heuristics in the evaluation of research. Journal of the Association for Information Science and Technology, 66(8):1738–1739, August 2015. CO-DEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Bornmann:2016:HMD

[Bor16a] Lutz Bornmann. How much does the expected number of citations for a publication change if it contains the address of a specific scientific institute? A new approach for the analysis of citation data on the institutional level based on regression models. Journal of the Association for Information Science and Technology, 67(9):2274–2282, September 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Bornmann:2016:LEW

[Bor16b] Lutz Bornmann. Letter to the editor: What do altmetrics counts mean? A plea for content analyses. Journal of the Association for Information Science and Technology, 67 (4):1016–1017, April 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Borrego:2016:MCS

[Bor16c] Ángel Borrego. Measuring compliance with a Spanish Government open access mandate. Journal of the Association for Information Science and Technology, 67(4):757–764, April 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Bornmann:2017:CAS

[Bor17] Lutz Bornmann. Is collaboration among scientists related to the citation impact of papers because their quality increases with collaboration? An analysis based on data from F1000Prime and normalized citation scores. Journal of the Association for Information Science and Technology, 68(4): 1036–1047, April 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Belitz:2023:CCM

[BOR^{+}23] Clara Belitz, Jaclyn Ocumpaugh, Steven Ritter, Ryan S. Baker, Stephen E. Fancsali, and Nigel Bosch. Constructing categories: Moving beyond protected classes in algorithmic fairness. Journal of the Association for Information Science and Technology, 74(6):663–668, June 2023. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Bowker:2019:BRR

[Bow19a] Lynne Bowker. Book review: Register Variation Online. Douglas Biber and Jesse Egbert. Cambridge, UK: Cambridge University Press. 2018. $110.00. 252 pp. (hardcover). (ISBN 978-1-107-12216-1). Journal of the Association for Information Science and Technology, 70(11):1293– 1295, November 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Bowler:2019:BRP

[Bow19b] Leanne Bowler. Book review: Protecting Children Online? Cyberbullying Policies of Social Media Companies. Milosevic, Tijana. Cambridge, MA: MIT Press, 2018. 296 pp. $35.00 (Hardcover). (ISBN 978-0-262-03709-9). Journal of the Association for Information Science and Technology, 70 (2):202–203, February 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Burgess:2021:MOP

[BPB21] Stephen Burgess, Craig M. Parker, and Scott Bingley. Mapping the online presence of small local sporting clubs. Journal of the Association for Information Science and Technology, 72(4):433–448, April 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Bawden:2015:AISb

[BR15a] David Bawden and Lyn Robinson. Advances in information science: “A few exciting words”: Information and entropy revisited. Journal of the Association for Information Science and Technology, 66(10):1965–1987, October 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Bawden:2015:AISc

[BR15b] David Bawden and Lyn Robinson. Advances in information science: “Waiting for Carnot”: Information and complexity. Journal of the Association for Information Science and Technology, 66(11):2177–2186, November 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Bodoff:2016:QTI

[BR16] David Bodoff and Daphne Raban. Question types and intermediary elicitations. Journal of the Association for Information Science and Technology, 67(2):289–304, February 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Bawden:2020:DOP

[BR20] David Bawden and Lyn Robinson. “The dearest of our possessions”: Applying Floridi’s information privacy concept in models of information behavior and information literacy. Journal of the Association for Information Science and Technology, 71(9):1030–1043, September 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Bodoff:2020:VIT

[BRL20] David Bodoff and Yaffa Richter-Levin. Viewpoints in indexing term assignment. Journal of the Association for Information Science and Technology, 71(4):450–461, April 2020. CO-DEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Bawden:2015:AISa

[BRS15] David Bawden, Lyn Robinson, and Tyabba Siddiqui. Advances in information science: “Potentialities or possibilities”: Towards quantum information science? Journal of the Association for Information Science and Technology, 66 (3):437–449, March 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Blank:2016:LMR

[BRS16] Ido Blank, Lior Rokach, and Guy Shani. Leveraging metadata to recommend keywords for academic papers. Journal of the Association for Information Science and Technology, 67(12):3073–3091, December 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Brunton:2014:BRS

[Bru14] Finn Brunton. Book review: Summa Technologiae — by Stanislaw Lem. Journal of the Association for Information Science and Technology, 65(12):2592–2594, December 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Brunner:2018:BRD

[Bru18] Robert J. Brunner. Book review: The Data Science Handbook. Field Cady. Hoboken, NJ: John Wiley & Sons, Inc., 2017. 416 pp. $59.95 (Hardcover). (ISBN 978-1-119-09294-0). Journal of the Association for Information Science and Technology, 69(6):861–863, June 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Berget:2016:DAF

[BS16] Gerd Berget and Frode Eika Sandnes. Do autocomplete functions reduce the impact of dyslexia on information-searching behavior? The case of Google. Journal of the Association for Information Science and Technology, 67(10):2320–2328, October 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Bensman:2017:LIS

[BS17] Stephen J. Bensman and Lawrence J. Smolinsky. Lotka’s inverse square law of scientific productivity: Its methods and statistics. Journal of the Association for Information Science and Technology, 68(7):1786–1791, July 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Borgman:2019:DDA

[BSG19] Christine L. Borgman, Andrea Scharnhorst, and Milena S. Golshan. Digital data archives as knowledge infrastructures: Mediating data sharing and reuse. Journal of the Association for Information Science and Technology, 70(8):888–904, August 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Bando:2015:QBS

[BST15] Lorena Leal Bando, Falk Scholer, and Andrew Turpin. Query-biased summary generation assisted by query expansion. Journal of the Association for Information Science and Technology, 66(5):961–979, May 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Bessa:2017:EIC

[BSVZ17] A. Bessa, R. L. T. Santos, A. Veloso, and N. Ziviani. Exploiting item co-utility to improve collaborative filtering recommendations. Journal of the Association for Information Science and Technology, 68(10):2380–2393, October 2017. CO-DEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Brandao:2014:LEQ

[BSZ^{+}14] Wladmir C. Brandão, Rodrygo L. T. Santos, Nivio Ziviani, Edleno S. de Moura, and Altigran S. da Silva. Learning to expand queries using entities. Journal of the Association for Information Science and Technology, 65(9):1870–1883, September 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Bornmann:2016:ABR

[BTMS16] Lutz Bornmann, Andreas Thor, Werner Marx, and Hermann Schier. The application of bibliometrics to research evaluation in the humanities and social sciences: an exploratory study using normalized Google Scholar data for the publications of a research institute. Journal of the Association for Information Science and Technology, 67(11):2778–2789, November 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Buckland:2015:BRP

[Buc15] Michael K. Buckland. Book reviews: Paper Knowledge: Toward a Media History of Documents — edited by Lisa Gitelman. Journal of the Association for Information Science and Technology, 66(5):1088–1089, May 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Buchanan:2020:BRF

[Buc20] Elizabeth A. Buchanan. Book review: Foundations of Information Ethics. Edited by John T. F. Burgess and Emily J. M. Knox. Chicago: ALA Neal-Schuman, 2019. 168 pp. $54.99 (paperback). (ISBN 978-0-8389-1722-0). Journal of the Association for Information Science and Technology, 71 (6):725–726, June 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Budd:2017:BRA

[Bud17] John M. Budd. Book review: Access to Scientific Research: Challenges Facing Communication in STM. David J. Brown. Berlin: Walter de Gruyter, 2016. 423 pp. (Hardcover 93.41 ). (hardcover). (ISBN: 978-3-11-037616-6). Journal of the Association for Information Science and Technology, 68(8): 2033–2034, August 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Burford:2014:GTP

[Bur14] Sally Burford. A grounded theory of the practice of web information architecture in large organizations. Journal of the Association for Information Science and Technology, 65 (10):2017–2034, October 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Burmeister:2017:LEF

[Bur17] Oliver Burmeister. Letter to the Editor: Further considerations in EMR adoption. Journal of the Association for Information Science and Technology, 68(7):1804, July 2017. CO-DEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Burke:2020:RBR

[Bur20] Colin B. Burke. Response to the book review of America’s Information Wars. Journal of the Association for Information Science and Technology, 71(12):1512, December 2020. CO-DEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Brown:2021:IBS

[BV21] Lindsay K. Brown and Tiffany C. Veinot. Information behavior and social control: Toward an understanding of conflictual information behavior in families managing chronic illness. Journal of the Association for Information Science and Technology, 72(1):66–82, January 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Bron:2016:MSR

[BVd16] Marc Bron, Jasmijn Van Gorp, and Maarten de Rijke. Media studies research in the data-driven age: How research questions evolve. Journal of the Association for Information Science and Technology, 67(7):1535–1554, July 2016. CO-DEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Brown:2015:TAD

[BVH15] Susan A. Brown, Viswanath Venkatesh, and Hartmut Hoehle. Technology adoption decisions in the household: a seven-model comparison. Journal of the Association for Information Science and Technology, 66(9):1933–1949, September 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Bergman:2014:SFR

[BWF14] Ofer Bergman, Steve Whittaker, and Noa Falk. Shared files: the retrieval perspective. Journal of the Association for Information Science and Technology, 65(10):1949–1963, October 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Borgman:2021:CQR

[BWGD21] Christine L. Borgman, Morgan F. Wofford, Milena S. Golshan, and Peter T. Darch. Collaborative qualitative research at scale: Reflections on 20 years of acquiring global data and making data global. Journal of the Association for Information Science and Technology, 72(6):667–682, June 2021. CO-DEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Bornmann:2015:BCB

[BWL15] Lutz Bornmann, Caroline Wagner, and Loet Leydesdorff. Brief communication: BRICS countries and scientific excellence: a bibliometric analysis of most frequently cited papers. Journal of the Association for Information Science and Technology, 66(7):1507–1513, July 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Bhattacharya:2016:PPC

[BYSB16] Sanmitra Bhattacharya, Chao Yang, Padmini Srinivasan, and Bob Boynton. Perceptions of Presidential candidates’ personalities in Twitter. Journal of the Association for Information Science and Technology, 67(2):249–267, February 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Chassanoff:2020:CIF

[CA20] Alexandra Chassanoff and Micah Altman. Curation as “Interoperability With the Future”: Preserving scholarly research software in academic libraries. Journal of the Association for Information Science and Technology, 71(3):325– 337, March 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Cabanac:2016:BLS

[Cab16] Guillaume Cabanac. Bibliogifts in LibGen? A study of a text-sharing platform driven by biblioleaks and crowdsourcing. Journal of the Association for Information Science and Technology, 67(4):874–884, April 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Carbo:2014:MBT

[Car14a] Toni Carbo. In memoriam: A brief tribute by Toni Carbo. Journal of the Association for Information Science and Technology, 65(8):1731, August 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Carlin:2014:BRT

[Car14b] Andrew P. Carlin. Book review: Theories of Information, Communication and Knowledge: A Multidisciplinary Approach — edited by Fidelia Ibekwe-SanJuan and Thomas M. Dousa. Journal of the Association for Information Science and Technology, 65(6):1299–1302, June 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Carter:2021:IAR

[CAS21] Daniel Carter, Amelia Acker, and Dan Sholler. Investigative approaches to researching information technology companies. Journal of the Association for Information Science and Technology, 72(6):655–666, June 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Collins:2014:LES

[CB14] Harry Collins and Lutz Bornmann. Letter to the Editor: On scientific misconduct. Journal of the Association for Information Science and Technology, 65(5):1089–1090, May 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Chua:2015:URH

[CB15] Alton Y. K. Chua and Snehasish Banerjee. Understanding review helpfulness as a function of reviewer reputation, review rating, and review depth. Journal of the Association for Information Science and Technology, 66(2):354–362, February 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Cole:2015:EGK

[CBAL15] Charles Cole, Jamshid Beheshti, Dhary Abuhimed, and Isabelle Lamoureux. The end game in Kuhlthau’s ISP model: Knowledge construction for grade 8 students researching an inquiry-based history project. Journal of the Association for Information Science and Technology, 66(11):2249–2266, November 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Chen:2016:TOI

[CBH16] Hongliang Chen, Christopher E. Beaudoin, and Traci Hong. Teen online information disclosure: Empirical testing of a protection motivation and social capital model. Journal of the Association for Information Science and Technology, 67 (12):2871–2881, December 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Chin:2019:DRP

[CBJ19] Jin Yao Chin, Sourav S. Bhowmick, and Adam Jatowt. Ondemand recent personal tweets summarization on mobile devices. Journal of the Association for Information Science and Technology, 70(6):547–562, June 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Cleverley:2017:EIS

[CBM17] Paul H. Cleverley, Simon Burnett, and Laura Muir. Exploratory information searching in the enterprise: a study of user satisfaction and task performance. Journal of the Association for Information Science and Technology, 68(1): 77–96, January 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Colavizza:2018:CBS

[CBvEW18] Giovanni Colavizza, Kevin W. Boyack, Nees Jan van Eck, and Ludo Waltman. The closer the better: Similarity of publication pairs at different cocitation levels. Journal of the Association for Information Science and Technology, 69 (4):600–609, April 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Chen:2016:VWS

[CC16] Rex H.-G. Chen and Chi-Ming Chen. Visualizing the world’s scientific publications. Journal of the Association for Information Science and Technology, 67(10):2477–2488, October 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Clement:2017:CTP

[CC17] Tanya E. Clement and Daniel Carter. Connecting theory and practice in digital humanities information work. Journal of the Association for Information Science and Technology, 68 (6):1385–1396, June 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Candela:2022:DET

[CC22] Gustavo Candela and Rafael C. Carrasco. Discovering emerging topics in textual corpora of galleries, libraries, archives, and museums institutions. Journal of the Association for Information Science and Technology, 73(6):820–833, June 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Chen:2014:CDB

[CCC14] Yen-Liang Chen, Ching-Hao Chuang, and Yu-Ting Chiu. Community detection based on social interactions in a social network. Journal of the Association for Information Science and Technology, 65(3):539–550, March 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Coughlin:2016:WAA

[CCJ16] Daniel M. Coughlin, Mark C. Campbell, and Bernard J. Jansen. A web analytics approach for appraising electronic resources in academic libraries. Journal of the Association for Information Science and Technology, 67(3):518–534, March 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Cheang:2014:MAE

[CCLL14] Brenda Cheang, Samuel Kai Wah Chu, Chongshou Li, and Andrew Lim. A multidimensional approach to evaluating management journals: Refining PageRank via the differentiation of citation types and identifying the roles that management journals play. Journal of the Association for Information Science and Technology, 65(12):2581–2591, December 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Cerda-Cosme:2023:ASR

[CCM23] Roxana Cerda-Cosme and Eva Méndez. Analysis of shared research data in Spanish scientific papers about COVID-19: a first approach. Journal of the Association for Information Science and Technology, 74(4):402–414, April 2023. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Candela:2015:AIS

[CCMT15] Leonardo Candela, Donatella Castelli, Paolo Manghi, and Alice Tani. Advances in information science: Data journals: a survey. Journal of the Association for Information Science and Technology, 66(9):1747–1762, September 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Chao:2015:DPC

[CCP15] Tiffany C. Chao, Melissa H. Cragin, and Carole L. Palmer. Data Practices and Curation Vocabulary (DPCVocab): an empirically derived framework of scientific data practices and curatorial processes. Journal of the Association for Information Science and Technology, 66(3):616–633, March 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Chavula:2024:SCH

[CCR24] Catherine Chavula, Yujin Choi, and Soo Young Rieh. Searching for creativity: How people search to generate new ideas. Journal of the Association for Information Science and Technology, 75(4):438–453, April 2024. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Cotelo:2014:DTR

[CCT14] Juan M. Cotelo, Fermin L. Cruz, and Jose A. Troyano. Dynamic topic-related tweet retrieval. Journal of the Association for Information Science and Technology, 65(3):513–523, March 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Colace:2015:IRF

[CDGN15] Francesco Colace, Massimo De Santo, Luca Greco, and Paolo Napoletano. Improving relevance feedback-based query expansion by the use of a weighted word pairs approach. Journal of the Association for Information Science and Technology, 66(11):2223–2234, November 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Cho:2018:AAT

[CDL18] Hyerim Cho, Andy Donovan, and Jin Ha Lee. Art in an algorithm: a taxonomy for describing video game visual styles. Journal of the Association for Information Science and Technology, 69(5):633–646, May 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Chen:2022:MCC

[CDL22] Liyue Chen, Jielan Ding, and Vincent Larivière. Measuring the citation context of national self-references. Journal of the Association for Information Science and Technology, 73(5):

671–686, May 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Costas:2021:HCO

[CdRM21] Rodrigo Costas, Sarah de Rijcke, and Noortje Marres. “Heterogeneous couplings”: Operationalizing network perspectives to study science-society interactions through social media metrics. Journal of the Association for Information Science and Technology, 72(5):595–610, May 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Casnici:2017:ARM

[CGGS17] Niccolò Casnici, Francisco Grimaldo, Nigel Gilbert, and Flaminio Squazzoni. Attitudes of referees in a multidisciplinary journal: an empirical analysis. Journal of the Association for Information Science and Technology, 68(7):1763– 1771, July 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Chang:2016:ERR

[CH16] Han-Wen Chang and Mu-Hsuan Huang. The effects of research resources on international collaboration in the astronomy community. Journal of the Association for Information Science and Technology, 67(10):2489–2510, October 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Chan:2014:BRT

[Cha14] Anita Say Chan. Book reviews: Technologies of Choice: ICTs, Development, and the Capabilities Approach — by Dorothea Kleine. Journal of the Association for Information Science and Technology, 65(11):2378–2379, November 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Chen:2015:IUI

[Che15] Annie T. Chen. Information use and illness representations: Understanding their connection in illness coping. Journal of the Association for Information Science and Technology, 66 (2):340–353, February 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Chen:2016:CBL

[Che16] Tsung Teng Chen. The congruity between linkage-based factors and content-based clusters — an experimental study using multiple document corpora. Journal of the Association for Information Science and Technology, 67(3):610–619, March 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Chen:2022:IBA

[Che22] Annie T. Chen. Interactions between affect, cognition, and information behavior in the context of fibromyalgia. Journal of the Association for Information Science and Technology, 73(1):31–44, January 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Cabanac:2014:SVC

[CHH14] Guillaume Cabanac, Gilles Hubert, and James Hartley. Solo versus collaborative writing: Discrepancies in the use of tables and graphs in academic articles. Journal of the Association for Information Science and Technology, 65(4):812–820, April 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Chen:2017:KDC

[CHH^{+}17] Liang Chen, Clyde W. Holsapple, Shih-Hui (Steven) Hsiao, Zhihong Ke, Jae-Young Oh, and Zhiguo Yang. Knowledgedissemination channels: Analytics of stature evaluation. Journal of the Association for Information Science and Technology, 68(4):911–930, April 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Chi:2019:FSR

[Chi19] Pei-Shan Chi. The field-specific reference patterns of periodical and nonserial publications. Journal of the Association for Information Science and Technology, 70(3):283–292, March 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Chi:2020:LSS

[CHJ20] Yu Chi, Daqing He, and Wei Jeng. Laypeople’s source selection in online health information-seeking process. Journal of the Association for Information Science and Technology, 71 (12):1484–1499, December 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Crespo:2014:ECI

[CHLRC14] Juan A. Crespo, Neus Herranz, Yunrong Li, and Javier Ruiz-Castillo. The effect on citation inequality of differences in citation practices at the Web of Science subject category level. Journal of the Association for Information Science and Technology, 65(6):1244–1256, June 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Choi:2015:CAT

[Cho15] Yunseon Choi. A complete assessment of tagging quality: a consolidated methodology. Journal of the Association for Information Science and Technology, 66(4):798–817, April 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Chowdhury:2016:HIS

[Cho16] Gobiinda G. Chowdhury. How to improve the sustainability of digital libraries and information Services? Journal of the Association for Information Science and Technology, 67(10): 2379–2391, October 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Chowdhury:2019:BRD

[Cho19] Gobinda Chowdhury. Book review: Digital Libraries for Cultural Heritage: Development, Outcomes, and Challenges from European Perspectives. Aparac-Jelusić Tatjana. San Raphael, CA: Morgan and Claypool, 2017. 175 pp. $59.95 (paperback). (ISBN 978-1-68173-083-7). Journal of the Association for Information Science and Technology, 70(1):96– 97, January 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Choi:2020:OAC

[Cho20] Wonchan Choi. Older adults’ credibility assessment of online health information: an exploratory study using an extended typology of web credibility. Journal of the Association for Information Science and Technology, 71(11):1295– 1307, November 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Choo:2023:CCI

[Cho23] Chun Wei Choo. Climate change information seeking. Journal of the Association for Information Science and Technology, 74(9):1086–1099, September 2023. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Christensen:2017:RII

[Chr17] Hans Dam Christensen. Rethinking image indexing? Journal of the Association for Information Science and Technology, 68(7):1782–1785, July 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Chen:2015:JMS

[CHT^{+}15] Ziyan Chen, Yu Huang, Jing Tian, Xiaoyan Liu, Kun Fu, and Tinglei Huang. Joint model for subsentence-level sentiment analysis with Markov logic. Journal of the Association for Information Science and Technology, 66(9):1913–1922, September 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Costa-jussa:2015:HMH

[Cj15] Marta R. Costa-jussà. How much hybridization does machine translation need? Journal of the Association for Information Science and Technology, 66(10):2160–2165, October 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Choi:2016:BWE

[CJ16a] Namjoo Choi and Soohyung Joo. Booklovers’ world: an examination of factors affecting continued usage of social cataloging sites. Journal of the Association for Information Science and Technology, 67(12):3022–3035, December 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Coughlin:2016:MJB

[CJ16b] Daniel M. Coughlin and Bernard J. Jansen. Modeling journal bibliometrics to predict downloads and inform purchase decisions at university research libraries. Journal of the Association for Information Science and Technology, 67(9): 2263–2273, September 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Ciesielska:2022:FDS

[CJ22] Malgorzata Ciesielska and Dariusz Jemielniak. Fairness in digital sharing legal professional attitudes toward digital piracy and digital commons. Journal of the Association for Information Science and Technology, 73(7):899–912, July 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Chen:2014:SMM

[CK14] Ya-Ning Chen and Hao-Ren Ke. A study on mental models of taggers and experts for article indexing based on analysis of keyword usage. Journal of the Association for Information Science and Technology, 65(8):1675–1694, August 2014. CO-DEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Chew:2016:CDI

[CK16a] Shu Wen Chew and Christopher S. G. Khoo. Comparison of drug information on consumer drug review sites versus authoritative health information websites. Journal of the Association for Information Science and Technology, 67(2): 333–349, February 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Colvin:2016:FRS

[CK16b] Erin Colvin and Donald H. Kraft. Fuzzy retrieval for software reuse. Journal of the Association for Information Science and Technology, 67(10):2454–2463, October 2016. CO-DEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Chowdhury:2017:IPS

[CK17] Gobinda Chowdhury and Kushwanth Koya. Information practices for sustainability: Role of iSchools in achieving the UN sustainable development goals (SDGs). Journal of the Association for Information Science and Technology, 68(9): 2128–2138, September 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Cheng:2021:ISS

[CK21a] Wei-Ning Cheng and Christopher S. G. Khoo. Information structures in sociology research papers: Modeling causeeffect and comparison relations in research objective and result statements. Journal of the Association for Information Science and Technology, 72(11):1367–1385, November 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Cooke:2021:OWL

[CK21b] Nicole A. Cooke and Vanessa L. Kitzie. Outsiders-within-library and information science: Reprioritizing the marginalized in critical sociocultural work. Journal of the Association for Information Science and Technology, 72(10):1285–1294, October 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Cushing:2022:PIM

[CK22] Amber L. Cushing and Páraic Kerrigan. Personal information management burden: a framework for describing nonwork personal information management in the context of inequality. Journal of the Association for Information Science and Technology, 73(11):1543–1558, November 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Chung:2016:USC

[CKL16] EunKyung Chung, Nahyun Kwon, and Jungyeoun Lee. Understanding scientific collaboration in the research life cycle: Bio- and nanoscientists’ motivations, information-sharing and communication practices, and barriers to collaboration. Journal of the Association for Information Science and Technology, 67(8):1836–1848, August 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Cox:2017:DRD

[CKLP17] Andrew M. Cox, Mary Anne Kennan, Liz Lyon, and Stephen Pinfield. Developments in research data management in academic libraries: Towards an understanding of research data service maturity. Journal of the Association for Information Science and Technology, 68(9):2182–2200, September 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Colebunders:2014:BCI

[CKR14] Robert Colebunders, Chris Kenyon, and Ronald Rousseau. Brief communication: Increase in numbers and proportions of review articles in Tropical Medicine, Infectious Diseases, and oncology. Journal of the Association for Information Science and Technology, 65(1):201–205, January 2014. CO-DEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Chen:2014:PCM

[CL14] Chaomei Chen and Loet Leydesdorff. Patterns of connections and movements in dual-map overlays: a new method of publication portfolio analysis. Journal of the Association for Information Science and Technology, 65(2):334–351, February 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Comins:2017:ILT

[CL17] Jordan A. Comins and Loet Leydesdorff. Identification of long-term concept-symbols among citations: Do common intellectual histories structure citation behavior? Journal of the Association for Information Science and Technology, 68 (5):1224–1233, May 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Cabanac:2021:PNA

[CL21] Guillaume Cabanac and Cyril Labbé. Prevalence of nonsensical algorithmically generated papers in the scientific literature. Journal of the Association for Information Science and Technology, 72(12):1461–1476, December 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Conde:2016:LCT

[CLA^{+}16] Angel Conde, Mikel Larrañaga, Ana Arruarte, Jon A. Elorriaga, and Dan Roth. LiTeWi: a combined term extraction and entity linking method for eliciting educational ontologies from textbooks. Journal of the Association for Information Science and Technology, 67(2):380–399, February 2016. CO-DEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Clement:2018:BRC

[Cle18] Tanya E. Clement. Book review: Coding Literacy: How Computer Programming Is Changing Writing. Annette Vee. Cambridge, MA: The MIT Press, 2017. $34.00 (hardcover). (ISBN 978-0-262-03624-5). Journal of the Association for Information Science and Technology, 69(9):1174– 1176, September 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Cheung:2015:HOS

[CLL15] Christy M. K. Cheung, Ivy L. B. Liu, and Matthew K. O. Lee. How online social interactions influence customer information contribution behavior in online social shopping communities: a social learning theory perspective. Journal of the Association for Information Science and Technology, 66(12): 2511–2521, December 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Cloonan:2018:BRR

[Clo18] Michèle V. Cloonan. Book review: Research in the Archival Multiverse. Anne J Gilliland, Sue McKemmish, and Andrew J. Lau, (Eds.). Clayton, VIC., Australia: Monash University Publishing, 2017. 1072 pp. $99.95 (paperback). (ISBN: 978-1-876924-67-6). Journal of the Association for Information Science and Technology, 69(4):623–624, April 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Chen:2014:GSO

[CLSS14] Wenhong Chen, Kye-Hyoung Lee, Joseph D. Straubhaar, and Jeremiah Spence. Getting a second opinion: Social capital, digital inequalities, and health information repertoires. Journal of the Association for Information Science and Technology, 65(12):2552–2563, December 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Choo:2023:IMH

[CM23] Chun Wei Choo and Marco Meyer. Information misbehavior: How organizations use information to deceive. Journal of the Association for Information Science and Technology, 74(9): 1081–1085, September 2023. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Costello:2017:ODI

[CMB17] Kaitlin L. Costello, John D. Martin III, and Ashlee Edwards Brinegar. Online disclosure of illicit information: Information behaviors in two drug forums. Journal of the Association for Information Science and Technology, 68(10):2439– 2448, October 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Cui:2015:IIA

[CML^{+}15] Chaoran Cui, Jun Ma, Tao Lian, Zhumin Chen, and Shuaiqiang Wang. Improving image annotation via rankingoriented neighbor search and learning-based keyword propagation. Journal of the Association for Information Science and Technology, 66(1):82–98, January 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Corbara:2023:SQP

[CMS23] Silvia Corbara, Alejandro Moreo, and Fabrizio Sebastiani. Syllabic quantity patterns as rhythmic features for Latin authorship attribution. Journal of the Association for Information Science and Technology, 74(1):128–141, January 2023. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Choo:2014:EWI

[CN14] Chun Wei Choo and Indrani Nadarajah. Early warning information seeking in the 2009 Victorian Bushfires. Journal of the Association for Information Science and Technology, 65(1):84–97, January 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Case:2016:WUM

[CO16] Donald O. Case and Lisa G. O’Connor. What’s the use? Measuring the frequency of studies of information outcomes. Journal of the Association for Information Science and Technology, 67(3):649–661, March 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Celebi:2018:SHA

[ÇÖ18] Arda Çelebi and Arzucan Özgür. Segmenting hashtags and analyzing their grammatical structure. Journal of the Association for Information Science and Technology, 69(5):675– 686, May 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Cocq:2022:RDH

[Coc22] Coppélie Cocq. Revisiting the digital humanities through the lens of indigenous studies — or how to question the cultural blindness of our technologies and practices. Journal of the Association for Information Science and Technology, 73(2):

333–344, February 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Colliander:2015:NAC

[Col15] Cristian Colliander. A novel approach to citation normalization: a similarity-based method for creating reference sets. Journal of the Association for Information Science and Technology, 66(3):489–500, March 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Cole:2017:BRI

[Col17a] Charles Cole. Book review: Introduction to Information Behaviour. Nigel Ford. London: Facet Publishing, 2015. 272 pp. £49.95 (paperback). (ISBN 978-1-85604-850-7). Journal of the Association for Information Science and Technology, 68 (2):530–532, February 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Cole:2017:BRL

[Col17b] Charles Cole. Book reviews: Looking for information: a survey of research on information seeking, needs, and behavior (4th edition). Donald O. Case and Lisa M. Given. Bingley, UK: Emerald Group Publishing, 2016. 528 pp. $82.95 (hardcover). (ISBN: 978-1-78560-968-8). Journal of the Association for Information Science and Technology, 68(9):2284– 2286, September 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Cole:2020:RBR

[Col20] Charles Cole. A rebuttal of the book review of the book titled “The Consciousness’ Drive: Information Need and the Search for Meaning”. Journal of the Association for Information Science and Technology, 71(2):242, February 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic). See [Jul20].

Conway:2020:BRP

[Con20] Paul Conway. Book review: The Politics of Mass Digitization. Nanna Bonde Thylstrup. Cambridge, MA: The MIT Press, 2018. 216 pp. $35.00 (hardcover). (ISBN 978-0-262-03687-0). Journal of the Association for Information Science and Technology, 71(2):245–247, February 2020. CO-DEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Cox:2014:BRH

[Cox14] Andrew M. Cox. Book reviews: How to Conduct a Practice-Based Study: Problems and Methods; Practice Theory, Work and Organization: An Introduction; The Dynamics of Social Practice: Everyday Life and How It Changes. Journal of the Association for Information Science and Technology, 65 (7):1511–1514, July 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Cox:2015:BRR

[Cox15] Richard J. Cox. Book reviews: Records Management and Information Culture: Tackling the People Problem — by Gillian Oliver and Fiorella Foscarini. Journal of the Association for Information Science and Technology, 66(8):1744– 1746, August 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Cox:2023:HAI

[Cox23] Andrew Cox. How artificial intelligence might change academic library work: Applying the competencies literature and the theory of the professions. Journal of the Association for Information Science and Technology, 74(3):367–380, March 2023. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Choi:2022:ABM

[CPL22] Wonchan Choi, Min Sook Park, and Yura Lee. Associations between mastery of life and everyday life information-seeking behavior among older adults: Analysis of the Pew Research Center’s information engaged and information wary survey data. Journal of the Association for Information Science and Technology, 73(3):393–406, March 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Chae:2016:LCA

[CPP^{+}16] Gunho Chae, Jaram Park, Juyong Park, Woon Seung Yeo, and Chungkon Shi. Linking and clustering artworks using social tags: Revitalizing crowd-sourced information on cultural collections. Journal of the Association for Information Science and Technology, 67(4):885–899, April 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Collins:2017:NCP

[CRGG17] Harry M. Collins, Luis Reyes-Galindo, and Paul Ginsparg. A note concerning primary source knowledge. Journal of the Association for Information Science and Technology, 68 (5):1105–1110, May 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Crispo:2015:BCN

[Cri15] Erika Crispo. Brief communications: A new index to use in conjunction with the h-index to account for an author’s relative contribution to publications with high impact. Journal of the Association for Information Science and Technology, 66(11):2381–2383, November 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Cronin:2014:EML

[Cro14] Blaise Cronin. Editorial: Meta-life. Journal of the Association for Information Science and Technology, 65(3):431–432, March 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Cronin:2015:ETN

[Cro15a] Blaise Cronin. Editorial: The ties that (no longer) bind. Journal of the Association for Information Science and Technology, 66(12):2397–2398, December 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Cronin:2015:EWW

[Cro15b] Blaise Cronin. Editorial: The writing on the wall. Journal of the Association for Information Science and Technology, 66(5):873–875, May 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Choi:2015:AIS

[CS15a] Wonchan Choi and Besiki Stvilia. Advances in information science: Web credibility assessment: Conceptualization, operationalization, variability, and models. Journal of the Association for Information Science and Technology, 66(12): 2399–2414, December 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Clarke:2015:SBP

[CS15b] Andrew Clarke and Robert Steele. Smartphone-based public health information systems: Anonymity, privacy and intervention. Journal of the Association for Information Science and Technology, 66(12):2596–2608, December 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Carter:2016:DSG

[CS16a] Daniel Carter and Dan Sholler. Data science on the ground: Hype, criticism, and everyday work. Journal of the Association for Information Science and Technology, 67(10):2309– 2319, October 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Choi:2016:UMA

[CS16b] Erik Choi and Chirag Shah. User motivations for asking questions in online Q&A services. Journal of the Association for Information Science and Technology, 67(5):1182–1197, May 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Choi:2016:CTB

[CS16c] Youngok Choi and Sue Yeon Syn. Characteristics of tagging behavior in digitized humanities online collections. Journal of the Association for Information Science and Technology, 67(5):1089–1104, May 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Chebil:2016:IBD

[CSOD16] Wiem Chebil, Lina Fatima Soualmia, Mohamed Nazih Omri, and Stéfan Jacques Darmoni. Indexing biomedical documents with a possibilistic network. Journal of the Association for Information Science and Technology, 67(4):928–941, April 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Cruz:2016:MLA

[CTM16] Noa P. Cruz, Maite Taboada, and Ruslan Mitkov. A machine-learning approach to negation and speculation detection for sentiment analysis. Journal of the Association for Information Science and Technology, 67(9):2118–2136, September 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Cushing:2015:BRP

[Cus15] Amber L. Cushing. Book reviews: Personal Archiving: Preserving Our Digital Heritage — edited by Donald Hawkins. Journal of the Association for Information Science and Technology, 66(8):1742–1743, August 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Costello:2020:SAH

[CV20] Kaitlin L. Costello and Tiffany C. Veinot. A spectrum of approaches to health information interaction: From avoidance to verification. Journal of the Association for Information Science and Technology, 71(8):871–886, August 2020. CO-DEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Cao:2018:ROC

[CW18] Xiongfei Cao and Dan Wang. The role of online communities in reducing urban-rural health disparities in China. Journal of the Association for Information Science and Technology, 69(7):890–899, July 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Cai:2017:BBP

[CWdR17] Fei Cai, Shuaiqiang Wang, and Maarten de Rijke. Behaviorbased personalization in web search. Journal of the Association for Information Science and Technology, 68(4):855–868, April 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Cui:2023:MSC

[CWL^{+}23] Yunxue Cui, Yongzhen Wang, Xiaozhong Liu, Xianwen Wang, and Xuhong Zhang. Multidimensional scholarly citations: Characterizing and understanding scholars’ citation behaviors. Journal of the Association for Information Science and Technology, 74(1):115–127, January 2023. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Chuang:2014:ISE

[CY14] Katherine Y. Chuang and Christopher C. Yang. Informational support exchanges using different computer-mediated communication formats in a social media alcoholism community. Journal of the Association for Information Science and Technology, 65(1):37–52, January 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Chipidza:2022:EFC

[CY22] Wallace Chipidza and Jie (Kevin) Yan. The effectiveness of flagging content belonging to prominent individuals: The case of Donald Trump on Twitter. Journal of the Association for Information Science and Technology, 73(11):1641–1658, November 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Choi:2015:THI

[CYP15] Sujin Choi, Joshua SungWoo Yang, and Han Woo Park. The triple helix and international collaboration in science. Journal of the Association for Information Science and Technology, 66(1):201–212, January 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Cohan:2017:TCS

[CYYG17] Arman Cohan, Sydney Young, Andrew Yates, and Nazli Goharian. Triaging content severity in online mental health forums. Journal of the Association for Information Science and Technology, 68(11):2675–2689, November 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Chung:2016:SMB

[CZ16] Wingyan Chung and Daniel Zeng. Social-media-based public policy informatics: Sentiment and network analyses of U.S. Immigration and Border Security. Journal of the Association for Information Science and Technology, 67(7):1588–1606, July 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic). See [Ano17n].

Costas:2015:DAC

[CZW15] Rodrigo Costas, Zohreh Zahedi, and Paul Wouters. Do “altmetrics” correlate with citations? Extensive comparison of altmetric indicators with citations from a multidisciplinary perspective. Journal of the Association for Information Science and Technology, 66(10):2003–2019, October 2015. CO-DEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Chang:2015:TEP

[CZZY15] Ke-Chiun Chang, Wei Zhou, Sifei Zhang, and Chien-Chung Yuan. Threshold effects of the patent H-index in the relationship between patent citations and market value. Journal of the Association for Information Science and Technology, 66 (12):2697–2703, December 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Daquino:2020:CAA

[Daq20] Marilena Daquino. A computational analysis of art historical linked data for assessing authoritativeness of attributions. Journal of the Association for Information Science and Technology, 71(7):757–769, July 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Douglass:2014:MSD

[DAT^{+}14] Kimberly Douglass, Suzie Allard, Carol Tenopir, Lei Wu, and Mike Frame. Managing scientific data as public assets: Data sharing practices and policies among full-time government employees. Journal of the Association for Information Science and Technology, 65(2):251–262, February 2014. CO-DEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Dessne:2015:ICI

[DB15] Karin Dessne and Katriina Byström. Imitating CoPs: Imposing formality on informality. Journal of the Association for Information Science and Technology, 66(11):2277–2284, November 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Daraio:2017:BUR

[DB17a] Cinzia Daraio and Andrea Bonaccorsi. Beyond university rankings? Generating new indicators on universities by linking data in open platforms. Journal of the Association for Information Science and Technology, 68(2):508–529, February 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Duretec:2017:FTL

[DB17b] Kresimir Duretec and Christoph Becker. Format technology lifecycle analysis. Journal of the Association for Information Science and Technology, 68(10):2484–2500, October 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Dobrota:2016:NAQ

[DBBJ16] Marina Dobrota, Milica Bulajic, Lutz Bornmann, and Veljko Jeremic. A new approach to the QS university ranking using the composite I-distance indicator: Uncertainty and sensitivity analyses. Journal of the Association for Information Science and Technology, 67(1):200–211, January 2016. CO-DEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Didegah:2018:DBC

[DBH18] Fereshteh Didegah, Timothy D. Bowman, and Kim Holmberg. On the differences between citations and altmetrics: an investigation of factors driving altmetrics versus citations for Finnish articles. Journal of the Association for Information Science and Technology, 69(6):832–843, June 2018. CO-DEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Ding:2014:DTD

[DC14] Wanying Ding and Chaomei Chen. Dynamic topic detection and tracking: a comparison of HDP, C-word, and cocitation methods. Journal of the Association for Information Science and Technology, 65(10):2084–2097, October 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Donaldson:2015:UCT

[DC15] Devan Ray Donaldson and Paul Conway. User conceptions of trustworthiness for digital archival documents. Journal of the Association for Information Science and Technology, 66 (12):2427–2444, December 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

deCastro:2016:LEM

[dC16] Alexandre de Castro. Letters to the Editor: Mental models may fail when faced with self-referential descriptors. Journal of the Association for Information Science and Technology, 67(2):490, February 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Dattolo:2022:ARB

[DC22] Antonina Dattolo and Marco Corbatto. Assisting researchers in bibliographic tasks: a new usable, real-time tool for analyzing bibliographies. Journal of the Association for Information Science and Technology, 73(6):757–776, June 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Du:2021:SDD

[DCLH21] Caifan Du, Johanna Cohoon, Patrice Lopez, and James Howison. Softcite dataset: a dataset of software mentions in biomedical and economic research publications. Journal of the Association for Information Science and Technology, 72(7):870–884, July 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Dobrota:2016:BCA

[DD16] Milan Dobrota and Marina Dobrota. Brief communications: ARWU ranking uncertainty and sensitivity: What if the award factor was excluded? Journal of the Association for Information Science and Technology, 67(2):480–482, February 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

DeRoo:2016:IFB

[DDB16] Berdien De Roo, Philippe De Maeyer, and Jean Bourgeois. Information flows as bases for archeology-specific geodata infrastructures: an exploratory study in Flanders. Journal of the Association for Information Science and Technology, 67(8):1928–1942, August 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Diaz-Faes:2014:ASP

[DFB14] Adrián A. Díaz-Faes and María Bordons. Acknowledgments in scientific publications: Presence in Spanish science and text patterns across disciplines. Journal of the Association for Information Science and Technology, 65(9):1834–1849, September 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

deFremery:2022:CRL

[dFB22a] Wayne de Fremery and Michael K. Buckland. Context, relevance, and labor. Journal of the Association for Information Science and Technology, 73(9):1268–1278, September 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

deFremery:2022:CT

[dFB22b] Wayne de Fremery and Michael K. Buckland. Copy theory. Journal of the Association for Information Science and Technology, 73(3):407–418, March 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Dalip:2017:GMF

[DGCC17] Daniel H. Dalip, Marcos André Gonçalves, Marco Cristo, and Pável Calado. A general multiview framework for assessing the quality of collaboratively created content on Web 2.0. Journal of the Association for Information Science and Technology, 68(2):286–308, February 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Dourado:2019:BTG

[DGGdST19] Ícaro Cavalcante Dourado, Renata Galante, Marcos André Gonçalves, and Ricardo da Silva Torres. Bag of textual graphs (BoTG): a general graph-based text representation model. Journal of the Association for Information Science and Technology, 70(8):817–829, August 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Doval:2019:CNG

[DGR19] Yerai Doval and Carlos Gómez-Rodríguez. Comparing neural- and N-gram-based language models for word segmentation. Journal of the Association for Information Science and Technology, 70(2):187–197, February 2019. CO-DEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

deHerrera:2017:SMC

[dHSM17] Alba G. Seco de Herrera, Roger Schaer, and Henning Müller. Shangri-La: a medical case-based retrieval tool. Journal of the Association for Information Science and Technology, 68 (11):2587–2601, November 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Dinneen:2020:UDF

[DJ20] Jesse David Dinneen and Charles-Antoine Julien. The ubiquitous digital file: a review of file management research. Journal of the Association for Information Science and Technology, 71(1):E1–E32, January 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Detlor:2022:CLD

[DJLS22] Brian Detlor, Heidi Julien, Tara La Rose, and Alexander Serenko. Community-led digital literacy training: Toward a conceptual framework. Journal of the Association for Information Science and Technology, 73(10):1387–1400, October 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Donaldson:2022:RTM

[DL22] Devan Ray Donaldson and Colin Bradley LeFevre. Records, trust, and misinformation: Using birtherism to understand the influence of conspiracy theories on human information interactions. Journal of the Association for Information Science and Technology, 73(11):1579–1593, November 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Dang:2014:BBW

[DLA14] Edward K. F. Dang, Robert W. P. Luk, and James Allan. Beyond bag-of-words: Bigram-enhanced context-dependent term weights. Journal of the Association for Information Science and Technology, 65(6):1134–1148, June 2014. CO-DEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Dang:2016:CDR

[DLA16] Edward Kai Fung Dang, Robert W. P. Luk, and James Allan. A context-dependent relevance model. Journal of the Association for Information Science and Technology, 67(3): 582–593, March 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Dang:2022:RMF

[DLA22] Edward Kai Fung Dang, Robert Wing Pong Luk, and James Allan. A retrieval model family based on the probability ranking principle for ad hoc retrieval. Journal of the Association for Information Science and Technology, 73(8):1140– 1154, August 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Du:2021:PCP

[DLD^{+}21] Qianzhou Du, Jing Li, Yanqing Du, G. Alan Wang, and Weiguo Fan. Predicting crowdfunding project success based on backers’ language preferences. Journal of the Association for Information Science and Technology, 72(12):1558–1574, December 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Pertile:2016:CCC

[dLPMR16] Solange de L. Pertile, Viviane P. Moreira, and Paolo Rosso. Comparing and combining content- and citation-based approaches for plagiarism detection. Journal of the Association for Information Science and Technology, 67(10):2511–2526, October 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Dougherty:2014:CTP

[DM14] Meghan Dougherty and Eric T. Meyer. Community, tools, and practices in web archiving: the state-of-the-art in relation to social science and humanities research needs. Journal of the Association for Information Science and Technology, 65(11):2195–2209, November 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

DeSordi:2016:TMT

[DMdO16] José Osvaldo De Sordi, Manuel Meireles, and Osvaldo Luiz de Oliveira. The Text Matrix as a tool to increase the cohesion of extensive texts. Journal of the Association for Information Science and Technology, 67(4):900–914, April 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Delgado:2017:PND

[DMMF17] Agustín D. Delgado, Raquel Martínez, Soto Montalvo, and Víctor Fresno. Person name disambiguation in the web using adaptive threshold clustering. Journal of the Association for Information Science and Technology, 68(7):1751–1762, July 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Deardorff:2017:PDA

[DMR^{+}17] Ariel Deardorff, Kate Masterton, Kirk Roberts, Halil Kilicoglu, and Dina Demner-Fushman. A protocol-driven approach to automatically finding authoritative answers to consumer health questions in online resources. Journal of the Association for Information Science and Technology, 68(7):

1724–1736, July 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Diekema:2014:TPI

[DO14] Anne R. Diekema and M. Whitney Olsen. Teacher personal information management (PIM) practices: Finding, keeping, and re-finding information. Journal of the Association for Information Science and Technology, 65(11):2261–2277, November 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Davison:2014:CEW

[DOM^{+}14] Robert M. Davison, Carol X. J. Ou, Maris G. Martinsons, Angela Y. Zhao, and Rong Du. The communicative ecology of Web 2.0 at work: Social networking in the workspace. Journal of the Association for Information Science and Technology, 65(10):2035–2047, October 2014. CO-DEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Donner:2016:ESC

[Don16] Paul Donner. Enhanced self-citation detection by fuzzy author name matching and complementary error estimates. Journal of the Association for Information Science and Technology, 67(3):662–670, March 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Dorsch:2019:BRB

[Dor19] Isabelle Dorsch. Book review: Becoming Metric-Wise: a Bibliometric Guide for Researchers. Rousseau, Ronald, Egghe, Leo, and Guns, Raf. Cambridge, MA: Chandos Publishing, 2018. 402 pp. $125.00 (paperback). (ISBN 978-0-08-102474-4). Journal of the Association for Information Science and Technology, 70(5):530–532, May 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Downey:2015:BRD

[Dow15] Gregory J. Downey. Book review: Digital Labor: The Internet as Playground and Factory — edited by Trebor Scholz. Journal of the Association for Information Science and Technology, 66(6):1303–1304, June 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

duPlessis:2020:SWC

[dP20] Jacques Christiaan du Plessis. Smart World Cities in the 21st Century. Agnes, Mainka. Boston: Walter de Gruyter, 2018. 266 pp. (ISBN 978-3-11-057525-5). Journal of the Association for Information Science and Technology, 71(2): 243–244, February 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Das:2023:GTN

[DP23] Sudeshna Das and Jiaul H. Paik. Gender tagging of named entities using retrieval-assisted multi-context aggregation: an unsupervised approach. Journal of the Association for Information Science and Technology, 74(4):461–475, April 2023. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Desrochers:2017:FDG

[DPHP17] Nadine Desrochers, Adèle Paul-Hus, and Jen Pecoskie. Five decades of gratitude: a meta-synthesis of acknowledgments research. Journal of the Association for Information Science and Technology, 68(12):2821–2833, December 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Denning:2016:RLP

[DPN16] Joel Denning, Maria Soledad Pera, and Yiu-Kai Ng. A readability level prediction tool for K–12 books. Journal of the Association for Information Science and Technology, 67(3): 550–565, March 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

DiIorio:2014:DSP

[DPPV14] Angelo Di Iorio, Silvio Peroni, Francesco Poggi, and Fabio Vitali. Dealing with structural patterns of XML documents. Journal of the Association for Information Science and Technology, 65(9):1884–1900, September 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Delgado-Quiros:2024:WPM

[DQAMM^{+}24] Lorena Delgado-Quirós, Isidro F. Aguillo, Alberto Martín-Martín, Emilio Delgado López-Cózar, Enrique Orduña-Malea, and José Luis Ortega. Why are these publications missing? Uncovering the reasons behind the exclusion of documents in free-access scholarly databases. Journal of the Association for Information Science and Technology, 75(1): 43–58, January 2024. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

deRijcke:2015:IIQ

[dRR15] Sarah de Rijcke and Alexander Rushforth. To intervene or not to intervene; is that the question? On the role of scientometrics in research evaluation. Journal of the Association for Information Science and Technology, 66(9):1954–1958, September 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Demarest:2015:AOA

[DS15] Bradford Demarest and Cassidy R. Sugimoto. Argue, observe, assess: Measuring disciplinary identities and differences through socio-epistemic discourse. Journal of the Association for Information Science and Technology, 66(7):1374– 1387, July 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Dumitrescu:2021:FCR

[DS21] Alexandra Dumitrescu and Simone Santini. Full coverage of a reader’s interests in context-based information filtering. Journal of the Association for Information Science and Technology, 72(8):1011–1027, August 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

daSilva:2019:MPD

[dSBG19] Luiz Fernando Cal da Silva, Marcelo Werneck Barbosa, and Rodrigo Richard Gomes. Measuring participation in distance education online discussion forums using social network analysis. Journal of the Association for Information Science and Technology, 70(2):140–150, February 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Darch:2020:LCD

[DSBG20] Peter T. Darch, Ashley E. Sands, Christine L. Borgman, and Milena S. Golshan. Library cultures of data curation: Adventures in astronomy. Journal of the Association for Information Science and Technology, 71(12):1470–1483, December 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Darch:2021:DSA

[DSBG21] Peter T. Darch, Ashley E. Sands, Christine L. Borgman, and Milena S. Golshan. Do the stars align?: Stakeholders and strategies in libraries’ curation of an astronomy dataset. Journal of the Association for Information Science and Technology, 72(2):239–252, February 2021. CO-DEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Motta:2015:PSA

[dSMdAGQ15] Gustavo da Silva Motta, Pauli Adriano de Almada Garcia, and Rogério Hermida Quintella. A patento-scientometric approach to venture capital investment prioritization. Journal of the Association for Information Science and Technology, 66(4):765–777, April 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Didegah:2018:CSC

[DT18] Fereshteh Didegah and Mike Thelwall. Co-saved, co-tweeted, and co-cited networks. Journal of the Association for Information Science and Technology, 69(8):959–973, August 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Dootson:2021:TPR

[DTDT21] Paula Dootson, Mary Tate, Kevin C. Desouza, and Peter Townson. Transforming public records management: Six key insights. Journal of the Association for Information Science and Technology, 72(5):643–648, May 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Du:2016:ERL

[DTW16] Jian Du, Xiaoli Tang, and Yishan Wu. The effects of research level and article type on the differences between citation metrics and F1000 recommendations. Journal of the Association for Information Science and Technology, 67(12): 3008–3021, December 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Du:2014:RME

[Du14a] Helen S. Du. The role of media-embedded heuristics in achieving online readership popularity. Journal of the Association for Information Science and Technology, 65(2):302– 312, February 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Du:2014:IJM

[Du14b] Jia Tina Du. The information journey of marketing professionals: Incorporating work task-driven information seeking, information judgments, information use, and information sharing. Journal of the Association for Information Science and Technology, 65(9):1850–1869, September 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Dutton:2020:BRI

[Dut20] William H. Dutton. Book review: Internet Daemons: Digital Communications Possessed. Fenwick, McKelvey. Minneapolis, MN: University of Minnesota Press, 2018. 336 pp. $28.00 (Paperback). (ISBN 978-1-5179-0154-7). Journal of the Association for Information Science and Technology, 71 (1):121–123, January 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Datta:2014:ZMT

[DYB14] Anwitaman Datta, Jackson Tan Teck Yong, and Stefano Braghin. The zen of multidisciplinary team recommendation. Journal of the Association for Information Science and Technology, 65(12):2518–2533, December 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Ding:2014:CBC

[DZC^{+}14] Ying Ding, Guo Zhang, Tamy Chambers, Min Song, Xiaolong Wang, and Chengxiang Zhai. Content-based citation analysis: the next generation of citation analysis. Journal of the Association for Information Science and Technology, 65 (9):1820–1833, September 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Erfani:2018:IUS

[EA18] Seyedezahra Shadi Erfani and Babak Abedin. Impacts of the use of social network sites on users’ psychological well-being: a systematic review. Journal of the Association for Information Science and Technology, 69(7):900–912, July 2018. CO-DEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Erfani:2017:ESN

[EAB17] Seyedezahra Shadi Erfani, Babak Abedin, and Yvette Blount. The effect of social network site use on the psychological well-being of cancer patients. Journal of the Association for Information Science and Technology, 68(5):1308– 1322, May 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Egbert:2015:DBU

[EBD15] Jesse Egbert, Douglas Biber, and Mark Davies. Developing a bottom-up, user-based method of web register classification. Journal of the Association for Information Science and Technology, 66(9):1817–1831, September 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Escolano:2021:BMN

[ECJF21] Carlos Escolano, Marta R. Costa-Jussà, and José A. R. Fonollosa. From bilingual to multilingual neural-based machine translation by incremental training. Journal of the Association for Information Science and Technology, 72(2): 190–203, February 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Eckerdal:2018:ERA

[Eck18] Johanna Rivano Eckerdal. Equipped for resistance: An agonistic conceptualization of the public library as a verb. Journal of the Association for Information Science and Technology, 69(12):1405–1413, December 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Egghe:2014:IES

[Egg14a] L. Egghe. Informetric explanation of some Leiden Ranking graphs. Journal of the Association for Information Science and Technology, 65(4):737–741, April 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Egghe:2014:BCG

[Egg14b] Leo Egghe. Brief communication: A good normalized impact and concentration measure. Journal of the Association for Information Science and Technology, 65(10):2152–2154, October 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Elsweiler:2015:EMS

[EH15] David Elsweiler and Morgan Harvey. Engaging and maintaining a sense of being informed: Understanding the tasks motivating Twitter search. Journal of the Association for Information Science and Technology, 66(2):264–281, February 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Einarsson:2021:HDM

[EH21] Árni Már Einarsson and Morten Hertzum. How do makers obtain information for their makerspace projects? Journal of the Association for Information Science and Technology, 72 (12):1528–1544, December 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Estrada:2017:TBT

[EHdBvO17] Liliana Melgar Estrada, Michiel Hildebrand, Victor de Boer, and Jacco van Ossenbruggen. Time-based tags for fiction movies: comparing experts to novices using a video labeling game. Journal of the Association for Information Science and Technology, 68(2):348–364, February 2017. CO-DEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Eschenfelder:2014:MDC

[EJ14] Kristin R. Eschenfelder and Andrew Johnson. Managing the data commons: Controlled sharing of scholarly data. Journal of the Association for Information Science and Technology, 65(9):1757–1774, September 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Emmons:2017:MVA

[ELB17] Scott R. Emmons, Robert P. Light, and Katy Börner. MOOC visual analytics: Empowering students, teachers, researchers, and platform developers of massively open online courses. Journal of the Association for Information Science and Technology, 68(10):2350–2363, October 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Ekbia:2015:AIS

[EMK^{+}15] Hamid Ekbia, Michael Mattioli, Inna Kouper, G. Arave, Ali Ghazinejad, Timothy Bowman, Venkata Ratandeep Suri, Andrew Tsou, Scott Weingart, and Cassidy R. Sugimoto.

Advances in information science: Big data, bigger dilemmas: a critical review. Journal of the Association for Information Science and Technology, 66(8):1523–1545, August 2015. CO-DEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Engerer:2017:CSV

[Eng17a] Volkmar Engerer. Control and syntagmatization: Vocabulary requirements in information retrieval thesauri and natural language lexicons. Journal of the Association for Information Science and Technology, 68(6):1480–1490, June 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Engerer:2017:EIR

[Eng17b] Volkmar Engerer. Exploring interdisciplinary relationships between linguistics and information retrieval from the 1960s to today. Journal of the Association for Information Science and Technology, 68(3):660–680, March 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Evans:2016:RRD

[EOG16] Heather K. Evans, Joycelyn Ovalle, and Stephen Green. Rockin’ robins: Do Congresswomen rule the roost in the Twittersphere? Journal of the Association for Information Science and Technology, 67(2):268–275, February 2016. CO-DEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Epifanio:2014:BCM

[Epi14] Irene Epifanio. Brief communication: Mapping the asymmetrical citation relationships between journals by h-plots. Journal of the Association for Information Science and Technology, 65(6):1293–1298, June 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Erdelez:2018:BRA

[Erd18] Sanda Erdelez. Book review: Accidental Information Discovery: Cultivating Serendipity in the Digital Age. Tammera M. Race & Stephann Makri. Cambridge, MA: Chandos Publishing. 2016. 136 pp. $80.00 (paperback). (ISBN 978-1-84334-750-7). Journal of the Association for Information Science and Technology, 69(5):753–756, May 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Eschenfelder:2022:FMS

[ESD22] Kristin R. Eschenfelder, Kalpana Shankar, and Greg Downey. The financial maintenance of social science data archives: Four case studies of long-term infrastructure work. Journal of the Association for Information Science and Technology, 73(12):1723–1740, December 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Eskens:2020:PIS

[Esk20] Sarah Eskens. The personal information sphere: an integral approach to privacy and related information and communication rights. Journal of the Association for Information Science and Technology, 71(9):1116–1128, September 2020. CO-DEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Ebrahimi:2018:TUG

[ESWC18] Mohammad Ebrahimi, Elaheh ShafieiBavani, Raymond Wong, and Fang Chen. Twitter user geolocation by filtering of highly mentioned users. Journal of the Association for Information Science and Technology, 69(7):879–889, July 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Erickson:2016:BBP

[EWX^{+}16] Lee B. Erickson, Pamela Wisniewski, Heng Xu, John M. Carroll, Mary Beth Rosson, and Daniel F. Perkins. The boundaries between: Parental involvement in a teen’s online world. Journal of the Association for Information Science and Technology, 67(6):1384–1403, June 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Fidler:2017:MIC

[FA17] Bradley Fidler and Amelia Acker. Metadata, infrastructure, and computer-mediated communication in historical perspective. Journal of the Association for Information Science and Technology, 68(2):412–422, February 2017. CO-DEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Feng:2019:RPI

[FA19] Yuanyuan Feng and Denise E. Agosto. Revisiting personal information management through information practices with activity tracking technology. Journal of the Association for Information Science and Technology, 70(12):1352–1367, December 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Fang:2018:DCP

[Fan18] Hui Fang. A discussion of citations from the perspective of the contribution of the cited paper to the citing paper. Journal of the Association for Information Science and Technology, 69(12):1513–1520, December 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Farrugia:2023:BRA

[Far23] Charles Farrugia. Book review: Architects of memory: Information and rhetoric in a networked archival age. Nathan R. Johnson. Tuscaloosa: University of Alabama Press. 2020. pp. 224. $49.95 (hardcover). ISBN: 978-0-8173-2060-7. Journal of the Association for Information Science and Technology, 74(10):1203–1206, October 2023. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Fassin:2018:NQR

[Fas18] Yves Fassin. A new qualitative rating system for scientific publications and a fame index for academics. Journal of the Association for Information Science and Technology, 69(11): 1396–1399, November 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Foster:2018:EAC

[FC18] Jonathan Foster and Paul Clough. Embedded, added, cocreated: Revisiting the value of information in an age of data. Journal of the Association for Information Science and Technology, 69(5):744–748, May 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Floegel:2022:MFT

[FC22] Diana Floegel and Kaitlin L. Costello. Methods for a feminist technoscience of information practice: Design justice and speculative futurities. Journal of the Association for Information Science and Technology, 73(4):625–634, April 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Fang:2021:HSC

[FCT^{+}21] Zhichao Fang, Rodrigo Costas, Wencan Tian, Xianwen Wang, and Paul Wouters. How is science clicked on Twitter? Click metrics for Bitly short links to scientific publications. Journal of the Association for Information Science and Technology, 72(7):918–932, July 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Fang:2020:STM

[FDC20] Zhichao Fang, Jonathan Dudek, and Rodrigo Costas. The stability of Twitter metrics: a study on unavailable Twitter mentions of scientific publications. Journal of the Association for Information Science and Technology, 71(12): 1455–1469, December 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Fang:2022:FVT

[FDC22] Zhichao Fang, Jonathan Dudek, and Rodrigo Costas. Facing the volatility of tweets in altmetric research. Journal of the Association for Information Science and Technology, 73(8): 1192–1195, August 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Ferreira:2018:WPM

[FdGPC18] Raoni S. Ferreira, Maria da Graça Pimentel, and Marco Cristo. A wikification prediction model based on the combination of latent, dyadic, and monadic features. Journal of the Association for Information Science and Technology, 69 (3):380–394, March 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Finegan-Dollak:2016:SSC

[FDR16] Catherine Finegan-Dollak and Dragomir R. Radev. Sentence simplification, compression, and disaggregation for summarization of sophisticated documents. Journal of the Association for Information Science and Technology, 67(10):2437– 2453, October 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Frings-Hessami:2023:API

[FHO23] Viviane Frings-Hessami and Gillian Oliver. Accessing and preserving information: Combining ICT4D and archival science to empower marginalized communities. Journal of the Association for Information Science and Technology, 74(12): 1350–1364, December 2023. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Fleischmann:2017:SRC

[FHW17] Kenneth R. Fleischmann, Cindy Hui, and William A. Wallace. The societal responsibilities of computational modelers: Human values and professional codes of ethics. Journal of the Association for Information Science and Technology, 68 (3):543–552, March 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Fiala:2014:BCC

[Fia14] Dalibor Fiala. Brief communication: Current index: A Proposal for a dynamic rating system for researchers. Journal of the Association for Information Science and Technology, 65(4):850–855, April 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Fields:2016:NNT

[Fie16] Chris Fields. Nobel numbers: Time-dependent centrality measures on coauthorship graphs. Journal of the Association for Information Science and Technology, 67(9):2212–2222, September 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Finnemann:2017:HCG

[Fin17] Niels Ole Finnemann. Hypertext configurations: Genres in networked digital media. Journal of the Association for Information Science and Technology, 68(4):845–854, April 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Flynn-Jones:2019:BRP

[FJ19] Emily Flynn-Jones. Book review: Playing With Feelings: Video Games and Affect. Anable, Aubrey. Minneapolis, MN: University of Minnesota Press, 2018. 200 pp. $25.00 (paperback). (ISBN 978-1-5179-0025-0). Journal of the Association for Information Science and Technology, 70(10):1165–1166, October 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Falavarjani:2021:CRB

[FJF^{+}21] Seyed Amin Mirlohi Falavarjani, Jelena Jovanovic, Hossein Fani, Ali A. Ghorbani, Zeinab Noorian, and Ebrahim Bagheri. On the causal relation between real world activities and emotional expressions of social media users. Journal of the Association for Information Science and Technology, 72(6):723–743, June 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Fan:2023:UDC

[FJT23] Wei-Min Fan, Wei Jeng, and Muh-Chyun Tang. Using data citation to define a knowledge domain: a case study of the Add-Health dataset. Journal of the Association for Information Science and Technology, 74(1):81–98, January 2023. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Faniel:2016:SSS

[FKY16] Ixchel M. Faniel, Adam Kriesberg, and Elizabeth Yakel. Social scientists’ satisfaction with data reuse. Journal of the Association for Information Science and Technology, 67 (6):1404–1416, June 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Fang:2023:HDF

[FLJD23] Zian Fang, Yongmei Liu, Fei Jiang, and Wei Dong. How does family support influence digital immigrants’ extended use of smartphones? An empirical study based on IT identity theory. Journal of the Association for Information Science and Technology, 74(12):1463–1481, December 2023. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Fernandez-Molina:2014:BRF

[FM14] J. Carlos Fernández-Molina. Book review: The Fight Over Digital Rights: The Politics of Copyright and Technology — by Bill D. Herman. Journal of the Association for Information Science and Technology, 65(4):862–865, April 2014. CO-DEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Fonseca:2019:CHS

[FMD19] Fred Fonseca, Michael Marcinkowski, and Clodoveu Davis. Cyber-human systems of thought and understanding. Journal of the Association for Information Science and Technology, 70(4):402–411, April 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Fleming-May:2016:DIP

[FMG16] Rachel A. Fleming-May and Harriett Green. Digital innovations in poetry: Practices of creative writing faculty in online literary publishing. Journal of the Association for Information Science and Technology, 67(4):859–873, April 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Foster:2018:DWC

[FMNG18] Jonathan Foster, Julie McLeod, Jan Nolin, and Elke Greifeneder. Data work in context: Value, risks, and governance. Journal of the Association for Information Science and Technology, 69(12):1414–1427, December 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Frittelli:2016:SRM

[FMP16] Marco Frittelli, Loriano Mancini, and Ilaria Peri. Scientific research measures. Journal of the Association for Information Science and Technology, 67(12):3051–3063, December 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Frandsen:2017:CBL

[FN17a] Tove Faber Frandsen and Jeppe Nicolaisen. Citation behavior: a large-scale test of the persuasion by name-dropping hypothesis. Journal of the Association for Information Science and Technology, 68(5):1278–1284, May 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Frandsen:2017:LER

[FN17b] Tove Faber Frandsen and Jeppe Nicolaisen. Letters to the Editor: Rejoinder: Noble prize effects in citation networks. Journal of the Association for Information Science and Technology, 68(12):2844–2845, December 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Fonseca:2021:WWQ

[Fon21] Fred Fonseca. Whether or when: The question on the use of theories in data science. Journal of the Association for Information Science and Technology, 72(12):1593–1604, December 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Fourie:2019:BRE

[Fou19] Ina Fourie. Book review: Exploring Context in Information Behavior: Seeker, Situation, Surroundings, and Shared Identities. Naresh Agarwal. Synthesis Lectures on Information Concepts, Retrieval, and Services. San Rafael, CA: Morgan and Claypool, 2018. 200 pp. $14.99 (e-book). (ISBN 978-1-68173-082-0). Journal of the Association for Information Science and Technology, 70(3):301–303, March 2019. CO-DEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Fox:2020:PMH

[Fox20] Grace Fox. “To protect my health or to protect my health privacy?” A mixed-methods investigation of the privacy paradox. Journal of the Association for Information Science and Technology, 71(9):1015–1029, September 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Fichman:2023:TCF

[FR23] Pnina Fichman and Maanvi Rathi. Trolling CNN and Fox News on Facebook, Instagram, and Twitter. Journal of the Association for Information Science and Technology, 74(5): 493–505, May 2023. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Frank:2021:BRC

[Fra21] Rebecca D. Frank. Book review: Cinderella’s stick: a fairy tale for digital preservation. Tzitzikas, Yiannis, Marketakis, Yiannis. Berlin: Springer, 2018, 249 pp., EUR 51.16 (paperback). (ISBN 978-3-319-98487-2). Journal of the Association for Information Science and Technology, 72(5):649–652, May 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Freund:2015:CIS

[Fre15] Luanne Freund. Contextualizing the information-seeking behavior of software engineers. Journal of the Association for Information Science and Technology, 66(8):1594–1605, August 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Freund:2018:BRS

[Fre18] Luanne Freund. Book review: Social Information Seeking: Leveraging the Wisdom of the Crowd. Chirag Shah. Cham, Switzerland: Springer International Publishing, 2017. 177 pp. $99.00 (Hardcover). (ISBN 978-3-319-56755-6). Journal of the Association for Information Science and Technology, 69(6):864–865, June 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Fricke:2015:BDE

[Fri15] Martin Frické. Big data and its epistemology. Journal of the Association for Information Science and Technology, 66(4): 651–661, April 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Fricke:2017:BRR

[Fri17] Martin Frické. Book review: Review of Building Ontologies with Basic Formal Ontology. Robert Arp, Barry Smith and Andrew D. Spear. Cambridge: The MIT Press, 2015. 248 pp. $30.00 (Paperback) (ISBN: 978-0-262-52781-1). Journal of the Association for Information Science and Technology, 68(3):801–804, March 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Froehlich:2021:BRA

[Fro21] Thomas J. Froehlich. Book review: Affelt, Amy. All that’s not fit to print: Fake news and the call to action for librarians and information professionals. London, UK: Emerald, 2019, 176 pp. £39.99 (paperback) (ISBN 978-1-78973-364-8). Journal of the Association for Information Science and Technology, 72(2):259–263, February 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Ferro:2018:TAI

[FS18] Nicola Ferro and Gianmaria Silvello. Toward an anatomy of IR system component performances. Journal of the Association for Information Science and Technology, 69(2):187–200, February 2018. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Ferro:2016:TMI

[FSK^{+}16] Nicola Ferro, Gianmaria Silvello, Heikki Keskustalo, Ari Pirkola, and Kalervo Järvelin. The twist measure for IR evaluation: Taking user’s effort into account. Journal of the Association for Information Science and Technology, 67(3): 620–648, March 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Fry:2016:TUR

[FSPC16] Jenny Fry, Valérie Spezi, Stephen Probets, and Claire Creaser. Towards an understanding of the relationship between disciplinary research cultures and open access repository behaviors. Journal of the Association for Information Science and Technology, 67(11):2710–2724, November 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Freund:2016:IAF

[FT16] Luanne Freund and Elaine G. Toms. Interacting with archival finding aids. Journal of the Association for Information Science and Technology, 67(4):994–1008, April 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Fafalios:2017:SRB

[FT17] Pavlos Fafalios and Yannis Tzitzikas. Stochastic reranking of biomedical search results based on extracted entities. Journal of the Association for Information Science and Technology, 68(11):2572–2586, November 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Furner:2020:DMB

[Fur20] Jonathan Furner. Definitions of “metadata”: A brief survey of international standards. Journal of the Association for Information Science and Technology, 71(6):E33–E42, June 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Fichman:2021:RBM

[FV21] Pnina Fichman and Matthew Vaughn. The relationships between misinformation and outrage trolling tactics on two Yahoo! Answers categories. Journal of the Association for Information Science and Technology, 72(12):1498–1510, December 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Fox:2022:EDA

[FvdWR^{+}22] Grace Fox, Lisa van der Werff, Pierangelo Rosati, Patricia Takako Endo, and Theo Lynn. Examining the determinants of acceptance and use of mobile contact tracing applications in Brazil: an extended privacy calculus perspective. Journal of the Association for Information Science and Technology, 73(7):944–967, July 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Ferreira:2014:STA

[FVGL14] Anderson A. Ferreira, Adriano Veloso, Marcos André Gonçalves, and Alberto H. F. Laender. Self-training author name disambiguation for information scarce scenarios. Journal of the Association for Information Science and Technology, 65(6):1257–1278, June 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Farys:2017:MCG

[FW17] Rudolf Farys and Tobias Wolbring. Matched control groups for modeling events in citation data: an illustration of Nobel Prize effects in citation networks. Journal of the Association for Information Science and Technology, 68(9):2201–2210, September 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Franssen:2019:SSO

[FW19] Thomas Franssen and Paul Wouters. Science and its significant other: Representing the humanities in bibliometric scholarship. Journal of the Association for Information Science and Technology, 70(10):1124–1137, October 2019. CO-DEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Golbeck:2018:CTU

[GAB^{+}18] Jennifer Golbeck, Brooke Auxier, Abigail Bickford, Lautaro Cabrera, Meaghan Conte McHugh, Stephani Moore, Jacquelyn Hart, Justin Resti, Anthony Rogers, and Jenna Zimmerman. Congressional Twitter use revisited on the platform’s 10-year anniversary. Journal of the Association for Information Science and Technology, 69(8):1067–1070, August 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Gonzalez-Agirre:2016:WSI

[GARA^{+}16] Aitor Gonzalez-Agirre, German Rigau, Eneko Agirre, Nikolaos Aletras, and Mark Stevenson. Why are these similar? Investigating item similarity types in a large digital library. Journal of the Association for Information Science and Technology, 67(7):1624–1638, July 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Gruss:2020:CBE

[GAS^{+}20] Richard Gruss, Alan Abrahams, Yuhyun Song, Daniel Berry, and Sultan M. Al-Daihani. Community building as an effective user engagement strategy: a case study in academic libraries. Journal of the Association for Information Science and Technology, 71(2):208–220, February 2020. CO-DEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Gauducheau:2016:ESI

[Gau16] Nadia Gauducheau. An exploratory study of the informationseeking activities of adolescents in a discussion forum. Journal of the Association for Information Science and Technology, 67(1):43–55, January 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Gorrell:2016:CTF

[GB16] Genevieve Gorrell and Kalina Bontcheva. Classifying Twitter favorites: Like, bookmark, or thanks? Journal of the Association for Information Science and Technology, 67(1): 17–25, January 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Genuis:2017:LNS

[GB17] Shelagh K. Genuis and Jenny Bronstein. Looking for “normal”: Sense making in the context of health disruption. Journal of the Association for Information Science and Technology, 68(3):750–761, March 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Girdhar:2019:CDS

[GB19a] Nancy Girdhar and K. K. Bharadwaj. Community detection in signed social networks using multiobjective genetic algorithm. Journal of the Association for Information Science and Technology, 70(8):788–804, August 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Gu:2019:DSC

[GB19b] Xin Gu and Karen L. Blackmore. Developing a scholar classification scheme from publication patterns in academic science: a cluster analysis approach. Journal of the Association for Information Science and Technology, 70(11):1262–1276, November 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Goyal:2018:DME

[GBL^{+}18] Nitesh Goyal, Marc Bron, Mounia Lalmas, Andrew Haines, and Henriette Cramer. Designing for mobile experience beyond the native ad click: Exploring landing page presentation style and media usage. Journal of the Association for Information Science and Technology, 69(7):913–923, July 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Gray:2015:IED

[GBS15] Peter Gray, Brian S. Butler, and Nikhil Sharma. The interacting effects of distributed work arrangements and individual dispositions on willingness to engage in sensemaking behaviors. Journal of the Association for Information Science and Technology, 66(10):2085–2097, October 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Guo:2023:LSL

[GBZ^{+}23] Teng Guo, Xiaomei Bai, Shihao Zhen, Shagufta Abid, and Feng Xia. Lost at starting line: Predicting maladaptation of university freshmen based on educational big data. Journal of the Association for Information Science and Technology, 74(1):17–32, January 2023. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Gantman:2018:ROI

[GD18] Ernesto R. Gantman and Marcelo P. Dabós. Research output and impact of the fields of management, economics, and sociology in Spain and France: an analysis using Google Scholar and Scopus. Journal of the Association for Information Science and Technology, 69(8):1054–1066, August 2018. CO-DEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Gao:2017:PEL

[GDO17] Ning Gao, Mark Dredze, and Douglas W. Oard. Person entity linking in email with NIL detection. Journal of the Association for Information Science and Technology, 68(10): 2412–2424, October 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Granikov:2022:FOC

[GEBP22] Vera Granikov, Reem El Sherif, France Bouthillier, and Pierre Pluye. Factors and outcomes of collaborative information seeking: a mixed studies review with a framework synthesis. Journal of the Association for Information Science and Technology, 73(4):542–560, April 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Gazni:2016:APC

[GG16] Ali Gazni and Zahra Ghaseminik. Author practices in citing other authors, institutions, and journals. Journal of the Association for Information Science and Technology, 67(10): 2536–2549, October 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Gregory:2019:SDR

[GGC^{+}19] Kathleen Gregory, Paul Groth, Helena Cousijn, Andrea Scharnhorst, and Sally Wyatt. Searching data: a review of observational data retrieval practices in selected disciplines. Journal of the Association for Information Science and Technology, 70(5):419–432, May 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Gao:2022:FFA

[GGS22] Ruoyuan Gao, Yingqiang Ge, and Chirag Shah. FAIR: Fairness-aware information retrieval evaluation. Journal of the Association for Information Science and Technology, 73 (10):1461–1473, October 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Gwizdka:2017:TDE

[GHCW17] Jacek Gwizdka, Rahilsadat Hosseini, Michael Cole, and Shouyi Wang. Temporal dynamics of eye-tracking and EEG during reading and relevance decisions. Journal of the Association for Information Science and Technology, 68(10): 2299–2312, October 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Gill:2017:IWS

[GHKB^{+}17] Alastair J. Gill, Saba Hinrichs-Krapels, Tobias Blanke, Jonathan Grant, Mark Hedges, and Simon Tanner. Insight workflow: Systematically combining human and computational methods to explore textual data. Journal of the Association for Information Science and Technology, 68(7):1671– 1686, July 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Giannella:2016:IAU

[Gia16] Chris Giannella. An improved algorithm for unsupervised decomposition of a multi-author document. Journal of the Association for Information Science and Technology, 67(2): 400–411, February 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Giannetti:2018:TCS

[Gia18] Francesca Giannetti. A Twitter case study for assessing digital sound. Journal of the Association for Information Science and Technology, 69(5):687–699, May 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Gonzalez-Ibanez:2019:CUM

[GIEVVGS19] Roberto González-Ibáñez, Aileen Esparza-Villamán, Juan Carlos Vargas-Godoy, and Chirag Shah. A comparison of unimodal and multimodal models for implicit detection of relevance in interactive IR. Journal of the Association for Information Science and Technology, 70(11):1223–1235, November 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Gonzalez-Ibanez:2015:CCM

[GISW15] Roberto González-Ibáñez, Chirag Shah, and Ryen W. White. Capturing Collabportunities: a method to evaluate collaboration opportunities in information search using pseudocollaboration. Journal of the Association for Information Science and Technology, 66(9):1897–1912, September 2015. CO-DEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Greyson:2016:RIH

[GJ16] Devon L. Greyson and Joy L. Johnson. The role of information in health behavior: a scoping study and discussion of major public health models. Journal of the Association for Information Science and Technology, 67(12):2831–2841, December 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Green:2022:CEM

[GJ22] Lucy Santos Green and Melissa P. Johnston. A contextualization of editorial misconduct in the library and information science academic information ecosystem. Journal of the Association for Information Science and Technology, 73(7): 913–928, July 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Ghosh:2014:IEC

[GK14] Jaideep Ghosh and Avinash Kshitij. An integrated examination of collaboration coauthorship networks through structural cohesion, holes, hierarchy, and percolating clusters. Journal of the Association for Information Science and Technology, 65(8):1639–1661, August 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Gok:2024:HII

[GK24] Abdullah Gök and Maria Karaulova. How “international” is international research collaboration? Journal of the Association for Information Science and Technology, 75(2):97–114, February 2024. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Gruda:2021:ACF

[GKM^{+}21] Dritjon Gruda, Dimitra Karanatsiou, Kanishka Mendhekar, Jennifer Golbeck, and Athena Vakali. I alone can fix it: Examining interactions between narcissistic leaders and anxious followers on Twitter using a machine learning approach. Journal of the Association for Information Science and Technology, 72(11):1323–1336, November 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Goh:2016:IAL

[GLR16] Dion Hoe-Lian Goh, Chei Sian Lee, and Khasfariyati Razikin. Interfaces for accessing location-based information on mobile devices: an empirical evaluation. Journal of the Association for Information Science and Technology, 67(12): 2882–2896, December 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Gu:2024:DLC

[GLZ^{+}24] Dongxiao Gu, Hu Liu, Huimin Zhao, Xuejie Yang, Min Li, and Changyong Liang. A deep learning and clustering-based topic consistency modeling framework for matching health information supply and demand. Journal of the Association for Information Science and Technology, 75(2):152–166, February 2024. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Gibson:2019:RSI

[GM19] Amelia N. Gibson and John D. Martin III. Re-situating information poverty: Information marginalization and parents of individuals with disabilities. Journal of the Association for Information Science and Technology, 70(5):476–487, May 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Gipp:2014:CBP

[GMB14] Bela Gipp, Norman Meuschke, and Corinna Breitinger. Citation-based Plagiarism Detection: Practicability on a Large-Scale Scientific Corpus. Journal of the Association for Information Science and Technology, 65(8):1527–1540, August 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Greenstein-Messica:2017:PDS

[GMRS17] Asnat Greenstein-Messica, Lior Rokach, and Asaf Shabtai. Personal-discount sensitivity prediction for mobile coupon conversion optimization. Journal of the Association for Information Science and Technology, 68(8):1940–1952, August 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Gwizdka:2019:ISI

[GMW19] Jacek Gwizdka, Yashar Moshfeghi, and Max L. Wilson. Introduction to the special issue on neuro-information science. Journal of the Association for Information Science and Technology, 70(9):911–916, September 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Gunaratne:2017:UIN

[GN17] Junius Gunaratne and Oded Nov. Using interactive “nutrition labels” for financial products to assist decision making under uncertainty. Journal of the Association for Information Science and Technology, 68(8):1836–1849, August 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Gomez-Nunez:2016:USJ

[GNVQdMA16] Antonio J. Gómez-Núñez, Benjamín Vargas-Quesada, and Félix de Moya-Anegón. Updating the SCImago journal and country rank classification: a new approach using Ward’s clustering and alternative combination of citation measures. Journal of the Association for Information Science and Technology, 67(1):178–190, January 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Godert:2016:OBM

[Göd16] Winfried Gödert. An ontology-based model for indexing and retrieval. Journal of the Association for Information Science and Technology, 67(3):594–609, March 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Graus:2018:BCM

[GOdR18] David Graus, Daan Odijk, and Maarten de Rijke. The birth of collective memories: analyzing emerging entities in text streams. Journal of the Association for Information Science and Technology, 69(6):773–786, June 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Gorichanaz:2019:IEP

[Gor19] Tim Gorichanaz. Information experience in personally meaningful activities. Journal of the Association for Information Science and Technology, 70(12):1302–1310, December 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Gorichanaz:2020:BRL

[Gor20a] Tim Gorichanaz. Book review: Lupton, Deborah. Data selves: More-than-human perspectives. New York, NY: Polity Press, 2019. 208 pp. $22.95 (paperback) (ISBN 978-1-5095-3642-9). Journal of the Association for Information Science and Technology, 71(12):1505–1506, December 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Gorichanaz:2020:UIW

[Gor20b] Tim Gorichanaz. Understanding and information in the work of visual artists. Journal of the Association for Information Science and Technology, 71(6):685–695, June 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Gorichanaz:2022:RIS

[Gor22] Tim Gorichanaz. Relating information seeking and use to intellectual humility. Journal of the Association for Information Science and Technology, 73(5):643–654, May 2022. CO-DEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Gorichanaz:2024:VSF

[Gor24] Tim Gorichanaz. Virtuous search: a framework for intellectual virtue in online search. Journal of the Association for Information Science and Technology, 75(5):538–549, May 2024. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Garechana:2015:CMW

[GRBCS15] Gaizka Garechana, Rosa Rio-Belver, Ernesto Cilleruelo, and Jaso Larruscain Sarasola. Clusterization and mapping of waste recycling science. Evolution of research from 2002 to 2012. Journal of the Association for Information Science and Technology, 66(7):1431–1446, July 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Giachanou:2021:IES

[GRC21] Anastasia Giachanou, Paolo Rosso, and Fabio Crestani. The impact of emotional signals on credibility assessment. Journal of the Association for Information Science and Technology, 72(9):1117–1132, September 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Greifeneder:2016:EDT

[Gre16] Elke Greifeneder. The effects of distraction on task completion scores in a natural environment test setting. Journal of the Association for Information Science and Technology, 67 (12):2858–2870, December 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Greyson:2018:ITC

[Gre18] Devon Greyson. Information triangulation: a complex and agentic everyday information practice. Journal of the Association for Information Science and Technology, 69(7):869– 878, July 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Greyson:2019:SII

[Gre19] Devon Greyson. The social informatics of ignorance. Journal of the Association for Information Science and Technology, 70(4):412–415, April 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Grassano:2017:FDP

[GRH^{+}17] Nicola Grassano, Daniele Rotolo, Joshua Hutton, Frédérique Lang, and Michael M. Hopkins. Funding data from publication acknowledgments: Coverage, uses, and limitations. Journal of the Association for Information Science and Technology, 68(4):999–1017, April 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Gok:2016:IRF

[GRS16] Abdullah Gök, John Rigby, and Philip Shapira. The impact of research funding on scientific outputs: Evidence from six smaller European countries. Journal of the Association for Information Science and Technology, 67(3):715–730, March 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Garcia:2015:ASR

[GRSFV15a] Jose A. García, Rosa Rodriguez-Sánchez, and Joaquín Fdez-Valdivia. Adverse selection of reviewers. Journal of the Association for Information Science and Technology, 66(6):1252– 1262, June 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Garcia:2015:BEP

[GRSFV15b] Jose A. García, Rosa Rodriguez-Sánchez, and Joaquín Fdez-Valdivia. Bias and effort in peer review. Journal of the Association for Information Science and Technology, 66(10): 2020–2030, October 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Garcia:2015:PAP

[GRSFV15c] Jose A. García, Rosa Rodriguez-Sánchez, and Joaquín Fdez-Valdivia. The principal-agent problem in peer review. Journal of the Association for Information Science and Technology, 66(2):297–308, February 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Garcia:2015:SIS

[GRSFV15d] Jose A. García, Rosa Rodriguez-Sánchez, and Joaquín Fdez-Valdivia. Social impact of scholarly articles in a citation network. Journal of the Association for Information Science and Technology, 66(1):117–127, January 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Gwak:2018:NAE

[GS18] Jae Ha Gwak and So Young Sohn. A novel approach to explore patent development paths for subfield technologies. Journal of the Association for Information Science and Technology, 69(3):410–419, March 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Ghosh:2022:AMF

[GS22] Isha Ghosh and Vivek Singh. “Not all my friends are friends”: Audience-group-based nudges for managing location privacy. Journal of the Association for Information Science and Technology, 73(6):797–810, June 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Golub:2016:AIS

[GSB^{+}16] Koraljka Golub, Dagobert Soergel, George Buchanan, Douglas Tudhope, Marianne Lykke, and Debra Hiom. Advances in information science: A framework for evaluating automatic indexing or classification in the context of retrieval. Journal of the Association for Information Science and Technology, 67(1):3–16, January 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Geras:2020:SWI

[GSG20] Agnieszka Geras, Grzegorz Siudem, and Marek Gagolewski. Should we introduce a dislike button for academic articles? Journal of the Association for Information Science and Technology, 71(2):221–229, February 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Geras:2022:TVT

[GSG22] Agnieszka Geras, Grzegorz Siudem, and Marek Gagolewski. Time to vote: Temporal clustering of user activity on Stack Overflow. Journal of the Association for Information Science and Technology, 73(12):1681–1691, December 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Gonzalez-Teruel:2022:DTT

[GTAS22] Aurora González-Teruel, Carlos-Alberto-Ávila Araújo, and Martha Sabelli. Diffusion of theories and theoretical models in the Ibero–American research on information behavior. Journal of the Association for Information Science and Technology, 73(4):561–578, April 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Golub:2014:TRK

[GTZŽ14] Koraljka Golub, Douglas Tudhope, Marcia Lei Zeng, and Maja Žumer. Terminology registries for knowledge organization systems: Functionality, use, and attributes. Journal of the Association for Information Science and Technology, 65(9):1901–1916, September 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Guimaraes:2017:BRI

[Gui17] José Augusto Guimarães. Book reviews: Interdisciplinary Knowledge Organization. Rick Szostak, Claudio Gnoli, and María López–Huertas. Cham, Switzerland: Springer. 2016. 227 pp. $129.00 (Hardcover) (ISBN 978-3-319-30148-8). Journal of the Association for Information Science and Technology, 68(9):2287–2289, September 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Gooch:2017:IIC

[GVB17] Daniel Gooch, Asimina Vasalou, and Laura Benton. Impact in interdisciplinary and cross-sector research: Opportunities and challenges. Journal of the Association for Information Science and Technology, 68(2):378–391, February 2017. CO-DEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Granitzer:2024:IDO

[GVF^{+}24] Michael Granitzer, Stefan Voigt, Noor Afshan Fathima, Martin Golasowski, Christian Guetl, Tobias Hecking, Gijs Hendriksen, Djoerd Hiemstra, Jan Martinovic, Jelena Mitrović, Izidor Mlakar, Stavros Moiras, Alexander Nussbaumer, Per Öster, Martin Potthast, Marjana Sencar Srdic, Sharikadze Megi, Katerina Slaninová, Benno Stein, Arjen P. de Vries, Vít Vondrák, Andreas Wagner, and Saber Zerhoudi. Impact and development of an Open Web Index for open web search. Journal of the Association for Information Science and Technology, 75(5):512–520, May 2024. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Given:2018:ITH

[GW18] Lisa M. Given and Rebekah Willson. Information technology and the humanities scholar: Documenting digital research practices. Journal of the Association for Information Science and Technology, 69(6):807–819, June 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Goldberg:2022:MLC

[GZBA22] David M. Goldberg, Nohel Zaman, Arin Brahma, and Mariano Aloiso. Are mortgage loan closing delay risks predictable? A predictive analysis using text mining on discussion threads. Journal of the Association for Information Science and Technology, 73(3):419–437, March 2022. CO-DEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Haider:2017:DTS

[H17] Jutta Haider and Fredrik ström. Dimensions of trust in scholarly communication: Problematizing peer review in the aftermath of John Bohannon’s “Sting” in science. Journal of the Association for Information Science and Technology, 68 (2):450–467, February 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Herceg:2018:CES

[HABT18] Paul M. Herceg, Timothy B. Allison, Robert S. Belvin, and Evelyne Tzoukermann. Collaborative exploratory search for information filtering and large-scale information triage. Journal of the Association for Information Science and Technology, 69(3):395–409, March 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Haigh:2016:BRT

[Hai16] Thomas Haigh. Book reviews: Turing: Pioneer of the Information Age. Jack Copeland. Oxford: Oxford University Press, 2013. 224 pp. $17.95 (paperback). (ISBN 978-0-19-871918-2). Journal of the Association for Information Science and Technology, 67(7):1787–1789, July 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Huvila:2017:BOI

[HAJ^{+}17] Isto Huvila, Theresa Dirndorfer Anderson, Eva Hourihan Jansen, Pam McKenzie, and Adam Worrall. Boundary objects in information science. Journal of the Association for Information Science and Technology, 68(8):1807–1822, August 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Hajibayova:2020:TCS

[Haj20] Lala Hajibayova. (un)theorizing citizen science: Investigation of theories applied to citizen science studies. Journal of the Association for Information Science and Technology, 71 (8):916–926, August 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Hakken:2014:BRE

[Hak14] David Hakken. Book reviews: Ethnography and Virtual Worlds: A Handbook of Methods — by Tom Boellstorff, Bonnie Nardi, Celia Pearce and T. L. Taylor. Journal of the Association for Information Science and Technology, 65(10): 2155–2157, October 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Haley:2014:BCR

[Hal14] M. Ryan Haley. Brief communication: Ranking top economics and finance journals using Microsoft Academic search versus Google Scholar: How does the new publish or perish option compare? Journal of the Association for Information Science and Technology, 65(5):1079–1084, May 2014. CO-DEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Haley:2017:NDA

[Hal17] M. Ryan Haley. On the normalization and distributional adjustment of journal ranking metrics: a simple parametric approach. Journal of the Association for Information Science and Technology, 68(6):1590–1593, June 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Halavais:2019:BRQ

[Hal19] Alexander Halavais. Book review: The Qualified Self: Social Media and the Accounting of Everyday Life. Humphreys, Lee. Cambridge, MA: The MIT Press, 2018. 200 pp. $28.00 (Hardcover). (ISBN 978-0-262-03785-3). Journal of the Association for Information Science and Technology, 70(9): 1043–1044, September 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Haley:2020:SPA

[Hal20] M. Ryan Haley. A simple paradigm for augmenting the Euclidean index to reflect journal impact and visibility. Journal of the Association for Information Science and Technology, 71(3):370–373, March 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Hansson:2021:BRL

[Han21] Joacim Hansson. Book review: Libraries, Archives and Museums as Democratic Spaces in a Digital Age. Audunson Ragnar, Andresen Herbjørn, Fagerlid Cicilie, Henningsen Erik, Hobohm Hans-Christoph, Jochumsen Henrik, Larsen Håkon, Vold Tonje (Eds.). Berlin: Walter de Gruyter GmbH, 2020. 370 pp. EUR 99.95 (hardcover). (ISBN 978-3-11-062954-5). Journal of the Association for Information Science and Technology, 72(12):1605–1608, December 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Hartel:2014:AIS

[Har14] Jenna Hartel. An arts-informed study of information using the draw-and-write technique. Journal of the Association for Information Science and Technology, 65(7):1349–1367, July 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Hawamdeh:2018:BRK

[Haw18] Suliman Hawamdeh. Book review: Knowledge Management Theory and Practice. (3rd ed.) Kimiz Dalkir. Cambridge, MA: The MIT Press, 2017. 552 pp. $65.00 (hardcover). (ISBN 978-0-262-03687-0). Journal of the Association for Information Science and Technology, 69(12):1521–1522, December 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Haythornthwaite:2018:BRS

[Hay18] Caroline Haythornthwaite. Book review: The SAGE Handbook of Social Media Research Methods. Luke Sloan and Anabel Quan-Haase (Eds.). Thousand Oaks, CA: SAGE Publishing, 2017. 728 pp. $175.00 (Hardcover). (ISBN 978-1-4739-1632-6). Journal of the Association for Information Science and Technology, 69(4):625–630, April 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Ho:2014:DAM

[HB14] Shuyuan Mary Ho and Izak Benbasat. Dyadic attribution model: a mechanism to assess trustworthiness in virtual organizations. Journal of the Association for Information Science and Technology, 65(8):1555–1576, August 2014. CO-DEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Hoenkamp:2015:HEL

[HB15] Eduard Hoenkamp and Peter Bruza. How everyday language can and will boost effective information retrieval. Journal of the Association for Information Science and Technology, 66 (8):1546–1558, August 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Howison:2016:SSL

[HB16] James Howison and Julia Bullard. Software in the scientific literature: Problems with seeing, finding, and using software mentioned in the biology literature. Journal of the Association for Information Science and Technology, 67(9): 2137–2155, September 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Hammache:2021:TPB

[HB21] Arezki Hammache and Mohand Boughanem. Term positionbased language model for information retrieval. Journal of the Association for Information Science and Technology, 72 (5):627–642, May 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Huang:2019:ZOP

[HBDL19] Yong Huang, Yi Bu, Ying Ding, and Wei Lu. From zero to one: a perspective on citing. Journal of the Association for Information Science and Technology, 70(10):1098–1107, October 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Huang:2019:OCV

[HBG19] Jianxiong Huang, Wai Fong Boh, and Kim Huat Goh. Opinion convergence versus polarization: examining opinion distributions in online word-of-mouth. Journal of the Association for Information Science and Technology, 70(11):1183– 1193, November 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Haustein:2016:BCT

[HBH^{+}16] Stefanie Haustein, Timothy D. Bowman, Kim Holmberg, Andrew Tsou, Cassidy R. Sugimoto, and Vincent Larivière. Brief communications: Tweets as impact indicators: Examining the implications of automated “bot” accounts on Twitter. Journal of the Association for Information Science and Technology, 67(1):232–238, January 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Huang:2015:HIP

[HBS15] Sheng-Cheng Huang, Randolph G. Bias, and David Schnyer. How are icons processed by the brain? Neuroimaging measures of four types of visual stimuli used in information systems. Journal of the Association for Information Science and Technology, 66(4):702–720, April 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Huang:2015:IBE

[HCC15] Hong Huang, Samuel Kai-Wah Chu, and Dora Yu-Ting Chen. Interactions between English-speaking and Chinesespeaking users and librarians on social networking sites. Journal of the Association for Information Science and Technology, 66(6):1150–1166, June 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Haimson:2021:MLE

[HCC^{+}21] Oliver L. Haimson, Albert J. Carter, Shanley Corvite, Brookelyn Wheeler, Lingbo Wang, Tianxiao Liu, and Alexxus Lige. The major life events taxonomy: Social readjustment, social media information sharing, and online network separation during times of life transition. Journal of the Association for Information Science and Technology, 72(7): 933–947, July 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Hu:2017:FEM

[HCD17] Xiao Hu, Kahyun Choi, and J. Stephen Downie. A framework for evaluating multimodal music mood classification. Journal of the Association for Information Science and Technology, 68(2):273–285, February 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Huvila:2015:PPT

[HCDÅ15] Isto Huvila, Åsa Cajander, Mats Daniels, and Rose-Mharie Åhlfeldt. Patients’ perceptions of their medical records from different subject positions. Journal of the Association for Information Science and Technology, 66(12):2456–2470, December 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Hartley:2015:BCR

[HCKH15] James Hartley, Guillaume Cabanac, Marcin Kozak, and Gilles Hubert. Brief communication: Research on tables and graphs in academic articles: Pitfalls and promises. Journal of the Association for Information Science and Technology, 66(2):428–431, February 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Huang:2021:RDM

[HCS21] Yingshen Huang, Andrew M. Cox, and Laura Sbaffi. Research data management policy and practice in Chinese university libraries. Journal of the Association for Information Science and Technology, 72(4):493–506, April 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Hansson:2022:ORD

[HD22] Karin Hansson and Anna Dahlgren. Open research data repositories: Practices, norms, and metadata for sharing images. Journal of the Association for Information Science and Technology, 73(2):303–316, February 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Haines:2023:CUI

[HDT23] Jelina Haines, Jia Tina Du, and Aunty Ellen Trevorrow. Cultural use of ICT4D to promote indigenous knowledge continuity of Ngarrindjeri stories and communal practices. Journal of the Association for Information Science and Technology, 74(12):1449–1462, December 2023. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Hu:2015:LLA

[HDZ^{+}15] Beibei Hu, Xianlei Dong, Chenwei Zhang, Timothy D. Bowman, Ying Ding, Staša Milojević, Chaoqun Ni, Erjia Yan, and Vincent Larivière. A lead-lag analysis of the topic evolution patterns for preprints and publications. Journal of the Association for Information Science and Technology, 66(12): 2643–2656, December 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

He:2018:BRM

[He18] Daqing He. Book review: Mobile Information Retrieval, Fabio Crestani, Stefano Mizzaro, and Ivan Scagnetto. Cham, Switzerland: Springer, 2017. 110 pp. $54.99 (softcover). (ISBN 978-3-319-60776-4). Journal of the Association for Information Science and Technology, 69(10):1283–1287, October 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Hasanain:2022:SEW

[HE22] Maram Hasanain and Tamer Elsayed. Studying effectiveness of Web search for fact checking. Journal of the Association for Information Science and Technology, 73(5):738–751, May 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Huvila:2022:IBP

[HEEB^{+}22] Isto Huvila, Heidi Enwald, Kristina Eriksson-Backa, Ying-Hsang Liu, and Noora Hirvonen. Information behavior and practices research informing information systems design. Journal of the Association for Information Science and Technology, 73(7):1043–1057, July 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Heneberg:2014:PWC

[Hen14] Petr Heneberg. Parallel worlds of citable documents and others: Inflated commissioned opinion articles enhance scientometric indicators. Journal of the Association for Information Science and Technology, 65(3):635–643, March 2014. CO-DEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Henriksen:2019:ACA

[Hen19] Dorte Henriksen. Alphabetic or contributor author order. What is the norm in Danish economics and political science and why? Journal of the Association for Information Science and Technology, 70(6):607–618, June 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Hertzum:2023:ISE

[Her23] Morten Hertzum. Information seeking by experimentation: Trying something out to discover what happens. Journal of the Association for Information Science and Technology, 74(4):383–387, April 2023. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Hernandez-Garcia:2016:SIM

[HGCKDR16] Yoscelina I. Hernandez-Garcia, José Antonio Chamizo, Mina Kleiche-Dray, and Jane M. Russell. The scientific impact of Mexican steroid research 1935–1965: a bibliometric and historiographic analysis. Journal of the Association for Information Science and Technology, 67(5):1245–1256, May 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Harvey:2014:PSI

[HH14] Melissa J. Harvey and Michael G. Harvey. Privacy and security issues for mobile health platforms. Journal of the Association for Information Science and Technology, 65(7): 1305–1318, July 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Hammarfelt:2018:CMV

[HH18] Björn Hammarfelt and Gaby Haddow. Conflicting measures and values: How humanities scholars in Australia and Sweden use and react to bibliometric indicators. Journal of the Association for Information Science and Technology, 69(7): 924–935, July 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Haddow:2019:QIQ

[HH19] Gaby Haddow and Björn Hammarfelt. Quality, impact, and quantification: Indicators and metrics use by social scientists. Journal of the Association for Information Science and Technology, 70(1):16–26, January 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Ho:2017:EDD

[HHB17] Shuyuan Mary Ho, Jeffrey T. Hancock, and Cheryl Booth. Ethical dilemma: Deception dynamics in computermediated group communication. Journal of the Association for Information Science and Technology, 68(12):2729–2742, December 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Huang:2014:GSP

[HHC^{+}14] Mu-Hsuan Huang, Wei-Tzu Huang, Cheng-Ching Chang, Dar-Zen Chen, and Chang-Pin Lin. The greater scattering phenomenon beyond Bradford’s law in patent citation. Journal of the Association for Information Science and Technology, 65(9):1917–1928, September 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Hiniker:2017:TOV

[HHK^{+}17] Alexis Hiniker, Sungsoo (Ray) Hong, Yea-Seul Kim, Nan-Chen Chen, Jevin D. West, and Cecilia Aragon. Toward the operationalization of visual metaphor. Journal of the Association for Information Science and Technology, 68(10): 2338–2349, October 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Hemphill:2021:SSM

[HHL21] Libby Hemphill, Margaret L. Hedstrom, and Susan Hautaniemi Leonard. Saving social media data: Understanding data management practices among social media researchers and their implications for archives. Journal of the Association for Information Science and Technology, 72(1):97–109, January 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Hyatt:2021:WWI

[HHPI21] Ed Hyatt, Morgan Harvey, Matthew Pointon, and Perla Innocenti. Whither wilderness? An investigation of technology use by long-distance backpackers. Journal of the Association for Information Science and Technology, 72(6):683–698, June 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Hoeber:2022:ESI

[HHSP22] Orland Hoeber, Morgan Harvey, Shaheed Ahmed Dewan Sagar, and Matthew Pointon. The effects of simulated interruptions on mobile search tasks. Journal of the Association for Information Science and Technology, 73(6):777–796, June 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Hilbert:2014:BCH

[Hil14a] Martin Hilbert. Brief communication: How much of the global information and communication explosion is driven by more, and how much by better technology? Journal of the Association for Information Science and Technology, 65 (4):856–861, April 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Hilbert:2014:TII

[Hil14b] Martin Hilbert. Technological information inequality as an incessantly moving target: the redistribution of information and communication capacities between 1986 and 2010. Journal of the Association for Information Science and Technology, 65(4):821–835, April 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Hjorland:2015:RTC

[Hjø15a] Birger Hjørland. Are relations in thesauri “context-free, definitional, and true in all possible worlds”? Journal of the Association for Information Science and Technology, 66(7): 1367–1373, July 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Hjorland:2015:BCP

[Hjø15b] Birger Hjørland. Brief communications: The phrase “information storage and retrieval” (IS&R): an historical note. Journal of the Association for Information Science and Technology, 66(6):1299–1302, June 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Hjorland:2015:CDK

[Hjø15c] Birger Hjørland. Classical databases and knowledge organization: a case for Boolean retrieval and human decisionmaking during searches. Journal of the Association for Information Science and Technology, 66(8):1559–1575, August 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Hjorland:2017:BRT

[Hjø17a] Birger Hjørland. Book review: Theory Development in the Information Sciences. Diane H. Sonnenwald. Austin, TX: University of Texas Press, 2016. 343 pp. $90.00 (Hardcover) (ISBN 978-1-4773-0824-0). Journal of the Association for Information Science and Technology, 68(7):1796–1801, July 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Hjorland:2017:LED

[Hjø17b] Birger Hjørland. Letters to the Editor: Does informetrics need a theory? A rejoinder to Professor Anthony van Raan. Journal of the Association for Information Science and Technology, 68(12):2846, December 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Ho:2014:BSH

[HK14] Yuh-Shan Ho and Michael Kahn. A bibliometric study of highly cited reviews in the Science Citation Index Expanded^{T M}. Journal of the Association for Information Science and Technology, 65(2):372–385, February 2014. CO-DEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Hu:2017:TCD

[HK17] Xiao Hu and Noriko Kando. Task complexity and difficulty in music information retrieval. Journal of the Association for Information Science and Technology, 68(7):1711–1723, July 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Huttunen:2021:MMG

[HK21] Aira Huttunen and Terttu Kortelainen. Meaning-making on gender: Deeply meaningful information in a significant life change among transgender people. Journal of the Association for Information Science and Technology, 72(7):799–810, July 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Ho:2018:TAI

[HKBB18] Shuyuan Mary Ho, Michelle Kaarst-Brown, and Izak Benbasat. Trustworthiness attribution: Inquiry into insider threat detection. Journal of the Association for Information Science and Technology, 69(2):271–280, February 2018. CO-DEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Hellsten:2016:CID

[HL16] Iina Hellsten and Loet Leydesdorff. The construction of interdisciplinarity: the development of the knowledge base and programmatic focus of the journal Climatic Change, 1977– 2013. Journal of the Association for Information Science and Technology, 67(9):2181–2193, September 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Hellsten:2020:AAA

[HL20] Iina Hellsten and Loet Leydesdorff. Automated analysis of actor-topic networks on Twitter: New approaches to the analysis of socio-semantic networks. Journal of the Association for Information Science and Technology, 71(1):3–15, January 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Hu:2017:MGC

[HLB^{+}17] Xiao Hu, Jin Ha Lee, David Bainbridge, Kahyun Choi, Peter Organisciak, and J. Stephen Downie. The MIREX grand challenge: a framework of holistic user-experience evaluation in music information retrieval. Journal of the Association for Information Science and Technology, 68(1):97–112, January 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Hung:2015:KSI

[HLC15] Shin-Yuan Hung, Hui-Min Lai, and Yu-Che Chou. Knowledgesharing intention in professional virtual communities: a comparison between posters and lurkers. Journal of the Association for Information Science and Technology, 66(12): 2494–2510, December 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Hoffmann:2016:RAN

[HLM16] Christian Pieter Hoffmann, Christoph Lutz, and Miriam Meckel. A relational altmetric? Network centrality on ResearchGate as an indicator of scientific impact. Journal of the Association for Information Science and Technology, 67 (4):765–775, April 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

He:2018:MCD

[HLW18] Zhongyang He, Zhen Lei, and Dashun Wang. Modeling citation dynamics of “atypical” articles. Journal of the Association for Information Science and Technology, 69(9):1148– 1160, September 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic). URL https://asistdl.

onlinelibrary.wiley.com/doi/10.1002/asi.24041.

Hardre:2014:EJT

[HM14] Patricia Hardré and Chad Mortensen. Education journals: Two decades of change and implications for the field. Journal of the Association for Information Science and Technology, 65(1):188–200, January 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Haley:2018:PPM

[HM18] M. Ryan Haley and M. Kevin McGee. A parametric “parent metric” approach for comparing maximum-normalized journal ranking metrics. Journal of the Association for Information Science and Technology, 69(1):172–176, January 2018. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Hu:2018:UCE

[HNX18] Xiao Hu, Jeremy Ng, and Shengping Xia. User-centered evaluation of metadata schema for nonmovable cultural heritage: Murals and stone cave temples. Journal of the Association for Information Science and Technology, 69(12):1476–1487, December 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Hodges:2021:FRB

[Hod21] James A. Hodges. Forensically reconstructing biomedical maintenance labor: PDF metadata under the epistemic conditions of COVID-19. Journal of the Association for Information Science and Technology, 72(11):1400–1414, November 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Hoffmann:2017:BDP

[Hof17] Anna Lauren Hoffmann. Beyond distributions and primary goods: Assessing applications of Rawls in information science and technology literature since 1990. Journal of the Association for Information Science and Technology, 68(7): 1601–1618, July 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Holmberg:2015:BRD

[Hol15] Kim Holmberg. Book reviews: Digital Methods — By Richard Rogers. Journal of the Association for Information Science and Technology, 66(12):2721–2722, December 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Hook:2017:UCS

[Hoo17] Peter A. Hook. Using Course-Subject Co-Occurrence (CSCO) to reveal the structure of an academic discipline: a framework to evaluate different inputs of a domain map. Journal of the Association for Information Science and Technology, 68(1):182–196, January 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Howarth:2017:BRM

[How17] Lynne C. Howarth. Book reviews: Metadata. Jeffrey Pomerantz. Cambridge, MA: MIT Press, 2015. 256 pp. $15.95 (paperback). (ISBN 978-0-262-52851-1). Journal of the Association for Information Science and Technology, 68(9):2271– 2274, September 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Hemphill:2022:HDP

[HPL^{+}22] Libby Hemphill, Amy Pienta, Sara Lafia, Dharma Akmon, and David A. Bleckley. How do properties of data, their curation, and their funding relate to reuse? Journal of the Association for Information Science and Technology, 73(10): 1432–1444, October 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

He:2019:PFA

[HPLC19] Jiangen He, Qing Ping, Wen Lou, and Chaomei Chen. PaperPoles: Facilitating adaptive visual exploration of scientific publications by citation links. Journal of the Association for Information Science and Technology, 70(8):843–857, August 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Hagen:2023:HSD

[HPLR23] Loni Hagen, Mihir Patel, and Luis Luna-Reyes. Humansupervised data science framework for city governments: a design science approach. Journal of the Association for Information Science and Technology, 74(8):923–936, August 2023. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Haustein:2014:TBA

[HPS^{+}14] Stefanie Haustein, Isabella Peters, Cassidy R. Sugimoto, Mike Thelwall, and Vincent Larivière. Tweeting biomedicine: an analysis of tweets and citations in the biomedical literature. Journal of the Association for Information Science and Technology, 65(4):656–669, April 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Huang:2022:DRB

[HQH^{+}22] Shengzhi Huang, Jiajia Qian, Yong Huang, Wei Lu, Yi Bu, Jinqing Yang, and Qikai Cheng. Disclosing the relationship between citation structure and future impact of a publication. Journal of the Association for Information Science and Technology, 73(7):1025–1042, July 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Hu:2019:DCC

[HR19] Xiaojun Hu and Ronald Rousseau. Do citation chimeras exist? The case of under-cited influential articles suffering delayed recognition. Journal of the Association for Information Science and Technology, 70(5):499–508, May 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Hasler:2014:UIG

[HRB14] Laura Hasler, Ian Ruthven, and Steven Buchanan. Using Internet groups in situations of information poverty: Topics and information needs. Journal of the Association for Information Science and Technology, 65(1):25–36, January 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Hottenrott:2021:RMI

[HRL21] Hanna Hottenrott, Michael E. Rose, and Cornelia Lawson. The rise of multiple institutional affiliations in academia. Journal of the Association for Information Science and Technology, 72(8):1039–1058, August 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Han:2023:ELM

[HRN^{+}23] Kanyao Han, Rezvaneh Rezapour, Katia Nakamura, Dikshya Devkota, Daniel C. Miller, and Jana Diesner. An expert-in-the-loop method for domain-specific document categorization based on small training data. Journal of the Association for Information Science and Technology, 74(6):669–684, June 2023. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Hills:2014:NAO

[HS14] Thomas Hills and Elad Segev. The news is American but our memories are … Chinese? Journal of the Association for Information Science and Technology, 65(9):1810–1819, September 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Hara:2017:ARE

[HS17] Noriko Hara and Madelyn Rose Sanfilippo. Analysis of roles in engaging contentious online discussions in science. Journal of the Association for Information Science and Technology, 68(8):1953–1966, August 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Hoeber:2022:SVL

[HS22] Orland Hoeber and Soumya Shukla. A study of visually linked keywords to support exploratory browsing in academic search. Journal of the Association for Information Science and Technology, 73(8):1171–1191, August 2022. CO-DEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Hsieh:2017:APP

[Hsi17] Ji-Lung Hsieh. Author publication preferences and journal competition. Journal of the Association for Information Science and Technology, 68(2):365–377, February 2017. CO-DEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Huang:2015:MAT

[HSK15] Xiaoli Huang, Dagobert Soergel, and Judith L. Klavans. Modeling and analyzing the topicality of art images. Journal of the Association for Information Science and Technology, 66(8):1616–1644, August 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Haussler:2024:GRS

[HSL24] Helena Häußler, Sebastian Schultheiß, and Dirk Lewandowski. Is googling risky? A study on risk perception and experiences of adverse consequences in web search. Journal of the Association for Information Science and Technology, 75(5): 567–580, May 2024. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Heinstrom:2020:DEM

[HSSE20] Jannica Heinström, Eero Sormunen, Reijo Savolainen, and Stefan Ek. Developing an empirical measure of everyday information mastering. Journal of the Association for Information Science and Technology, 71(7):729–741, July 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Hamari:2016:SEW

[HSU16] Juho Hamari, Mimmi Sjöklint, and Antti Ukkonen. The sharing economy: Why people participate in collaborative consumption. Journal of the Association for Information Science and Technology, 67(9):2047–2059, September 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Hemminger:2014:AWA

[HT14] Bradley M. Hemminger and Julia TerMaat. Annotating for the world: Attitudes toward sharing scholarly annotations. Journal of the Association for Information Science and Technology, 65(11):2278–2292, November 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

He:2017:LSU

[HT17] Wu He and Xin Tian. A longitudinal study of user queries and browsing requests in a case-based reasoning retrieval system. Journal of the Association for Information Science and Technology, 68(5):1124–1136, May 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Huvila:2020:BRI

[Huv20] Isto Huvila. Book review: Information at Work: Information Management in the Workplace. Edited by Katriina Byström, Jannica Heinström, and Ian Ruthven. London: Facet Publishing, 2019. 200 pp. £69.95 (paperback). (ISBN 978-1-78330-275-8). Journal of the Association for Information Science and Technology, 71(4):497–499, April 2020. CO-DEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Huvila:2022:BRF

[Huv22a] Isto Huvila. Book review: The Filing Cabinet: a Vertical History of Information. Craig Robertson. Minneapolis, MN: University of Minnesota Press, 2021. 312 pp. $34.95 (paperback). (ISBN 978-1-5179-0946-8). Journal of the Association for Information Science and Technology, 73(8):1196–1198, August 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Huvila:2022:MTI

[Huv22b] Isto Huvila. Making and taking information. Journal of the Association for Information Science and Technology, 73(4): 528–541, April 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Huang:2017:IOA

[HWC17] Hen-Hsen Huang, Jun-Jie Wang, and Hsin-Hsi Chen. Implicit opinion analysis: Extraction and polarity labelling. Journal of the Association for Information Science and Technology, 68(9):2076–2087, September 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Huang:2015:SRC

[HWW15] Mu-Hsuan Huang, Ling-Ling Wu, and Yi-Chen Wu. A study of research collaboration in the pre-web and post-web stages: a coauthorship analysis of the information systems discipline. Journal of the Association for Information Science and Technology, 66(4):778–797, April 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Haraty:2017:DSE

[HWW^{+}17] Mona Haraty, Zhongyuan Wang, Helen Wang, Shamsi Iqbal, and Jaime Teevan. Design and in-situ evaluation of a mixedinitiative approach to information organization. Journal of the Association for Information Science and Technology, 68 (9):2211–2224, September 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

He:2020:RLF

[HWZ20] Chaocheng He, Jiang Wu, and Qingpeng Zhang. Research leadership flow determinants and the role of proximity in research collaborations. Journal of the Association for Information Science and Technology, 71(11):1341–1356, November 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

He:2022:PAR

[HWZ22] Chaocheng He, Jiang Wu, and Qingpeng Zhang. Proximityaware research leadership recommendation in research collaboration via deep neural networks. Journal of the Association for Information Science and Technology, 73(1):70–89, January 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Hu:2017:MCP

[HY17] Xiao Hu and Yi-Hsuan Yang. The mood of Chinese pop music: Representation and recognition. Journal of the Association for Information Science and Technology, 68(8):1899– 1910, August 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Hong:2020:CCC

[HYD^{+}20] Hong Hong, Qiang Ye, Qianzhou Du, G. Alan Wang, and Weiguo Fan. Crowd characteristics and crowd wisdom: Evidence from an online investment community. Journal of the Association for Information Science and Technology, 71(4): 423–435, April 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Ibanez:2016:GAG

[IABL16] Alfonso Ibáñez, Rubén Armañanzas, Concha Bielza, and Pedro Larrañaga. Genetic algorithms and Gaussian Bayesian networks to uncover the predictive core set of bibliometric indices. Journal of the Association for Information Science and Technology, 67(7):1703–1721, July 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Isah:2016:PLW

[IB16] Esther Ebole Isah and Katriina Byström. Physicians’ learning at work through everyday access to information. Journal of the Association for Information Science and Technology, 67(2):318–332, February 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Ferrer-i-Cancho:2018:OZM

[iCV18] Ramon Ferrer i Cancho and Michael S. Vitevitch. The origins of Zipf’s meaning-frequency law. Journal of the Association for Information Science and Technology, 69(11):1369–1379, November 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Introne:2020:DSO

[IESG20] Joshua Introne, Ingrid Erickson, Bryan Semaan, and Sean Goggins. Designing sustainable online support: Examining the effects of design change in 49 online health support communities. Journal of the Association for Information Science and Technology, 71(4):379–394, April 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Iorio:2014:WAW

[Ior14] L. Iorio. Withdrawn article: Withdrawal: ‘A new type of misconduct in the field of the physical sciences: the case of the pseudonyms used by I. Ciufolini to anonymously criticize other people’s works on arXiv’ by L. Iorio. Journal of the Association for Information Science and Technology, 65 (11):2375, November 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Iandoli:2016:OCC

[IQDS16] Luca Iandoli, Ivana Quinto, Anna De Liddo, and Simon Buckingham Shum. On online collaboration and construction of shared knowledge: Assessing mediation capability in computer supported argument visualization tools. Journal of the Association for Information Science and Technology, 67(5):1052–1067, May 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Ikae:2022:GIT

[IS22] Catherine Ikae and Jacques Savoy. Gender identification on Twitter. Journal of the Association for Information Science and Technology, 73(1):58–69, January 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Jemielniak:2016:OBG

[JA16] Dariusz Jemielniak and Eduard Aibar. Opinion: Bridging the gap between Wikipedia and academia. Journal of the Association for Information Science and Technology, 67(7): 1773–1776, July 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Jones:2020:WBT

[JAG^{+}20] Kyle M. L. Jones, Andrew Asher, Abigail Goben, Michael R. Perry, Dorothea Salo, Kristin A. Briney, and M. Brooke Robertshaw. “We’re being tracked at all times”: Student perspectives of their privacy in relation to learning analytics in higher education. Journal of the Association for Information Science and Technology, 71(9):1044–1059, September 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

James:2017:FPF

[Jam17] Jack E. James. Free-to-publish, free-to-read, or both? Cost, equality of access, and integrity in science publishing. Journal of the Association for Information Science and Technology, 68(6):1584–1589, June 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

James:2020:POA

[Jam20] Jack E. James. Pirate open access as electronic civil disobedience: Is it ethical to breach the paywalls of monetized academic publishing? Journal of the Association for Information Science and Technology, 71(12):1500–1504, December 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Jana:2015:BCS

[Jan15] Siladitya Jana. Brief communication: Sister Nivedita’s influence on J. C. Bose’s writings. Journal of the Association for Information Science and Technology, 66(3):645–650, March 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Jacucci:2019:INR

[JBD^{+}19] Giulio Jacucci, Oswald Barral, Pedram Daee, Markus Wenzel, Baris Serim, Tuukka Ruotsalo, Patrik Pluchino, Jonathan Freeman, Luciano Gamberini, Samuel Kaski, and Benjamin Blankertz. Integrating neurophysiologic relevance feedback in intent modeling for information retrieval. Journal of the Association for Information Science and Technology, 70(9):917–930, September 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Jeng:2017:IEA

[JDHL17] Wei Jeng, Spencer DesAutels, Daqing He, and Lei Li. Information exchange on an academic social networking site: a multidiscipline comparison on ResearchGate Q&A. Journal of the Association for Information Science and Technology, 68(3):638–652, March 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Jaffe:2017:PCD

[JdR17] Adam B. Jaffe and Gaétan de Rassenfosse. Patent citation data in social science research: Overview and best practices. Journal of the Association for Information Science and Technology, 68(6):1360–1374, June 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Jia:2024:IPO

[JDZ24] Romy Menghao Jia, Jia Tina Du, and Yuxiang (Chris) Zhao. Interaction with peers online: LGBTQIA+ individuals’ information seeking and meaning-making during the life transitions of identity construction. Journal of the Association for Information Science and Technology, 75(1):24–42, January 2024. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Jiang:2015:RST

[JGCR15] Shan Jiang, Qiang Gao, Hsinchun Chen, and Mihail C. Roco. The roles of sharing, transfer, and public funding in nanotechnology knowledge-diffusion networks. Journal of the Association for Information Science and Technology, 66 (5):1017–1029, May 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Jiang:2019:GTV

[JGY^{+}19] Zhiwei Jiang, Qing Gu, Yafeng Yin, Jianxiang Wang, and Daoxu Chen. GRAW+: a two-view graph propagation method with word coupling for readability assessment. Journal of the Association for Information Science and Technology, 70(5):433–447, May 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Jeng:2015:UPA

[JHJ15] Wei Jeng, Daqing He, and Jiepu Jiang. User participation in an academic social networking service: a survey of open group users on Mendeley. Journal of the Association for Information Science and Technology, 66(5):890–904, May 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Jin:2014:CIA

[JJ14] Tao Jin and Boryung Ju. The corporate information agency: Do competitive intelligence practitioners utilize It? Journal of the Association for Information Science and Technology, 65(3):589–608, March 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Jiang:2019:CHR

[JJC19] Jian-De Jiang, Jyun-Yu Jiang, and Pu-Jen Cheng. Cocluster hypothesis and ranking consistency for relevance ranking in web search. Journal of the Association for Information Science and Technology, 70(6):535–546, June 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Jarvelin:2016:IRH

[JKS^{+}16] Anni Järvelin, Heikki Keskustalo, Eero Sormunen, Miamaria Saastamoinen, and Kimmo Kettunen. Information retrieval from historical newspaper collections in highly inflectional languages: a query expansion approach. Journal of the Association for Information Science and Technology, 67(12): 2928–2946, December 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Jia:2018:DCA

[JL18] Yanli Jia and Ivy L. B. Liu. Do consumers always follow “useful” reviews? The interaction effect of review valence and review usefulness on consumers’ purchase decisions. Journal of the Association for Information Science and Technology, 69(11):1304–1317, November 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Jiang:2023:EEB

[JL23] Xiaorui Jiang and Junjun Liu. Extracting the evolutionary backbone of scientific domains: The semantic main path network analysis approach based on citation context analysis. Journal of the Association for Information Science and Technology, 74(5):546–569, May 2023. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Jarrahi:2023:AIW

[JLB^{+}23] Mohammad Hossein Jarrahi, Christoph Lutz, Karen Boyd, Carsten Oesterlund, and Matthew Willis. Artificial intelligence in the work context. Journal of the Association for Information Science and Technology, 74(3):303–310, March 2023. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Jiang:2016:RUC

[JLC16] Zhuoren Jiang, Xiaozhong Liu, and Yan Chen. Recovering uncaptured citations in a scholarly network: a two-step citation analysis to estimate publication importance. Journal of the Association for Information Science and Technology, 67(7):1722–1735, July 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Jahn:2022:TTH

[JML22] Najko Jahn, Lisa Matthias, and Mikael Laakso. Toward transparency of hybrid open access through publisherprovided metadata: an article-level study of Elsevier. Journal of the Association for Information Science and Technology, 73(1):104–118, January 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Juola:2019:CAD

[JMV19] Patrick Juola, George K. Mikros, and Sean Vinsick. A comparative assessment of the difficulty of authorship attribution in Greek and in English. Journal of the Association for Information Science and Technology, 70(1):61–70, January 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Johnson:2017:RCW

[Joh17] Nathan R. Johnson. Rhetoric and the Cold War politics of information science. Journal of the Association for Information Science and Technology, 68(6):1375–1384, June 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Jones:2017:BRS

[Jon17] William Jones. Book reviews: The Science of Managing our Digital Stuff. Ofer Bergman and Steve Whittaker. Cambridge, MA: MIT Press, 2016. 296 pp. $29.95 (Hardcover). (ISBN 978-0-262-03517-0). Journal of the Association for Information Science and Technology, 68(12):2834–2840, December 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Jordan:2019:SMP

[Jor19] Katy Jordan. Separating and merging professional and personal selves online: The structure and processes that shape academics’ ego-networks on academic social networking sites and Twitter. Journal of the Association for Information Science and Technology, 70(8):830–842, August 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Johnson:2016:BPC

[JPF16] Robert Johnson, Stephen Pinfield, and Mattia Fosci. Business process costs of implementing “gold” and “green” open access in institutional and national contexts. Journal of the Association for Information Science and Technology, 67(9): 2283–2295, September 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Jarrahi:2019:PKP

[JPS^{+}19] Mohammad Hossein Jarrahi, Gabriela Philips, Will Sutherland, Steve Sawyer, and Ingrid Erickson. Personalization of knowledge, personal knowledge ecology, and digital nomadism. Journal of the Association for Information Science and Technology, 70(4):313–324, April 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Jiao:2024:DED

[JQMY24] Hong Jiao, Yuhong Qiu, Xiaowei Ma, and Bo Yang. Dissemination effect of data papers on scientific datasets. Journal of the Association for Information Science and Technology, 75 (2):115–131, February 2024. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Jones:2020:MTH

[JRL20] Kyle M. L. Jones, Alan Rubel, and Ellen LeClere. A matter of trust: Higher education institutions as information fiduciaries in an age of educational data mining and learning analytics. Journal of the Association for Information Science and Technology, 71(10):1227–1241, October 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Johnson:2016:EII

[JRS16] Frances Johnson, Jennifer Rowley, and Laura Sbaffi. Exploring information interactions in the context of Google. Journal of the Association for Information Science and Technology, 67(4):824–840, April 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Jarrahi:2015:TTS

[JS15] Mohammad Hossein Jarrahi and Steve Sawyer. Theorizing on the take-up of social technologies, organizational policies and norms, and consultants’ knowledge-sharing practices. Journal of the Association for Information Science and Technology, 66(1):162–179, January 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Jett:2016:CMV

[JSLC16] Jacob Jett, Simone Sacchi, Jin Ha Lee, and Rachel Ivy Clarke. A conceptual model for video games and interactive media. Journal of the Association for Information Science and Technology, 67(3):505–517, March 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Jorgensen:2014:ARA

[JSW14] Corinne Jörgensen, Besiki Stvilia, and Shuheng Wu. Assessing the relationships among tag syntax, semantics, and perceived usefulness. Journal of the Association for Information Science and Technology, 65(4):836–849, April 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Jiang:2016:EHS

[JSY^{+}16] Xiaorui Jiang, Xiaoping Sun, Zhe Yang, Hai Zhuge, and Jianmin Yao. Exploiting heterogeneous scientific literature networks to combat ranking bias: Evidence from the computational linguistics area. Journal of the Association for Information Science and Technology, 67(7):1679–1702, July 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Jarrahi:2017:IBI

[JT17] Mohammad Hossein Jarrahi and Leslie Thomson. The interplay between information practices and information context: the case of mobile knowledge workers. Journal of the Association for Information Science and Technology, 68(5):1073– 1089, May 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Julien:2020:BRC

[Jul20] Heidi Julien. Book review: The Consciousness’ Drive: Information Need and the Search for Meaning. Cole, Charles. Cham, Switzerland: Springer International Publishing, 2018. 247 pp. $79.99 (hardcover). (ISBN 978-3-319-92455-7). Journal of the Association for Information Science and Technology, 71(1):118–120, January 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic). See rebuttal [Col20].

Johnson:2017:AA

[JV17] Deborah G. Johnson and Mario Verdicchio. AI anxiety. Journal of the Association for Information Science and Technology, 68(9):2267–2270, September 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Jackson:2019:IID

[JVF19] Stephen Jackson, Nadia Vanteeva, and Colm Fearon. An investigation of the impact of data breach severity on the readability of mandatory data breach notification letters: Evidence from U.S. firms. Journal of the Association for Information Science and Technology, 70(11):1277–1289, November 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Jiang:2015:PJI

[JW15] Ling Jiang and Christian Wagner. Perceptions of justice or injustice as determinants of contributor defections from online communities. Journal of the Association for Information Science and Technology, 66(7):1477–1493, July 2015. CO-DEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Jemielniak:2017:CDQ

[JW17] Dariusz Jemielniak and Maciej Wilamowski. Cultural diversity of quality of information on Wikipedias. Journal of the Association for Information Science and Technology, 68(10): 2460–2470, October 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Jones:2019:RBH

[JWJ^{+}19] Lenette M. Jones, Kathy D. Wright, Anthony I. Jack, Jared P. Friedman, David M. Fresco, Tiffany Veinot, Wei Lu, and Shirley M. Moore. The relationships between health information behavior and neural processing in African Americans with prehypertension. Journal of the Association for Information Science and Technology, 70(9):968–980, September 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Jones:2018:TOW

[JWWV18] Lenette M. Jones, Kathy D. Wright, McKenzie K. Wallace, and Tiffany Veinot. “Take an opportunity whenever you get it”: Information sharing among African–American women with hypertension. Journal of the Association for Information Science and Technology, 69(1):168–171, January 2018. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Jiang:2020:MPA

[JZC20] Xiaorui Jiang, Xinghao Zhu, and Jingqiang Chen. Main path analysis on cyclic citation networks. Journal of the Association for Information Science and Technology, 71(5): 578–595, May 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Kriesberg:2022:SUP

[KA22] Adam Kriesberg and Amelia Acker. The second US presidential social media transition: How private platforms impact the digital preservation of public records. Journal of the Association for Information Science and Technology, 73(11): 1529–1542, November 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Kozlowski:2024:DUA

[KAL24] Diego Kozlowski, Jens Peter Andersen, and Vincent Larivière. The decrease in uncited articles and its effect on the concentration of citations. Journal of the Association for Information Science and Technology, 75(2):188–197, February 2024. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Kang:2022:MAS

[Kan22] Minhyung Kang. Motivational affordances and survival of new askers on social QA sites: The case of Stack Exchange network. Journal of the Association for Information Science and Technology, 73(1):90–103, January 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Klavans:2017:WTC

[KB17] Richard Klavans and Kevin W. Boyack. Which type of citation analysis generates the most accurate taxonomy of scientific and technical knowledge? Journal of the Association for Information Science and Technology, 68(4):984–998, April 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Kipp:2017:MIF

[KBC17] Margaret E. I. Kipp, Jihee Beak, and Inkyung Choi. Motivations and intentions of Flickr users in enriching Flick records for Library of Congress photos. Journal of the Association for Information Science and Technology, 68(10):2364–2379, October 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Kajanan:2014:EAS

[KBD^{+}14] Sangaralingam Kajanan, Yang Bao, Anindya Datta, Debra VanderMeer, and Kaushik Dutta. Efficient automatic search query formulation using phrase-level analysis. Journal of the Association for Information Science and Technology, 65 (5):1058–1075, May 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Koho:2022:HPH

[KBH^{+}22] Mikko Koho, Toby Burrows, Eero Hyvönen, Esko Ikkala, Kevin Page, Lynn Ransom, Jouni Tuominen, Doug Emery, Mitch Fraas, Benjamin Heller, David Lewis, Andrew Morrison, Guillaume Porte, Emma Thomson, Athanasios Velios, and Hanno Wijsman. Harmonizing and publishing heterogeneous premodern manuscript metadata as Linked Open Data. Journal of the Association for Information Science and Technology, 73(2):240–257, February 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Kim:2018:TDA

[KBS18] Munui Kim, Injun Baek, and Min Song. Topic diffusion analysis of a weighted citation network in biomedical literature. Journal of the Association for Information Science and Technology, 69(2):329–342, February 2018. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Koya:2020:CHI

[KC20] Kushwanth Koya and Gobinda Chowdhury. Cultural heritage information practices and iSchools education for achieving sustainable development. Journal of the Association for Information Science and Technology, 71(6):696–710, June 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Kobsa:2016:EPP

[KCK16] Alfred Kobsa, Hichang Cho, and Bart P. Knijnenburg. The effect of personalization provider characteristics on privacy attitudes and behaviors: an Elaboration Likelihood Model approach. Journal of the Association for Information Science and Technology, 67(11):2587–2606, November 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Kim:2014:DFI

[KCL14] Daejoong Kim, Heasun Chun, and Hyunjoo Lee. Determining the factors that influence college students’ adoption of smartphones. Journal of the Association for Information Science and Technology, 65(3):578–588, March 2014. CO-DEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Ku:2014:TMS

[KCZ^{+}14] Yungchang Ku, Chaochang Chiu, Yulei Zhang, Hsinchun Chen, and Handsome Su. Text mining self-disclosing health information for public health service. Journal of the Association for Information Science and Technology, 65(5):928–947, May 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Kim:2015:CND

[KD15] Jinseok Kim and Jana Diesner. Coauthorship networks: a directed network approach considering the order and number of coauthors. Journal of the Association for Information Science and Technology, 66(12):2685–2696, December 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Kim:2016:DEI

[KD16] Jinseok Kim and Jana Diesner. Distortive effects of initialbased name disambiguation on measurements of large-scale coauthorship networks. Journal of the Association for Information Science and Technology, 67(6):1446–1461, June 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Kholghi:2017:CIE

[KDS^{+}17] Mahnoosh Kholghi, Lance De Vine, Laurianne Sitbon, Guido Zuccon, and Anthony Nguyen. Clinical information extraction using small data: an active learning approach based on sequence representations and word embeddings. Journal of the Association for Information Science and Technology, 68(11):2543–2556, November 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Kearns:2022:BRI

[Kea22] Jodi Kearns. Book review: Information: a Historical Companion. Ann Blair, Paul Duguid, Anja-Silvia Goeing, and Anthony Grafton Princeton: Princeton University Press, 2021. 904 pp. $65.00 (hardcover). (ISBN: 978-0-691-17954-4). Journal of the Association for Information Science and Technology, 73(3):485–488, March 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Kelly:2021:EEB

[Kel21] Matthew Kelly. Epistemology, epistemic belief, personal epistemology, and epistemics: a review of concepts as they impact information behavior research. Journal of the Association for Information Science and Technology, 72(4):507– 519, April 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Kanan:2016:AAT

[KF16] Tarek Kanan and Edward A. Fox. Automated Arabic text classification with P-Stemmer, machine learning, and a tailored news article taxonomy. Journal of the Association for Information Science and Technology, 67(11):2667–2683, November 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Kulczycki:2020:MPS

[KGP^{+}20] Emanuel Kulczycki, Raf Guns, Janne Pölönen, Tim C. E. Engels, Ewa A. Rozkosz, Alesia A. Zuccala, Kasper Bruun, Olli Eskola, Andreja Istenic Starcic, Michal Petr, and Gunnar Sivertsen. Multilingual publishing in the social sciences and humanities: a seven-country European study. Journal of the Association for Information Science and Technology, 71(11):1371–1385, November 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Karaulova:2019:IAH

[KGS19] Maria Karaulova, Abdullah Gök, and Philip Shapira. Identifying author heritage using surname data: an application for Russian surnames. Journal of the Association for Information Science and Technology, 70(5):488–498, May 2019. CO-DEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Kurtz:2017:MMY

[KH17] Michael J. Kurtz and Edwin A. Henneken. Measuring metrics — a 40-year longitudinal cross-validation of citations, downloads, and peer review in astrophysics. Journal of the Association for Information Science and Technology, 68(3): 695–708, March 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Khazraee:2019:ANT

[Kha19] Emad Khazraee. Assembling narratives: Tensions in collaborative construction of knowledge. Journal of the Association for Information Science and Technology, 70(4):325–337, April 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Kulczycki:2022:UJI

[KHZ^{+}22] Emanuel Kulczycki, Ying Huang, Alesia A. Zuccala, Tim C. E. Engels, Antonio Ferrara, Raf Guns, Janne Pölönen, Gunnar Sivertsen, Zehra Taskin, and Lin Zhang. Uses of the journal impact factor in national journal rankings in China and Europe. Journal of the Association for Information Science and Technology, 73(12):1741–1754, December 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Kozak:2016:SSP

[KIH16] Marcin Kozak, Olesia Iefremova, and James Hartley. Spamming in scholarly publishing: a case study. Journal of the Association for Information Science and Technology, 67(8): 2009–2015, August 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Kim:2015:EKD

[Kim15] Sung Un Kim. Exploring the knowledge development process of English language learners at a high school: How do English language proficiency and the nature of research task influence student Learning? Journal of the Association for Information Science and Technology, 66(1):128–143, January 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Kim:2019:ABA

[Kim19a] Jinseok Kim. Author-based analysis of conference versus journal publication in computer science. Journal of the Association for Information Science and Technology, 70(1):71– 82, January 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Kim:2019:SFC

[Kim19b] Jinseok Kim. Scale-free collaboration networks: an author name disambiguation perspective. Journal of the Association for Information Science and Technology, 70(7):685–700, July 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Kurata:2017:PDR

[KIMM17] Keiko Kurata, Emi Ishita, Yosuke Miyata, and Yukiko Minami. Print or digital? Reading behavior and preferences in Japan. Journal of the Association for Information Science and Technology, 68(4):884–894, April 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Kozak:2015:BCD

[KISS15] Marcin Kozak, Olesia Iefremova, Jarosław Szkoła, and Daniel Sas. Brief communications: Do researchers provide public or institutional E-mail accounts as correspondence Emails in scientific articles? Journal of the Association for Information Science and Technology, 66(10):2149–2154, October 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Kitzie:2019:LLM

[Kit19] Vanessa Kitzie. “That looks like me or something I can do”: Affordances and constraints in the online identity work of US LGBTQ+ millennials. Journal of the Association for Information Science and Technology, 70(12):1340–1351, December 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Kim:2016:GSS

[KK16] Hyun Hee Kim and Yong Ho Kim. Generic speech summarization of transcribed lecture videos: Using tags and their semantic relations. Journal of the Association for Information Science and Technology, 67(2):366–379, February 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Kim:2019:EMA

[KK19a] Hyun Hee Kim and Yong Ho Kim. ERP/MMR algorithm for classifying topic-relevant and topic-irrelevant visual shots of documentary videos. Journal of the Association for Information Science and Technology, 70(9):931–941, September 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Kim:2019:VSU

[KK19b] Hyun Hee Kim and Yong Ho Kim. Video summarization using event-related potential responses to shot boundaries in real-time video watching. Journal of the Association for Information Science and Technology, 70(2):164–175, February 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Kim:2020:EFS

[KK20] Jinseok Kim and Jenna Kim. Effect of forename string on author name disambiguation. Journal of the Association for Information Science and Technology, 71(7):839–855, July 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Korkeamaki:2024:TDT

[KKK24] Laura Korkeamäki, Heikki Keskustalo, and Sanna Kumpulainen. Types of domain and task-solving information in media scholars’ data interaction. Journal of the Association for Information Science and Technology, 75(4):454–468, April 2024. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Kim:2015:CLS

[KKO15] Sungho Kim, Youngjoong Ko, and Douglas W. Oard. Combining lexical and statistical translation evidence for crosslanguage information retrieval. Journal of the Association for Information Science and Technology, 66(1):23–39, January 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Kim:2021:EBN

[KKOS21] Jinseok Kim, Jenna Kim, and Jason Owen-Smith. Ethnicitybased name partitioning for author name disambiguation using supervised machine learning. Journal of the Association for Information Science and Technology, 72(8):979–994, August 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Kumpulainen:2020:HRA

[KKZS20] Sanna Kumpulainen, Heikki Keskustalo, Boyang Zhang, and Kostas Stefanidis. Historical reasoning in authentic research tasks: Mapping cognitive and document spaces. Journal of the Association for Information Science and Technology, 71 (2):230–241, February 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Kumpulainen:2022:SDH

[KL22] Sanna Kumpulainen and Elina Late. Struggling with digitized historical newspapers: Contextual barriers to information interaction in history research activities. Journal of the Association for Information Science and Technology, 73 (7):1012–1024, July 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Kazmer:2014:DKO

[KLC^{+}14] Michelle M. Kazmer, Mia Liza A. Lustria, Juliann Cortese, Gary Burnett, Ji-Hyun Kim, Jinxuan Ma, and Jeana Frost. Distributed knowledge in an online patient support community: Authority and discovery. Journal of the Association for Information Science and Technology, 65(7):1319–1334, July 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Klavans:2014:SMC

[KLG14] Judith L. Klavans, Rebecca LaPlante, and Jennifer Golbeck. Subject matter categorization of tags applied to digital images from art museums. Journal of the Association for Information Science and Technology, 65(1):3–12, January 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Kauchak:2017:MTD

[KLH17] David Kauchak, Gondy Leroy, and Alan Hogue. Measuring text difficulty using parse-tree frequency. Journal of the Association for Information Science and Technology, 68(9): 2088–2100, September 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Km:2020:GCE

[KMC20] Pooja Km, Samrat Mondal, and Joydeep Chandra. A graph combination with edge pruning-based approach for author name disambiguation. Journal of the Association for Information Science and Technology, 71(1):69–83, January 2020.

CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Kljun:2016:TUS

[KMD16] Matjaž Kljun, John Mariani, and Alan Dix. Toward understanding short-term personal information preservation: a study of backup strategies of end users. Journal of the Association for Information Science and Technology, 67(12): 2947–2963, December 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Kronegger:2015:CSD

[KMFD15] Luka Kronegger, Franc Mali, Anuška Ferligoj, and Patrick Doreian. Classifying scientific disciplines in Slovenia: a study of the evolution of collaboration structures. Journal of the Association for Information Science and Technology, 66(2): 321–339, February 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Korman:2018:DTE

[KMJR18] Daniel Z. Korman, Eric Mack, Jacob Jett, and Allen H. Renear. Defining textual entailment. Journal of the Association for Information Science and Technology, 69(6):763–772, June 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Kless:2015:TOS

[KMKL15] Daniel Kless, Simon Milton, Edmund Kazmierczak, and Jutta Lindenthal. Thesaurus and ontology structure: Formal and pragmatic differences and similarities. Journal of the Association for Information Science and Technology, 66 (7):1348–1366, July 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Kodua-Ntim:2023:NRO

[KN23] Kwame Kodua-Ntim. Narrative review on open access institutional repositories and knowledge sharing in South Africa. Journal of the Association for Information Science and Technology, 74(9):1118–1123, September 2023. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Knott:2024:BRB

[Kno24] Cheryl Knott. Book review: Birth of modern facts: How the information revolution transformed academic research, governments, and businesses, Cortada, James W., Rowman & Littlefield. 2023. 462 pp. $125.00 (hardcover). (ISBN: 978-1-5381-7390-9). Journal of the Association for Information Science and Technology, 75(4):496–499, April 2024. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Kay:2014:POM

[KNY^{+}14] Luciano Kay, Nils Newman, Jan Youtie, Alan L. Porter, and Ismael Rafols. Patent overlay mapping: Visualizing technological distance. Journal of the Association for Information Science and Technology, 65(12):2432–2443, December 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Ko:2015:NTW

[Ko15] Youngjoong Ko. A new term-weighting scheme for text classification using the odds of positive and negative class probabilities. Journal of the Association for Information Science and Technology, 66(12):2553–2565, December 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Konieczny:2016:AIS

[Kon16] Piotr Konieczny. Advances in information science: Teaching with Wikipedia in a 21^{st}-century classroom: Perceptions of Wikipedia and its educational benefits. Journal of the Association for Information Science and Technology, 67(7): 1523–1534, July 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Kosciejew:2020:BRM

[Kos20] Marc Kosciejew. Book review: The Misinformation Age: How False Ideas Spread. Cailin O’Connor and James Owen Weatherall. New Haven, CT: Yale University Press, 2019. 280 pp. $26.00 (hardcover). (ISBN 978-0-300-23401-5). Journal of the Association for Information Science and Technology, 71(5):612–615, May 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Kosciejew:2021:BRI

[Kos21] Marc Kosciejew. Book review: The information manifold: Why computers cannot solve algorithmic bias and fake news. Badia, Antonio Cambridge, UK: The MIT press, 2019. 352 pp. $50.00 (hardcover). (ISBN 978-0-262-04303-8). Journal of the Association for Information Science and Technology, 72(3):357–361, March 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Kvasny:2018:MHH

[KP18] Lynette Kvasny and Fay Cobb Payton. Managing hypervisibility in the HIV prevention information-seeking practices of black female college students. Journal of the Association for Information Science and Technology, 69(6):798–806, June 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Kelty:2015:SDC

[KPC^{+}15] Christopher Kelty, Aaron Panofsky, Morgan Currie, Roderic Crooks, Seth Erickson, Patricia Garcia, Michael Wartenbe, and Stacy Wood. Seven dimensions of contemporary participation disentangled. Journal of the Association for Information Science and Technology, 66(3):474–488, March 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Kim:2020:SJS

[KPWS20] Lanu Kim, Jason H. Portenoy, Jevin D. West, and Katherine W. Stovel. Scientific journals still matter in the era of academic search engines and preprint archives. Journal of the Association for Information Science and Technology, 71 (10):1218–1226, October 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Kalman:2015:FPE

[KR15] Yoram M. Kalman and Gilad Ravid. Filing, piling, and everything in between: the dynamics of E-mail inbox management. Journal of the Association for Information Science and Technology, 66(12):2540–2552, December 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Kraft:2014:MLL

[Kra14] Donald H. Kraft. In memoriam: Lois Lunin. Journal of the Association for Information Science and Technology, 65 (9):1948, September 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Krauskopf:2017:CCU

[Kra17] Erwin Krauskopf. Call for caution in the use of bibliometric data. Journal of the Association for Information Science and Technology, 68(8):2029–2032, August 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Koppel:2014:MDI

[KS14a] Moshe Koppel and Nadav Schweitzer. Measuring direct and indirect authorial influence in historical corpora. Journal of the Association for Information Science and Technology, 65 (10):2138–2144, October 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Koumenides:2014:AIS

[KS14b] Christos L. Koumenides and Nigel R. Shadbolt. Advances in information science: Ranking methods for entity-oriented semantic web search. Journal of the Association for Information Science and Technology, 65(6):1091–1106, June 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Kim:2016:IIF

[KS16] Youngseek Kim and Jeffrey M. Stanton. Institutional and individual factors affecting scientists’ data-sharing behaviors: a multilevel analysis. Journal of the Association for Information Science and Technology, 67(4):776–799, April 2016. CO-DEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Kocher:2017:SEA

[KS17] Mirco Kocher and Jacques Savoy. A simple and efficient algorithm for authorship verification. Journal of the Association for Information Science and Technology, 68(1):259–269, January 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Khosrowjerdi:2020:CPI

[KSB20] Mahmood Khosrowjerdi, Anneli Sundqvist, and Katriina Byström. Cultural patterns of information source use: A global study of 47 countries. Journal of the Association for Information Science and Technology, 71(6):711–724, June 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Koh:2019:PML

[KSL19] Kyungwon Koh, John T. Snead, and Kun Lu. The processes of maker learning and information behavior in a technologyrich high school class. Journal of the Association for Information Science and Technology, 70(12):1395–1412, December 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Kaba:2014:UIC

[KT14a] Bangaly Kaba and Bakary Touré. Understanding information and communication technology behavioral intention to use: applying the UTAUT model to social networking site adoption by young people in a least developed country. Journal of the Association for Information Science and Technology, 65(8):1662–1674, August 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Kousha:2014:DRW

[KT14b] Kayvan Kousha and Mike Thelwall. Disseminating research with web CV hyperlinks. Journal of the Association for Information Science and Technology, 65(8):1615–1626, August 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Kousha:2015:AME

[KT15] Kayvan Kousha and Mike Thelwall. An automatic method for extracting citations from Google Books. Journal of the Association for Information Science and Technology, 66(2): 309–320, February 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Kousha:2016:AMA

[KT16a] Kayvan Kousha and Mike Thelwall. An automatic method for assessing the teaching impact of books from online academic syllabi. Journal of the Association for Information Science and Technology, 67(12):2993–3007, December 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Kousha:2016:CAC

[KT16b] Kayvan Kousha and Mike Thelwall. Can amazon.com reviews help to assess the wider impacts of books? Journal of the Association for Information Science and Technology, 67(3):566–581, March 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Kousha:2017:WCI

[KT17a] Kayvan Kousha and Mike Thelwall. Are Wikipedia citations important evidence of the impact of scholarly articles and books? Journal of the Association for Information Science and Technology, 68(3):762–779, March 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Kousha:2017:NSE

[KT17b] Kayvan Kousha and Mike Thelwall. News stories as evidence for research? BBC citations from articles, books, and Wikipedia. Journal of the Association for Information Science and Technology, 68(8):2017–2028, August 2017. CO-DEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Kousha:2017:PCA

[KT17c] Kayvan Kousha and Mike Thelwall. Patent citation analysis with Google. Journal of the Association for Information Science and Technology, 68(1):48–61, January 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Kousha:2017:GRA

[KTA17] Kayvan Kousha, Mike Thelwall, and Mahshid Abdoli. Goodreads reviews to assess the wider impacts of books. Journal of the Association for Information Science and Technology, 68(8):2004–2016, August 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Kinley:2014:MUW

[KTPE14] Khamsum Kinley, Dian Tjondronegoro, Helen Partridge, and Sylvia Edwards. Modeling users’ web search behavior and their cognitive styles. Journal of the Association for Information Science and Technology, 65(6):1107–1123, June 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Kim:2015:ETA

[KTS^{+}15] Jaewon Kim, Paul Thomas, Ramesh Sankaranarayana, Tom Gedeon, and Hwan-Jin Yoon. Eye-tracking analysis of user behavior and performance in web search on large and small screens. Journal of the Association for Information Science and Technology, 66(3):526–544, March 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Kim:2016:UEM

[KTS^{+}16] Jaewon Kim, Paul Thomas, Ramesh Sankaranarayana, Tom Gedeon, and Hwan-Jin Yoon. Understanding eye movements on mobile devices for better presentation of search results. Journal of the Association for Information Science and Technology, 67(11):2607–2619, November 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Kang:2020:DCA

[KTZ20] Lele Kang, Chuan-Hoo Tan, and J. Leon Zhao. Do cognitive and affective expressions matter in purchase conversion? A live chat perspective. Journal of the Association for Information Science and Technology, 71(4):436–449, April 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Kuehn:2023:IEC

[Kue23] Evan F. Kuehn. The information ecosystem concept in information literacy: a theoretical approach and definition. Journal of the Association for Information Science and Technology, 74(4):434–443, April 2023. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Karanam:2017:CMA

[KvOS^{+}17] Saraschandra Karanam, Herre van Oostendorp, Mylene Sanchiz, Aline Chevalier, Jessie Chin, and Wai-Tat Fu. Cognitive modeling of age-related differences in information search behavior. Journal of the Association for Information Science and Technology, 68(10):2328–2337, October 2017.

CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Koppel:2014:DIT

[KW14] Moshe Koppel and Yaron Winter. Determining if two documents are written by the same author. Journal of the Association for Information Science and Technology, 65(1):178– 187, January 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Kitzie:2022:AIP

[KWLV22] Vanessa L. Kitzie, Travis L. Wagner, Valerie Lookingbill, and Nicolas Vera. Advancing information practices theoretical discourses centered on marginality, community, and embodiment: Learning from the experiences of lesbian, gay, bisexual, transgender, queer, intersex, and asexual (LGBTQIA+) communities. Journal of the Association for Information Science and Technology, 73(4):494–510, April 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Kang:2020:TAC

[KWR20] Xin Kang, Yunong Wu, and Fuji Ren. Toward action comprehension for searching: Mining actionable intents in query entities. Journal of the Association for Information Science and Technology, 71(2):143–157, February 2020. CO-DEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Kim:2017:SDR

[KY17] Youngseek Kim and Ayoung Yoon. Scientists’ data reuse behaviors: a multilevel analysis. Journal of the Association for Information Science and Technology, 68(12):2709–2719, December 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Kong:2015:TPP

[KYFZ15] Shoubin Kong, Fei Ye, Ling Feng, and Zhe Zhao. Towards the prediction problems of bursting hashtags on Twitter. Journal of the Association for Information Science and Technology, 66(12):2566–2579, December 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Karanasios:2023:PRS

[KZ23] Stan Karanasios and Aljona Zorina. From participation roles to socio-emotional information roles: Insights from the closure of an online community. Journal of the Association for Information Science and Technology, 74(1):33–49, January 2023. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Koopman:2017:WME

[KZB17] Bevan Koopman, Guido Zuccon, and Peter Bruza. What makes an effective clinical query and querier? Journal of the Association for Information Science and Technology, 68(11): 2557–2571, November 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Leydesdorff:2014:SNT

[LA14] Loet Leydesdorff and Petra Ahrweiler. In search of a network theory of innovations: Relations, positions, and perspectives. Journal of the Association for Information Science and Technology, 65(11):2359–2374, November 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Lai:2015:AKF

[Lai15] Chin-Hui Lai. Applying knowledge flow mining to group recommendation methods for task-based groups. Journal of the Association for Information Science and Technology, 66 (3):545–563, March 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Lanzara:2015:BRM

[Lan15] Giovan Francesco Lanzara. Book reviews: Materiality and Organizing: Social Interaction in a Technological World — Edited by Paul M. Leonardi, Bonnie A. Nardi and Jannis Kallinikos. Journal of the Association for Information Science and Technology, 66(12):2717–2720, December 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Liu:2015:PIR

[LB15] Jingjing Liu and Nicholas J. Belkin. Personalizing information retrieval for multi-session tasks: Examining the roles of task stage, task type, and topic knowledge on the interpretation of dwell time as an indicator of document usefulness. Journal of the Association for Information Science and Technology, 66(1):58–81, January 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Leydesdorff:2016:OFW

[LB16] Loet Leydesdorff and Lutz Bornmann. The operationalization of “fields” as WoS subject categories (WCs) in evaluative bibliometrics: the cases of “library and information science” and “science & technology studies”. Journal of the Association for Information Science and Technology, 67(3): 707–714, March 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Lee:2019:FIC

[LB19a] Danielle H. Lee and Peter Brusilovsky. The first impression of conference papers: Does it matter in predicting future citations? Journal of the Association for Information Science and Technology, 70(1):83–95, January 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Lee:2019:HID

[LB19b] Myeong Lee and Brian S. Butler. How are information deserts created? A theory of local information landscapes. Journal of the Association for Information Science and Technology, 70(2):101–116, February 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Lin:2020:HDC

[LB20] Yan Lin and Wai Fong Boh. How different are crowdfunders? Examining archetypes of crowdfunders. Journal of the Association for Information Science and Technology, 71(11): 1357–1370, November 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Luhmann:2022:DHD

[LB22] Jan Luhmann and Manuel Burghardt. Digital humanities — a discipline in its own right? An analysis of the role and position of digital humanities in the academic landscape. Journal of the Association for Information Science and Technology, 73(2):148–171, February 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Lhadj:2016:EIR

[LBA16] Lynda Said Lhadj, Mohand Boughanem, and Karima Amrouche. Enhancing information retrieval through conceptbased language modeling and semantic smoothing. Journal of the Association for Information Science and Technology, 67(12):2909–2927, December 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Liu:2022:PCS

[LBC^{+}22] Meijun Liu, Yi Bu, Chongyan Chen, Jian Xu, Daifeng Li, Yan Leng, Richard B. Freeman, Eric T. Meyer, Wonjin Yoon, Mujeen Sung, Minbyul Jeong, Jinhyuk Lee, Jaewoo Kang, Chao Min, Min Song, Yujia Zhai, and Ying Ding. Pandemics are catalysts of scientific novelty: Evidence from COVID-19. Journal of the Association for Information Science and Technology, 73(8):1065–1078, August 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Lee:2021:AMN

[LBG^{+}21] Seungyoon Lee, Bailey C. Benedict, Yue ‘Gurt’ Ge, Pamela Murray-Tuite, and Satish V. Ukkusuri. An application of media and network multiplexity theory to the structure and perceptions of information environments in hurricane evacuation. Journal of the Association for Information Science and Technology, 72(7):885–900, July 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Liu:2024:MCD

[LBLL24] Xin Liu, Yi Bu, Ming Li, and Jiang Li. Monodisciplinary collaboration disrupts science more than multidisciplinary collaboration. Journal of the Association for Information Science and Technology, 75(1):59–78, January 2024. CO-DEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Leydesdorff:2019:SSE

[LBM19a] Loet Leydesdorff, Lutz Bornmann, and John Mingers. Statistical significance and effect sizes of differences among research universities at the level of nations and worldwide based on the Leiden rankings. Journal of the Association for Information Science and Technology, 70(5):509–525, May 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Lueg:2019:CEF

[LBM^{+}19b] Christopher Lueg, Bridgette Banks, Josephine Michalek, Jack Dimsey, and Dominic Oswin. Close encounters of the fifth kind: Recognizing system-initiated engagement as interaction type. Journal of the Association for Information Science and Technology, 70(6):634–637, June 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Leydesdorff:2019:RIG

[LBW19a] Loet Leydesdorff, Lutz Bornmann, and Caroline S. Wagner. The relative influences of government funding and international collaboration on citation impact. Journal of the Association for Information Science and Technology, 70(2):198– 201, February 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Lu:2019:ESW

[LBW^{+}19b] Chao Lu, Yi Bu, Jie Wang, Ying Ding, Vetle Torvik, Matthew Schnaars, and Chengzhi Zhang. Examining scientific writing styles from the perspective of linguistic complexity. Journal of the Association for Information Science and Technology, 70(5):462–475, May 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Laparra:2023:ASH

[LBWE^{+}23] Egoitz Laparra, Alex Binford-Walsh, Kirk Emerson, Marc L. Miller, Laura López-Hoffman, Faiz Currim, and Steven Bethard. Addressing structural hurdles for metadata extraction from environmental impact statements. Journal of the Association for Information Science and Technology, 74(9): 1124–1139, September 2023. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Liu:2021:ANE

[LC21] Xiaomei Liu and Xiaotian Chen. Authors’ noninstitutional emails and their correlation with retraction. Journal of the Association for Information Science and Technology, 72(4): 473–477, April 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

LaBella:2018:APO

[LCB^{+}18] Agostino La Bella, Andrea Fronzetti Colladon, Elisa Battistoni, Silvia Castellan, and Matteo Francucci. Assessing perceived organizational leadership styles through Twitter text mining. Journal of the Association for Information Science and Technology, 69(1):21–31, January 2018. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Li:2014:PCA

[LCD^{+}14] Rui Li, Tamy Chambers, Ying Ding, Guo Zhang, and Liansheng Meng. Patent citation analysis: Calculating science linkage based on citing motivation. Journal of the Association for Information Science and Technology, 65(5):1007– 1017, May 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Labib:2022:ANB

[LCH^{+}22] Ashraf Labib, Salem Chakhar, Lorraine Hope, John Shimell, and Mark Malinowski. Analysis of noise and bias errors in intelligence information systems. Journal of the Association for Information Science and Technology, 73(12):1755–1775, December 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Liu:2014:CDL

[LCHL14] John S. Liu, Hsiao-Hui Chen, Mei Hsiu-Ching Ho, and Yu-Chen Li. Citations with different levels of relevancy: Tracing the main paths of legal opinions. Journal of the Association for Information Science and Technology, 65(12):2479–2488, December 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Lee:2016:UMI

[LCK16] Jin Ha Lee, Hyerim Cho, and Yea-Seul Kim. Users’ music information needs and behaviors: Design implications for music information retrieval systems. Journal of the Association for Information Science and Technology, 67(6):1301– 1330, June 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Liu:2016:USM

[LCL16] Ivy L. B. Liu, Christy M. K. Cheung, and Matthew K. O. Lee. User satisfaction with microblogging: Information dissemination versus social networking. Journal of the Association for Information Science and Technology, 67(1):56–70, January 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Lehmann:2017:SFR

[LCLBY17] Janette Lehmann, Carlos Castillo, Mounia Lalmas, and Ricardo Baeza-Yates. Story-focused reading in online news and its potential for user engagement. Journal of the Association for Information Science and Technology, 68(4):869–883, April 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Lian:2018:TPO

[LCLM18] Tao Lian, Zhumin Chen, Yujie Lin, and Jun Ma. Temporal patterns of the online video viewing behavior of smart TV viewers. Journal of the Association for Information Science and Technology, 69(5):647–659, May 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Larosiliere:2017:HDW

[LCM17] Gregory D. Larosiliere, Lemuria D. Carter, and Christian Meske. How does the world connect? Exploring the global diffusion of social network sites. Journal of the Association for Information Science and Technology, 68(8):1875–1885, August 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Laut:2017:ICS

[LCNP17] Jeffrey Laut, Francesco Cappa, Oded Nov, and Maurizio Porfiri. Increasing citizen science contribution using a virtual peer. Journal of the Association for Information Science and Technology, 68(3):583–593, March 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Lee:2015:EEM

[LCP15] Jin Ha Lee, Rachel Ivy Clarke, and Andrew Perti. Empirical evaluation of metadata for video games and interactive media. Journal of the Association for Information Science and Technology, 66(12):2609–2625, December 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Lopez-Cozar:2014:GSE

[LCRGTS14] Emilio Delgado López-Cózar, Nicolás Robinson-García, and Daniel Torres-Salinas. The Google Scholar experiment: How to index false papers and manipulate bibliometric indicators. Journal of the Association for Information Science and Technology, 65(3):446–454, March 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Liu:2015:CQN

[LCSL15] Duen-Ren Liu, Yu-Hsuan Chen, Minxin Shen, and Pei-Jung Lu. Complementary QA network analysis for QA retrieval in social question-answering websites. Journal of the Association for Information Science and Technology, 66(1):99–116, January 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Li:2023:TIB

[LCWC23] Yuan Li, Anita Crescenzi, Austin R. Ward, and Rob Capra. Thinking inside the box: an evaluation of a novel searchassisting tool for supporting (meta)cognition during exploratory search. Journal of the Association for Information Science and Technology, 74(9):1049–1066, September 2023. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Luo:2018:ESI

[LCWL18] Peng Luo, Kun Chen, Chong Wu, and Yongli Li. Exploring the social influence of multichannel access in an online health community. Journal of the Association for Information Science and Technology, 69(1):98–109, January 2018. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Liu:2014:AND

[LDK^{+}14] Wanli Liu, Rezarta Islamaj Doğan, Sun Kim, Donald C. Comeau, Won Kim, Lana Yeganova, Zhiyong Lu, and W. John Wilbur. Author name disambiguation for PubMed. Journal of the Association for Information Science and Technology, 65(4):765–781, April 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Leydesdorff:2016:AJJ

[LdMAdN16] Loet Leydesdorff, Félix de Moya-Anegón, and Wouter de Nooy. Aggregated journal–journal citation relations in Scopus and Web of Science matched and compared in terms of networks, maps, and interactive overlays. Journal of the Association for Information Science and Technology, 67(9): 2194–2211, September 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Leydesdorff:2015:JMI

[LdMAGB15] Loet Leydesdorff, Félix de Moya-Anegón, and Vicente P. Guerrero-Bote. Journal maps, interactive overlays, and the measurement of interdisciplinarity on the basis of Scopus data (1996–2012). Journal of the Association for Information Science and Technology, 66(5):1001–1016, May 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Leydesdorff:2017:CHS

[LdN17] Loet Leydesdorff and Wouter de Nooy. Can “hot spots” in the sciences be mapped using the dynamics of aggregated journal–journal citation relations? Journal of the Association for Information Science and Technology, 68(1):197–213, January 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Liu:2019:ISR

[LDS^{+}19] Yezheng Liu, Fei Du, Jianshan Sun, Thushari Silva, Yuanchun Jiang, and Tingting Zhu. Identifying social roles using heterogeneous features in online social networks. Journal of the Association for Information Science and Technology, 70(7):660–674, July 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Levine:2015:BRI

[Lev15] Emil Levine. Book reviews: Information and Intrigue: From Index Cards to Dewey Decimal to Alger Hiss — by Colin Burke. Journal of the Association for Information Science and Technology, 66(10):2168–2170, October 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Levine:2020:BRA

[Lev20] Emil Levine. Book review: America’s information wars: The untold story of information systems in America’s conflicts and politics from World War II to the Internet age. Burke, Colin B. Lanham, MD: Rowman & Littlefield, 2018. 390 pp. $85.00 (hardcover). (ISBN 978-1-5381-1245-8). Journal of the Association for Information Science and Technology, 71 (8):991–994, August 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Lewandowski:2015:ERE

[Lew15] Dirk Lewandowski. Evaluating the retrieval effectiveness of web search engines using a representative query sample. Journal of the Association for Information Science and Technology, 66(9):1763–1775, September 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Lai:2021:UAA

[LF21] Shiyang Lai and Ningyuan Fan. Understanding the attenuation of the accommodation recommendation spillover effect in view of spatial distance. Journal of the Association for Information Science and Technology, 72(11):1448–1453, November 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Leydesdorff:2014:IJS

[LG14] Loet Leydesdorff and Robert L. Goldstone. Interdisciplinarity at the journal and specialty level: the changing knowledge bases of the journal cognitive science. Journal of the Association for Information Science and Technology, 65(1):164– 177, January 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Li:2020:DOD

[LGD20] Kai Li, Jane Greenberg, and Jillian Dunic. Data objects and documenting scientific processes: an analysis of data events in biodiversity data papers. Journal of the Association for Information Science and Technology, 71(2):172–182, February 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Lariviere:2015:TSM

[LGST15] Vincent Larivière, Yves Gingras, Cassidy R. Sugimoto, and Andrew Tsou. Team size matters: Collaboration and scientific impact since 1900. Journal of the Association for Information Science and Technology, 66(7):1323–1332, July 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Liu:2014:SMK

[LGZ14] Xiaozhong Liu, Chun Guo, and Lin Zhang. Scholar metadata and knowledge generation with human and artificial intelligence. Journal of the Association for Information Science and Technology, 65(6):1187–1201, June 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Lutz:2017:OSC

[LHM17] Christoph Lutz, Christian Pieter Hoffmann, and Miriam Meckel. Online serendipity: a contextual differentiation of antecedents and outcomes. Journal of the Association for Information Science and Technology, 68(7):1698–1710, July 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Leydesdorff:2016:BCJ

[LHR16] Loet Leydesdorff, Gaston Heimeriks, and Daniele Rotolo. Brief communication: Journal portfolio analysis for countries, cities, and organizations: Maps and comparisons. Journal of the Association for Information Science and Technology, 67(3):741–748, March 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Lewandowski:2024:JSI

[LHS24] Dirk Lewandowski, Jutta Haider, and Olof Sundin. JASIST special issue editorial: Re-orienting search engine research in information science. Journal of the Association for Information Science and Technology, 75(5):503–511, May 2024. CO-DEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Liu:2022:MDA

[LHXS22] Xiaoxiao Liu, Mingye Hu, Bo Sophia Xiao, and Jingbo Shao. Is my doctor around me? Investigating the impact of doctors’ presence on patients’ review behaviors on an online health platform. Journal of the Association for Information Science and Technology, 73(9):1279–1296, September 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Leydesdorff:2014:MRI

[LI14a] Loet Leydesdorff and Inga A. Ivanova. Mutual redundancies in interhuman communication systems: Steps toward a calculus of processing meaning. Journal of the Association for Information Science and Technology, 65(2):386–399, February 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Li:2014:IGF

[Li14b] Rowena Li. The influence of geospatial factors on democracy: Its representation on web interface design. Journal of the Association for Information Science and Technology, 65(2): 313–333, February 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Leydesdorff:2021:MIS

[LI21a] Loet Leydesdorff and Inga Ivanova. The measurement of “interdisciplinarity” and “synergy” in scientific and extrascientific collaborations. Journal of the Association for Information Science and Technology, 72(4):387–402, April 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Li:2021:YPI

[Li21b] Xiaofeng Li. Young people’s information practices in library makerspaces. Journal of the Association for Information Science and Technology, 72(6):744–758, June 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Louvier:2022:HAM

[LI22] Kahina Le Louvier and Perla Innocenti. Heritage as an affective and meaningful information literacy practice: an interdisciplinary approach to the integration of asylum seekers and refugees. Journal of the Association for Information Science and Technology, 73(5):687–701, May 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Liao:2017:ESE

[Lia17] Chien Hsiang Liao. Exploring the social effect of outstanding scholars on future research accomplishments. Journal of the Association for Information Science and Technology, 68(10): 2449–2459, October 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Lindenthal:2014:VWP

[Lin14a] Thies Lindenthal. Valuable words: the price dynamics of Internet domain names. Journal of the Association for Information Science and Technology, 65(5):869–881, May 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Ling:2014:BRN

[Lin14b] Richard Ling. Book review: Networked: The New Social Operating System — by Lee Rainie and Barry Wellman. Journal of the Association for Information Science and Technology, 65(9):1944–1947, September 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Lingel:2015:IPU

[Lin15] Jessa Lingel. Information practices of urban newcomers: an analysis of habits and wandering. Journal of the Association for Information Science and Technology, 66(6):1239–1251, June 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Liu:2017:ATA

[LJ17] Zhe Liu and Bernard J. Jansen. ASK: a taxonomy of accuracy, social, and knowledge information seeking posts in social question and answering. Journal of the Association for Information Science and Technology, 68(2):333–347, February 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Li:2022:DPS

[LJ22] Kai Li and Chenyue Jiao. The data paper as a sociolinguistic epistemic object: a content analysis on the rhetorical moves used in data paper abstracts. Journal of the Association for Information Science and Technology, 73(6):834–846, June 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Leydesdorff:2018:TCR

[LJI18] Loet Leydesdorff, Mark W. Johnson, and Inga Ivanova. Toward a calculus of redundancy: Signification, codification, and anticipation in cultural evolution. Journal of the Association for Information Science and Technology, 69(10): 1181–1192, October 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Lee:2021:DUS

[LJK21] Jun Lee, Adam Jatowt, and Kyoung-Sook Kim. Discovering underlying sensations of human emotions based on social media. Journal of the Association for Information Science and Technology, 72(4):417–432, April 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Lu:2017:FIQ

[LJLH17] Kun Lu, Soohyung Joo, Taehun Lee, and Rong Hu. Factors that influence query reformulations and search performance in health information retrieval: a multilevel modeling approach. Journal of the Association for Information Science and Technology, 68(8):1886–1898, August 2017. CO-DEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Lee:2015:MBB

[LJT^{+}15] Lung-Hao Lee, Yen-Cheng Juan, Wei-Lin Tseng, Hsin-Hsi Chen, and Yuen-Hsien Tseng. Mining browsing behaviors for objectionable content filtering. Journal of the Association for Information Science and Technology, 66(5):930–942, May 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Lu:2014:URE

[LK14] Kun Lu and Margaret E. I. Kipp. Understanding the retrieval effectiveness of collaborative tags and author keywords in different retrieval environments: an experimental study on medical collections. Journal of the Association for Information Science and Technology, 65(3):483–500, March 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Liu:2016:NAM

[LK16] John S. Liu and Chung-Huei Kuan. A new approach for main path analysis: Decay in knowledge diffusion. Journal of the Association for Information Science and Technology, 67(2): 465–476, February 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Li:2020:CPC

[LK20] Yao Li and Alfred Kobsa. Context and privacy concerns in friend request decisions. Journal of the Association for Information Science and Technology, 71(6):632–643, June 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Lavranos:2016:ISM

[LKKP16] Charilaos Lavranos, Petros Kostagiolas, Nikolaos Korfiatis, and Joseph Papadatos. Information seeking for musical creativity: a systematic literature review. Journal of the Association for Information Science and Technology, 67(9):2105– 2117, September 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Lee:2017:CEB

[LKKS17] Keeheon Lee, SuYeon Kim, Erin Hea-Jin Kim, and Min Song. Comparative evaluation of bibliometric content networks by tomographic content analysis: an application to Parkinson’s disease. Journal of the Association for Information Science and Technology, 68(5):1295–1307, May 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Lewandowski:2018:EIS

[LKRS18] Dirk Lewandowski, Friederike Kerkmann, Sandra Rümmele, and Sebastian Sünkler. An empirical investigation on search engine ad disclosure. Journal of the Association for Information Science and Technology, 69(3):420–437, March 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Lee:2020:CLW

[LKY^{+}20] Yang-Yin Lee, Hao Ke, Ting-Yu Yen, Hen-Hsen Huang, and Hsin-Hsi Chen. Combining and learning word embedding with WordNet for semantic relatedness and similarity measurement. Journal of the Association for Information Science and Technology, 71(6):657–670, June 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

LaFata:2017:CFS

[LL17] Concetta Manuela La Fata and Toni Lupo. A combined fuzzy-SEM evaluation approach to identify the key drivers of the academic library service quality in the digital technology era: an empirical study. Journal of the Association for Information Science and Technology, 68(10):2425–2438, October 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Loehrlein:2014:CFP

[LLB14] Aaron J. Loehrlein, Victoria L. Lemieux, and Michael Bennett. The classification of financial products. Journal of the Association for Information Science and Technology, 65(2): 263–280, February 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Liu:2016:PIS

[LLB16] Jingjing Liu, Chang Liu, and Nicholas J. Belkin. Predicting information searchers’ topic knowledge at different search stages. Journal of the Association for Information Science and Technology, 67(11):2652–2666, November 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Liu:2020:PTI

[LLB20] Jingjing Liu, Chang Liu, and Nicholas J. Belkin. Personalization in text information retrieval: a survey. Journal of the Association for Information Science and Technology, 71 (3):349–369, March 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Li:2017:ULM

[LLD^{+}17] Daifeng Li, Zhipeng Luo, Ying Ding, Jie Tang, Gordon Guo-Zheng Sun, Xiaowen Dai, John Du, Jingwei Zhang, and Shoubin Kong. User-level microblogging recommendation incorporating social influence. Journal of the Association for Information Science and Technology, 68(3):553–568, March 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Lariviere:2014:EJD

[LLG14] Vincent Larivière, George A. Lozano, and Yves Gingras. Are elite journals declining? Journal of the Association for Information Science and Technology, 65(4):649–655, April 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Liu:2019:SSS

[LLH19] Jingjing Liu, Yuan Li, and Samantha K. Hastings. Simplified scheme of search task difficulty reasons. Journal of the Association for Information Science and Technology, 70(5): 526–529, May 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Liu:2015:FMB

[LLHF15] Yu Liu, Weijia Li, Zhen Huang, and Qiang Fang. A fast method based on multiple clustering for name disambiguation in bibliographic citations. Journal of the Association for Information Science and Technology, 66(3):634–644, March 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Lin:2015:IID

[lLTBH15] Yi ling Lin, Christoph Trattner, Peter Brusilovsky, and Daqing He. The impact of image descriptions on user tagging behavior: a study of the nature and functionality of crowdsourced tags. Journal of the Association for Information Science and Technology, 66(9):1785–1798, September 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Laing:2015:SRV

[LM15a] Simon Laing and Masood Masoodian. A study of the role of visual information in supporting ideation in graphic design. Journal of the Association for Information Science and Technology, 66(6):1199–1211, June 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Lu:2015:AAW

[LM15b] Kun Lu and Jin Mao. An automatic approach to weighted subject indexing — an empirical study in the biomedical domain. Journal of the Association for Information Science and Technology, 66(9):1776–1784, September 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Lindau:2022:SIA

[LMA^{+}22] Stacy Tessler Lindau, Jennifer A. Makelarski, Emily M. Abramsohn, David G. Beiser, Kelly Boyd, Elbert S. Huang, Kelsey Paradise, and Elizabeth L. Tung. Sharing information about health-related resources: Observations from a community resource referral intervention trial in a predominantly African American/Black community. Journal of the Association for Information Science and Technology, 73(3):438– 448, March 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Laakso:2021:OFS

[LMJ21a] Mikael Laakso, Lisa Matthias, and Najko Jahn. Open is not forever: a study of vanished open access journals. Journal of the Association for Information Science and Technology, 72 (9):1099–1112, September 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Laakso:2021:RCO

[LMJ21b] Mikael Laakso, Lisa Matthias, and Najko Jahn. Response to comment on “Open is not forever: a study of vanished open access journals”. Journal of the Association for Information Science and Technology, 72(9):1115–1116, September 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Lin:2016:USS

[LML16] Yu-Ru Lin, Drew Margolin, and David Lazer. Uncovering social semantics from textual traces: a theory-driven approach and evidence from public statements of U.S. Members of Congress. Journal of the Association for Information Science and Technology, 67(9):2072–2089, September 2016. CO-DEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Lu:2018:TEA

[LML18] Kun Lu, Jin Mao, and Gang Li. Toward effective automated weighted subject indexing: a comparison of different approaches in different environments. Journal of the Association for Information Science and Technology, 69(1):121– 133, January 2018. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Liang:2023:BAS

[LML23] Zhentao Liang, Jin Mao, and Gang Li. Bias against scientific novelty: a prepublication perspective. Journal of the Association for Information Science and Technology, 74(1):99–114, January 2023. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Lopez:2022:MNN

[LMM^{+}22] Marisela Gutierrez Lopez, Stephann Makri, Andrew Mac-Farlane, Colin Porlezza, Glenda Cooper, and Sondess Missaoui. Making newsworthy news: The integral role of creativity and verification in the human information behavior that drives news story creation. Journal of the Association for Information Science and Technology, 73(10):1445–1460, October 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Lawrie:2015:CLP

[LMMO15] Dawn Lawrie, James Mayfield, Paul McNamee, and Douglas W. Oard. Cross-language person-entity linking from 20 languages. Journal of the Association for Information Science and Technology, 66(6):1106–1123, June 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Lemke:2021:CAR

[LMP21] Steffen Lemke, Athanasios Mazarakis, and Isabella Peters. Conjoint analysis of researchers’ hidden preferences for bibliometrics, altmetrics, and usage metrics. Journal of the Association for Information Science and Technology, 72(6): 777–792, June 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Leydesdorff:2017:CWM

[LN17] Loet Leydesdorff and Adina Nerghes. Co-word maps and topic modeling: a comparison using small and medium-sized corpora (N < 1 000). Journal of the Association for Information Science and Technology, 68(4):1024–1035, April 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Luo:2019:LGU

[LN19] Margaret Meiling Luo and Diane Nahl. Let’s Google: Uncertainty and bilingual search. Journal of the Association for Information Science and Technology, 70(9):1014–1025, September 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Lloyd:2019:UKI

[LO19] Annemaree Lloyd and Michael Olsson. Untangling the knot: The information practices of enthusiast car restorers. Journal of the Association for Information Science and Technology, 70(12):1311–1323, December 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Lee:2019:MSA

[LOD^{+}19] Jongwook Lee, Sanghee Oh, Hang Dong, Fang Wang, and Gary Burnett. Motivations for self-archiving on an academic social networking site: a study on ResearchGate. Journal of the Association for Information Science and Technology, 70 (6):563–574, June 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Lee:2022:IBP

[LOM22] Lo Lee, Melissa G. Ocepek, and Stephann Makri. Information behavior patterns: a new theoretical perspective from an empirical study of naturalistic information acquisition. Journal of the Association for Information Science and Technology, 73(4):594–608, April 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Lopatovska:2014:TME

[Lop14] Irene Lopatovska. Toward a model of emotions and mood in the online information search process. Journal of the Association for Information Science and Technology, 65(9): 1775–1793, September 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Losee:2016:TSD

[Los16] Robert Losee. Thesaurus structure, descriptive parameters, and scale. Journal of the Association for Information Science and Technology, 67(9):2156–2165, September 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Lee:2016:UEC

[LP16] Jin Ha Lee and Rachel Price. User experience with commercial music services: an empirical exploration. Journal of the Association for Information Science and Technology, 67(4): 800–811, April 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Losada:2019:WSM

[LPB19] David E. Losada, Javier Parapar, and Alvaro Barreiro. When to stop making relevance judgments? A study of stopping methods for building information retrieval test collections. Journal of the Association for Information Science and Technology, 70(1):49–60, January 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Lachance:2014:BCK

[LPL14] Christian Lachance, Steve Poirier, and Vincent Larivière. Brief communication: The kiss of death? The effect of being cited in a review on subsequent citations. Journal of the Association for Information Science and Technology, 65(7): 1501–1505, July 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Leydesdorff:2015:MTH

[LPU15] Loet Leydesdorff, Evgeniy Perevodchikov, and Alexander Uvarov. Measuring triple-helix synergy in the Russian innovation systems at regional, provincial, and national levels. Journal of the Association for Information Science and Technology, 66(6):1229–1238, June 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Leydesdorff:2014:ICR

[LPW14] Loet Leydesdorff, Han Woo Park, and Caroline Wagner. International coauthorship relations in the Social Sciences Citation Index: Is internationalization leading the Network? Journal of the Association for Information Science and Technology, 65(10):2111–2126, October 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Liu:2014:IMM

[LQ14] Xiaozhong Liu and Jian Qin. An interactive metadata model for structural, descriptive, and referential representation of scholarly output. Journal of the Association for Information Science and Technology, 65(5):964–983, May 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Liu:2014:CAD

[LR14] Yuxian Liu and Ronald Rousseau. Citation analysis and the development of science: a case study using articles by some Nobel prize winners. Journal of the Association for Information Science and Technology, 65(2):281–289, February 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Lee:2019:MRB

[LRB19] Deborah Lee, Lyn Robinson, and David Bawden. Modeling the relationship between scientific and bibliographic classification for music. Journal of the Association for Information Science and Technology, 70(3):230–241, March 2019. CO-DEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Lee:2021:ODM

[LRB21] Deborah Lee, Lyn Robinson, and David Bawden. Orthogonality, dependency, and music: an exploration of the relationships between music facets. Journal of the Association for Information Science and Technology, 72(5):570–582, May 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Li:2014:EIP

[LS14] Ting Li and Thomas Slee. The effects of information privacy concerns on digitizing personal health records. Journal of the Association for Information Science and Technology, 65(8): 1541–1554, August 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Li:2016:SBG

[LS16] Jiang Li and Dongbo Shi. Sleeping beauties in genius work: When were they awakened? Journal of the Association for Information Science and Technology, 67(2):432–440, February 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Li:2017:EFG

[LS17] Chenliang Li and Aixin Sun. Extracting fine-grained location with temporal awareness in tweets: a two-stage approach. Journal of the Association for Information Science and Technology, 68(7):1652–1670, July 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Le:2018:RPI

[LS18] Long T. Le and Chirag Shah. Retrieving people: Identifying potential answerers in Community Question–Answering. Journal of the Association for Information Science and Technology, 69(10):1246–1258, October 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Lee:2021:USM

[LS21] Grace E. Lee and Aixin Sun. Understanding the stability of medical concept embeddings. Journal of the Association for Information Science and Technology, 72(3):346–356, March 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Lee:2022:CMM

[LS22] Hanseul Stephanie Lee and Catherine Arnott Smith. A comparative mixed methods study on health information seeking among US-born/US-dwelling, Korean-born/US-dwelling, and Korean-born/Korean-dwelling mothers. Journal of the Association for Information Science and Technology, 73(7):929–943, July 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Liu:2018:PCE

[LSC18] Yunzhong Liu, Jinhe Shi, and Yi Chen. Patient-centered and experience-aware mining for effective adverse drug reaction discovery in online health forums. Journal of the Association for Information Science and Technology, 69(2):215–228, February 2018. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Lee:2023:SPR

[LSHH23] Dong Joon Lee, Besiki Stvilia, Seungyeon Ha, and Douglas Hahn. The structure and priorities of researchers’ scholarly profile maintenance activities: a case of institutional research information management system. Journal of the Association for Information Science and Technology, 74(2): 186–204, February 2023. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Li:2023:EKG

[LSL23] Guangtong Li, L. Siddharth, and Jianxi Luo. Embedding knowledge graph of patent metadata to measure knowledge proximity. Journal of the Association for Information Science and Technology, 74(4):476–490, April 2023. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Lariviere:2014:APJ

[LSM^{+}14] Vincent Larivière, Cassidy R. Sugimoto, Benoit Macaluso, Staša Milojević, Blaise Cronin, and Mike Thelwall. arXiv e-prints and the journal of record: an analysis of roles and relationships. Journal of the Association for Information Science and Technology, 65(6):1157–1169, June 2014. CO-DEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Li:2018:DDD

[LSX18] Jing Li, Aixin Sun, and Zhenchang Xing. To do or not to do: Distill crowdsourced negative caveats to augment API documentation. Journal of the Association for Information Science and Technology, 69(12):1460–1475, December 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Li:2015:TLO

[LTD^{+}15] Daifeng Li, Jie Tang, Ying Ding, Xin Shuai, Tamy Chambers, Guozheng Sun, Zhipeng Luo, and Jingwei Zhang. Topic-level opinion influence model (TOIM): an investigation using Tencent microblogging. Journal of the Association for Information Science and Technology, 66(12):2657–2673, December 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Lueg:2015:OPM

[Lue15] Christopher P. Lueg. Opinion piece: The missing link: Information behavior research and its estranged relationship with embodiment. Journal of the Association for Information Science and Technology, 66(12):2704–2707, December 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Lueg:2020:EQ

[Lue20] Christopher Lueg. To be or not to be (embodied): That is not the question. Journal of the Association for Information Science and Technology, 71(1):114–117, January 2020. CO-DEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Luyt:2016:WCM

[Luy16] Brendan Luyt. Wikipedia, collective memory, and the Vietnam War. Journal of the Association for Information Science and Technology, 67(8):1956–1961, August 2016. CO-DEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Lee:2017:UUC

[LWA^{+}17] Jin Ha Lee, Rachel Wishkoski, Lara Aase, Perry Meas, and Chris Hubbles. Understanding users of cloud music services: Selection factors, management and access behavior, and perceptions. Journal of the Association for Information Science and Technology, 68(5):1186–1200, May 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Leydesdorff:2016:LER

[LWB16] Loet Leydesdorff, Caroline Wagner, and Lutz Bornmann. Letters to the Editor: Replicability and the public/private divide. Journal of the Association for Information Science and Technology, 67(7):1777–1778, July 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Li:2018:PPS

[LWL^{+}18] Hanhui Li, Hefeng Wu, Donghui Li, Shujin Lin, Zhuo Su, and Xiaonan Luo. PSI: A probabilistic semantic interpretable framework for fine-grained image ranking. Journal of the Association for Information Science and Technology, 69(12): 1488–1501, December 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Lim:2018:CVP

[LWLH18] Sesil Lim, JongRoul Woo, Jongsu Lee, and Sung-Yoon Huh. Consumer valuation of personal information in the age of big data. Journal of the Association for Information Science and Technology, 69(1):60–71, January 2018. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Li:2019:ASM

[LWM^{+}19] Daifeng Li, Yintian Wang, Andrew Madden, Ying Ding, Jie Tang, Gordon Guozheng Sun, Ning Zhang, and Enguo Zhou. Analyzing stock market trends using social media user moods and social influence. Journal of the Association for Information Science and Technology, 70(9):1000–1013, September 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Lund:2023:CNA

[LWM^{+}23] Brady D. Lund, Ting Wang, Nishith Reddy Mannuru, Bing Nie, Somipam Shimray, and Ziang Wang. ChatGPT and a new academic reality: Artificial intelligence-written research papers and the ethics of the large language models in scholarly publishing. Journal of the Association for Information Science and Technology, 74(5):570–581, May 2023. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Leydesdorff:2019:SKB

[LWPG^{+}19] Loet Leydesdorff, Caroline S. Wagner, Igone Porto-Gomez, Jordan A. Comins, and Fred Phillips. Synergy in the knowledge base of U.S. innovation systems at national, state, and regional levels: The contributions of high-tech manufacturing and knowledge-intensive services. Journal of the Association for Information Science and Technology, 70(10):1108– 1123, October 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Liu:2017:LCA

[LX17] Yang Liu and Songhua Xu. A local context-aware LDA model for topic modeling in a document network. Journal of the Association for Information Science and Technology, 68(6):1429–1448, June 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Li:2014:NAP

[LXLL14] Yuhua Li, Songhua Xu, Xiaonan Luo, and Shujin Lin. A new algorithm for product image search based on salient edge characterization. Journal of the Association for Information Science and Technology, 65(12):2534–2551, December 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Lu:2018:IUP

[LXW18] Tian Lu, Yunjie (Calvin) Xu, and Scott Wallace. Internet usage and patient’s trust in physician during diagnoses: a knowledge power perspective. Journal of the Association for Information Science and Technology, 69(1):110–120, January 2018. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Liu:2022:MBE

[LYC^{+}22] Qian Liu, Zhengfa Yang, Xiaofang Cai, Qianzhou Du, and Weiguo Fan. The more, the better? The effect of feedback and user’s past successes on idea implementation in open innovation communities. Journal of the Association for Information Science and Technology, 73(3):376–392, March 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Liu:2014:WYP

[LYCX14] Bo Liu, Quan Yuan, Gao Cong, and Dong Xu. Where your photo is taken: Geolocation prediction for social images. Journal of the Association for Information Science and Technology, 65(6):1232–1243, June 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Liew:2021:DIK

[LYL21] Chern Li Liew, Jamie Yeates, and Spencer Charles Lilley. Digitized indigenous knowledge collections: Impact on cultural knowledge transmission, social connections, and cultural identity. Journal of the Association for Information Science and Technology, 72(12):1575–1592, December 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Luo:2015:STI

[LYOW15] Zhunchen Luo, Yang Yu, Miles Osborne, and Ting Wang. Structuring Tweets for improving Twitter search. Journal of the Association for Information Science and Technology, 66 (12):2522–2539, December 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Lin:2014:DSR

[LYWF14] Weijane Lin, Hsiu-Ping Yueh, Hsin-Ying Wu, and Li-Chen Fu. Developing a service robot for a children’s library: a design-based research approach. Journal of the Association for Information Science and Technology, 65(2):290–301, February 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Liu:2019:RDK

[LZ19] Jingjing Liu and Xiangmin Zhang. The role of domain knowledge in document selection from search results. Journal of the Association for Information Science and Technology, 70(11):1236–1247, November 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Liu:2021:MTP

[LZ21] Jun Liu and Jingyi Zhao. More than plain text: Censorship deletion in the Chinese social media. Journal of the Association for Information Science and Technology, 72(1): 18–31, January 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Lee:2022:SME

[LZ22] Chang Heon Lee and J. Leon Zhao. Social media engagement and crowdfunding performance: The moderating role of product type and entrepreneurs’ characteristics. Journal of the Association for Information Science and Technology, 73(11):1559–1578, November 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Lu:2020:CCN

[LZA^{+}20] Chao Lu, Yingyi Zhang, Yong-Yeol Ahn, Ying Ding, Chenwei Zhang, and Dandan Ma. Co-contributorship network and division of labor in individual scientific collaborations. Journal of the Association for Information Science and Technology, 71(10):1162–1178, October 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Leginus:2016:PGW

[LZD16] Martin Leginus, ChengXiang Zhai, and Peter Dolog. Personalized generation of word clouds from tweets. Journal of the Association for Information Science and Technology, 67 (5):1021–1032, May 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Liu:2023:ASM

[LZG^{+}23] Jiayue Liu, Ziyao Zhou, Ming Gao, Jiafu Tang, and Weiguo Fan. Aspect sentiment mining of short bullet screen comments from online TV series. Journal of the Association for Information Science and Technology, 74(8):1026–1045, August 2023. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Li:2017:UEM

[LZSW17] Jingfei Li, Peng Zhang, Dawei Song, and Yue Wu. Understanding an enriched multidimensional user relevance model by analyzing query logs. Journal of the Association for Information Science and Technology, 68(12):2743–2754, December 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Li:2021:CMR

[LZZ^{+}21] Wenjie Li, Yi Zheng, Yuejie Zhang, Rui Feng, Tao Zhang, and Weiguo Fan. Cross-modal retrieval with dual multi-angle self-attention. Journal of the Association for Information Science and Technology, 72(1):46–65, January 2021. CO-DEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Mayernik:2018:TTC

[MA18a] Matthew S. Mayernik and Amelia Acker. Tracing the traces: the critical role of metadata within networked communications. Journal of the Association for Information Science and Technology, 69(1):177–180, January 2018. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Mir:2018:BLS

[MA18b] Tariq Ahmad Mir and Marcel Ausloos. Benford’s law: a “sleeping beauty” sleeping in the dirty pages of logarithmic tables. Journal of the Association for Information Science and Technology, 69(3):349–358, March 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Ma:2021:UIA

[Ma21] Yuanye Ma. Understanding information: Adding a nonindividualistic lens. Journal of the Association for Information Science and Technology, 72(10):1295–1305, October 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Ma:2022:BRIb

[Ma22a] Lai Ma. Book review: Information: a reader. Eric Hayot, Anatoly Detwyler, Lea Pao. New York: Columbia University Press, 2021. 384 pp. $110.00 (hardcover). (ISBN: 978-0-231-18620-9). Journal of the Association for Information Science and Technology, 73(11):1659–1662, November 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Ma:2022:BRIa

[Ma22b] Lai Ma. Book review: Information: Keywords. Michele Kennerly, Samuel Frederick, and Jonathan E. Abel New York: Columbia University Press, 2021. 232 pp. $110.00 (hardcover). (ISBN: 978-0-231-19876-9). Journal of the Association for Information Science and Technology, 73(7):1058– 1061, July 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Ma:2023:IP

[Ma23a] Lai Ma. Information, platformized. Journal of the Association for Information Science and Technology, 74(2):273–282, February 2023. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Ma:2023:RCC

[Ma23b] Yuanye Ma. Relatedness and compatibility: The concept of privacy in Mandarin Chinese and American English corpora. Journal of the Association for Information Science and Technology, 74(2):249–272, February 2023. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Meseguer-Artola:2016:FIT

[MAAL^{+}16] Antoni Meseguer-Artola, Eduard Aibar, Josep Lladós, Julià Minguillón, and Maura Lerga. Factors that influence the teaching use of Wikipedia in higher education. Journal of the Association for Information Science and Technology, 67 (5):1224–1232, May 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Madsen:2016:LIM

[Mad16] Dorte Madsen. Liberating interdisciplinarity from myth. An exploration of the discursive construction of identities in information studies. Journal of the Association for Information Science and Technology, 67(11):2697–2709, November 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Mahesh:2017:LEA

[Mah17] G. Mahesh. Letter to the Editor: Academia’s recognition of Wikipedia. Journal of the Association for Information Science and Technology, 68(6):1596, June 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Mahony:2022:TOT

[Mah22] Simon Mahony. Toward openness and transparency to better facilitate knowledge creation. Journal of the Association for Information Science and Technology, 73(10):1474–1488, October 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Mak:2014:AD

[Mak14] Bonnie Mak. Archaeology of a digitization. Journal of the Association for Information Science and Technology, 65(8): 1515–1526, August 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Makri:2020:IID

[Mak20] Stephann Makri. Information informing design: Information Science research with implications for the design of digital information environments. Journal of the Association for Information Science and Technology, 71(11):1402–1412, November 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Malesios:2015:BCS

[Mal15] Chrisovalantis Malesios. Brief communications: Some variations on the standard theoretical models for the h-index: a comparative analysis. Journal of the Association for Information Science and Technology, 66(11):2384–2388, November 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Mukta:2019:TMB

[MAM19] Md. Saddam Hossain Mukta, Mohammed Eunus Ali, and Jalal Mahmud. Temporal modeling of basic human values from social network usage. Journal of the Association for Information Science and Technology, 70(2):151–163, February 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Manley:2022:LER

[Man22] Stewart Manley. Letters to the editor and the race for publication metrics. Journal of the Association for Information Science and Technology, 73(5):702–707, May 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Mishra:2015:ISU

[MAP15] Jyoti Mishra, David Allen, and Alan Pearman. Information seeking, use, and decision making. Journal of the Association for Information Science and Technology, 66(4):662–673, April 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Marino:2015:BRS

[Mar15] Mark C. Marino. Book reviews: Speaking Code: Coding as Aesthetic and Political Expression — by Geoff Cox and Alex McLean. Journal of the Association for Information Science and Technology, 66(9):1959–1962, September 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Marcinkowski:2016:DID

[Mar16] Michael Marcinkowski. Data, ideology, and the developing critical program of social informatics. Journal of the Association for Information Science and Technology, 67(5):1266– 1275, May 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Martin:2023:PPE

[Mar23] Kirsten Martin. Predatory predictions and the ethics of predictive analytics. Journal of the Association for Information Science and Technology, 74(5):531–545, May 2023. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

McNamara:2015:PPP

[MAS^{+}15] Mary McNamara, Corey Arnold, Karthik Sarma, Denise Aberle, Edward Garon, and Alex A. T. Bui. Patient portal preferences: Perspectives on imaging information. Journal of the Association for Information Science and Technology, 66 (8):1606–1615, August 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Mathiesen:2015:AIS

[Mat15] Kay Mathiesen. Advances in information science: Human rights as a topic and guide for LIS research and practice. Journal of the Association for Information Science and Technology, 66(7):1305–1322, July 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Matusiak:2020:BRM

[Mat20] Krystyna K. Matusiak. Book review: Machine translation and global research: Towards improved machine translation literacy in the scholarly community. Bowker, Lynne and Ciro, Jairo B.. Bingley, UK: Emerald Publishing, 2019. 128 pp. $95.00 (hardcover). (ISBN 978-1-78756-722-1). Journal of the Association for Information Science and Technology, 71(10):1275–1278, October 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Mathieu:2022:DKW

[Mat22] Camille Mathieu. Defining knowledge workers’ creation, description, and storage practices as impact on enterprise content management strategy. Journal of the Association for Information Science and Technology, 73(3):472–484, March 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Mayernik:2016:RDM

[May16] Matthew S. Mayernik. Research data and metadata curation as institutional issues. Journal of the Association for Information Science and Technology, 67(4):973–993, April 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Marx:2014:LEP

[MB14] Werner Marx and Lutz Bornmann. Letter to the Editor: On the problems of dealing with bibliometric data. Journal of the Association for Information Science and Technology, 65(4):866–867, April 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Makri:2020:RHI

[MB20] Stephann Makri and Lily Buckley. Down the rabbit hole: Investigating disruption of the information encountering process. Journal of the Association for Information Science and Technology, 71(2):127–142, February 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Mason:2023:TNP

[MB23] Tom Mason and David Bawden. Times new plural: The multiple temporalities of contemporary life and the infosphere. Journal of the Association for Information Science and Technology, 74(10):1159–1169, October 2023. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Martins:2016:CSA

[MBAG16] Eder F. Martins, Fabiano M. Belém, Jussara M. Almeida, and Marcos A. Gonçalves. On cold start for associative tag recommendation. Journal of the Association for Information Science and Technology, 67(1):83–105, January 2016. CO-DEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Marx:2014:DHR

[MBBL14] Werner Marx, Lutz Bornmann, Andreas Barth, and Loet Leydesdorff. Detecting the historical roots of research fields by Reference Publication Year Spectroscopy (RPYS). Journal of the Association for Information Science and Technology, 65(4):751–764, April 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Mutz:2015:TFP

[MBD15] Rüdiger Mutz, Lutz Bornmann, and Hans-Dieter Daniel. Testing for the fairness and predictive validity of research funding decisions: a multilevel multiple imputation for missing data approach using ex-ante and ex-post peer evaluation data from the Austrian science fund. Journal of the Association for Information Science and Technology, 66(11): 2321–2339, November 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Meng:2024:NCA

[MBML24] Kai Meng, Zhichao Ba, Yaxue Ma, and Gang Li. A network coupling approach to detecting hierarchical linkages between science and technology. Journal of the Association for Information Science and Technology, 75(2):167–187, February 2024. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Maruping:2017:GBI

[MBVB17] Likoebe M. Maruping, Hillol Bala, Viswanath Venkatesh, and Susan A. Brown. Going beyond intention: Integrating behavioral expectation into the unified theory of acceptance and use of technology. Journal of the Association for Information Science and Technology, 68(3):623–637, March 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Makri:2014:MMO

[MBW^{+}14] Stephann Makri, Ann Blandford, Mel Woods, Sarah Sharples, and Deborah Maxwell. “Making my own luck”: Serendipity strategies and how to support them in digital information environments. Journal of the Association for Information Science and Technology, 65(11):2179–2194, November 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Morrison:2022:CGO

[MBZ^{+}22] Heather Morrison, Luan Borges, Xuan Zhao, Tanoh Laurent Kakou, and Amit Nataraj Shanbhoug. Change and growth in open access journal publishing and charging trends 2011– 2021. Journal of the Association for Information Science and Technology, 73(12):1793–1805, December 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Mao:2018:IBB

[MC18] Jin Mao and Hong Cui. Identifying bacterial biotope entities using sequence labeling: Performance and feature analysis. Journal of the Association for Information Science and Technology, 69(9):1134–1147, September 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

McCain:2015:MFT

[McC15] Katherine W. McCain. Mining full-text journal articles to assess obliteration by incorporation: Herbert A. Simon’s concepts of bounded rationality and satisficing in economics, management, and psychology. Journal of the Association for Information Science and Technology, 66(11):2187–2201, November 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

McDowell:2021:SWS

[McD21] Kate McDowell. Storytelling wisdom: Story, information, and DIKW. Journal of the Association for Information Science and Technology, 72(10):1223–1233, October 2021. CO-DEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Mena-Chalco:2014:BBC

[MCDMM14] Jesús Pascual Mena-Chalco, Luciano Antonio Digiampietri, Fabrício Martins Lopes, and Roberto Marcondes Cesar Junior. Brazilian bibliometric coauthorship networks. Journal of the Association for Information Science and Technology, 65(7):1424–1445, July 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

McKay:2019:TWT

[MCSB19] Dana McKay, Shanton Chang, Wally Smith, and George Buchanan. The things we talk about when we talk about browsing: an empirical typology of library browsing behavior. Journal of the Association for Information Science and Technology, 70(12):1383–1394, December 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Min:2021:CCE

[MCY^{+}21] Chao Min, Qingyu Chen, Erjia Yan, Yi Bu, and Jianjun Sun. Citation cascade and the evolution of topic relevance. Journal of the Association for Information Science and Technology, 72(1):110–127, January 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Mutz:2015:BCW

[MD15] Rüdiger Mutz and Hans-Dieter Daniel. Brief communication: What is behind the curtain of the Leiden Ranking? Journal of the Association for Information Science and Technology, 66(9):1950–1953, September 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

McEwen:2016:III

[MD16] Rhonda McEwen and Adam K. Dubé. Intuitive or idiomatic: an interdisciplinary study of child-tablet computer interaction. Journal of the Association for Information Science and Technology, 67(5):1169–1181, May 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

McKeown:2016:PIS

[MDC^{+}16] Kathy McKeown, Hal Daume III, Snigdha Chaturvedi, John Paparrizos, Kapil Thadani, Pablo Barrio, Or Biran, Suvarna Bothe, Michael Collins, Kenneth R. Fleischmann, Luis Gravano, Rahul Jha, Ben King, Kevin McInerney, Taesun Moon, Arvind Neelakantan, Diarmuid O’Seaghdha, Dragomir Radev, Clay Templeton, and Simone Teufel. Predicting the impact of scientific concepts using full-text features. Journal of the Association for Information Science and Technology, 67(11):2684–2696, November 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Melo:2019:SSA

[MDJ^{+}19] Philipe F. Melo, Daniel H. Dalip, Manoel M. Junior, Marcos A. Gonçalves, and Fabrício Benevenuto. 10SENT: A stable sentiment analysis method based on the combination of off-the-shelf approaches. Journal of the Association for Information Science and Technology, 70(3):242–255, March 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Min:2018:IID

[MDL^{+}18] Chao Min, Ying Ding, Jiang Li, Yi Bu, Lei Pei, and Jianjun Sun. Innovation or imitation: The diffusion of citations. Journal of the Association for Information Science and Technology, 69(10):1271–1282, October 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Martins:2023:IOR

[MdSLdO^{+}23] Dalton Lopes Martins, Daniela Lucas da Silva Lemos, Luis Felipe Rosa de Oliveira, Joyce Siqueira, Danielle do Carmo, and Vinicius Nunes Medeiros. Information organization and representation in digital cultural heritage in Brazil: Systematic mapping of information infrastructure in digital collections for data science applications. Journal of the Association for Information Science and Technology, 74(6):707–726, June 2023. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Medina:2015:BRC

[Med15] Eden Medina. Book review: Conflicts in the Knowledge Society: The Contentious Politics of Intellectual Property — by Sebastian Haunss. Journal of the Association for Information Science and Technology, 66(4):869–871, April 2015. CO-DEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Mehra:2022:BRD

[Meh22] Bharat Mehra. Book review: The Digitally Disposed-Racial Capitalism and the Informatics of Value. Seb Franklin. Minneapolis: University of Minnesota Press, 2021. 280 pp. $27.00 (paperback). (ISBN: 978-1-5179-0715-0). Journal of the Association for Information Science and Technology, 73 (9):1356–1361, September 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Mehra:2023:TID

[Meh23] Bharat Mehra. Toward an impact-driven framework to operationalize social justice and implement ICT4D in the field of information. Journal of the Association for Information Science and Technology, 74(12):1419–1436, December 2023. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Menin:2021:CMI

[Men21] Boris Menin. Construction of a model as an information channel between the physical phenomenon and observer. Journal of the Association for Information Science and Technology, 72(9):1198–1210, September 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Marcinkowski:2016:CPE

[MF16] Michael Marcinkowski and Fred Fonseca. The conditions of peak empiricism in big data and interaction design. Journal of the Association for Information Science and Technology, 67(6):1279–1288, June 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Matthews:2021:GAM

[MG21] Paul Matthews and Kathrina Glitre. Genre analysis of movies using a topic model of plot summaries. Journal of the Association for Information Science and Technology, 72(12): 1511–1527, December 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Macias-Galindo:2015:EDM

[MGCTW15] Daniel Macias-Galindo, Lawrence Cavedon, John Thangarajah, and Wilson Wong. Effects of domain on measures of semantic relatedness. Journal of the Association for Information Science and Technology, 66(10):2116–2131, October 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Muresan:2016:INL

[MGIGW16] Smaranda Muresan, Roberto Gonzalez-Ibanez, Debanjan Ghosh, and Nina Wacholder. Identification of nonliteral language in social media: a case study on sarcasm. Journal of the Association for Information Science and Technology, 67 (11):2725–2737, November 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Menkov:2020:RPA

[MGK20] Vladimir Menkov, Paul Ginsparg, and Paul B. Kantor. Recommendations and privacy in the arXiv system: a simulation experiment using historical data. Journal of the Association for Information Science and Technology, 71(3):300– 313, March 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Mu:2016:DDC

[MGKA16] Tingting Mu, John Y. Goulermas, Ioannis Korkontzelos, and Sophia Ananiadou. Descriptive document clustering via discriminant learning in a co-embedded space of multilevel similarities. Journal of the Association for Information Science and Technology, 67(1):106–133, January 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Moed:2015:MAS

[MH15] Henk F. Moed and Gali Halevi. Multidimensional assessment of scholarly research impact. Journal of the Association for Information Science and Technology, 66(10):1988–2002, October 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Moed:2016:FTD

[MH16] Henk F. Moed and Gali Halevi. On full text download and citation distributions in scientific-scholarly journals. Journal of the Association for Information Science and Technology, 67(2):412–431, February 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Mostafa:2018:PQM

[MH18] Javed Mostafa and Deanna Morrow Hall. A proposed quantitative methodology to characterize the corporate library universe. Journal of the Association for Information Science and Technology, 69(5):760, May 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Makri:2019:III

[MHJ19] Stephann Makri, Tsui-Ling Hsueh, and Sara Jones. Ideation as an intellectual information acquisition and use context: Investigating game designers’ information-based ideation behavior. Journal of the Association for Information Science and Technology, 70(8):775–787, August 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Mayernik:2017:ATO

[MHMW17] Matthew S. Mayernik, David L. Hart, Keith E. Maull, and Nicholas M. Weber. Assessing and tracing the outcomes and impact of research infrastructures. Journal of the Association for Information Science and Technology, 68(6):1341– 1359, June 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Milard:2014:SCB

[Mil14] Béatrice Milard. The social circles behind scientific references: Relationships between citing and cited authors in chemistry publications. Journal of the Association for Information Science and Technology, 65(12):2459–2468, December 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Min:2016:PIC

[Min16] Jinyoung Min. Personal information concerns and provision in social network sites: Interplay between secure preservation and true presentation. Journal of the Association for Information Science and Technology, 67(1):26–42, January 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Miroiu:2014:LEH

[Mir14] Adrian Miroiu. Letter to the Editor: How many athletes qualified for the London Olympic Games? Journal of the Association for Information Science and Technology, 65(8): 1728, August 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Mehra:2023:DSG

[MJ23] Bharat Mehra and Baheya S. Jaber. “Don’t Say Gay” in Alabama: a taxonomic framework of LGBTQ+ information support services in public libraries — an exploratory website content analysis of critical resistance. Journal of the Association for Information Science and Technology, 74(8):954– 970, August 2023. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Min:2015:HPE

[MK15] Jinyoung Min and Byoungsoo Kim. How are people enticed to disclose personal information despite privacy concerns in social network sites? The calculus between benefit and cost. Journal of the Association for Information Science and Technology, 66(4):839–857, April 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Minkov:2017:GBR

[MKK17] Einat Minkov, Keren Kahanov, and Tsvi Kuflik. Graphbased recommendation integrating rating history and domain knowledge: Application to on-site guidance of museum visitors. Journal of the Association for Information Science and Technology, 68(8):1911–1924, August 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Mongeon:2016:CCI

[ML16] Philippe Mongeon and Vincent Larivière. Costly collaborations: the impact of scientific fraud on co-authors’ careers. Journal of the Association for Information Science and Technology, 67(3):535–542, March 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Ma:2021:ESR

[ML21] Jinxuan Ma and Brady Lund. The evolution and shift of research topics and methods in library and information science. Journal of the Association for Information Science and Technology, 72(8):1059–1074, August 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Ma:2022:DHC

[ML22] Rongqian Ma and Kai Li. Digital humanities as a crossdisciplinary battleground: an examination of inscriptions in journal publications. Journal of the Association for Information Science and Technology, 73(2):172–187, February 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

MacRoberts:2018:MSC

[MM18] Michael H. MacRoberts and Barbara R. MacRoberts. The mismeasure of science: Citation analysis. Journal of the Association for Information Science and Technology, 69(3): 474–482, March 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Manzuch:2020:GRR

[MM20] Zinaida Manzuch and Elena Maceviciute. Getting ready to reduce the digital divide: Scenarios of Lithuanian public libraries. Journal of the Association for Information Science and Technology, 71(10):1205–1217, October 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Martin-Martin:2015:LEN

[MMALCOM15] Alberto Martín-Martín, Juan M. Ayllón, Emilio Delgado López-Cózar, and Enrique Orduna-Malea. Letters to the Editor: Nature’s top 100 re-revisited. Journal of the Association for Information Science and Technology, 66(12):2714, December 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Maemura:2017:OAF

[MMB17] Emily Maemura, Nathan Moles, and Christoph Becker. Organizational assessment frameworks for digital preservation: a literature review and mapping. Journal of the Association for Information Science and Technology, 68(7):1619–1637, July 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Mann:2020:BBT

[MMFA20] Monique Mann, Peta Mitchell, Marcus Foth, and Irina Anastasiu. #BlockSidewalk to Barcelona: Technological sovereignty and the social license to operate smart cities. Journal of the Association for Information Science and Technology, 71(9):1103–1115, September 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Montalvo:2015:ENE

[MMFD15] Soto Montalvo, Raquel Martínez, Víctor Fresno, and Agustín Delgado. Exploiting named entities for bilingual news clustering. Journal of the Association for Information Science and Technology, 66(2):363–376, February 2015. CO-DEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Melo:2023:DFM

[MMHA23] Marijel Melo, Laura March, Kimberly Hirsh, and Emily Arnsberg. Description framework of makerspaces: Examining the relationship between spatial arrangement and diverse user populations. Journal of the Association for Information Science and Technology, 74(5):506–516, May 2023. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Moilanen:2015:VXD

[MNNK15] Katja Moilanen, Timo Niemi, Turkka Näppilä, and Mikko Kuru. A visual XML dataspace approach for satisfying ad hoc information needs. Journal of the Association for Information Science and Technology, 66(11):2304–2320, November 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Moksness:2020:PQS

[MO20] Lars Moksness and Svein Ottar Olsen. Perceived quality and self-identity in scholarly publishing. Journal of the Association for Information Science and Technology, 71(3):338–348, March 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Mesgari:2015:AIS

[MOM^{+}15] Mostafa Mesgari, Chitu Okoli, Mohamad Mehdi, Finn Årup Nielsen, and Arto Lanamäki. Advances in information science: “The sum of all human knowledge”: a systematic review of scholarly research on the content of Wikipedia. Journal of the Association for Information Science and Technology, 66(2):219–245, February 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Moore:2020:RDS

[Moo20] Samuel A. Moore. Revisiting “the 1990s debutante”: Scholar-led publishing and the prehistory of the open access movement. Journal of the Association for Information Science and Technology, 71(7):856–866, July 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Morrison:2024:MIV

[Mor24] Renee Morrison. Making the invisible visible: Critical discourse analysis as a tool for search engine research. Journal of the Association for Information Science and Technology, 75(5):600–612, May 2024. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Mostafa:2017:ESS

[Mos17] Javed Mostafa. Editorial: Sanitizing signals in scholarship and mass media: Integrity informatics I. Journal of the Association for Information Science and Technology, 68(1):3–4, January 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Mostafa:2018:EDM

[Mos18] Javed Mostafa. Editorial: Documents and (as) machines. Journal of the Association for Information Science and Technology, 69(1):3–5, January 2018. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Moshfeghi:2019:NMR

[MP19] Yashar Moshfeghi and Frank E. Pollick. Neuropsychological model of the realization of information need. Journal of the Association for Information Science and Technology, 70(9): 954–967, September 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Milard:2023:ECN

[MP23] Béatrice Milard and Yoann Pitarch. Egocentric cocitation networks and scientific papers destinies. Journal of the Association for Information Science and Technology, 74(4):415– 433, April 2023. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Md:2020:BRC

[MPB^{+}20] Paul Kudlow Md, PhD, Devin Bissky Dziadyk Bsc, Alan Rutledge Bsc, Aviv Shachak PhD, Gunther Eysenbach Md, and Mph. Book review: The citation advantage of promoted articles in a cross-publisher distribution platform: a 12-month randomized controlled trial. Journal of the Association for Information Science and Technology, 71(10): 1257–1274, October 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

McCay-Peet:2015:ISH

[MPT15] Lori McCay-Peet and Elaine G. Toms. Investigating serendipity: How it unfolds and what may influence it. Journal of the Association for Information Science and Technology, 66(7):1463–1476, July 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Mayor:2014:ORC

[MR14] Charlie Mayor and Lyn Robinson. Ontological realism and classification: Structures and concepts in the Gene Ontology. Journal of the Association for Information Science and Technology, 65(4):686–697, April 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

McNeese:2017:RTC

[MR17] Nathan J. McNeese and Madhu C. Reddy. The role of team cognition in collaborative information seeking. Journal of the Association for Information Science and Technology, 68 (1):129–140, January 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

McCoy:2019:UUS

[MR19] Chase McCoy and Howard Rosenbaum. Uncovering unintended and shadow practices of users of decision support system dashboards in higher education institutions. Journal of the Association for Information Science and Technology, 70(4):370–384, April 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Martinez-Romo:2016:SEK

[MRAF16] Juan Martinez-Romo, Lourdes Araujo, and Andres Duque Fernandez. SemGraph: Extracting keyphrases following a novel semantic graph-based approach. Journal of the Association for Information Science and Technology, 67(1):71– 82, January 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Moohebat:2015:III

[MRKT15] Mohammadreza Moohebat, Ram Gopal Raj, Sameem Binti Abdul Kareem, and Dirk Thorleuchter. Identifying ISI-indexed articles by their lexical usage: a text analysis approach. Journal of the Association for Information Science and Technology, 66(3):501–511, March 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Montejo-Raez:2014:KBA

[MRMCMVUL14] Arturo Montejo-Ráez, Eugenio Martínez-Cámara, M. Teresa Martín-Valdivia, and L. Alfonso Ureña-López. A knowledgebased approach for polarity classification in Twitter. Journal of the Association for Information Science and Technology, 65(2):414–425, February 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Martin:2016:WEM

[MS16] Kirsten Martin and Katie Shilton. Why experience matters to privacy: How context-based experience moderates consumer privacy expectations for mobile applications. Journal of the Association for Information Science and Technology, 67(8):1871–1882, August 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Mizzaro:2019:BRM

[MS19] Stefano Mizzaro and Ivan Scagnetto. Book review: Mobile Search Behaviors: an In-Depth Analysis Based on Contexts, APPs, and Devices. Dan Wu and Shaobo Liang. Synthesis Lectures on Information Concepts, Retrieval, and Services. San Rafael, CA: Morgan & Claypool, 2018. 159 pp. $51.96 (ebook). (ISBN 978-1-68173-300-5). Journal of the Association for Information Science and Technology, 70(11):1290–1292, November 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Malik:2021:PUP

[MSB21] Nishant Malik, David Spencer, and Quang Neo Bui. Power in the U.S. political economy: a network analysis. Journal of the Association for Information Science and Technology, 72(7):811–823, July 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Machado:2019:SWW

[MSdGS19] Luís Miguel Oliveira Machado, Renato Rocha Souza, and Maria da Graça Simões. Semantic Web or Web of Data? A diachronic study (1999 to 2017) of the publications of Tim Berners-Lee and the World Wide Web Consortium. Journal of the Association for Information Science and Technology, 70(7):701–714, July 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Muneer:2019:CCL

[MSI^{+}19] Iqra Muneer, Muhammad Sharjeel, Muntaha Iqbal, Rao Muhammad Adeel Nawab, and Paul Rayson. CLEU — a cross-language English–Urdu corpus and benchmark for text reuse experiments. Journal of the Association for Information Science and Technology, 70(7):729–741, July 2019. CO-DEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Malo:2014:GDB

[MSK^{+}14] Pekka Malo, Ankur Sinha, Pekka Korhonen, Jyrki Wallenius, and Pyry Takala. Good debt or bad debt: Detecting semantic orientations in economic texts. Journal of the Association for Information Science and Technology, 65(4): 782–796, April 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

McDonald:2021:DOM

[MSM^{+}21] Claire McDonald, Marc Schmalz, Allee Monheim, Stephen Keating, Kelsey Lewin, Frank Cifaldi, and Jin Ha Lee. Describing, organizing, and maintaining video game development artifacts. Journal of the Association for Information Science and Technology, 72(5):540–553, May 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

MacKrill:2021:WMI

[MSPP21] Kate MacKrill, Connor Silvester, James W. Pennebaker, and Keith J. Petrie. What makes an idea worth spreading? language markers of popularity in TED talks by academics and other speakers. Journal of the Association for Information Science and Technology, 72(8):1028–1038, August 2021. CO-DEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Morrow:2022:ESC

[MSTP^{+}22] Garrett Morrow, Briony Swire-Thompson, Jessica Montgomery Polny, Matthew Kopec, and John P. Wihbey. The emerging science of content labeling: Contextualizing social media content moderation. Journal of the Association for Information Science and Technology, 73(10):1365–1386, October 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Meyer:2019:SIK

[MSW^{+}19] Eric T. Meyer, Kalpana Shankar, Matthew Willis, Sarika Sharma, and Steve Sawyer. The social informatics of knowledge. Journal of the Association for Information Science and Technology, 70(4):307–312, April 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Mohammadi:2014:MRA

[MT14] Ehsan Mohammadi and Mike Thelwall. Mendeley readership altmetrics for the social sciences and humanities: Research evaluation and knowledge flows. Journal of the Association for Information Science and Technology, 65(8):1627–1638, August 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Maflahi:2016:WRC

[MT16] Nabeil Maflahi and Mike Thelwall. When are readership counts as useful as citation counts? Scopus versus Mendeley for LIS journals. Journal of the Association for Information Science and Technology, 67(1):191–199, January 2016. CO-DEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Maflahi:2018:HQD

[MT18a] Nabeil Maflahi and Mike Thelwall. How quickly do publications get read? The evolution of Mendeley reader counts for new articles. Journal of the Association for Information Science and Technology, 69(1):158–167, January 2018. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Milard:2018:CST

[MT18b] Béatrice Milard and Ludovic Tanguy. Citations in scientific texts: Do social relations matter? Journal of the Association for Information Science and Technology, 69(11):1380–1395, November 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Makri:2020:CEM

[MT20] Stephann Makri and Sophie Turner. “I can’t express my thanks enough”: The “gratitude cycle” in online communities. Journal of the Association for Information Science and Technology, 71(5):503–515, May 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Mohammadi:2015:WRR

[MTHL15] Ehsan Mohammadi, Mike Thelwall, Stefanie Haustein, and Vincent Larivière. Who reads research articles? An altmetrics analysis of Mendeley user categories. Journal of the Association for Information Science and Technology, 66(9):

1832–1846, September 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Mohammadi:2016:CMB

[MTK16] Ehsan Mohammadi, Mike Thelwall, and Kayvan Kousha. Can Mendeley bookmarks reflect readership? A survey of user motivations. Journal of the Association for Information Science and Technology, 67(5):1198–1209, May 2016. CO-DEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Moulahi:2014:IMR

[MTY14] Bilel Moulahi, Lynda Tamine, and Sadok Ben Yahia. iAggregator: Multidimensional relevance aggregation based on a fuzzy operator. Journal of the Association for Information Science and Technology, 65(10):2062–2083, October 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Miah:2022:SMA

[MVA22] Shah Jahan Miah, Huy Quan Vu, and Damminda Alahakoon. A social media analytics perspective for humanoriented smart city planning and management. Journal of the Association for Information Science and Technology, 73 (1):119–135, January 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Meadowbrooke:2014:IBH

[MVL^{+}14] Chrysta C. Meadowbrooke, Tiffany C. Veinot, Jimena Loveluck, Andrew Hickok, and José A. Bauermeister. Information behavior and HIV testing intentions among young men at risk for HIV/AIDS. Journal of the Association for Information Science and Technology, 65(3):609–620, March 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Mutz:2017:EVI

[MWD17] Rüdiger Mutz, Tobias Wolbring, and Hans-Dieter Daniel. The effect of the “very important paper” (VIP) designation in Angewandte Chemie International Edition on citation impact: a propensity score matching analysis. Journal of the Association for Information Science and Technology, 68(9): 2139–2153, September 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Maemura:2018:ICC

[MWMB18] Emily Maemura, Nicholas Worby, Ian Milligan, and Christoph Becker. If these crawls could talk: Studying and documenting web archives provenance. Journal of the Association for Information Science and Technology, 69(10): 1223–1233, October 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Moskovitch:2017:EJS

[MWPF17] Robert Moskovitch, Fei Wang, Jian Pei, and Carol Friedman. Editorial: JASIST special issue on biomedical information retrieval. Journal of the Association for Information Science and Technology, 68(11):2525–2528, November 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Miah:2015:CIP

[MYK15] Md. Waliur Rahman Miah, John Yearwood, and Siddhivinayak Kulkarni. Constructing an inter-post similarity measure to differentiate the psychological stages in offensive chats. Journal of the Association for Information Science and Technology, 66(5):1065–1081, May 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Ming:2022:ROA

[MZ22] Wei Ming and Zhenyue Zhao. Rethinking the open access citation advantage: Evidence from the “reverse-flipping” journals. Journal of the Association for Information Science and Technology, 73(11):1608–1620, November 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Mercun:2017:PBF

[MŽA17] Tanja Merčun, Maja Žumer, and Trond Aalberg. Presenting bibliographic families using information visualization: Evaluation of FRBR-based prototype and hierarchical visualizations. Journal of the Association for Information Science and Technology, 68(2):392–411, February 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Nobarany:2015:UPS

[NB15] Syavash Nobarany and Kellogg S. Booth. Use of politeness strategies in signed open peer review. Journal of the Association for Information Science and Technology, 66(5):1048– 1064, May 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Nobarany:2017:USA

[NB17] Syavash Nobarany and Kellogg S. Booth. Understanding and supporting anonymity policies in peer review. Journal of the Association for Information Science and Technology, 68(4): 957–971, April 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Nakash:2023:ECI

[NB23] Maayan Nakash and Dan Bouhnik. The effects of COVID-19 on information management in remote and hybrid work environments. Journal of the Association for Information Science and Technology, 74(9):1067–1080, September 2023. CO-DEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Nobarany:2016:WMP

[NBH16] Syavash Nobarany, Kellogg S. Booth, and Gary Hsieh. What motivates people to review articles? The case of the humancomputer interaction community. Journal of the Association for Information Science and Technology, 67(6):1358–1371, June 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Nelson:2017:CMM

[NE17] Gregory M. Nelson and Dennis L. Eggett. Citations, mandates, and money: Author motivations to publish in chemistry hybrid open access journals. Journal of the Association for Information Science and Technology, 68(10):2501–2510, October 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Newell:2023:SIP

[New23] Bryce Clayton Newell. Surveillance as information practice. Journal of the Association for Information Science and Technology, 74(4):444–460, April 2023. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Nicolaisen:2015:BCB

[NF15] Jeppe Nicolaisen and Tove Faber Frandsen. Brief communication: Bibliometric evolution: Is the journal of the association for information science and technology transforming into a specialty Journal? Journal of the Association for Information Science and Technology, 66(5):1082–1085, May 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Najaftorkaman:2015:TAU

[NGTKR15] Mohammadreza Najaftorkaman, Amir Hossein Ghapanchi, Amir Talaei-Khoei, and Pradeep Ray. A taxonomy of antecedents to user adoption of health information systems: a synthesis of thirty years of research. Journal of the Association for Information Science and Technology, 66(3):576–598, March 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Niu:2015:AIP

[NH15] Xi Niu and Bradley Hemminger. Analyzing the interaction patterns in a faceted search interface. Journal of the Association for Information Science and Technology, 66(5):1030– 1047, May 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Ntuli:2015:DRO

[NILCP15] Hamilton Ntuli, Roula Inglesi-Lotz, Tsangyao Chang, and Anastassios Pouris. Does research output cause economic growth or vice versa? Evidence from 34 OECD countries. Journal of the Association for Information Science and Technology, 66(8):1709–1716, August 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Nikolov:2019:QBO

[NLFM19] Dimitar Nikolov, Mounia Lalmas, Alessandro Flammini, and Filippo Menczer. Quantifying biases in online information exposure. Journal of the Association for Information Science and Technology, 70(3):218–229, March 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Nov:2016:BCU

[NLP16] Oded Nov, Jeffrey Laut, and Maurizio Porfiri. Brief communications: Using targeted design interventions to encourage extra-role crowdsourcing behavior. Journal of the Association for Information Science and Technology, 67(2):483–489, February 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Nori:2020:WSN

[NPG^{+}20] Raffaella Nori, Massimiliano Palmiero, Fiorella Giusberti, Elisa Gambetti, and Laura Piccardi. Web searching and navigation: Age, intelligence, and familiarity. Journal of the Association for Information Science and Technology, 71 (8):902–915, August 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Naik:2017:LVP

[NSK17] Jatothu Brahmaiah Naik, Chanamallu Srinivasarao, and Giri Babu Kande. Local vector pattern with global index angles for a content-based image retrieval system. Journal of the Association for Information Science and Technology, 68 (12):2755–2770, December 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Nemer:2017:PEI

[NT17a] David Nemer and Michail Tsikerdekis. Political engagement and ICTs: Internet use in marginalized communities. Journal of the Association for Information Science and Technology, 68(6):1539–1550, June 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Nichols:2017:MO

[NT17b] David M. Nichols and Michael B. Twidale. Metrics for openness. Journal of the Association for Information Science and Technology, 68(4):1048–1060, April 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Nakayama:2019:SIM

[NTNP19] Shinnosuke Nakayama, Tyrone J. Tolbert, Oded Nov, and Maurizio Porfiri. Social information as a means to enhance engagement in citizen science-based telerehabilitation. Journal of the Association for Information Science and Technology, 70(6):587–595, June 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Ninkov:2017:WAO

[NV17] Anton Ninkov and Liwen Vaughan. A webometric analysis of the online vaccination debate. Journal of the Association for Information Science and Technology, 68(5):1285–1294, May 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Ortega:2014:MAS

[OA14] José Luis Ortega and Isidro F. Aguillo. Microsoft Academic search and Google Scholar citations: Comparative analysis of author profiles. Journal of the Association for Information Science and Technology, 65(6):1149–1156, June 2014. CO-DEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Ozdemiray:2015:ESR

[OA15] Ahmet Murat Ozdemiray and Ismail Sengor Altingovde. Explicit search result diversification using score and rank aggregation methods. Journal of the Association for Information Science and Technology, 66(6):1212–1228, June 2015. CO-DEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

OBrien:2017:ALO

[O’B17] Heather L. O’Brien. Antecedents and learning outcomes of online news engagement. Journal of the Association for Information Science and Technology, 68(12):2809–2820, December 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Oh:2019:SWD

[OB19] Chi Young Oh and Brian Butler. Small worlds in a distant land: International newcomer students’ local information behaviors in unfamiliar environments. Journal of the Association for Information Science and Technology, 70(10):1060– 1073, October 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Oberbichler:2022:IIW

[OBD^{+}22] Sarah Oberbichler, Emanuela Boros, Antoine Doucet, Jani Marjanen, Eva Pfanzelter, Juha Rautiainen, Hannu Toivonen, and Mikko Tolonen. Integrated interdisciplinary workflows for research on historical newspapers: Perspectives from humanities scholars, computer scientists, and librarians. Journal of the Association for Information Science and Technology, 73(2):225–239, February 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

OConnor:2016:BRB

[O’C16] Daniel O’Connor. Book reviews: Beyond Bibliometrics: Harnessing Multidimensional Indicators of Scholarly Intent — Edited by Blaise Cronin and Cassidy R. Sugimoto. Journal of the Association for Information Science and Technology, 67(7):1780–1783, July 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Osterlund:2019:DAK

[ØC19] Carsten Østerlund and Kevin Crowston. Documentation and access to knowledge in online communities: Know your audience and write appropriately? Journal of the Association for Information Science and Technology, 70(6):619–633, June 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Ou:2016:SGN

[OD16] Carol X. J. Ou and Robert M. Davison. Shaping guanxi networks at work through instant messaging. Journal of the Association for Information Science and Technology, 67 (5):1153–1168, May 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Ostermaier-Grabow:2019:CSP

[OGL19] Anika Ostermaier-Grabow and Stephanie B. Linek. Communication and self-presentation behavior on academic social networking sites: an exploratory case study on profiles and discussion threads on ResearchGate. Journal of the Association for Information Science and Technology, 70(10): 1153–1164, October 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Oh:2017:TPI

[Oh17] Kyong Eun Oh. Types of personal information categorization: Rigid, fuzzy, and flexible. Journal of the Association for Information Science and Technology, 68(6):1491–1504, June 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

OHanlon:2018:BRD

[O’H18] Shane O’Hanlon. Book review: Disability, Human Rights, and Information Technology. Edited by Jonathan Lazar and Michael Ashley Stein. Philadelphia, PA: University of Pennsylvania Press, 2017. 360 pp. $69.95 (hardcover). (ISBN 978-0-8122-4923-1). Journal of the Association for Information Science and Technology, 69(7):953–955, July 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Osatuyi:2016:SBL

[OHP16] Babajide Osatuyi, Starr Roxanne Hiltz, and Katia Passerini. Seeing is believing (or at least changing your mind): the influence of visibility and task complexity on preference changes in computer-supported team decision making. Journal of the Association for Information Science and Technology, 67(9):2090–2104, September 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Oh:2016:BWI

[OHT^{+}16] Kyong Eun Oh, Daniel Halpern, Marilyn Tremaine, James Chiang, Deborah Silver, and Karen Bemis. Blocked: When the information is hidden by the visualization. Journal of the Association for Information Science and Technology, 67 (5):1033–1051, May 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Osterlund:2021:AIW

[ØJW^{+}21] Carsten Østerlund, Mohammad Hossein Jarrahi, Matthew Willis, Karen Boyd, and Christine T. Wolf. Artificial intelligence and the world of work, a co-constitutive relationship. Journal of the Association for Information Science and Technology, 72(1):128–135, January 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Oguz:2016:BCU

[OK16] Fatih Oguz and Wallace Koehler. Brief communications: URL decay at year 20: a research note. Journal of the Association for Information Science and Technology, 67(2):477– 479, February 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

ONeill:2017:AW

[OKB17] Edward T. O’Neill, Kerre A. Kammerer, and Rick Bennett. The aboutness of words. Journal of the Association for Information Science and Technology, 68(10):2471–2483, October 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Oliphant:2021:EIR

[Oli21] Tami Oliphant. Emerging (information) realities and epistemic injustice. Journal of the Association for Information Science and Technology, 72(8):951–962, August 2021. CO-DEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Olson:2016:BRI

[Ols16] Hope A. Olson. Book reviews: Indexing It All: The [Subject] in the Age of Documentation, Information, and Data. Ronald E. Day. Cambridge, MA: MIT Press, 2014. 184 pp. $30 (hardcover). (ISBN 978-0-262-02821-9). Journal of the Association for Information Science and Technology, 67(7): 1784–1786, July 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Oltmann:2015:DCP

[Olt15] Shannon M. Oltmann. Data, censorship, and politics: Analyzing the restricted flow of information in federal scientific policy development. Journal of the Association for Information Science and Technology, 66(1):144–161, January 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Orduna-Malea:2015:BRW

[OM15] Enrique Orduña-Malea. Book review: Web Metrics for Library and Information Professionals — Edited by David Stuart. Journal of the Association for Information Science and Technology, 66(11):2392–2395, November 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Okoli:2014:AIS

[OMM^{+}14] Chitu Okoli, Mohamad Mehdi, Mostafa Mesgari, Finn Årup Nielsen, and Arto Lanamäki. Advances in information science: Wikipedia in the eyes of its beholders: a systematic review of scholarly research on Wikipedia readers and readership. Journal of the Association for Information Science and Technology, 65(12):2381–2403, December 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Orduna-Malea:2017:WCP

[OMTK17] Enrique Orduna-Malea, Mike Thelwall, and Kayvan Kousha. Web citations in patents: Evidence of technological impact? Journal of the Association for Information Science and Technology, 68(8):1967–1974, August 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Orduna-Malea:2015:HET

[OMTSLC15] Enrique Orduña-Malea, Daniel Torres-Salinas, and Emilio Delgado López-Cózar. Hyperlinks embedded in Twitter as a proxy for total external in-links to international university websites. Journal of the Association for Information Science and Technology, 66(7):1447–1462, July 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Oh:2023:SAS

[ONZ23] Hanseok Oh, Seojin Nam, and Yongjun Zhu. Structured abstract summarization of scientific articles: Summarization using full-text section information. Journal of the Association for Information Science and Technology, 74(2):234–248, February 2023. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Ofek:2015:BCC

[OR15] Nir Ofek and Lior Rokach. Brief communication: A classifier to determine which Wikipedia biographies will be accepted. Journal of the Association for Information Science and Technology, 66(1):213–218, January 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Oduntan:2021:PPB

[OR21] Olubukola Oduntan and Ian Ruthven. People and places: Bridging the information gaps in refugee integration. Journal of the Association for Information Science and Technology, 72(1):83–96, January 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Ortega:2022:CAP

[Ort22] José Luis Ortega. Classification and analysis of PubPeer comments: How a web journal club is used. Journal of the Association for Information Science and Technology, 73(5):

655–670, May 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Oh:2015:MSI

[OS15] Sanghee Oh and Sue Yeon Syn. Motivations for sharing information and social support in social media: a comparative analysis of Facebook, Twitter, Delicious, YouTube, and Flickr. Journal of the Association for Information Science and Technology, 66(10):2045–2060, October 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Organisciak:2022:GSL

[OSD22] Peter Organisciak, Benjamin M. Schmidt, and J. Stephen Downie. Giving shape to large digital libraries through exploratory data analysis. Journal of the Association for Information Science and Technology, 73(2):317–332, February 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Olensky:2016:ECM

[OSvE16] Marlies Olensky, Marion Schmidt, and Nees Jan van Eck. Evaluation of the citation matching algorithms of CWTS and iFQ in comparison to the Web of Science. Journal of the Association for Information Science and Technology, 67 (10):2550–2564, October 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Ossenblok:2014:CJA

[OVE14] Truyken L. B. Ossenblok, Frederik T. Verleysen, and Tim C. E. Engels. Coauthorship of journal articles and book chapters in the social sciences and humanities (2000–2010). Journal of the Association for Information Science and Technology, 65(5):882–897, May 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Owen:2021:BRW

[Owe21] Victoria Owen. Book review: Warner, Julian. Copyright, data and creativity in the digital age: a journey through Feist. Abingdon: Routledge, 2020, 178 pp. £96.00 (hardback). (ISBN 978-0-367-90285-8). Journal of the Association for Information Science and Technology, 72(9):1211–1214, September 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Onodera:2015:FAC

[OY15] Natsuo Onodera and Fuyuki Yoshikane. Factors affecting citation rates of research articles. Journal of the Association for Information Science and Technology, 66(4):739–764, April 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Paltoglou:2016:SBE

[Pal16] Georgios Paltoglou. Sentiment-based event detection in Twitter. Journal of the Association for Information Science and Technology, 67(7):1576–1587, July 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Pan:2021:BRS

[Pan21] Xuwei Pan. Book review: Social tagging in a linked data environment. Edited by Diane Rasmussen Pennington and Louise F. Spiteri. London, UK: Facet Publishing, 2018. 240 pp. £74.95 (paperback). (ISBN 978-1-78330-338-0). Journal of the Association for Information Science and Technology, 72(2):264–265, February 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Park:2019:UCS

[Par19] Min Sook Park. Understanding characteristics of semantic associations in health consumer generated knowledge representation in social media. Journal of the Association for Information Science and Technology, 70(11):1210–1222, November 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Pontis:2015:UIE

[PB15] Sheila Pontis and Ann Blandford. Understanding “influence:” an exploratory study of academics’ processes of knowledge construction through iterative and interactive information seeking. Journal of the Association for Information Science and Technology, 66(8):1576–1593, August 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Pontis:2016:UIE

[PB16] Sheila Pontis and Ann Blandford. Understanding “influence”: an empirical test of the Data-Frame Theory of Sensemaking. Journal of the Association for Information Science and Technology, 67(4):841–858, April 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Pontis:2017:KDA

[PBG^{+}17] Sheila Pontis, Ann Blandford, Elke Greifeneder, Hesham Attalla, and David Neal. Keeping up to date: an academic researcher’s information journey. Journal of the Association for Information Science and Technology, 68(1):22–35, January 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Pekar:2020:EDH

[PBN^{+}20] Viktor Pekar, Jane Binner, Hossein Najafi, Chris Hale, and Vincent Schmidt. Early detection of heterogeneous disaster events using social media. Journal of the Association for Information Science and Technology, 71(1):43–54, January 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Pozo:2022:BTF

[PdlRR^{+}22] Álvaro Pérez Pozo, Javier de la Rosa, Salvador Ros, Elena González-Blanco, Laura Hernández, and Mirella de Sisto. A bridge too far for artificial intelligence?: Automatic classification of stanzas in Spanish poetry. Journal of the Association for Information Science and Technology, 73(2):258–267, February 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Pech:2022:CPS

[PDS22] Gerson Pech, Catarina Delgado, and Silvio Paolo Sorella. Classifying papers into subfields using Abstracts, Titles, Keywords and KeyWords Plus through pattern detection and optimization procedures: an application in physics. Journal of the Association for Information Science and Technology, 73(11):1513–1528, November 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Podgornik:2016:DTV

[PDŠB16] Bojana Boh Podgornik, Danica Dolničar, Andrej Šorgo, and Tomaž Bartol. Development, testing, and validation of an information literacy test (ILT) for higher education. Journal of the Association for Information Science and Technology, 67(10):2420–2436, October 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Pluye:2019:HOO

[PEG^{+}19] Pierre Pluye, Reem El Sherif, Vera Granikov, Quan Nha Hong, Isabelle Vedel, Maria Cristiane Barbosa Galvao, Francesca E. Y. Frati, Sophie Desroches, Carol Repchinsky, Benoît Rihoux, France Légaré, Bernard Burnand, Mathieu Bujold, and Roland Grad. Health outcomes of online consumer health information: a systematic mixed studies review with framework synthesis. Journal of the Association for Information Science and Technology, 70(7):643–659, July 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Peterson:2019:ECR

[Pet19] Gabriel M. Peterson. The effectiveness of correction and republication as quality control in scholarly communication — a bibliometric analysis. Journal of the Association for Information Science and Technology, 70(3):293–295, March 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Poole:2020:DDM

[PG20] Alex H. Poole and Deborah A. Garwood. Digging into data management in public-funded, international research in digital humanities. Journal of the Association for Information Science and Technology, 71(1):84–97, January 2020. CO-DEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Polkinghorne:2021:HIR

[PG21] Sarah Polkinghorne and Lisa M. Given. Holistic information research: From rhetoric to paradigm. Journal of the Association for Information Science and Technology, 72(10): 1261–1271, October 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Peset:2020:SAA

[PGFG^{+}20] F. Peset, F. Garzón-Farinós, Lm González, X. García-Massó, A. Ferrer-Sapena, Jl Toca-Herrera, and Ea Sánchez-Pérez. Survival analysis of author keywords: An application to the library and information sciences area. Journal of the Association for Information Science and Technology, 71(4):462– 473, April 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Pardi:2024:IKT

[PGK24] Georg Pardi, Steffen Gottschling, and Yvonne Kammerer. The influence of knowledge type and source reputation on preferences for website or video search results. Journal of the Association for Information Science and Technology, 75(5): 521–537, May 2024. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Potnis:2022:IPA

[PH22] Devendra Potnis and Macy Halladay. Information practices of administrators for controlling information in an online community of new mothers in rural America. Journal of the Association for Information Science and Technology, 73(11): 1621–1640, November 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Ping:2017:HMW

[PHC17] Qing Ping, Jiangen He, and Chaomei Chen. How many ways to use CiteSpace? A study of user interactive events over 14 months. Journal of the Association for Information Science and Technology, 68(5):1234–1256, May 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

PhD:2020:SCR

[PhD20] Rebecca D. Frank PhD. The social construction of risk in digital preservation. Journal of the Association for Information Science and Technology, 71(4):474–484, April 2020. CO-DEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Pan:2020:SKC

[PHH^{+}20] Min Pan, Jimmy Xiangji Huang, Tingting He, Zhiming Mao, Zhiwei Ying, and Xinhui Tu. A simple kernel co-occurrence-based enhancement for pseudo-relevance feedback. Journal of the Association for Information Science and Technology, 71(3):264–281, March 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Potnis:2023:CIE

[PHJ23] Devendra Potnis, Macy Halladay, and Sara-Elizabeth Jones. Consequences of information exchanges of vulnerable women on Facebook: an “information grounds” study informing value co-creation and ICT4D research. Journal of the Association for Information Science and Technology, 74(12): 1365–1383, December 2023. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Pan:2016:HUE

[PHZF16] Xuwei Pan, Shenglan He, Xiyong Zhu, and Qingmiao Fu. How users employ various popular tags to annotate resources in social tagging: an empirical study. Journal of the Association for Information Science and Technology, 67(5):1121– 1137, May 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Pilerot:2014:MDR

[Pil14] Ola Pilerot. Making design researchers’ information sharing visible through material objects. Journal of the Association for Information Science and Technology, 65(10):2006–2016, October 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Pang:2020:EIW

[PKA20] Natalie Pang, Stan Karanasios, and Misita Anwar. Exploring the information worlds of older persons during disasters. Journal of the Association for Information Science and Technology, 71(6):619–631, June 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Pontis:2016:ARE

[PKB^{+}16] Sheila Pontis, Genovefa Kefalidou, Ann Blandford, Jamie Forth, Stephann Makri, Sarah Sharples, Geraint Wiggins, and Mel Woods. Academics’ responses to encountered information: Context matters. Journal of the Association for Information Science and Technology, 67(8):1883–1903, August 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Paramita:2024:TIU

[PKK^{+}24] Monica Lestari Paramita, Maria Kasinidou, Styliani Kleanthous, Paolo Rosso, Tsvi Kuflik, and Frank Hopfgartner. Towards improving user awareness of search engine biases: a participatory design approach. Journal of the Association for Information Science and Technology, 75(5):581–599, May 2024. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Parapar:2020:USD

[PLPQB20] Javier Parapar, David E. Losada, Manuel A. Presedo-Quindimil, and Alvaro Barreiro. Using score distributions to compare statistical significance tests for information retrieval evaluation. Journal of the Association for Information Science and Technology, 71(1):98–113, January 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Perez-Molina:2018:RPC

[PM18] Eduardo Perez-Molina. The role of patent citations as a footprint of technology. Journal of the Association for Information Science and Technology, 69(4):610–618, April 2018. CO-DEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Pal:2014:IQE

[PMD14] Dipasree Pal, Mandar Mitra, and Kalyankumar Datta. Improving query expansion using WordNet. Journal of the Association for Information Science and Technology, 65(12): 2469–2478, December 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Platas:2021:MSC

[PMGB^{+}21] M^{a} Luisa Díez Platas, Salvador Ros Muñoz, Elena González-Blanco, Pablo Ruiz Fabo, and Elena Álvarez Mellado. Medieval Spanish (12th–15th centuries) named entity recognition and attribute annotation system based on contextual information. Journal of the Association for Information Science and Technology, 72(2):224–238, February 2021. CO-DEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Palavitsinis:2014:MQD

[PMSA14] Nikos Palavitsinis, Nikos Manouselis, and Salvador Sanchez-Alonso. Metadata quality in digital repositories: Empirical results from the cross-domain transfer of a quality assurance process. Journal of the Association for Information Science and Technology, 65(6):1202–1216, June 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Pera:2018:RBE

[PN18] Maria Soledad Pera and Yiu-Kai Ng. Recommending books to be exchanged online in the absence of wish lists. Journal of the Association for Information Science and Technology, 69(4):541–552, April 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Park:2015:EEI

[POK15] Namkee Park, Hyun Sook Oh, and Naewon Kang. Effects of ego involvement and social norms on individuals’ uploading intention on Wikipedia: a comparative study between the United States and South Korea. Journal of the Association for Information Science and Technology, 66(7):1494–1506, July 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Poole:2017:GUA

[Poo17] Alex H. Poole. “A greatly unexplored area”: Digital curation and innovation in digital humanities. Journal of the Association for Information Science and Technology, 68(7): 1772–1781, July 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Pee:2019:AIH

[PPC19] L. G. Pee, Shan L. Pan, and Lili Cui. Artificial intelligence in healthcare robots: A social informatics study of knowledge embodiment. Journal of the Association for Information Science and Technology, 70(4):351–369, April 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Pee:2020:SII

[PPLJ20] Loo Geok Pee, Shan L. Pan, Mingwei Li, and Suling Jia. Social informatics of information value cocreation: A case study of xiaomi’s online user community. Journal of the Association for Information Science and Technology, 71(4): 409–422, April 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Pepermans:2016:DSJ

[PR16] Guido Pepermans and Sandra Rousseau. The decision to submit to a journal: Another example of a valence-consistent shift? Journal of the Association for Information Science and Technology, 67(6):1372–1383, June 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Prathap:2014:BCT

[Pra14a] Gangan Prathap. Brief communication: A three-class, threedimensional bibliometric performance indicator. Journal of the Association for Information Science and Technology, 65 (7):1506–1508, July 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Prathap:2014:BCM

[Pra14b] Gangan Prathap. Brief communication: Measures for impact, consistency, and the h- and g-indices. Journal of the Association for Information Science and Technology, 65 (5):1076–1078, May 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Prathap:2014:BCZ

[Pra14c] Gangan Prathap. Brief communication: The zynergy-index and the formula for the h-index. Journal of the Association for Information Science and Technology, 65(2):426–427, February 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Prathap:2014:LEQ

[Pra14d] Gangan Prathap. Letter to the Editor: Quantity, quality, and consistency as bibliometric indicators. Journal of the Association for Information Science and Technology, 65(1): 214, January 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Prathap:2015:LEM

[Pra15a] Gangan Prathap. Letters to the Editor: Measures for bibliometric size, impact, and concentration. Journal of the Association for Information Science and Technology, 66(8): 1740–1741, August 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Pratt:2015:LEB

[Pra15b] Travis C. Pratt. Letters to the Editor: a bibliometric for the publication of inconsequential work. Journal of the Association for Information Science and Technology, 66(10): 2167, October 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Paris:2022:SOC

[PRM22] Britt Paris, Rebecca Reynolds, and Catherine McGowan. Sins of omission: Critical informatics perspectives on privacy in e-learning systems in higher education. Journal of the Association for Information Science and Technology, 73(5): 708–725, May 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Perianes-Rodriguez:2016:UCD

[PRRC16] Antonio Perianes-Rodriguez and Javier Ruiz-Castillo. University citation distributions. Journal of the Association for Information Science and Technology, 67(11):2790–2804, November 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Perianes-Rodriguez:2018:ICS

[PRRC18] Antonio Perianes-Rodriguez and Javier Ruiz-Castillo. The impact of classification systems in the evaluation of the research performance of the Leiden Ranking universities. Journal of the Association for Information Science and Technology, 69(8):1046–1053, August 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Parsons:2014:APV

[PS14a] Paul Parsons and Kamran Sedig. Adjustable properties of visual representations: Improving the quality of humaninformation interaction. Journal of the Association for Information Science and Technology, 65(3):455–482, March 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Pislyakov:2014:MER

[PS14b] Vladimir Pislyakov and Elena Shukshina. Measuring excellence in Russia: Highly cited papers, leading institutions, patterns of national and international collaboration. Journal of the Association for Information Science and Technology, 65(11):2321–2330, November 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Potha:2019:IAV

[PS19] Nektaria Potha and Efstathios Stamatatos. Improving author verification based on topic modeling. Journal of the Association for Information Science and Technology, 70(10): 1074–1088, October 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Phan:2020:CNE

[PS20] Minh C. Phan and Aixin Sun. Collective named entity recognition in user comments via parameterized label propagation. Journal of the Association for Information Science and Technology, 71(5):568–577, May 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Pinfield:2014:OAR

[PSB^{+}14] Stephen Pinfield, Jennifer Salter, Peter A. Bath, Bill Hubbard, Peter Millington, Jane H. S. Anders, and Azhar Hussain. Open-access repositories worldwide, 2005–2012: Past growth, current characteristics, and future possibilities. Journal of the Association for Information Science and Technology, 65(12):2404–2421, December 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Pinfield:2016:TCP

[PSB16] Stephen Pinfield, Jennifer Salter, and Peter A. Bath. The “total cost of publication” in a hybrid open-access environment: Institutional approaches to funding journal articleprocessing charges in combination with subscriptions. Journal of the Association for Information Science and Technology, 67(7):1751–1766, July 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Pinfield:2017:GCI

[PSB17] Stephen Pinfield, Jennifer Salter, and Peter A. Bath. A “gold-centric” implementation of open access: Hybrid journals, the “total cost of publication,” and policy development in the UK and beyond. Journal of the Association for Information Science and Technology, 68(9):2248–2263, September 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Purpura:2022:LRR

[PSS22] Alberto Purpura, Gianmaria Silvello, and Gian Antonio Susto. Learning to rank from relevance judgments distributions. Journal of the Association for Information Science and Technology, 73(9):1236–1252, September 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Patin:2021:IEP

[PSY^{+}21] Beth Patin, Melinda Sebastian, Jieun Yeon, Danielle Bertolini, and Alexandra Grimm. Interrupting epistemicide: a practical framework for naming, identifying, and ending epistemic injustice in the information professions. Journal of the Association for Information Science and Technology, 72(10):1306–1318, October 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Potnis:2021:HGI

[PT21] Devendra Potnis and Iman Tahamtan. Hashtags for gatekeeping of information on social media. Journal of the Association for Information Science and Technology, 72(10): 1234–1246, October 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Pe-Than:2017:DIM

[PTGL17] Ei Pa Pa Pe-Than, Dion Hoe-Lian Goh, and Chei Sian Lee. Does it matter how you play? The effects of collaboration and competition among players of human computation games. Journal of the Association for Information Science and Technology, 68(8):1823–1835, August 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Purves:2019:PIS

[PWK19] Ross S. Purves, Stephan Winter, and Werner Kuhn. Places in information science. Journal of the Association for Information Science and Technology, 70(11):1173–1182, November 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Pan:2016:SCD

[PYH16] Xuelian Pan, Erjia Yan, and Weina Hua. Science communication and dissemination in different cultures: an analysis of the audience for TED videos in China and abroad. Journal of the Association for Information Science and Technology, 67(6):1473–1486, June 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Park:2018:IDC

[PYW18] Hyoungjoo Park, Sukjin You, and Dietmar Wolfram. Informal data citation for data sharing and reuse is more common than formal data citation in biomedical fields. Journal of the Association for Information Science and Technology, 69(11): 1346–1354, November 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Quan-Haase:2018:BRF

[QH18] Anabel Quan-Haase. Book review: Facets of Facebook: Use and Users. Knautz Kathrin and Baran Katsiaryna S. (eds). Berlin: De Gruyter, 2016. 328 pp. $126 (Hardcover). Open access (Ebook). (ISBN 978-3-11-041816-3). Journal of the Association for Information Science and Technology, 69(5): 757–759, May 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Qiao:2020:JNN

[QH20] Chen Qiao and Xiao Hu. A joint neural network model for combining heterogeneous user data sources: an example of at-risk student prediction. Journal of the Association for Information Science and Technology, 71(10):1192–1204, October 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Quan-Haase:2020:OPC

[QHH20] Anabel Quan-Haase and Dennis Ho. Online privacy concerns and privacy protection strategies among older adults in East York, Canada. Journal of the Association for Information Science and Technology, 71(9):1089–1102, September 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Qin:2022:DSO

[QLM^{+}22] Chunxiu Qin, Yaxi Liu, Xubu Ma, Jiangping Chen, and Huigang Liang. Designing for serendipity in online knowledge communities: an investigation of tag presentation formats and openness to experience. Journal of the Association for Information Science and Technology, 73(10):1401–1417, October 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Qiu:2014:SLM

[QSH14] Xin Ying Qiu, Padmini Srinivasan, and Yong Hu. Supervised learning models to predict firm performance with annual reports: an empirical study. Journal of the Association for Information Science and Technology, 65(2):400–413, February 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Qiu:2021:KOO

[QZWC21] Jiangnan Qiu, Min Zuo, Jingxian Wang, and Chengjie Cai. Knowledge order in an online knowledge community: Group heterogeneity and two paths mediated by group interaction. Journal of the Association for Information Science and Technology, 72(10):1075–1091, August 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Rodrigues:2014:SJB

[RA14] Rosângela Schwarz Rodrigues and Ernest Abadal. Scientific journals in Brazil and Spain: Alternative publishing models. Journal of the Association for Information Science and Technology, 65(10):2145–2151, October 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Reynolds:2022:DDC

[RAMP22] Rebecca Reynolds, Julie Aromi, Catherine McGowan, and Britt Paris. Digital divide, critical-, and crisis-informatics perspectives on K–12 emergency remote teaching during the pandemic. Journal of the Association for Information Science and Technology, 73(12):1665–1680, December 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Riley:2022:WPE

[RAW22] Franklin Riley, David K. Allen, and Thomas Daniel Wilson. When politicians and the experts collide: Organization and the creation of information spheres. Journal of the Association for Information Science and Technology, 73(8):1127– 1139, August 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Rubel:2014:FAC

[RB14] Alan Rubel and Ryan Biava. A framework for analyzing and comparing privacy states. Journal of the Association for Information Science and Technology, 65(12):2422–2431, December 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Ruthven:2018:IOW

[RBJ18a] Ian Ruthven, Steven Buchanan, and Cara Jardine. Isolated, overwhelmed, and worried: Young first-time mothers asking for information and support online. Journal of the Association for Information Science and Technology, 69(9): 1073–1083, September 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Ruthven:2018:REH

[RBJ18b] Ian Ruthven, Steven Buchanan, and Cara Jardine. Relationships, environment, health and development: the information needs expressed online by young first-time mothers. Journal of the Association for Information Science and Technology, 69(8):985–995, August 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Roy:2022:IAW

[RBJ22] Dwaipayan Roy, Sumit Bhatia, and Prateek Jain. Information asymmetry in Wikipedia across different languages: a statistical analysis. Journal of the Association for Information Science and Technology, 73(3):347–361, March 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Reynolds:2016:RDD

[RC16] Rebecca Reynolds and Ming Ming Chiu. Reducing digital divide effects through student engagement in coordinated game design, online resource use, and social computing activities in school. Journal of the Association for Information Science and Technology, 67(8):1822–1835, August 2016. CO-DEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Radford:2017:SVN

[RCM^{+}17a] Marie L. Radford, Lynn Silipigni Connaway, Stephanie Mikitish, Mark Alpert, Chirag Shah, and Nicole A. Cooke. Shared values, new vision: Collaboration and communities of practice in virtual reference and SQA. Journal of the Association for Information Science and Technology, 68(2): 438–449, February 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Ren:2017:DTP

[RCM^{+}17b] Pengjie Ren, Zhumin Chen, Jun Ma, Zhiwei Zhang, Luo Si, and Shuaiqiang Wang. Detecting temporal patterns of user queries. Journal of the Association for Information Science and Technology, 68(1):113–128, January 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Radford:2022:DSE

[RCM22] Marie L. Radford, Laura Costello, and Kaitlin E. Montague. “Death of social encounters”: Investigating COVID-19’s initial impact on virtual reference services in academic libraries. Journal of the Association for Information Science and Technology, 73(11):1594–1607, November 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Rousseau:2016:BCD

[RD16] Ronald Rousseau and Jielan Ding. Brief communication: Does international collaboration yield a higher citation potential for US scientists publishing in highly visible interdisciplinary journals? Journal of the Association for Information Science and Technology, 67(4):1009–1013, April 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Reichenbacher:2016:AGR

[RDPF16] Tumasch Reichenbacher, Stefano De Sabbata, Ross S. Purves, and Sara I. Fabrikant. Assessing geographic relevance for mobile search: a computational model and its validation via crowdsourcing. Journal of the Association for Information Science and Technology, 67(11):2620–2634, November 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Ren:2021:HDS

[RDPN21] Jie Ren, Hang Dong, Balaji Padmanabhan, and Jeffrey V. Nickerson. How does social media sentiment impact mass media sentiment? a study of news in the financial markets. Journal of the Association for Information Science and Technology, 72(9):1183–1197, September 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Rieks:2020:RBC

[RDS20] Angela Ramnarine Rieks, Jason Dedrick, and Jeffrey Stanton. Risks, benefits, and control of information: Two studies of smart electric meter privacy. Journal of the Association for Information Science and Technology, 71(9):1060–1073, September 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Rodriguez-Esteban:2022:RDS

[REVR22] Raul Rodriguez-Esteban, Dina Vishnyakova, and Fabio Rinaldi. Revisiting the decay of scientific email addresses. Journal of the Association for Information Science and Technology, 73(1):136–139, January 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Reynolds:2021:BRD

[Rey21] Rebecca Reynolds. Book review: Dijk, Jan. (2020). The digital divide. Cambridge, UK: Polity, 208 pp. £17.99 (paperback) (ISBN 978-1-5095-3445-6). Journal of the Association for Information Science and Technology, 72(1):136–138, January 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Rockmore:2018:CEN

[RFF^{+}18] Daniel N. Rockmore, Chen Fang, Nicholas J. Foti, Tom Ginsburg, and David C. Krakauer. The cultural evolution of national constitutions. Journal of the Association for Information Science and Technology, 69(3):483–494, March 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Robinson-Garcia:2016:ADC

[RGJCTS16] Nicolas Robinson-García, Evaristo Jiménez-Contreras, and Daniel Torres-Salinas. Analyzing data citation practices using the data citation index. Journal of the Association for Information Science and Technology, 67(12):2964–2975, December 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Rozas:2021:LCL

[RH21] David Rozas and Steven Huckle. Loosen control without losing control: Formalization and decentralization within commons-based peer production. Journal of the Association for Information Science and Technology, 72(2):204–223, February 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Rheinberger:2017:BRR

[Rhe17] Hans-Jörg Rheinberger. Book review: Representation in Scientific Practice Revisited. Catelijne Coopmans, Janet Vertesi, Michael Lynch, and Steve Woolgar (Eds.). Cambridge, MA: MIT Press, 2014. 384 pp. $38.00 (paperback). (ISBN: 978-0-262-52538-1). Journal of the Association for Information Science and Technology, 68(4):1068–1069, April 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Rotolo:2023:DFS

[RHG23] Daniele Rotolo, Michael Hopkins, and Nicola Grassano. Do funding sources complement or substitute? Examining the impact of cancer research publications. Journal of the Association for Information Science and Technology, 74(1):50– 66, January 2023. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Richards:2018:RMC

[Ric18] Lorraine L. Richards. Records management in the cloud: From system design to resource ownership. Journal of the Association for Information Science and Technology, 69(2): 281–289, February 2018. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Riviera:2015:TSN

[Riv15] Emanuela Riviera. Testing the strength of the normative approach in citation theory through relational bibliometrics: the case of Italian sociology. Journal of the Association for Information Science and Technology, 66(6):1178–1188, June 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Radev:2016:BNA

[RJGM16] Dragomir R. Radev, Mark Thomas Joseph, Bryan Gibson, and Pradeep Muthukrishnan. A bibliometric and network analysis of the field of computational linguistics. Journal of the Association for Information Science and Technology, 67 (3):683–706, March 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Ruotsalo:2020:IFQ

[RJK20] Tuukka Ruotsalo, Giulio Jacucci, and Samuel Kaski. Interactive faceted query suggestion for exploratory search: Wholesession effectiveness and interaction engagement. Journal of the Association for Information Science and Technology, 71(7):742–756, July 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Rowley:2017:GIO

[RJS17a] Jennifer Rowley, Frances Johnson, and Laura Sbaffi. Gender as an influencer of online health information-seeking and evaluation behavior. Journal of the Association for Information Science and Technology, 68(1):36–47, January 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Rowley:2017:ABA

[RJS^{+}17b] Jennifer Rowley, Frances Johnson, Laura Sbaffi, Will Frass, and Elaine Devine. Academics’ behaviors and attitudes towards open access publishing in scholarly journals. Journal of the Association for Information Science and Technology, 68(5):1201–1211, May 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Rahmi:2019:AND

[RJS19] Rahmi Rahmi, Hideo Joho, and Tetsuya Shirai. An analysis of natural disaster-related information-seeking behavior using temporal stages. Journal of the Association for Information Science and Technology, 70(7):715–728, July 2019. CO-DEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Reichmann:2021:BAR

[RKHMRH21] Stefan Reichmann, Thomas Klebel, Ilire Hasani-Mavriqi, and Tony Ross-Hellauer. Between administration and research: Understanding data management practices in an institutional context. Journal of the Association for Information Science and Technology, 72(11):1415–1431, November 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Ravindran:2014:AES

[RKL14] Thara Ravindran, Alton Chua Yeow Kuan, and Dion Goh Hoe Lian. Antecedents and effects of social network fatigue. Journal of the Association for Information Science and Technology, 65(11):2306–2320, November 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Rotolo:2015:BCM

[RL15a] Daniele Rotolo and Loet Leydesdorff. Brief communications: Matching Medline/PubMed data with Web of Science: a routine in R language. Journal of the Association for Information Science and Technology, 66(10):2155–2159, October 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Rubin:2015:TDR

[RL15b] Victoria L. Rubin and Tatiana Lukoianova. Truth and deception at the rhetorical structure level. Journal of the Association for Information Science and Technology, 66(5):905–917, May 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Rowland:2024:SIK

[RLAB^{+}24] Jussara Rowland, Sergi López-Asensio, Ataberk Bagci, Ana Delicado, and Ana Prades. Shaping information and knowledge on climate change technologies: a cross-country qualitative analysis of carbon capture and storage results on Google search. Journal of the Association for Information Science and Technology, 75(5):625–639, May 2024. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Rae:2023:NLM

[RMDF23] Alastair R. Rae, James G. Mork, and Dina Demner-Fushman. The National Library of Medicine indexer assignment dataset: a new large-scale dataset for reviewer assignment research. Journal of the Association for Information Science and Technology, 74(2):205–218, February 2023. CO-DEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Rodriguez-Navarro:2016:RAB

[RN16] Alonso Rodríguez-Navarro. Research assessment based on infrequent achievements: a comparison of the United States and Europe in terms of highly cited papers and Nobel Prizes. Journal of the Association for Information Science and Technology, 67(3):731–740, March 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Robinson:2015:OMP

[Rob15] Lyn Robinson. Opinion: Multisensory, pervasive, immersive: Towards a new generation of documents. Journal of the Association for Information Science and Technology, 66 (8):1734–1737, August 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Rogers:2015:BRA

[Rog15] Yvonne Rogers. Book review: Adaptive Interaction: A Utility Maximization Approach to Understanding Human Interaction with Technology — by Stephen J. Payne and Andrew Howes. Journal of the Association for Information Science and Technology, 66(7):1520–1521, July 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Rohman:2021:EPA

[Roh21] Abdul Rohman. The emergence, peak, and abeyance of an online information ground: The lifecycle of a Facebook group for verifying information during violence. Journal of the Association for Information Science and Technology, 72 (3):302–314, March 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Rousseau:2015:LEE

[Rou15] Ronald Rousseau. Letter to the Editor: Egghe’s g-index is not a proper concentration measure. Journal of the Association for Information Science and Technology, 66(7):1518– 1519, July 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Rousseau:2016:CDP

[Rou16] Ronald Rousseau. Citation data as a proxy for quality or scientific influence are at best PAC (probably approximately correct). Journal of the Association for Information Science and Technology, 67(12):3092–3094, December 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Rousseau:2017:BRS

[Rou17] Ronald Rousseau. Book review: Scholarly metrics under the microscope: From citation analysis to academic auditing. Edited by Blaise Cronin and Cassidy R. Sugimoto. Medford, NJ: Information Today, Inc., 2015. 976 pp. $149.50 (hardcover). (ISBN 978-1-57387-499-1). Journal of the Association for Information Science and Technology, 68(2):537–538, February 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Ronda-Pupo:2016:PLR

[RPK16] Guillermo Armando Ronda-Pupo and J. Sylvan Katz. The power-law relationship between citation-based performance and collaboration in articles in management journals: a scale-independent approach. Journal of the Association for Information Science and Technology, 67(10):2565–2572, October 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Ronda-Pupo:2017:SRB

[RPK17] Guillermo Armando Ronda-Pupo and J. Sylvan Katz. The scaling relationship between citation-based performance and coauthorship patterns in natural sciences. Journal of the Association for Information Science and Technology, 68(5): 1257–1265, May 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Rousseau:2015:LEM

[RR15] Sandra Rousseau and Ronald Rousseau. Letters to the Editor: Metric-wiseness. Journal of the Association for Information Science and Technology, 66(11):2389, November 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Robert:2017:IDE

[RR17] Lionel P. Robert, Jr. and Daniel M. Romero. The influence of diversity and experience on the effects of crowd size. Journal of the Association for Information Science and Technology, 68(2):321–332, February 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Rusho:2020:HIE

[RR20] Yonit Rusho and Daphne R. Raban. Hands on: Information experiences as sources of value. Journal of the Association for Information Science and Technology, 71(6):671–684, June 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Rusho:2021:JCP

[RR21] Yonit Rusho and Daphne R. Raban. Join the club? Peer effects on information value perception. Journal of the Association for Information Science and Technology, 72(2):156– 172, February 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Rotolo:2017:SIE

[RRHL17] Daniele Rotolo, Ismael Rafols, Michael M. Hopkins, and Loet Leydesdorff. Strategic intelligence on emerging technologies: Scientometric overlay mapping. Journal of the Association for Information Science and Technology, 68(1): 214–233, January 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Rajabi:2014:BCA

[RSAS14] Enayat Rajabi, Salvador Sanchez-Alonso, and Miguel-Angel Sicilia. Brief communication: Analyzing broken links on the web of data: an experiment with DBpedia. Journal of the Association for Information Science and Technology, 65(8):

1721–1727, August 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Ren:2017:AWB

[RTS^{+}17] Yongli Ren, Martin Tomko, Flora Dilys Salim, Kevin Ong, and Mark Sanderson. Analyzing Web behavior in indoor retail spaces. Journal of the Association for Information Science and Technology, 68(1):62–76, January 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Rubenstein:2015:TAT

[Rub15] Ellen L. Rubenstein. “They are always there for me”: the convergence of social support and information in an online breast cancer community. Journal of the Association for Information Science and Technology, 66(7):1418–1430, July 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Ruthven:2021:RER

[Rut21] Ian Ruthven. Resonance and the experience of relevance. Journal of the Association for Information Science and Technology, 72(5):554–569, May 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Ruthven:2022:IBT

[Rut22] Ian Ruthven. An information behavior theory of transitions. Journal of the Association for Information Science and Technology, 73(4):579–593, April 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Ruthven:2024:ISb

[Rut24a] Ian Ruthven. Information sculpting. Journal of the Association for Information Science and Technology, 75(4):483–495, April 2024. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Ruthven:2024:ISa

[Rut24b] Ian Ruthven. Information shaping. Journal of the Association for Information Science and Technology, 75(4):469–482, April 2024. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Rodriguez-Vidal:2021:APS

[RVdAGP21] Javier Rodríguez-Vidal, Jorge Carrillo de Albornoz, Julio Gonzalo, and Laura Plaza. Authority and priority signals in automatic summary generation for online reputation management. Journal of the Association for Information Science and Technology, 72(5):583–594, May 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Rodriguez-Vidal:2019:ADI

[RVGPS19] Javier Rodríguez-Vidal, Julio Gonzalo, Laura Plaza, and Henry Anaya Sánchez. Automatic detection of influencers in social networks: Authority versus domain signals. Journal of the Association for Information Science and Technology, 70(7):675–684, July 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Robert:2018:YSY

[RY18] Lionel P. Robert, Jr. and Sangseok You. Are you satisfied yet? Shared leadership, individual trust, autonomy, and satisfaction in virtual teams. Journal of the Association for Information Science and Technology, 69(4):503–513, April 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Ren:2018:OCR

[RYEP18] Jie Ren, William Yeoh, Mong Shan Ee, and Ales Popovic. Online consumer reviews and sales: Examining the chicken– egg relationships. Journal of the Association for Information Science and Technology, 69(3):449–460, March 2018. CO-DEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Skouvig:2015:UIH

[SA15] Laura Skouvig and Jack Andersen. Understanding information history from a genre-theoretical perspective. Journal of the Association for Information Science and Technology, 66 (10):2061–2070, October 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Safder:2023:NMT

[SAA^{+}23] Iqra Safder, Momin Ali, Naif Radi Aljohani, Raheel Nawaz, and Saeed-Ul Hassan. Neural machine translation for in-text citation classification. Journal of the Association for Information Science and Technology, 74(10):1229–1240, October 2023. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Sarigil:2018:CPH

[SAB^{+}18] Erdem Sarigil, Ismail Sengor Altingovde, Roi Blanco, B. Barla Cambazoglu, Rifat Ozcan, and Özgür Ulusoy. Characterizing, predicting, and handling web search queries that match very few or no results. Journal of the Association for Information Science and Technology, 69(2):256–270, February 2018. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Shah:2021:IAD

[SAHZ21] Chirag Shah, Theresa Anderson, Loni Hagen, and Yin Zhang. An iSchool approach to data science: Humancentered, socially responsible, and context-driven. Journal of the Association for Information Science and Technology, 72(6):793–796, June 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Salvatore:2020:BRM

[Sal20] Cecilia Lizama Salvatore. Book review: The monumental challenge of preservation: The past in a volatile world. Cloonan, Michele Valerie Cambridge, MA: The MIT press, 2018. 280 pp. $30.00. (ISBN: 978-0-262-03773-0). Journal of the Association for Information Science and Technology, 71(7): 867–868, July 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Sanfilippo:2022:BRA

[San22] Madelyn Rose Sanfilippo. Book review: Algorithms and autonomy: The ethics of automated decision systems. Rubel, Alan Castro, Clinton Pham, Adam Cambridge University Press, 2021. 206 pp. £29.99 (paperback). (978-1-108-79539-5). Journal of the Association for Information Science and Technology, 73(10):1506–1509, October 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Saracevic:2016:BRH

[Sar16] Tefko Saracevic. Book review: Handbook of Information Science — By Wolfgang G. Stock and Mechtild Stock. Journal of the Association for Information Science and Technology, 67(3):749–750, March 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Savolainen:2015:PIS

[Sav15a] Reijo Savolainen. Providing informational support in an online discussion group and a Q&A site: the case of travel planning. Journal of the Association for Information Science and Technology, 66(3):450–461, March 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Savoy:2015:TCA

[Sav15b] Jacques Savoy. Text clustering: an application with the State of the Union addresses. Journal of the Association for Information Science and Technology, 66(8):1645–1654, August 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Savoy:2016:EPA

[Sav16a] Jacques Savoy. Estimating the probability of an authorship attribution. Journal of the Association for Information Science and Technology, 67(6):1462–1472, June 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Savoy:2016:TRS

[Sav16b] Jacques Savoy. Text representation strategies: an example with the State of the Union addresses. Journal of the Association for Information Science and Technology, 67(8):1858– 1870, August 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Savolainen:2017:CCG

[Sav17] Reijo Savolainen. Contributions to conceptual growth: the elaboration of Ellis’s model for information-seeking behavior. Journal of the Association for Information Science and Technology, 68(3):594–608, March 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Savolainen:2018:ISP

[Sav18a] Reijo Savolainen. Information-seeking processes as temporal developments: Comparison of stage-based and cyclic approaches. Journal of the Association for Information Science and Technology, 69(6):787–797, June 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Savoy:2018:BRW

[Sav18b] Jacques Savoy. Book review: Working with Text. Tools, Techniques and Approaches for Text Mining. Emme L. Tonkin & Gregory J. L. Tourte. Chandos Publisher, Cambridge (MA). 2016. 330 pp. (ISBN 978-1-84334-749-1). Journal of the Association for Information Science and Technology, 69(1):181–184, January 2018. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Savoy:2019:APE

[Sav19] Jacques Savoy. Authorship of Pauline epistles revisited. Journal of the Association for Information Science and Technology, 70(10):1089–1097, October 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Sawyer:2019:BRI

[Saw19] Steve Sawyer. Book review: The Internet of Things. Bunz, Mercedes and Meikle, Graham. New York: Polity/Digital Media and Society Press, 2018. 165 pp. $64.95 hardcover, $15.95 Kindle, $21.53 (paperback). (ISBN-13: 978-1-5095-1746-6). Journal of the Association for Information Science and Technology, 70(6):638–639, June 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Sjogaarde:2021:ALH

[SAW21] Peter Sjögårde, Per Ahlgren, and Ludo Waltman. Algorithmic labeling in hierarchical classifications of publications: Evaluation of bibliographic fields and term weighting approaches. Journal of the Association for Information Science and Technology, 72(7):853–869, July 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Sonnenwald:2014:GE

[SB14] Diane H. Sonnenwald and Harry Bruce. Guest editorial. Journal of the Association for Information Science and Technology, 65(1):1–2, January 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Sanchez:2016:CSP

[SB16] David Sánchez and Montserrat Batet. C-sanitized: a privacy model for document redaction and sanitization. Journal of the Association for Information Science and Technology, 67 (1):148–163, January 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Serenko:2018:CEE

[SB18] Alexander Serenko and Nick Bontis. A critical evaluation of expert survey-based journal rankings: the role of personal research interests. Journal of the Association for Information Science and Technology, 69(5):749–752, May 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Sturges:2015:RDS

[SBA^{+}15] Paul Sturges, Marianne Bamkin, Jane H. S. Anders, Bill Hubbard, Azhar Hussain, and Melanie Heeley. Research data sharing: Developing a stakeholder-driven model for journal policies. Journal of the Association for Information Science and Technology, 66(12):2445–2455, December 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Soedring:2021:MPS

[SBH21] Thomas Soedring, Pia Borlund, and Markus Helfert. The migration and preservation of six Norwegian municipality record-keeping systems: Lessons learned. Journal of the Association for Information Science and Technology, 72(6): 699–709, June 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Shema:2014:DBC

[SBIT14] Hadas Shema, Judit Bar-Ilan, and Mike Thelwall. Do blog citations correlate with a higher number of future citations? Research blogs as a potential source for alternative metrics. Journal of the Association for Information Science and Technology, 65(5):1018–1027, May 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Shema:2015:HRB

[SBIT15] Hadas Shema, Judit Bar-Ilan, and Mike Thelwall. How is research blogged? A content analysis approach. Journal of the Association for Information Science and Technology, 66 (6):1136–1149, June 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Shahbazi:2023:CSS

[SBS23] Maryam Shahbazi, Deborah Bunker, and Tania C. Sorrell. Communicating shared situational awareness in times of chaos: Social media and the COVID-19 pandemic. Journal of the Association for Information Science and Technology, 74(10):1185–1202, October 2023. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Symonds:2014:AQE

[SBZ^{+}14] Michael Symonds, Peter Bruza, Guido Zuccon, Bevan Koopman, Laurianne Sitbon, and Ian Turner. Automatic query expansion: a structural linguistic perspective. Journal of the Association for Information Science and Technology, 65(8): 1577–1596, August 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Schneider:2017:IPB

[SC17] Jesper W. Schneider and Rodrigo Costas. Identifying potential “breakthrough” publications using refined citation analyses: Three related explorative approaches. Journal of the Association for Information Science and Technology, 68(3): 709–723, March 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Stephens:2021:KCT

[SC21] Bryan Stephens and Jonathon N. Cummings. Knowledge creation through collaboration: The role of shared institutional affiliations and physical proximity. Journal of the Association for Information Science and Technology, 72(11): 1337–1353, November 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Schreiber:2015:BCV

[Sch15] Michael Schreiber. Brief communications: A variant of the h-index to measure recent performance. Journal of the Association for Information Science and Technology, 66(11): 2373–2380, November 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Schmidt:2018:AVR

[Sch18] Marion Schmidt. An analysis of the validity of retraction annotation in PubMed and the Web of Science. Journal of the Association for Information Science and Technology, 69 (2):318–328, February 2018. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Singh:2020:FLM

[SCIF20] Vivek K. Singh, Mary Chayko, Raj Inamdar, and Diana Floegel. Female librarians and male computer programmers? Gender bias in occupational images on digital media platforms. Journal of the Association for Information Science and Technology, 71(11):1281–1294, November 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Sun:2022:DDB

[SDT22] Mengyi Sun, Jainabou Barry Danfa, and Misha Teplitskiy. Does double-blind peer review reduce bias? Evidence from a top computer science conference. Journal of the Association for Information Science and Technology, 73(6):811–819, June 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Shahzad:2024:HTK

[SDUDL24] Khuram Shahzad, Marco De Sisto, Shajara Ul-Durar, and Wei Liu. How technological knowledge management capability compliments knowledge-intensive human resource management practices to enhance team outcomes: a moderated mediation analysis. Journal of the Association for Information Science and Technology, 75(4):377–394, April 2024. CO-DEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Suri:2016:SMH

[SE16] Venkata Ratnadeep Suri and Hamid R. Ekbia. Spatial mediations in historical understanding: GIS and epistemic practices of history. Journal of the Association for Information Science and Technology, 67(9):2296–2306, September 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Seadle:2020:FRG

[Sea20] Michael Seadle. Film revew: The Great Hack (documentary film). Produced and directed by Karim Amer and Jehane Noujaim. Netflix, 2019. 1 hour 54 minutes. Journal of the Association for Information Science and Technology, 71(12): 1507–1511, December 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Seewang:2017:BRM

[See17] Laila Seewang. Book review: Mechanization Takes Command: A Contribution to Anonymous History. Sigfried Giedion. Minneapolis, MN: University of Minnesota Press, 2013. 785 pp. $29.95 (softcover) (ISBN: 978-0-8166-9043-5). Journal of the Association for Information Science and Technology, 68(3):805–808, March 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Suissa:2022:TAU

[SEZG22] Omri Suissa, Avshalom Elmalech, and Maayan Zhitomirsky-Geffet. Text analysis using deep neural networks in digital humanities and information science. Journal of the Association for Information Science and Technology, 73(2):268–287, February 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Soo:2015:BCS

[SF15] Jason Soo and Ophir Frieder. Brief communications: On searching misspelled collections. Journal of the Association for Information Science and Technology, 66(6):1294–1298, June 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Subelj:2017:PBW

[ŠF17] Lovro Šubelj and Dalibor Fiala. Publication boost in Web of Science journals and its effect on citation distributions. Journal of the Association for Information Science and Technology, 68(4):1018–1023, April 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Sun:2020:CTL

[SF20] Lia H. Sun and Pnina Fichman. The collective trolling lifecycle. Journal of the Association for Information Science and Technology, 71(7):770–783, July 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Slota:2023:LWA

[SFG^{+}23] Stephen C. Slota, Kenneth R. Fleischmann, Sherri Greenberg, Nitin Verma, Brenna Cummings, Lan Li, and Chris Shenefiel. Locating the work of artificial intelligence ethics. Journal of the Association for Information Science and Technology, 74(3):311–322, March 2023. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Slota:2023:FDH

[SFL^{+}23] Stephen C. Slota, Kenneth R. Fleischmann, Min Kyung Lee, Sherri R. Greenberg, Ishan Nigam, Tara Zimmerman, Sarah Rodriguez, and James Snow. A feeling for the data: How government and nonprofit stakeholders negotiate value conflicts in data science approaches to ending homelessness. Journal of the Association for Information Science and Technology, 74(6):727–741, June 2023. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Santana:2017:IAN

[SGLF17] Alan Filipe Santana, Marcos André Gonçalves, Alberto H. F. Laender, and Anderson A. Ferreira. Incremental author name disambiguation by exploiting domain-specific heuristics. Journal of the Association for Information Science and Technology, 68(4):931–945, April 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Singh:2021:DFN

[SGS21] Vivek K. Singh, Isha Ghosh, and Darshan Sonagara. Detecting fake news stories via multimodal analysis. Journal of the Association for Information Science and Technology, 72(1): 3–17, January 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Silva:2014:TSA

[SGV14] Rodrigo M. Silva, Marcos A. Gonçalves, and Adriano Veloso. A Two-stage active learning method for learning to rank. Journal of the Association for Information Science and Technology, 65(1):109–128, January 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Suorsa:2014:KCC

[SH14] Anna Suorsa and Maija-Leena Huotari. Knowledge creation and the concept of a human being: a phenomenological approach. Journal of the Association for Information Science and Technology, 65(5):1042–1057, May 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Suh:2019:MEK

[SH19] Minhyang (Mia) Suh and Gary Hsieh. The “had mores”: Exploring Korean immigrants’ information behavior and ICT usage when settling in the United States. Journal of the Association for Information Science and Technology, 70(1): 38–48, January 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Shah:2014:AIS

[Sha14] Chirag Shah. Advances in information science: Collaborative information seeking. Journal of the Association for Information Science and Technology, 65(2):215–236, February 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Shankar:2021:BRP

[Sha21a] Kalpana Shankar. Book review: The pandemic information gap: The brutal economics of COVID-19, Gans, Joshua, Cambridge, MA: The MIT Press, 2020. 249. pp. EUR 51.16 (paperback). (ISBN 978-3-319-98487-2). Journal of the Association for Information Science and Technology, 72(11): 1454–1457, November 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Shaw:2021:BRE

[Sha21b] Ryan Shaw. Book reviews: Engines of order: a mechanology of algorithmic techniques. Rieder, Bernhard. Amsterdam: Amsterdam University Press, 2020, 352 pp., EUR 115,00 (hardcover). (ISBN 978-94-6298-619-0). EUR 0,00 (PDF). (ISBN 978-90-485-3741-9). Journal of the Association for Information Science and Technology, 72(10):1092–1095, August 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Sheble:2016:RSM

[She16] Laura Sheble. Research synthesis methods and library and information science: Shared problems, limited diffusion. Journal of the Association for Information Science and Technology, 67(8):1990–2008, August 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Sheble:2017:ARM

[She17] Laura Sheble. AIS review: Macro-level diffusion of a methodological knowledge innovation: Research synthesis methods, 1972–2011. Journal of the Association for Information Science and Technology, 68(12):2693–2708, December 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Shen:2019:CSI

[She19] Yi Shen. Create synergies and inspire collaborations around the development of intelligent infrastructure for humancentered communities. Journal of the Association for Information Science and Technology, 70(6):596–606, June 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Shelomi:2021:COF

[She21] Matan Shelomi. Comment on “Open is not forever: a study of vanished open access journals”. Journal of the Association for Information Science and Technology, 72(9):1113–1114, September 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Shah:2016:RSF

[SHGI16] Chirag Shah, Chathra Hendahewa, and Roberto González-Ibáñez. Rain or shine? Forecasting search process performance in exploratory search tasks. Journal of the Association for Information Science and Technology, 67(7):1607– 1623, July 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Siren-Heikel:2023:CLA

[SHKL23] Stefanie Sirén-Heikel, Martin Kjellman, and Carl-Gustav Lindén. At the crossroads of logics: Automating newswork with artificial intelligence-(Re)defining journalistic logics from the perspective of technologists. Journal of the Association for Information Science and Technology, 74(3): 354–366, March 2023. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Shmargad:2018:SDT

[Shm18] Yotam Shmargad. Structural diversity and tie strength in the purchase of a social networking app. Journal of the Association for Information Science and Technology, 69(5): 660–674, May 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Sidone:2017:SPC

[SHMC17] Otávio José Guerci Sidone, Eduardo Amaral Haddad, and Jesús Pascual Mena-Chalco. Scholarly publication and collaboration in Brazil: the role of geography. Journal of the Association for Information Science and Technology, 68(1): 243–258, January 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Song:2016:DHT

[SHQ^{+}16] Jun Song, Yu Huang, Xiang Qi, Yuheng Li, Feng Li, Kun Fu, and Tinglei Huang. Discovering hierarchical topic evolution in time-stamped documents. Journal of the Association for Information Science and Technology, 67(4):915–927, April 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Stvilia:2015:RPT

[SHW^{+}15] Besiki Stvilia, Charles C. Hinnant, Shuheng Wu, Adam Worrall, Dong Joon Lee, Kathleen Burnett, Gary Burnett, Michelle M. Kazmer, and Paul F. Marty. Research project tasks, data, and perceptions of data quality in a condensed matter physics community. Journal of the Association for Information Science and Technology, 66(2):246–263, February 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Silvello:2017:LCF

[Sil17] Gianmaria Silvello. Learning to cite framework: How to automatically construct citations for hierarchical data. Journal of the Association for Information Science and Technology, 68(6):1505–1524, June 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Silvello:2018:ART

[Sil18] Gianmaria Silvello. AIS review: Theory and practice of data citation. Journal of the Association for Information Science and Technology, 69(1):6–20, January 2018. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Siler:2020:DSP

[Sil20] Kyle Siler. Demarcating spectrums of predatory publishing: Economic and institutional sources of academic legitimacy. Journal of the Association for Information Science and Technology, 71(11):1386–1401, November 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Sin:2016:SMP

[Sin16] Sei-Ching Joanna Sin. Social media and problematic everyday life information-seeking outcomes: Differences across use frequency, gender, and problem-solving styles. Journal of the Association for Information Science and Technology, 67(8): 1793–1807, August 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Saastamoinen:2017:STF

[SJ17] Miamaria Saastamoinen and Kalervo Järvelin. Search task features in work tasks of varying types and complexity. Journal of the Association for Information Science and Technology, 68(5):1111–1123, May 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Song:2015:ITK

[SJK15] Min Song, Yoo Kyung Jeong, and Ha Jin Kim. Identifying the topology of the K-pop video community on YouTube: a combined co-comment analysis approach. Journal of the Association for Information Science and Technology, 66(12): 2580–2595, December 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Shu:2019:DWS

[SJL19] Fei Shu, Charles-Antoine Julien, and Vincent Larivière. Does the Web of Science accurately represent Chinese scientific performance? Journal of the Association for Information Science and Technology, 70(10):1138–1152, October 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Salminen:2014:LCM

[SJN14] Airi Salminen, Eliisa Jauhiainen, and Reija Nurmeksela. A life cycle model of XML documents. Journal of the Association for Information Science and Technology, 65(12):2564– 2580, December 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Shankar:2021:DCC

[SJT^{+}21] Kalpana Shankar, Wei Jeng, Andrea Thomer, Nicholas Weber, and Ayoung Yoon. Data curation as collective action during COVID-19. Journal of the Association for Information Science and Technology, 72(3):280–284, March 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Seberger:2014:BRC

[SK14a] John S. Seberger and Cory Knobel. Book review: On Computing: The Fourth Great Scientific Domain — by P. S. Rosenbloom. Journal of the Association for Information Science and Technology, 65(3):644–646, March 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Stern:2014:IPS

[SK14b] Tziporah Stern and Nanda Kumar. Improving privacy settings control in online social networks with a wheel interface. Journal of the Association for Information Science and Technology, 65(3):524–538, March 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Skulimowski:2023:FOA

[SK23] Andrzej M. J. Skulimowski and Thomas Köhler. A futureoriented approach to the selection of artificial intelligence technologies for knowledge platforms. Journal of the Association for Information Science and Technology, 74(8):905– 922, August 2023. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Song:2018:IDD

[SKA18] Min Song, Keunyoung Kang, and Ju Young An. Investigating drug-disease interactions in drug-symptom-disease triples via citation relations. Journal of the Association for Information Science and Technology, 69(11):1355–1368, November 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Syazillah:2018:ATV

[SKC18] Nordin Hani Syazillah, K. Kiran, and G. Chowdhury. Adaptation, translation, and validation of information literacy assessment instrument. Journal of the Association for Information Science and Technology, 69(8):996–1006, August 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Sinha:2022:SEA

[SKKM22] Ankur Sinha, Satishwar Kedas, Rishu Kumar, and Pekka Malo. SEntFiN 1.0: Entity-aware sentiment analysis for financial news. Journal of the Association for Information Science and Technology, 73(9):1314–1335, September 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Song:2017:EAT

[SKL17] Min Song, SuYeon Kim, and Keeheon Lee. Ensemble analysis of topical journal ranking in bioinformatics. Journal of the Association for Information Science and Technology, 68 (6):1564–1583, June 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Skold:2018:UEN

[Skö18] Olle Sköld. Understanding the “expanded notion” of videogames as archival objects: a review of priorities, methods, and conceptions. Journal of the Association for Information Science and Technology, 69(1):134–145, January 2018. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Skouvig:2023:BRL

[Sko23] Laura Skouvig. Book review: Literary information in China: a history. By Jack W. Chen, Anatoly Detwyler, Xiao Lui, Christopher M. B. Nugent, Bruce Rusk (Eds.), New York: Columbia University Press. 2021. pp. 672. $90.00 (hbk). ISBN: 978-0-231-19552-2. Journal of the Association for Information Science and Technology, 74(8):937–940, August 2023. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Schultz:2024:DSA

[SKO24] Carsten D. Schultz, Christian Koch, and Rainer Olbrich. Dark sides of artificial intelligence: The dangers of automated decision-making in search engine advertising. Journal of the Association for Information Science and Technology, 75(5):550–566, May 2024. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Song:2014:PIB

[SKZ^{+}14] Min Song, SuYeon Kim, Guo Zhang, Ying Ding, and Tamy Chambers. Productivity and influence in bioinformatics: a bibliometric analysis using PubMed central. Journal of the Association for Information Science and Technology, 65(2): 352–371, February 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Schultheiss:2021:HUK

[SL21a] Sebastian Schultheiß and Dirk Lewandowski. How users’ knowledge of advertisements influences their viewing and selection behavior in search engines. Journal of the Association for Information Science and Technology, 72(3):285–301, March 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Steichen:2021:HDM

[SL21b] Ben Steichen and Ryan Lowe. How do multilingual users search? An investigation of query and result list language choices. Journal of the Association for Information Science and Technology, 72(6):759–776, June 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Siler:2024:VDA

[SL24] Kyle Siler and Vincent Larivière. Varieties of diffusion in academic publishing: How status and legitimacy influence growth trajectories of new innovations. Journal of the Association for Information Science and Technology, 75(2):132– 151, February 2024. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Su:2018:LAT

[SLCS18] Sen Su, Xiao Li, Xiang Cheng, and Chenna Sun. Locationaware targeted influence maximization in social networks. Journal of the Association for Information Science and Technology, 69(2):229–241, February 2018. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Soni:2021:FLD

[SLE21] Sandeep Soni, Kristina Lerman, and Jacob Eisenstein. Follow the leader: Documents on the leading edge of semantic change get more citations. Journal of the Association for Information Science and Technology, 72(4):478–492, April 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Stvilia:2021:SYO

[SLeH21] Besiki Stvilia, Dong Joon Lee, and Na eun Han. “Striking out on your own” — a study of research information management problems on university campuses. Journal of the Association for Information Science and Technology, 72 (8):963–978, August 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Sundin:2022:WRW

[SLH22] Olof Sundin, Dirk Lewandowski, and Jutta Haider. Whose relevance? Web search engines as multisided relevance machines. Journal of the Association for Information Science and Technology, 73(5):637–642, May 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Sela:2015:PNC

[SLI^{+}15] Michal Sela, Talia Lavie, Ohad Inbar, Ilit Oppenheim, and Joachim Meyer. Personalizing news content: an experimental study. Journal of the Association for Information Science and Technology, 66(1):1–12, January 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Su:2017:BAR

[SLLL17] Zhuo Su, Donghui Li, Hanhui Li, and Xiaonan Luo. Boosting attribute recognition with latent topics by matrix factorization. Journal of the Association for Information Science and Technology, 68(7):1737–1750, July 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Son:2020:UUD

[SLLW20] Jaebong Son, Jintae Lee, Kai R. Larsen, and Jiyoung Woo. Understanding the uncertainty of disaster tweets and its effect on retweeting: The perspectives of uncertainty reduction theory and information entropy. Journal of the Association for Information Science and Technology, 71(10):1145–1161, October 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Smolinsky:2015:TTP

[SLM15] Lawrence Smolinsky, Aaron Lercher, and Andrew McDaniel. Testing theories of preferential attachment in random networks of citations. Journal of the Association for Information Science and Technology, 66(10):2132–2145, October 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Shen:2019:KWO

[SLS^{+}19] Xiao-Liang Shen, Yang-Jun Li, Yongqiang Sun, Jun Chen, and Feng Wang. Knowledge withholding in online knowledge spaces: Social deviance behavior and secondary control perspective. Journal of the Association for Information Science and Technology, 70(4):385–401, April 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Siler:2020:DNM

[SLS20] Kyle Siler, Vincent Larivière, and Cassidy R. Sugimoto. The diverse niches of megajournals: Specialism within generalism. Journal of the Association for Information Science and Technology, 71(7):800–816, July 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Sugimoto:2018:NCC

[SM18] Cassidy R. Sugimoto and Javed Mostafa. A note of concern and context: on careful use of terminologies. Journal of the Association for Information Science and Technology, 69(3): 347–348, March 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Siqueira:2022:WMA

[SM22] Joyce Siqueira and Dalton Lopes Martins. Workflow models for aggregating cultural heritage data on the web: a systematic literature review. Journal of the Association for Information Science and Technology, 73(2):204–224, February 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Salager-Meyer:2009:BSS

[SMAB09] Françoise Salager-Meyer, María Ángeles Alcaraz Ariza, and Maryelis Pabón Berbesí. “Backstage solidarity” in Spanishand English-written medical research papers: Publication context and the acknowledgment paratext. Journal of the American Society for Information Science and Technology: JASIST, 60(2):307–317, February 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic). See [Ano14].

Smith:2015:DIS

[Smi15] Catherine L. Smith. Domain-independent search expertise: a description of procedural knowledge gained during guided instruction. Journal of the Association for Information Science and Technology, 66(7):1388–1405, July 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Smith:2017:DIS

[Smi17a] Catherine L. Smith. Domain-independent search expertise: Gaining knowledge in query formulation through guided practice. Journal of the Association for Information Science and Technology, 68(6):1462–1479, June 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Smith:2017:IRS

[Smi17b] Catherine L. Smith. Investigating the role of semantic priming in query expression: a framework and two experiments. Journal of the Association for Information Science and Technology, 68(1):168–181, January 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Smolinsky:2017:DPL

[Smo17] Lawrence Smolinsky. Discrete power law with exponential cutoff and Lotka’s law. Journal of the Association for Information Science and Technology, 68(7):1792–1795, July 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Salmeron:2016:MVI

[SMRR16] Ladislao Salmerón, Mônica Macedo-Rouet, and Jean-François Rouet. Multiple viewpoints increase students’ attention to source features in social question and answer forum messages. Journal of the Association for Information Science and Technology, 67(10):2404–2419, October 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Shen:2014:EST

[SMW14] Cuihua Shen, Peter Monge, and Dmitri Williams. The evolution of social ties online: a longitudinal study in a massively multiplayer online game. Journal of the Association for Information Science and Technology, 65(10):2127–2137, October 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Silva:2015:PBR

[SMYL15] Thushari Silva, Jian Ma, Chen Yang, and Haidan Liang. A profile-boosted research analytics framework to recommend journals for manuscripts. Journal of the Association for Information Science and Technology, 66(1):180–200, January 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Snyder:2014:VRI

[Sny14] Jaime Snyder. Visual representation of information as communicative practice. Journal of the Association for Information Science and Technology, 65(11):2233–2247, November 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Soderstrom:2023:GRR

[Söd23] Kristofer Rolf Söderström. Global reach, regional strength: Spatial patterns of a big science facility. Journal of the Association for Information Science and Technology, 74(9): 1140–1156, September 2023. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Soergel:2018:BRS

[Soe18] Dagobert Soergel. Book review: The Selected Works of Marcia J. Bates. Volume I Information and the Information Professions. Bates, Marcia J. Berkeley, CA: Ketchikan Press, 2016. 372 pp. $18.00 (Paperback). (ISBN 978-0-9817584-1-1) The Selected Works of Marcia J. Bates. Volume II Information Searching Theory and Practice. Bates, Marcia J.Berkeley, CA: Ketchikan Press, 2016. 392 pp. $18.00 (Paperback). (ISBN 978-0-9817584-2-8) The Selected Works of Marcia J. Bates. Volume III Information Users and Information System Design. Bates, Marcia J. Berkeley, CA: Ketchikan Press, 2016. 384 pp. $18.00 (Paperback). (ISBN 978-0-9817584-3-5). Journal of the Association for Information Science and Technology, 69(2):343–344, February 2018. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Sobak:2017:DSI

[SP17] Veslemøy Søbak and Nils Pharo. Decentralized subject indexing of television programs: the effects of using a semicontrolled indexing language. Journal of the Association for Information Science and Technology, 68(3):739–749, March 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Shen:2022:CES

[SP22] Zixing Shen and Michael J. Pritchard. Cognitive engagement on social media: a study of the effects of visual cueing in educational videos. Journal of the Association for Information Science and Technology, 73(9):1253–1267, September 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Stanton:2014:BRB

[SR14] Jeffrey M. Stanton and Ben Rothke. Book reviews: Burdens of Proof: Cryptographic Culture and Evidence Law in the Age of Electronic Documents — by Jean-François Blanchette. Journal of the Association for Information Science and Technology, 65(7):1509–1510, July 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Simkin:2015:WDA

[SR15] Mikhail V. Simkin and Vwani P. Roychowdhury. Why does attention to web articles fall with Time? Journal of the Association for Information Science and Technology, 66(9): 1847–1856, September 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Splendiani:2016:AGS

[SR16] Bruno Splendiani and Mireia Ribera. Accessibility of graphics in STEM research articles: Analysis and proposals for improvement. Journal of the Association for Information Science and Technology, 67(6):1507–1520, June 2016. CO-DEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Salles:2016:QAT

[SRG^{+}16] Thiago Salles, Leonardo Rocha, Marcos André Gonçalves, Jussara M. Almeida, Fernando Mourão, Wagner Meira Jr., and Felipe Viegas. A quantitative analysis of the temporal effects on automatic text classification. Journal of the Association for Information Science and Technology, 67(7): 1639–1667, July 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Shuai:2017:MIE

[SRM^{+}17] Xin Shuai, Jason Rollins, Isabelle Moulinier, Tonya Custis, Mathilda Edmunds, and Frank Schilder. A multidimensional investigation of the effects of publication retraction on scholarly impact. Journal of the Association for Information Science and Technology, 68(9):2225–2236, September 2017. CO-DEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Saunders:2016:OHI

[SRPS16] Carol Saunders, Anne F. Rutkowski, Jon Pluyter, and Ronald Spanjers. Opinion: Health information technologies: From hazardous to the dark side. Journal of the Association for Information Science and Technology, 67(7):1767–1772, July 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Shi:2017:JIF

[SRYL17] Dongbo Shi, Ronald Rousseau, Liu Yang, and Jiang Li. A journal’s impact factor is influenced by changes in publication delays of citing journals. Journal of the Association for Information Science and Technology, 68(3):780–789, March 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Seddon:2014:IOC

[SS14a] Jessica Seddon and Ramesh Srinivasan. Information and ontologies: Challenges in scaling knowledge for development. Journal of the Association for Information Science and Technology, 65(6):1124–1133, June 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Sinn:2014:HUD

[SS14b] Donghee Sinn and Nicholas Soares. Historians’ use of digital archival collections: the web, historical scholarship, and archival research. Journal of the Association for Information Science and Technology, 65(9):1794–1809, September 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Stieger:2014:TUI

[SS14c] Stefan Stieger and Viren Swami. Twitter users’ interest in asteroid 2012 DA_{14} mirrored the asteroid’s trajectory during its Earth flyby. Journal of the Association for Information Science and Technology, 65(7):1409–1415, July 2014. CO-DEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Sedhai:2017:AMT

[SS17] Surendra Sedhai and Aixin Sun. An analysis of 14 million tweets on hashtag-oriented spamming. Journal of the Association for Information Science and Technology, 68(7):1638– 1651, July 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Stark:2020:DWS

[SSA20] Luke Stark, Amanda Stanhaus, and Denise L. Anthony. “I Don’t Want Someone to Watch Me While I’m Working”: Gendered views of facial recognition technology in workplace surveillance. Journal of the Association for Information Science and Technology, 71(9):1074–1088, September 2020. CO-DEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Shvartzshnaider:2022:GCU

[SSA22] Yan Shvartzshnaider, Madelyn Rose Sanfilippo, and Noah Apthorpe. GKC-CI: a unifying framework for contextual norms and information governance. Journal of the Association for Information Science and Technology, 73(9):1297– 1313, September 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Saltz:2017:PDS

[SSC17] Jeffrey Saltz, Ivan Shamshurin, and Colin Connors. Predicting data science sociotechnical execution challenges by categorizing data science projects. Journal of the Association for Information Science and Technology, 68(12):2720–2728, December 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Soares-Silva:2020:ELU

[SSdMCM20] Danilo Soares-Silva, Gustavo Hermínio Salati Marcondes de Moraes, Alexandre Cappellozza, and Cristiano Morini. Explaining library user loyalty through perceived service quality: What is wrong? Journal of the Association for Information Science and Technology, 71(8):954–967, August 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Sanfilippo:2020:DPP

[SSR^{+}20] Madelyn R. Sanfilippo, Yan Shvartzshnaider, Irwin Reyes, Helen Nissenbaum, and Serge Egelman. Disaster privacy/ privacy disaster. Journal of the Association for Information Science and Technology, 71(9):1002–1014, September 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Schlagwein:2020:CSC

[SSS20] Daniel Schlagwein, Detlef Schoder, and Kai Spindeldreher. Consolidated, systemic conceptualization, and definition of the “sharing economy”. Journal of the Association for Information Science and Technology, 71(7):817–838, July 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Seo:2016:MMG

[ST16a] Hyunjin Seo and Stuart Thorson. A mixture model of global Internet capacity distributions. Journal of the Association for Information Science and Technology, 67(8):2032–2044, August 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Sud:2016:AIC

[ST16b] Pardeep Sud and Mike Thelwall. Not all international collaboration is beneficial: the Mendeley readership and citation impact of biochemical research collaboration. Journal of the Association for Information Science and Technology, 67(8): 1849–1857, August 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Suominen:2016:MST

[ST16c] Arho Suominen and Hannes Toivanen. Map of science with topic modeling: Comparison of unsupervised learning and human-assigned subject classification. Journal of the Association for Information Science and Technology, 67(10): 2464–2476, October 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

StJean:2017:FMD

[St.17] Beth St. Jean. Factors motivating, demotivating, or impeding information seeking and use by people with type 2 diabetes: a call to work toward preventing, identifying, and addressing incognizance. Journal of the Association for Information Science and Technology, 68(2):309–320, February 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Savolainen:2022:ATP

[ST22] Reijo Savolainen and Leslie Thomson. Assessing the theoretical potential of an expanded model for everyday information practices. Journal of the Association for Information Science and Technology, 73(4):511–527, April 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Stamatatos:2018:MTR

[Sta18] Efstathios Stamatatos. Masking topic-related information to enhance authorship attribution. Journal of the Association for Information Science and Technology, 69(3):461–473, March 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Stahlman:2022:NKC

[Sta22] Gretchen R. Stahlman. From nostalgia to knowledge: Considering the personal dimensions of data lifecycles. Journal of the Association for Information Science and Technology, 73 (12):1692–1705, December 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Sormunen:2014:SGW

[STAH14] Eero Sormunen, Mikko Tanni, Tuulikki Alamettälä, and Jannica Heinström. Students’ group work strategies in sourcebased writing assignments. Journal of the Association for Information Science and Technology, 65(6):1217–1231, June 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Spina:2017:EAS

[STCS17] Damiano Spina, Johanne R. Trippas, Lawrence Cavedon, and Mark Sanderson. Extracting audio summaries to support effective spoken document search. Journal of the Association for Information Science and Technology, 68(9):2101– 2115, September 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Shklovski:2015:MTS

[STD15] Irina Shklovski, Emily Troshynski, and Paul Dourish. Mobile technologies and the spatiotemporal configurations of institutional practice. Journal of the Association for Information Science and Technology, 66(10):2098–2115, October 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Song:2019:CBL

[STF19] Lianlian Song, Geoffrey Tso, and Yelin Fu. Click behavior and link prioritization: Multiple demand theory application for web improvement. Journal of the Association for Information Science and Technology, 70(8):805–816, August 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Stock:2020:BRM

[Sto20] Wolfgang G. Stock. Book review: Minty, Ella. Social media and the Islamic state: Can public relations succeed where conventional diplomacy failed?. London, NY: Routledge, 2020, XV, 98 pp. £29.99 (paperback) (ISBN 978-0-367-18816-0). Journal of the Association for Information Science and Technology, 71(11):1413–1416, November 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Sahib:2014:IBV

[STS14] Nuzhah Gooda Sahib, Anastasios Tombros, and Tony Stockman. Investigating the behavior of visually impaired users for multi-session search tasks. Journal of the Association for Information Science and Technology, 65(1):69–83, January 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Sahib:2015:ESI

[STS15] Nuzhah Gooda Sahib, Anastasios Tombros, and Tony Stockman. Evaluating a search interface for visually impaired searchers. Journal of the Association for Information Science and Technology, 66(11):2235–2248, November 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Sebescen:2017:SHE

[SV17] Nina Sebescen and Jessica Vitak. Securing the human: Employee security vulnerability risk in organizational settings. Journal of the Association for Information Science and Technology, 68(9):2237–2247, September 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Smutny:2020:SIR

[SV20] Zdenek Smutny and Vasja Vehovar. Social informatics research: Schools of thought, methodological basis, and thematic conceptualization. Journal of the Association for Information Science and Technology, 71(5):529–539, May 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Smets:2022:SCU

[SVB22] Annelien Smets, Jorre Vannieuwenhuyze, and Pieter Ballon. Serendipity in the city: User evaluations of urban recommender systems. Journal of the Association for Information Science and Technology, 73(1):19–30, January 2022. CO-DEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Seeber:2022:CDD

[SVC22] Marco Seeber, Jef Vlegels, and Mattia Cattaneo. Conditions that do or do not disadvantage interdisciplinary research proposals in project evaluation. Journal of the Association for Information Science and Technology, 73(8):1106–1126, August 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Svenonius:2015:BRD

[Sve15] Elaine Svenonius. Book reviews: The Discipline of Organizing — edited by Robert J. Glushko. Journal of the Association for Information Science and Technology, 66(9): 1963–1964, September 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Sappelli:2017:ECA

[SVK17] Maya Sappelli, Suzan Verberne, and Wessel Kraaij. Evaluation of context-aware recommendation systems for information re-finding. Journal of the Association for Information Science and Technology, 68(4):895–910, April 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Shi:2018:RCV

[SW18] Wei Shi and Matthew S. Weber. Rethinking the complexity of virtual work and knowledge sharing. Journal of the Association for Information Science and Technology, 69(11): 1318–1329, November 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Swain:2022:BRK

[Swa22] Deborah Swain. Book review: Knowledge architectures: Structures and semantics. Denise Bedford, Abingdon-on-Thames: Routledge, 2021. 544 pp. £120.00 (hardcover). (ISBN 978-0-367-21944-4). Journal of the Association for Information Science and Technology, 73(6):892–896, June 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Stover:2016:BCC

[SWKK16] Justin Anthony Stover, Yaron Winter, Moshe Koppel, and Mike Kestemont. Brief communications: Computational authorship verification method attributes a new work to a major 2nd century African author. Journal of the Association for Information Science and Technology, 67(1):239–242, January 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Stvilia:2018:RUD

[SWL18] Besiki Stvilia, Shuheng Wu, and Dong Joon Lee. Researchers’ uses of and disincentives for sharing their research identity information in research information management systems. Journal of the Association for Information Science and Technology, 69(8):1035–1045, August 2018. CO-DEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Sugimoto:2017:ARS

[SWLH17] Cassidy R. Sugimoto, Sam Work, Vincent Larivière, and Stefanie Haustein. AIS review: Scholarly use of social media and altmetrics: a review of the literature. Journal of the Association for Information Science and Technology, 68(9): 2037–2062, September 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Sun:2019:BES

[SWSZ19] Yongqiang Sun, Nan Wang, Xiao-Liang Shen, and Xi Zhang. Bias effects, synergistic effects, and information contingency effects: Developing and testing an extended information adoption model in social Q&A. Journal of the Association for Information Science and Technology, 70(12):1368–1382, December 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Sa:2020:EUP

[SY20] Ning Sa and Xiaojun (Jenny) Yuan. Examining users’ partial query modification patterns in voice search. Journal of the Association for Information Science and Technology, 71(3): 251–263, March 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Sa:2022:IEV

[SY22] Ning Sa and Xiaojun (Jenny) Yuan. Improving the effectiveness of voice search systems through partial query modification. Journal of the Association for Information Science and Technology, 73(8):1092–1105, August 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Sanfilippo:2017:THT

[SYF17] Madelyn Sanfilippo, Shengnan Yang, and Pnina Fichman. Trolling here, there, and everywhere: Perceptions of trolling behaviors in context. Journal of the Association for Information Science and Technology, 68(10):2313–2327, October 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Soldaini:2017:LRL

[SYG17] Luca Soldaini, Andrew Yates, and Nazli Goharian. Learning to reformulate long queries for clinical decision support. Journal of the Association for Information Science and Technology, 68(11):2602–2619, November 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Schwartz:2017:NCO

[SYS17] David G. Schwartz, Inbal Yahav, and Gahl Silverman. News censorship in online social networks: a study of circumvention in the commentsphere. Journal of the Association for Information Science and Technology, 68(3):569–582, March 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Sbaffi:2020:MOH

[SZ20] Laura Sbaffi and Chen Zhao. Modeling the online health information seeking process: Information channel selection among university students. Journal of the Association for Information Science and Technology, 71(2):196–207, February 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Sun:2021:HAM

[SZJ^{+}21] Jianshan Sun, Mingyue Zhu, Yuanchun Jiang, Yezheng Liu, and Le Wu. Hierarchical attention model for personalized tag recommendation. Journal of the Association for Information Science and Technology, 72(2):173–189, February 2021. CO-DEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Sun:2022:IIE

[SZWS22] Xu Sun, Xiaosong Zhou, Qingfeng Wang, and Sarah Sharples. Investigating the impact of emotions on perceiving serendipitous information encountering. Journal of the Association for Information Science and Technology, 73(1): 3–18, January 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Sarnikar:2014:QPP

[SZZ14] Surendra Sarnikar, Zhu Zhang, and J. Leon Zhao. Queryperformance prediction for effective query routing in domainspecific repositories. Journal of the Association for Information Science and Technology, 65(8):1597–1614, August 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Shen:2016:HBC

[SZZ16] Xiao-Liang Shen, Kem Z. K. Zhang, and Sesia J. Zhao. Herd behavior in consumers’ adoption of online reviews. Journal of the Association for Information Science and Technology, 67(11):2754–2765, November 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Tabak:2014:JBC

[Tab14] Edin Tabak. Jumping between context and users: a difficulty in tracing information practices. Journal of the Association for Information Science and Technology, 65(11):2223–2232, November 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Tang:2015:VUF

[TBP^{+}15] David Li Tang, France Bouthillier, Pierre Pluye, Roland Grad, and Carol Repchinsky. The value of user feedback: Healthcare professionals’ comments to the health information provider. Journal of the Association for Information Science and Technology, 66(2):377–391, February 2015. CO-DEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Toraman:2017:DSC

[TC17] Cagri Toraman and Fazli Can. Discovering story chains: a framework based on zigzagged search and news actors. Journal of the Association for Information Science and Technology, 68(12):2795–2808, December 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Tian:2024:UCC

[TCF^{+}24] Wencan Tian, Ruonan Cai, Zhichao Fang, Yu Geng, Xianwen Wang, and Zhigang Hu. Understanding cocorresponding authorship: a bibliometric analysis and detailed overview. Journal of the Association for Information Science and Technology, 75(1):3–23, January 2024. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Tamine:2015:ABH

[TCP15] Lynda Tamine, Cécile Chouquet, and Thomas Palmer. Analysis of biomedical and health queries: Lessons learned from TREC and CLEF evaluation benchmarks. Journal of the Association for Information Science and Technology, 66(12): 2626–2642, December 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

TeixeiradaSilva:2015:LED

[TD15] Jaime A. Teixeira da Silva and Judit Dobránszki. Letters to the Editor: Do open access data files represent an academic risk? Journal of the Association for Information Science and Technology, 66(11):2390–2391, November 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Teich:2016:LCD

[TDOF^{+}16] Elke Teich, Stefania Degaetano-Ortlieb, Peter Fankhauser, Hannah Kermes, and Ekaterina Lapshinova-Koltunski. The linguistic construal of disciplinarity: a data-mining approach using register features. Journal of the Association for Information Science and Technology, 67(7):1668–1678, July 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

TeixeiradaSilva:2016:LEI

[Tei16] Jaime A. Teixeira da Silva. Letter to the Editor: The importance of the anonymous voice in postpublication peer review. Journal of the Association for Information Science and Technology, 67(1):243, January 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

TeixeiradaSilva:2017:FDJ

[Tei17] Jaime A. Teixeira da Silva. The false Donald J. Trump article and the ethics of misleading journalism. Journal of the Association for Information Science and Technology, 68 (4):1061–1063, April 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Tenopir:2016:BRB

[Ten16] Carol Tenopir. Book review: Big Data, Little Data, No Data: Scholarship in the Networked World — By Christine L. Borgman. Journal of the Association for Information Science and Technology, 67(3):751–753, March 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Thelwall:2021:MFG

[TF21] Mike Thelwall and David Foster. Male or female genderpolarized YouTube videos are less viewed. Journal of the Association for Information Science and Technology, 72(12): 1545–1557, December 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Trace:2015:VCC

[TFR15] Ciaran B. Trace and Luis Francisco-Revilla. The value and complexity of collection arrangement for evidentiary work. Journal of the Association for Information Science and Technology, 66(9):1857–1882, September 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Tan:2015:SSI

[TG15] Esther Meng-Yoke Tan and Dion Hoe-Lian Goh. A study of social interaction during mobile information seeking. Journal of the Association for Information Science and Technology, 66(10):2031–2044, October 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Tommasel:2017:LAU

[TG17] Antonela Tommasel and Daniela Godoy. Learning and adapting user criteria for recommending followees in social networks. Journal of the Association for Information Science and Technology, 68(8):1863–1874, August 2017. CO-DEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Tian:2021:IPL

[TGC^{+}21] Yubing Tian, Ricardo Gomez, Marika Cifor, James Wilson, and Henry Morgan. The information practices of law enforcement: Passive and active collaboration and its implication for sanctuary laws in Washington state. Journal of the Association for Information Science and Technology, 72(11): 1354–1366, November 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Thelwall:2016:CTP

[TGV^{+}16] Mike Thelwall, Olga Goriunova, Farida Vis, Simon Faulkner, Anne Burns, Jim Aulich, Amalia Mas-Bleda, Emma Stuart, and Francesco D’Orazio. Chatting through pictures? A classification of images tweeted in one week in the UK and USA. Journal of the Association for Information Science and Technology, 67(11):2575–2586, November 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Tu:2016:CCT

[TH16] Yi-Ning Tu and Shu-Lan Hsu. Constructing conceptual trajectory maps to trace the development of research fields. Journal of the Association for Information Science and Technology, 67(8):2016–2031, August 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Tsai:2015:FAR

[THC15] Chih-Fong Tsai, Ya-Han Hu, and Zong-Yao Chen. Factors affecting Rocchio-based pseudorelevance feedback in image retrieval. Journal of the Association for Information Science and Technology, 66(1):40–57, January 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Thelwall:2017:BGA

[The17] Mike Thelwall. Book genre and author gender: Romance > paranormal-romance to autobiography > memoir. Journal of the Association for Information Science and Technology, 68(5):1212–1223, May 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Thelwall:2020:FCI

[The20] Mike Thelwall. Female citation impact superiority 1996– 2018 in six out of seven English-speaking nations. Journal of the Association for Information Science and Technology, 71(8):979–990, August 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Tang:2017:FAA

[THL17] Li Tang, Guangyuan Hu, and Weishu Liu. Funding acknowledgment analysis: Queries and caveats. Journal of the Association for Information Science and Technology, 68(3):790– 794, March 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Thomer:2022:IDR

[Tho22] Andrea K. Thomer. Integrative data reuse at scientifically significant sites: Case studies at Yellowstone National Park and the La Brea Tar Pits. Journal of the Association for Information Science and Technology, 73(8):1155–1170, August 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Torres:2014:QRI

[THWS14] Sergio Duarte Torres, Djoerd Hiemstra, Ingmar Weber, and Pavel Serdyukov. Query recommendation in the information domain of children. Journal of the Association for Information Science and Technology, 65(7):1368–1384, July 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Tang:2020:MDR

[TJ20] Muh-Chyun Tang and Pei-Syuan Jhang. Music discovery and revisiting behaviors of individuals with different preference characteristics: An experience sampling approach. Journal of the Association for Information Science and Technology, 71(5):540–552, May 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Tuomaala:2014:ELI

[TJV14] Otto Tuomaala, Kalervo Järvelin, and Pertti Vakkari. Evolution of library and information science, 1965–2005: Content analysis of journal articles. Journal of the Association for Information Science and Technology, 65(7):1446–1462, July 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Thelwall:2014:AES

[TK14] Mike Thelwall and Kayvan Kousha. Academia.edu: Social network or Academic Network? Journal of the Association for Information Science and Technology, 65(4):721–731, April 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Thelwall:2015:RDC

[TK15] Mike Thelwall and Kayvan Kousha. ResearchGate: Disseminating, communicating, and measuring Scholarship? Journal of the Association for Information Science and Technology, 66(5):876–889, May 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Thelwall:2017:GSN

[TK17a] Mike Thelwall and Kayvan Kousha. Goodreads: a social network site for book readers. Journal of the Association for Information Science and Technology, 68(4):972–983, April 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Thelwall:2017:RAA

[TK17b] Mike Thelwall and Kayvan Kousha. ResearchGate articles: Age, discipline, audience size, and impact. Journal of the Association for Information Science and Technology, 68(2): 468–479, February 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Thelwall:2017:SPC

[TK17c] Mike Thelwall and Kayvan Kousha. SlideShare presentations, citations, users, and trends: a professional site with academic and educational uses. Journal of the Association for Information Science and Technology, 68(8):1989–2003, August 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Trace:2017:IMH

[TK17d] Ciaran B. Trace and Unmil P. Karadkar. Information management in the humanities: Scholarly processes, tools, and the construction of personal collections. Journal of the Association for Information Science and Technology, 68(2):491– 507, February 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Thelwall:2023:DAS

[TKA^{+}23] Mike Thelwall, Kayvan Kousha, Mahshid Abdoli, Emma Stuart, Meiko Makita, Paul Wilson, and Jonathan Levitt. Do altmetric scores reflect article quality? Evidence from the UK Research Excellence Framework 2021. Journal of the Association for Information Science and Technology, 74(5): 582–593, May 2023. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Thakker:2017:OCV

[TKB^{+}17] Dhavalkumar Thakker, Stan Karanasios, Emmanuel Blanchard, Lydia Lau, and Vania Dimitrova. Ontology for cultural variations in interpersonal communication: Building on theoretical models and crowdsourced knowledge. Journal of the Association for Information Science and Technology, 68(6):1411–1428, June 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Thelwall:2023:WFC

[TKS^{+}23] Mike Thelwall, Kayvan Kousha, Emma Stuart, Meiko Makita, Mahshid Abdoli, Paul Wilson, and Jonathan Levitt. In which fields are citations indicators of research quality? Journal of the Association for Information Science and Technology, 74(8):941–953, August 2023. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Thelwall:2018:NSP

[TL18] Mike Thelwall and Jonathan M. Levitt. National scientific performance evolution patterns: retrenchment, successful expansion, or overextension. Journal of the Association for Information Science and Technology, 69(5):720–727, May 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Tang:2022:PDO

[TL22] Muh-Chyun Tang and I-Han Liao. Preference diversity and openness to novelty: Scales construction from the perspective of movie recommendation. Journal of the Association for Information Science and Technology, 73(9):1222–1235, September 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Tenopir:2016:TAS

[TLA^{+}16] Carol Tenopir, Kenneth Levine, Suzie Allard, Lisa Christian, Rachel Volentine, Reid Boehm, Frances Nichols, David Nicholas, Hamid R. Jamali, Eti Herman, and Anthony Watkinson. Trustworthiness and authority of scholarly information in a digital age: Results of an international questionnaire. Journal of the Association for Information Science and Technology, 67(10):2344–2361, October 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Teplitskiy:2017:AIO

[TLD17] Misha Teplitskiy, Grace Lu, and Eamon Duede. Amplifying the impact of open access: Wikipedia and the diffusion of science. Journal of the Association for Information Science and Technology, 68(9):2116–2127, September 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Thomas:2023:MSE

[TLS^{+}23] Manoj A. Thomas, Yan Li, Vera Sistenich, Ken Ngoy Diango, and Diulu Kabongo. A multi-stakeholder engagement framework for knowledge management in ICT4D. Journal of the Association for Information Science and Technology, 74 (12):1384–1400, December 2023. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Thelwall:2015:SAD

[TM15] Mike Thelwall and Nabeil Maflahi. Are scholarly articles disproportionately read in their own country? An analysis of Mendeley readers. Journal of the Association for Information Science and Technology, 66(6):1124–1135, June 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Thelwall:2016:GRA

[TM16] Mike Thelwall and Nabeil Maflahi. Guideline references and academic citations as evidence of the clinical value of health research. Journal of the Association for Information Science and Technology, 67(4):960–966, April 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Thelwall:2020:ACR

[TM20] Mike Thelwall and Nabeil Maflahi. Academic collaboration rates and citation associations vary substantially between countries and fields. Journal of the Association for Information Science and Technology, 71(8):968–978, August 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Thompson:2020:CFR

[TMBA20] Nik Thompson, Tanya McGill, Anna Bunn, and Rukshan Alexander. Cultural factors and the role of privacy concerns in acceptance of government surveillance. Journal of the Association for Information Science and Technology, 71(9): 1129–1142, September 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Tang:2021:FDP

[TMDZ21a] Rong Tang, Bharat Mehra, Jia Tina Du, and Yuxiang (Chris) Zhao. Framing a discussion on paradigm shift(s) in the field of information. Journal of the Association for Information Science and Technology, 72(2):253–258, February 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Tang:2021:PSF

[TMDZ21b] Rong Tang, Bharat Mehra, Jia Tina Du, and Yuxiang (Chris) Zhao. Paradigm shift in the field of information special issue editorial. Journal of the Association for Information Science and Technology, 72(10):1217–1222, October 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Tan:2015:IDD

[TND15] Luke Kien-Weng Tan, Jin-Cheon Na, and Ying Ding. Influence diffusion detection using the influence style (INFUSE) model. Journal of the Association for Information Science and Technology, 66(8):1717–1733, August 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Twidale:2021:EED

[TNL21] Michael B. Twidale, David M. Nichols, and Christopher P. Lueg. Everyone everywhere: a distributed and embedded paradigm for usability. Journal of the Association for Information Science and Technology, 72(10):1272–1284, October 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

TeixeiraLopes:2017:ELT

[TPR17] Carla Teixeira Lopes, Dagmara Paiva, and Cristina Ribeiro. Effects of language and terminology of query suggestions on medical accuracy considering different user characteristics. Journal of the Association for Information Science and Technology, 68(9):2063–2075, September 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Tan:2015:HCI

[TPZ15] Barney Tan, Shan L. Pan, and Meiyun Zuo. Harnessing collective IT resources for sustainability: Insights from the green leadership strategy of China mobile. Journal of the Association for Information Science and Technology, 66(4): 818–838, April 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Tinto:2016:SHI

[TR16] Fiona Tinto and Ian Ruthven. Sharing “happy” information. Journal of the Association for Information Science and Technology, 67(10):2329–2343, October 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Thompson:2017:PWT

[TR17] Santi Thompson and Michele Reilly. “A picture is worth a thousand words”: Reverse image lookup and digital library assessment. Journal of the Association for Information Science and Technology, 68(9):2264–2266, September 2017. CO-DEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Thelwall:2016:MRC

[TS16] Mike Thelwall and Pardeep Sud. Mendeley readership counts: an investigation of temporal and disciplinary differences. Journal of the Association for Information Science and Technology, 67(12):3036–3050, December 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Thelwall:2021:DNR

[TS21] Mike Thelwall and Pardeep Sud. Do new research issues attract more citations? A comparison between 25 Scopus subject categories. Journal of the Association for Information Science and Technology, 72(3):269–279, March 2021. CO-DEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Tsikerdekis:2016:PCN

[Tsi16] Michail Tsikerdekis. Personal communication networks and their positive effects on online collaboration and outcome quality on Wikipedia. Journal of the Association for Information Science and Technology, 67(4):812–823, April 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Torres-Salinas:2016:LET

[TSRG16] Daniel Torres-Salinas and Nicolás Robinson-García. Letter to the editor: The time for bibliometric applications. Journal of the Association for Information Science and Technology, 67(4):1014–1015, April 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Tang:2015:TCE

[TSY15] Li Tang, Philip Shapira, and Jan Youtie. Is there a clubbing effect underlying Chinese research citation Increases? Journal of the Association for Information Science and Technology, 66(9):1923–1932, September 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Tao:2017:HCM

[TT17] Yihan Tao and Anastasios Tombros. How collaborators make sense of tasks together: a comparative analysis of collaborative sensemaking behavior in collaborative informationseeking tasks. Journal of the Association for Information Science and Technology, 68(3):609–622, March 2017. CO-DEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Tang:2016:DVO

[TTH16] Kuo-Hao Tang, Li-Chen Tsai, and Sheue-Ling Hwang. The development and validation of a one-bit comparison for evaluating the maturity of tag distributions in a Web 2.0 environment. Journal of the Association for Information Science and Technology, 67(6):1430–1445, June 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Thelwall:2016:DRS

[TW16a] Mike Thelwall and Paul Wilson. Does research with statistics have more impact? The citation rank advantage of structural equation modeling. Journal of the Association for Information Science and Technology, 67(5):1233–1244, May 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Thelwall:2016:MRA

[TW16b] Mike Thelwall and Paul Wilson. Mendeley readership altmetrics for medical articles: an analysis of 45 fields. Journal of the Association for Information Science and Technology, 67(8):1962–1972, August 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Thomer:2018:DPN

[TWB^{+}18] Andrea K. Thomer, Karen M. Wickett, Karen S. Baker, Bruce W. Fouke, and Carole L. Palmer. Documenting provenance in noncomputational workflows: Research process models based on geobiology fieldwork in Yellowstone National Park. Journal of the Association for Information Science and Technology, 69(10):1234–1245, October 2018. CO-DEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Tao:2023:MSB

[TZH23] Jie Tao, Lina Zhou, and Kevin Hickey. Making sense of the black-boxes: Toward interpretable text classification using deep learning models. Journal of the Association for Information Science and Technology, 74(6):685–700, June 2023. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Tay:2020:BMR

[TZK20] Wenyi Tay, Xiuzhen Zhang, and Sarvnaz Karimi. Beyond mean rating: Probabilistic aggregation of star ratings based on helpfulness. Journal of the Association for Information Science and Technology, 71(7):784–799, July 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Tavakoli:2022:ACA

[TZS^{+}22] Leila Tavakoli, Hamed Zamani, Falk Scholer, William Bruce Croft, and Mark Sanderson. Analyzing clarification in asynchronous information-seeking conversations. Journal of the Association for Information Science and Technology, 73(3): 449–471, March 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Unkel:2017:ECC

[UH17] Julian Unkel and Alexander Haas. The effects of credibility cues on the selection of search engine results. Journal of the Association for Information Science and Technology, 68(8): 1850–1862, August 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Urs:2023:EDS

[UM23] Shalini R. Urs and Mohamed Minhaj. Evolution of data science and its education in iSchools: an impressionistic study using curriculum analysis. Journal of the Association for Information Science and Technology, 74(6):606–622, June 2023. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Unsworth:2018:BRI

[Uns18] Kristene Unsworth. Book review: Information Cultures in the Digital Age: A Festschrift in Honor of Rafael Capurro. Matthew Kelly and Jared Bielby. Wiesbaden, Germany: Springer VS, 2016. 479 pp. $129.00 (Paperback). (ISBN 978-3-658-14679-5). Journal of the Association for Information Science and Technology, 69(3):495–497, March 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Vilares:2015:ULS

[VAGR15] David Vilares, Miguel A. Alonso, and Carlos Gómez-Rodríguez. On the usefulness of lexical and syntactic processing in polarity classification of Twitter messages. Journal of the Association for Information Science and Technology, 66(9):1799–1816, September 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Vasileiadou:2015:OCS

[Vas15] Eleftheria Vasileiadou. Opinion: Crowd science: It is not just a matter of time (or funding). Journal of the Association for Information Science and Technology, 66(7):1514–1517, July 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Vaughan:2016:UIS

[Vau16] Liwen Vaughan. Uncovering information from social media hyperlinks: an investigation of Twitter. Journal of the Association for Information Science and Technology, 67(5):1105– 1120, May 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Vaughan:2015:DMW

[VC15] Liwen Vaughan and Yue Chen. Data mining from web search queries: a comparison of Google Trends and Baidu Index. Journal of the Association for Information Science and Technology, 66(1):13–22, January 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Vieira:2014:DMB

[VCG14] Elizabeth S. Vieira, José A. S. Cabral, and José A. N. F. Gomes. Definition of a model based on bibliometric indicators for assessing applicants to academic positions. Journal of the Association for Information Science and Technology, 65(3):560–577, March 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Vezzani:2017:LEC

[VCG17] Antonio Vezzani, Alex Coad, and Petros Gkotsis. Letter to the Editor: Concerns about the consequences of patenting on scientometric research. Journal of the Association for Information Science and Technology, 68(9):2293–2295, September 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Vakkari:2022:DCR

[VCJ22] Pertti Vakkari, Yu-Wei Chang, and Kalervo Järvelin. Disciplinary contributions to research topics and methodology in Library and Information Science — leading to fragmentation? Journal of the Association for Information Science and Technology, 73(12):1706–1722, December 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

vandenBesselaar:2016:LEC

[vdB16] Peter van den Besselaar. Letters to the Editor: Correct assumptions? Journal of the Association for Information Science and Technology, 67(7):1779, July 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

vanderSluis:2023:FBA

[vdSvdB23] Frans van der Sluis and Egon L. van den Broek. Feedback beyond accuracy: Using eye-tracking to detect comprehensibility and interest during reading. Journal of the Association for Information Science and Technology, 74(1):3–16, January 2023. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

vanderSluis:2014:WCB

[vdSvdBG^{+}14] Frans van der Sluis, Egon L. van den Broek, Richard J. Glassey, Elisabeth M. A. G. van Dijk, and Franciska M. G. de Jong. When complexity becomes interesting. Journal of the Association for Information Science and Technology, 65 (7):1478–1500, July 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

vanderVegt:2021:DBS

[vdVZK21] Anton van der Vegt, Guido Zuccon, and Bevan Koopman. Do better search engines really equate to better clinical decisions? If not, why not? Journal of the Association for Information Science and Technology, 72(2):141–155, February 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Vechtomova:2014:MAE

[Vec14] Olga Vechtomova. A method for automatic extraction of multiword units representing business aspects from user reviews. Journal of the Association for Information Science and Technology, 65(7):1463–1477, July 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Verma:2022:TCP

[VFZ^{+}22] Nitin Verma, Kenneth R. Fleischmann, Le Zhou, Bo Xie, Min Kyung Lee, Kate Rich, Kristina Shiroma, Chenyan Jia, and Tara Zimmerman. Trust in COVID-19 public health information. Journal of the Association for Information Science and Technology, 73(12):1776–1792, December 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Vani:2018:ISS

[VG18] K Vani and Deepa Gupta. Integrating syntax-semantic-based text analysis with structural and citation information for scientific plagiarism detection. Journal of the Association for Information Science and Technology, 69(11):1330–1345, November 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Varathan:2017:COM

[VGC17] Kasturi Dewi Varathan, Anastasia Giachanou, and Fabio Crestani. Comparative opinion mining: a review. Journal of the Association for Information Science and Technology, 68(4):811–829, April 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Vannini:2020:MFG

[VGN20] Sara Vannini, Ricardo Gomez, and Bryce Clayton Newell. “Mind the five”: Guidelines for data privacy and security in humanitarian work with undocumented migrants and other vulnerable populations. Journal of the Association for Information Science and Technology, 71(8):927–938, August 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Vinkler:2017:CIP

[Vin17] Péter Vinkler. Core indicators and professional recognition of scientometricians. Journal of the Association for Information Science and Technology, 68(1):234–242, January 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Vasalou:2015:PFC

[VJH15] Asimina Vasalou, Adam Joinson, and David Houghton. Privacy as a fuzzy concept: a new conceptualization of privacy for practitioners. Journal of the Association for Information Science and Technology, 66(5):918–929, May 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Vincent-Lamarre:2016:EOA

[VLBG^{+}16] Philippe Vincent-Lamarre, Jade Boivin, Yassine Gargouri, Vincent Larivière, and Stevan Harnad. Estimating open access mandate effectiveness: The MELIBEA score. Journal of the Association for Information Science and Technology, 67 (11):2815–2828, November 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

vanLeeuwen:2018:EPU

[vLTW18] Thed N. van Leeuwen, Clifford Tatum, and Paul F. Wouters. Exploring possibilities to use bibliometric data to monitor gold open access publishing at the national level. Journal of the Association for Information Science and Technology, 69 (9):1161–1173, September 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Vaughan:2018:NAW

[VN18] Liwen Vaughan and Anton Ninkov. A new approach to web co-link analysis. Journal of the Association for Information Science and Technology, 69(6):820–831, June 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Vasquez:2023:IPA

[VO23] Fatima Espinoza Vasquez and Shannon M. Oltmann. Information precarity and the agentic practices of marginalized communities: Puerto rican activists addressing the crisis before, during, and after hurricane Maria. Journal of the Association for Information Science and Technology, 74(5): 517–530, May 2023. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Vasalou:2015:UEP

[VOBB15] Asimina Vasalou, Anne-Marie Oostveen, Chris Bowers, and Russell Beale. Understanding engagement with the privacy domain through design research. Journal of the Association for Information Science and Technology, 66(6):1263–1273, June 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Veinot:2019:MIE

[VP19] Tiffany C. Veinot and Casey S. Pierce. Materiality in information environments: Objects, spaces, and bodies in three outpatient hemodialysis facilities. Journal of the Association for Information Science and Technology, 70(12):1324–1339, December 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

vanRaan:2017:BRT

[vR17a] Anthony F. J. van Raan. Book reviews: Theories of Informetrics and Scholarly Communication. Cassidy R. Sugimoto, Editor. Berlin: de Gruyter Mouton, 2016. 426 pp. $112.00 (hardcover). (ISBN: 978-3-11-029803-1). Journal of the Association for Information Science and Technology, 68 (9):2275–2283, September 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Vilchez-Roman:2017:LEB

[VR17b] Carlos Vílchez-Román. Letter to the Editor: BRICS’ scientific excellence and the search for relevance and replicability. Journal of the Association for Information Science and Technology, 68(2):539, February 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Vaughan:2014:WSV

[VRF14] Liwen Vaughan and Esteban Romero-Frías. Web search volume as a predictor of academic fame: an exploration of Google trends. Journal of the Association for Information Science and Technology, 65(4):707–720, April 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Viejo:2014:PSN

[VS14] Alexandre Viejo and David Sánchez. Profiling social networks to provide useful and privacy-preserving web search. Journal of the Association for Information Science and Technology, 65(12):2444–2458, December 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Vrettas:2015:CVJ

[VS15] George Vrettas and Mark Sanderson. Conferences versus journals in computer science. Journal of the Association for Information Science and Technology, 66(12):2674–2684, December 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Vuong:2019:UUB

[VSJR19] Tung Vuong, Miamaria Saastamoinen, Giulio Jacucci, and Tuukka Ruotsalo. Understanding user behavior in naturalistic information search tasks. Journal of the Association for Information Science and Technology, 70(11):1248–1261, November 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Vlachidis:2016:KBA

[VT16] Andreas Vlachidis and Douglas Tudhope. A knowledgebased approach to Information Extraction for semantic interoperability in the archaeology domain. Journal of the Association for Information Science and Technology, 67(5):1138– 1152, May 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Vakkari:2021:PEQ

[VVP^{+}21] Pertti Vakkari, Michael Völske, Martin Potthast, Matthias Hagen, and Benno Stein. Predicting essay quality from search and writing behavior. Journal of the Association for Information Science and Technology, 72(7):839–852, July 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Vishwanath:2018:HPP

[VXN18] Arun Vishwanath, Weiai Xu, and Zed Ngoh. How people protect their privacy on Facebook: a cost-benefit view. Journal of the Association for Information Science and Technology, 69(5):700–709, May 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Vydiswaran:2015:OBL

[VZRP15] V. G. Vinod Vydiswaran, ChengXiang Zhai, Dan Roth, and Peter Pirolli. Overcoming bias to learn about controversial topics. Journal of the Association for Information Science and Technology, 66(8):1655–1672, August 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Wang:2018:BMI

[Wan18] Qi Wang. A bibliometric model for identifying emerging research topics. Journal of the Association for Information Science and Technology, 69(2):290–304, February 2018. CO-DEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Warner:2015:BRR

[War15] Julian Warner. Book reviews: “Raw Data” is an Oxymoron — edited by Lisa Gitelman. Journal of the Association for Information Science and Technology, 66(5):1086–1087, May 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Watson:2014:ESS

[Wat14] Curtis Watson. An exploratory study of secondary students’ judgments of the relevance and reliability of information. Journal of the Association for Information Science and Technology, 65(7):1385–1408, July 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Wijnhoven:2015:IIT

[WB15] Fons Wijnhoven and Michel Brinkhuis. Internet information triangulation: Design theory and prototype evaluation. Journal of the Association for Information Science and Technology, 66(4):684–701, April 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Winkler:2021:VSM

[WB21] Christa E. Winkler and Rebecca Fay Berenbon. Validation of a survey for measuring scientists’ attitudes toward data reuse. Journal of the Association for Information Science and Technology, 72(4):449–453, April 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Weissenberger:2018:EBB

[WBH18] Lynnsey K. Weissenberger, John M. Budd, and Ken R. Herold. Epistemology beyond the brain. Journal of the Association for Information Science and Technology, 69(5): 710–719, May 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Waltman:2014:FRP

[WC14] Ludo Waltman and Rodrigo Costas. F1000 recommendations as a potential new data source for research evaluation: A comparison with citations. Journal of the Association for Information Science and Technology, 65(3):433–445, March 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Wakeling:2017:UUG

[WCC^{+}17] Simon Wakeling, Paul Clough, Lynn Silipigni Connaway, Barbara Sen, and David Tomás. Users and uses of a global union catalog: a mixed-methods study of World-Cat.org. Journal of the Association for Information Science and Technology, 68(9):2166–2181, September 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Walsh:2022:DHI

[WCdF^{+}22] John A. Walsh, Peter J. Cobb, Wayne de Fremery, Koraljka Golub, Humphrey Keah, Jeonghyun Kim, Joseph Kiplang’at, Ying-Hsang Liu, Simon Mahony, Sam G. Oh, Chris Alen Sula, Ted Underwood, and Xiaoguang Wang. Digital humanities in the iSchool. Journal of the Association for Information Science and Technology, 73(2):188–203, February 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Worrall:2021:ISE

[WCO21] Adam Worrall, Alicia Cappello, and Rachel Osolen. The importance of socio-emotional considerations in online communities, social informatics, and information science. Journal of the Association for Information Science and Technology, 72(10):1247–1260, October 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Wakeling:2019:MUE

[WCP^{+}19] Simon Wakeling, Claire Creaser, Stephen Pinfield, Jenny Fry, Valérie Spezi, Peter Willett, and Monica Paramita. Motivations, understandings, and experiences of open-access mega-journal authors: Results of a large-scale survey. Journal of the Association for Information Science and Technology, 70(7):754–768, July 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Weinberger:2015:TFU

[WCTS15] Hadas Weinberger, Jonathan Cohen, Boaz Tadmor, and Pierre Singer. Towards a framework for untangling complexity: the interprofessional decision-making model for the complex patient. Journal of the Association for Information Science and Technology, 66(2):392–407, February 2015. CO-DEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Wang:2021:UPD

[WDL21] Xiaoguang Wang, Qingyu Duan, and Mengli Liang. Understanding the process of data reuse: an extensive review. Journal of the Association for Information Science and Technology, 72(9):1161–1182, September 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Wu:2020:UTP

[WDTC20] Dan Wu, Jing Dong, Yuan Tang, and Rob Capra. Understanding task preparation and resumption behaviors in crossdevice search. Journal of the Association for Information Science and Technology, 71(8):887–901, August 2020. CO-DEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Weber:2020:BRA

[Web20] Matthew S. Weber. Book review: The Archived Web: Doing History in the Digital Age. Brügger, Niels. Cambridge, MA: The MIT Press, 2018. 200 pp. $30.00 (hardcover). (ISBN 978-0-262-03902-4). Journal of the Association for Information Science and Technology, 71(3):374–376, March 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Weinberg:2014:BRP

[Wei14a] Bella Hass Weinberg. Book review: Powering Search: The Role of Thesauri in New Information Environments — by Ali Shiri. Journal of the Association for Information Science and Technology, 65(5):1085–1088, May 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Weiss:2014:BRD

[Wei14b] Martin B. H. Weiss. Book reviews: Digital Crossroads: Telecommunications Law and Policy in the Internet Age (2nd edition) — by Jonathan E. Nuechterlein and Philip J. Weiser. Journal of the Association for Information Science and Technology, 65(10):2158–2159, October 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Westbrook:2015:IPV

[Wes15] Lynn Westbrook. Intimate partner violence online: Expectations and agency in question and answer websites. Journal of the Association for Information Science and Technology, 66(3):599–615, March 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Workman:2014:FSI

[WFRN14] T. Elizabeth Workman, Marcelo Fiszman, Thomas C. Rindflesch, and Diane Nahl. Framing serendipitous informationseeking behavior for facilitating literature-based discovery: a proposed model. Journal of the Association for Information Science and Technology, 65(3):501–512, March 2014. CO-DEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Ward:2019:AIC

[WG19] Wesley S. Ward and Lisa M. Given. Assessing intercultural communication: Testing technology tools for information sharing in multinational research teams. Journal of the Association for Information Science and Technology, 70(4): 338–350, April 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Williams:2015:EWP

[WH15] Peter Williams and Christian Hennig. Effect of web page menu orientation on retrieving information by people with learning disabilities. Journal of the Association for Information Science and Technology, 66(4):674–683, April 2015. CO-DEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

White:2014:BDW

[Whi14] Ryen W. White. Belief dynamics in web search. Journal of the Association for Information Science and Technology, 65 (11):2165–2178, November 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

White:2017:BRA

[Whi17] Howard D. White. Book review: Atlas of Knowledge: Anyone Can Map. Katy Börner. Cambridge, MA, MIT Press, 2015. 224 pp. $39.95 (hardcover). (ISBN 978-0-262-02881-3). Journal of the Association for Information Science and Technology, 68(2):533–536, February 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Wen:2017:MST

[WHvdZvdB17] Bei Wen, Edwin Horlings, Mariëlle van der Zouwen, and Peter van den Besselaar. Mapping science through bibliometric triangulation: an experimental approach applied to water research. Journal of the Association for Information Science and Technology, 68(3):724–738, March 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Wang:2021:PUC

[WHWZ21] Xiangyu Wang, Andrew High, Xi Wang, and Kang Zhao. Predicting users’ continued engagement in online health communities from the quantity and quality of received support. Journal of the Association for Information Science and Technology, 72(6):710–722, June 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Wang:2014:RJM

[WHX^{+}14] Xiaolun Wang, Zhijuan Hong, Yunjie (Calvin) Xu, Chenghong Zhang, and Hong Ling. Relevance judgments of mobile commercial information. Journal of the Association for Information Science and Technology, 65(7):1335–1348, July 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Wang:2017:LSE

[WHYJ17] Pengwei Wang, Tianyong Hao, Jun Yan, and Lianwen Jin. Large-scale extraction of drug–disease pairs from the medical literature. Journal of the Association for Information Science and Technology, 68(11):2649–2661, November 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Wang:2023:SJP

[WHZ23] Jing Wang, Willem Halffman, and Yuehong Helen Zhang. Sorting out journals: The proliferation of journal lists in China. Journal of the Association for Information Science and Technology, 74(10):1207–1228, October 2023. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Williams:2014:MAK

[Wil14a] James G. Williams. In memoriam: Allen Kent (1921– 2014). Journal of the Association for Information Science and Technology, 65(10):2162–2163, October 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Williams:2014:MAD

[Wil14b] James G. Williams. In memoriam: Anthony Debons (1916– 2013). Journal of the Association for Information Science and Technology, 65(8):1729–1730, August 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Willson:2022:JSI

[WJB22] Rebekah Willson, Heidi Julien, and Gary Burnett. JASIS&T special issue on information behavior and information practices theory. Journal of the Association for Information Science and Technology, 73(4):491–493, April 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Wojdynski:2016:TDW

[WK16] Bartosz W. Wojdynski and Sriram Kalyanaraman. The three dimensions of website navigability: Explication and effects. Journal of the Association for Information Science and Technology, 67(2):454–464, February 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Wooldridge:2019:ASE

[WK19] Jenny Wooldridge and Mike B. King. Altmetric scores: An early indicator of research impact. Journal of the Association for Information Science and Technology, 70(3):271–282, March 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Wang:2014:ENE

[WKC14] Zhonghong Wang, Christopher S. G. Khoo, and Abdus Sattar Chaudhry. Evaluation of the navigation effectiveness of an organizational taxonomy built on a general classification scheme and domain thesauri. Journal of the Association for Information Science and Technology, 65(5):948–963, May 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Williamson:2016:DSA

[WKJW16] Kirsty Williamson, Mary Anne Kennan, Graeme Johanson, and John Weckert. Data sharing for the advancement of science: Overcoming barriers for citizen scientists. Journal of the Association for Information Science and Technology, 67(10):2392–2403, October 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Wan:2014:ALC

[WL14a] Xiaojun Wan and Fang Liu. Are all literature citations equally important? Automatic citation strength estimation and its applications. Journal of the Association for Information Science and Technology, 65(9):1929–1938, September 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Wan:2014:WIL

[WL14b] Xiaojun Wan and Fang Liu. WL-index: Leveraging citation mention number to quantify an individual’s scientific impact. Journal of the Association for Information Science and Technology, 65(12):2509–2517, December 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Wang:2020:AQI

[WL20] Ping Wang and Xiaodan Li. Assessing the quality of information on Wikipedia: a deep-learning approach. Journal of the Association for Information Science and Technology, 71 (1):16–28, January 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Wei:2014:ETC

[WLC^{+}14] Chih-Ping Wei, Yen-Hsien Lee, Yu-Sheng Chiang, Chun-Ta Chen, and Christopher C. C. Yang. Exploiting temporal characteristics of features for effectively discovering event episodes from news corpora. Journal of the Association for Information Science and Technology, 65(3):621–634, March 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Wang:2016:ILD

[WLC16] Yanshan Wang, Jae-Sung Lee, and In-Chan Choi. Indexing by Latent Dirichlet Allocation and an Ensemble Model.

Journal of the Association for Information Science and Technology, 67(7):1736–1750, July 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Wu:2023:UUG

[WLG23] Qian Wu, Chei Sian Lee, and Dion Hoe-Lian Goh. Understanding user-generated questions in social Q&A: a goalframing approach. Journal of the Association for Information Science and Technology, 74(8):990–1009, August 2023. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Wang:2023:AIC

[WLS23] Xuequn Wang, Xiaolin Lin, and Bin Shao. Artificial intelligence changes the way we work: a close look at innovating with chatbots. Journal of the Association for Information Science and Technology, 74(3):339–353, March 2023. CO-DEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Wu:2019:IRE

[WLTY19] Yingying Wu, Yiqun Liu, Yen-Hsi Richard Tsai, and Shing-Tung Yau. Investigating the role of eye movements and physiological signals in search satisfaction prediction using geometric analysis. Journal of the Association for Information Science and Technology, 70(9):981–999, September 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Wu:2017:IEC

[WLW^{+}17] Stephen Wu, Sijia Liu, Yanshan Wang, Tamara Timmons, Harsha Uppili, Steven Bedrick, William Hersh, and Hongfang Liu. Intrainstitutional EHR collections for patient-level information retrieval. Journal of the Association for Information Science and Technology, 68(11):2636–2648, November 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Wu:2020:USS

[WLZ^{+}20] Zongda Wu, Renchao Li, Zhifeng Zhou, Junfang Guo, Jionghui Jiang, and Xinning Su. A user sensitive subject protection approach for book search service. Journal of the Association for Information Science and Technology, 71(2):

183–195, February 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Wu:2014:ADF

[WLZB14] Shengli Wu, Jieyu Li, Xiaoqin Zeng, and Yaxin Bi. Adaptive data fusion methods in information retrieval. Journal of the Association for Information Science and Technology, 65(10): 2048–2061, October 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Wang:2023:QSB

[WMM^{+}23] Shiyun Wang, Yaxue Ma, Jin Mao, Yun Bai, Zhentao Liang, and Gang Li. Quantifying scientific breakthroughs by a novel disruption indicator based on knowledge entities. Journal of the Association for Information Science and Technology, 74 (2):150–167, February 2023. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Wasserman:2015:CBU

[WMS^{+}15] Max Wasserman, Satyam Mukherjee, Konner Scott, Xiao Han T. Zeng, Filippo Radicchi, and Luís A. N. Amaral. Correlations between user voting data, budget, and box office for films in the Internet movie database. Journal of the Association for Information Science and Technology, 66(4):858–868, April 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Wang:2022:TGH

[WMX^{+}22] Pianran Wang, Yue Ma, Huan Xie, Hanqing Wang, Jinyi Lu, and Jianhua Xu. “There is a gorilla holding a key on the book cover”: Young children’s known picture book search strategies. Journal of the Association for Information Science and Technology, 73(1):45–57, January 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Wu:2015:EAP

[WN15] I-Chin Wu and Yun-Fang Niu. Effects of anchoring process under preference stabilities for interactive movie recommendations. Journal of the Association for Information Science and Technology, 66(8):1673–1695, August 2015. CO-DEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Walhout:2017:ETC

[WOJBG17] Jaap Walhout, Paola Oomen, Halszka Jarodzka, and Saskia Brand-Gruwel. Effects of task complexity on online search behavior of adolescents. Journal of the Association for Information Science and Technology, 68(6):1449–1461, June 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Worrall:2019:CAB

[Wor19] Adam Worrall. “Connections above and beyond”: Information, translation, and community boundaries in Library-Thing and Goodreads. Journal of the Association for Information Science and Technology, 70(7):742–753, July 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Wang:2022:AVS

[WS22] Yiwei Wang and Chirag Shah. Authentic versus synthetic: an investigation of the influences of study settings and task configurations on search behaviors. Journal of the Association for Information Science and Technology, 73(3):362–375, March 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Wirth:2016:SOS

[WSvPK16] Werner Wirth, Katharina Sommer, Thilo von Pape, and Veronika Karnowski. Success in online searches: Differences between evaluation and finding tasks. Journal of the Association for Information Science and Technology, 67(12): 2897–2908, December 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Wang:2020:MMW

[WSWH20] Hong Wang, Yong-Qiang Song, Lu-Tong Wang, and Xiao-Hong Hu. Memory model for web ad effect based on multimodal features. Journal of the Association for Information Science and Technology, 71(1):29–42, January 2020. CO-DEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Wang:2022:RAS

[WSZC22] Xiaoguang Wang, Ningyuan Song, Huimin Zhou, and Hanghang Cheng. The representation of argumentation in scientific papers: a comparative analysis of two research areas. Journal of the Association for Information Science and Technology, 73(6):863–878, June 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Wu:2014:CBE

[WTT^{+}14] Mingfang Wu, Andrew Turpin, James A. Thom, Falk Scholer, and Ross Wilkinson. Cost and benefit estimation of experts’ mediation in an enterprise search. Journal of the Association for Information Science and Technology, 65 (1):146–163, January 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Winnink:2016:TCB

[WTvR16] Jos J. Winnink, Robert J. W. Tijssen, and Anthony F. J. van Raan. Theory-changing breakthroughs in science: the impact of research teamwork on scientific discoveries. Journal of the Association for Information Science and Technology, 67(5):1210–1223, May 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Wu:2019:PPC

[Wu19] Philip Fei Wu. The privacy paradox in the context of online social networking: A self-identity perspective. Journal of the Association for Information Science and Technology, 70 (3):207–217, March 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Wu:2023:VVV

[Wu23] Philip Fei Wu. Veni, vidi, vici? On the rise of scrape-and-report scholarship in online reviews research. Journal of the Association for Information Science and Technology, 74(2): 145–149, February 2023. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Wolf:2015:SSF

[WV15] Christine T. Wolf and Tiffany C. Veinot. Struggling for space and finding my place: an interactionist perspective on everyday use of biomedical information. Journal of the Association for Information Science and Technology, 66(2):282–296, February 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Wu:2018:ESO

[WV18] I-Chin Wu and Pertti Vakkari. Effects of subject-oriented visualization tools on search by novices and intermediates. Journal of the Association for Information Science and Technology, 69(12):1428–1445, December 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Woudstra:2016:QVA

[WvdHS16] Lilian Woudstra, Bart van den Hooff, and Alexander Schouten. The quality versus accessibility debate revisited: a contingency perspective on human information source selection. Journal of the Association for Information Science and Technology, 67(9):2060–2071, September 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Wiggers:2023:BEL

[WVvLZ23] Gineke Wiggers, Suzan Verberne, Wouter van Loon, and Gerrit-Jan Zwenne. Bibliometric-enhanced legal information retrieval: Combining usage and citations as flavors of impact relevance. Journal of the Association for Information Science and Technology, 74(8):1010–1025, August 2023. CO-DEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Wu:2020:CAI

[WVZ20] Philip Fei Wu, Jessica Vitak, and Michael T. Zimmer. A contextual approach to information privacy research. Journal of the Association for Information Science and Technology, 71(4):485–490, April 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Wang:2015:AJS

[WW15] Feifei Wang and Dietmar Wolfram. Assessment of journal similarity based on citing discipline analysis. Journal of the Association for Information Science and Technology, 66(6):1189–1198, June 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Walters:2016:DND

[WW16] William H. Walters and Esther Isabelle Wilder. Disciplinary, national, and departmental contributions to the literature of library and information science, 2007–2012. Journal of the Association for Information Science and Technology, 67 (6):1487–1506, June 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Wong:2021:UIS

[WWB21] Kevin Wong, Geoff Walton, and Gavin Bailey. Using information science to enhance educational preventing violent extremism programs. Journal of the Association for Information Science and Technology, 72(3):362–376, March 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Wei:2016:ESP

[WWL16] Jiuchang Wei, Fei Wang, and Michael K. Lindell. The evolution of stakeholders’ perceptions of disaster: a model of information flow. Journal of the Association for Information Science and Technology, 67(2):441–453, February 2016. CO-DEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Wisniewski:2015:FAT

[WXLBO15] Pamela Wisniewski, Heng Xu, Heather Lipford, and Emmanuel Bello-Ogunu. Facebook apps and tagging: the tradeoff between personal privacy and engaging with friends. Journal of the Association for Information Science and Technology, 66(9):1883–1896, September 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Wu:2023:WSW

[WXSL23] Dan Wu, Hao Xu, Yaqi Sun, and Siyu Lv. What should we teach? A human-centered data science graduate curriculum model design for iField schools. Journal of the Association for Information Science and Technology, 74(6):623–640, June 2023. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Yim:2017:CTE

[wYKY17] Wen wai Yim, Sharon W. Kwan, and Meliha Yetisgen. Classifying tumor event attributes in radiology reports. Journal of the Association for Information Science and Technology, 68(11):2662–2674, November 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Wu:2021:GIP

[WYZL21] Chaojiang Wu, Erjia Yan, Yongjun Zhu, and Kai Li. Gender imbalance in the productivity of funded projects: a study of the outputs of national institutes of health R01 grants. Journal of the Association for Information Science and Technology, 72(11):1386–1399, November 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

White:2018:LWC

[WZ18] Howard D. White and Alesia A. Zuccala. Libcitations, WorldCat, cultural impact, and fame. Journal of the Association for Information Science and Technology, 69(12): 1502–1512, December 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Wang:2022:USC

[WZFZ22] Xiangyu Wang, Min Zhang, Weiguo Fan, and Kang Zhao. Understanding the spread of COVID-19 misinformation on social media: The effects of topics and a political leader’s nudge. Journal of the Association for Information Science and Technology, 73(5):726–737, May 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Wang:2018:PMA

[WZK^{+}18] Chong Wang, Shuai Zhao, Achir Kalra, Cristian Borcea, and Yi Chen. Predictive models and analysis for webpage depthlevel dwell time. Journal of the Association for Information Science and Technology, 69(8):1007–1022, August 2018. CO-DEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Wang:2021:CLP

[WZL21] Chuang Wang, Rongxin Zhou, and Matthew K. O. Lee. Can loyalty be pursued and achieved? an extended RFD model to understand and predict user loyalty to mobile apps. Journal of the Association for Information Science and Technology, 72(7):824–838, July 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Xiao:2014:WIO

[XA14] Lu Xiao and Nicole Askin. What influences online deliberation? A Wikipedia study. Journal of the Association for Information Science and Technology, 65(5):898–910, May 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Xie:2018:IFA

[XBCH18] Iris Xie, Rakesh Babu, Melissa Davey Castillo, and Hyejung Han. Identification of factors associated with blind users’ help-seeking situations in interacting with digital libraries. Journal of the Association for Information Science and Technology, 69(4):514–527, April 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Xie:2021:OTA

[XBL^{+}21] Iris Xie, Rakesh Babu, Hyun Seung Lee, Shengang Wang, and Tae Hee Lee. Orientation tactics and associated factors in the digital library environment: Comparison between blind and sighted users. Journal of the Association for Information Science and Technology, 72(8):995–1010, August 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Xu:2022:TPD

[XBL^{+}22] Huimin Xu, Yi Bu, Meijun Liu, Chenwei Zhang, Mengyi Sun, Yi Zhang, Eric Meyer, Eduardo Salas, and Ying Ding. Team power dynamics and team impact: New perspectives on scientific collaboration using career age as a proxy for team power. Journal of the Association for Information Science and Technology, 73(10):1489–1505, October 2022. CO-DEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Xiao:2017:DRR

[XC17] Lu Xiao and Niall Conroy. Discourse relations in rationalecontaining text-segments. Journal of the Association for Information Science and Technology, 68(12):2783–2794, December 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Xiang:2021:LDA

[XCL^{+}21] Rong Xiang, Emmanuele Chersoni, Qin Lu, Chu-Ren Huang, Wenjie Li, and Yunfei Long. Lexical data augmentation for sentiment analysis. Journal of the Association for Information Science and Technology, 72(11):1432–1447, November 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Xu:2018:IFA

[XDDG18] Xiaoying Xu, Kaushik Dutta, Anindya Datta, and Chunmian Ge. Identifying functional aspects from user reviews for functionality-based mobile app recommendation. Journal of the Association for Information Science and Technology, 69(2):242–255, February 2018. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Xu:2016:ACA

[XDSC16] Jian Xu, Ying Ding, Min Song, and Tamy Chambers. Author credit-assignment schemas: a comparison and analysis. Journal of the Association for Information Science and Technology, 67(8):1973–1989, August 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Xia:2015:WPP

[XHC^{+}15] Jingfeng Xia, Jennifer L. Harmon, Kevin G. Connolly, Ryan M. Donnelly, Mary R. Anderson, and Heather A. Howard. Who publishes in “predatory” journals? Journal of the Association for Information Science and Technology, 66(7):1406–1417, July 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Xie:2020:GHC

[XHM^{+}20] Bo Xie, Daqing He, Tim Mercer, Youfa Wang, Dan Wu, Kenneth R. Fleischmann, Yan Zhang, Linda H. Yoder, Keri K. Stephens, Michael Mackert, and Min Kyung Lee. Global health crises are also information crises: a call to action. Journal of the Association for Information Science and Technology, 71(12):1419–1423, December 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Xiao:2014:ERA

[Xia14] Lu Xiao. Effects of rationale awareness in online ideation crowdsourcing tasks. Journal of the Association for Information Science and Technology, 65(8):1707–1720, August 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Xia:2023:WSI

[Xia23] Huichuan Xia. What scholars and IRBs talk when they talk about the Belmont principles in crowd work-based research. Journal of the Association for Information Science and Technology, 74(1):67–80, January 2023. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Xie:2017:UIS

[XJBK17] Iris Xie, Soohyung Joo, and Renee Bennett-Kapusniak. User involvement and system support in applying search tactics. Journal of the Association for Information Science and Technology, 68(5):1165–1185, May 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Xie:2022:CCE

[XLKC22] Juan Xie, Hongru Lu, Lele Kang, and Ying Cheng. Citing criteria and its effects on researcher’s intention to cite: a mixed-method study. Journal of the Association for Information Science and Technology, 73(8):1079–1091, August 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Xu:2016:ALR

[XLL16] Bo Xu, Hongfei Lin, and Yuan Lin. Assessment of learning to rank methods for query expansion. Journal of the Association for Information Science and Technology, 67(6):1345– 1357, June 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Xie:2018:MTP

[XOL^{+}18] Z. Xie, Z. Ouyang, J. Li, E. Dong, and D. Yi. Modelling transition phenomena of scientific coauthorship networks. Journal of the Association for Information Science and Technology, 69(2):305–317, February 2018. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Xu:2019:DFS

[XZ19] Chen Xu and Qin Zhang. The dominant factor of social tags for users’ decision behavior on e-commerce websites: Color or text. Journal of the Association for Information Science and Technology, 70(9):942–953, September 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Xu:2019:NMT

[XZW^{+}19] Shuo Xu, Dongsheng Zhai, Feifei Wang, Xin An, Hongshen Pang, and Yirong Sun. A novel method for topic linkages between scientific publications and patents. Journal of the Association for Information Science and Technology, 70(9): 1026–1042, September 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Yan:2014:FKP

[Yan14] Erjia Yan. Finding knowledge paths among scientific disciplines. Journal of the Association for Information Science and Technology, 65(11):2331–2347, November 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Yan:2015:RDI

[Yan15] Erjia Yan. Research dynamics, impact, and dissemination: a topic-level analysis. Journal of the Association for Information Science and Technology, 66(11):2357–2372, November 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Yan:2016:DKP

[Yan16] Erjia Yan. Disciplinary knowledge production and diffusion in science. Journal of the Association for Information Science and Technology, 67(9):2223–2245, September 2016. CO-DEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Ye:2018:SGV

[YB18] Fred Y. Ye and Lutz Bornmann. “Smart girls” versus “sleeping beauties” in the sciences: the identification of instant and delayed recognition by using the citation angle. Journal of the Association for Information Science and Technology, 69 (3):359–367, March 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Yoon:2020:TCI

[YC20] Ayoung Yoon and Andrea Copeland. Toward communityinclusive data ecosystems: Challenges and opportunities of open data for community-based organizations. Journal of the Association for Information Science and Technology, 71(12): 1439–1454, December 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Yi:2016:CAT

[YCK16] Kwan Yi, Namjoo Choi, and Yung Soo Kim. A content analysis of Twitter hyperlinks and their application in web resource indexing. Journal of the Association for Information Science and Technology, 67(8):1808–1821, August 2016. CO-DEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Yan:2020:ASP

[YCL20] Erjia Yan, Zheng Chen, and Kai Li. Authors’ status and the perceived quality of their work: Measuring citation sentiment change in Nobel articles. Journal of the Association for Information Science and Technology, 71(3):314–324, March 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Ye:2016:LRA

[YH16] Zheng Ye and Jimmy Xiangji Huang. A learning to rank approach for quality-aware pseudo-relevance feedback. Journal of the Association for Information Science and Technology, 67(4):942–959, April 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Yang:2021:FTG

[YHL^{+}21] Ting-Hao Yang, Yu-Lun Hsieh, Shih-Hung Liu, Yung-Chun Chang, and Wen-Lian Hsu. A flexible template generation and matching method with applications for publication reference metadata extraction. Journal of the Association for Information Science and Technology, 72(1):32–45, January 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Yanovsky:2021:OSD

[YHLG21] Stav Yanovsky, Nicholas Hoernle, Omer Lev, and Kobi Gal. One size does not fit all: a study of badge behavior in Stack Overflow. Journal of the Association for Information Science and Technology, 72(3):331–345, March 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Yulianti:2016:TBS

[YHS16] Evi Yulianti, Sharin Huspi, and Mark Sanderson. Tweetbiased summarization. Journal of the Association for Information Science and Technology, 67(6):1289–1300, June 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Ye:2014:ATP

[YL14] Fred Y. Ye and Loet Leydesdorff. The “academic trace” of the performance matrix: a mathematical synthesis of the h-index and the integrated impact indicator (I3). Journal of the Association for Information Science and Technology, 65(4):742–750, April 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Yan:2017:FPM

[YL17a] Bowen Yan and Jianxi Luo. Filtering patent maps for visualization of diversification paths of inventors and organizations. Journal of the Association for Information Science and Technology, 68(6):1551–1563, June 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Yan:2017:MTD

[YL17b] Bowen Yan and Jianxi Luo. Measuring technological distance for patent mapping. Journal of the Association for Information Science and Technology, 68(2):423–437, February 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Yan:2018:WDD

[YL18] Erjia Yan and Kai Li. Which domains do open-access journals do best in? A 5-year longitudinal study. Journal of the Association for Information Science and Technology, 69(6): 844–856, June 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Yang:2021:DIS

[YLHW21] Xuan Yang, Xiao Li, Daning Hu, and Harry Jiannan Wang. Differential impacts of social influence on initial and sustained participation in open source software projects. Journal of the Association for Information Science and Technology, 72(9):1133–1147, September 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Yip:2020:DPP

[YLL20] Jason C. Yip, Kung Jin Lee, and Jin Ha Lee. Design partnerships for participatory librarianship: a conceptual model for understanding librarians co designing with digital youth. Journal of the Association for Information Science and Technology, 71(10):1242–1256, October 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Yoon:2017:DRT

[Yoo17] Ayoung Yoon. Data reusers’ trust development. Journal of the Association for Information Science and Technology, 68(4):946–956, April 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Yeoh:2016:EUC

[YP16] William Yeoh and Aleš Popovič. Extending the understanding of critical success factors for implementing business intelligence systems. Journal of the Association for Information Science and Technology, 67(1):134–147, January 2016. CO-DEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Yablowitz:2016:IDP

[YR16] Michal Gaziel Yablowitz and Daphne R. Raban. Investment decision paths in the information age: the effect of online journalism. Journal of the Association for Information Science and Technology, 67(6):1417–1429, June 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

You:2023:SFH

[YR23] Sangseok You and Lionel P. Robert. Subgroup formation in human-robot teams: a multi-study mixed-method approach with implications for theory and practice. Journal of the Association for Information Science and Technology, 74(3): 323–338, March 2023. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Yan:2016:CCA

[YRH16] Sulan Yan, Ronald Rousseau, and Shuiqing Huang. Contributions of Chinese authors in PLOS One. Journal of the Association for Information Science and Technology, 67(3): 543–549, March 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Yang:2018:HIS

[YRWH18] Bo Yang, Ronald Rousseau, Xue Wang, and Shuiqing Huang. How important is scientific software in bioinformatics research? A comparative study between international and Chinese research communities. Journal of the Association for Information Science and Technology, 69(9):1122–1133, September 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Yigit-Sert:2021:EDS

[YSAM^{+}21] Sevgi Yigit-Sert, Ismail Sengor Altingovde, Craig Macdonald, Iadh Ounis, and Özgür Ulusoy. Explicit diversification of search results across multiple dimensions for educational search. Journal of the Association for Information Science and Technology, 72(3):315–330, March 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Yu:2014:ECT

[Yu14] Ning Yu. Exploring co-training strategies for opinion detection. Journal of the Association for Information Science and Technology, 65(10):2098–2110, October 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Yu:2023:LSE

[YXAL23] Chuanming Yu, Haodong Xue, Lu An, and Gang Li. A lightweight semantic-enhanced interactive network for efficient short-text matching. Journal of the Association for Information Science and Technology, 74(2):283–300, February 2023. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Yuan:2018:NMR

[YXJ18] Qingni Yuan, Songhua Xu, and Lv Jian. A new method for retrieving batik shape patterns. Journal of the Association for Information Science and Technology, 69(4):578–599, April 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Yan:2016:UPB

[YY16] Erjia Yan and Qi Yu. Using path-based approaches to examine the dynamic structure of discipline-level citation networks: 1997–2011. Journal of the Association for Information Science and Technology, 67(8):1943–1955, August 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Yu:2016:IIC

[YZ16] Liangzhi Yu and Wenjie Zhou. Information inequality in contemporary Chinese urban society: the results of a cluster analysis. Journal of the Association for Information Science and Technology, 67(9):2246–2262, September 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Yan:2017:ADK

[YZ17] Erjia Yan and Yongjun Zhu. Adding the dimension of knowledge trading to source impact assessment: Approaches, indicators, and implications. Journal of the Association for Information Science and Technology, 68(5):1090–1104, May 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Yao:2020:GVV

[YZQT20] Xinlin Yao, Cheng Zhang, Zhe Qu, and Bernard C. Y. Tan. Global village or virtual Balkans? Evolution and performance of scientific collaboration in the information age. Journal of the Association for Information Science and Technology, 71(4):395–408, April 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Yin:2021:HOR

[YZYR21] Hang Yin, Shuang Zheng, William Yeoh, and Jie Ren. How online review richness impacts sales: an attribute substitution perspective. Journal of the Association for Information Science and Technology, 72(7):901–917, July 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Zheng:2019:SMP

[ZAE^{+}19] Han Zheng, Htet Htet Aung, Mojisola Erdt, Tai-Quan Peng, Aravind Sesagiri Raamkumar, and Yin-Leng Theng. Social media presence of scholarly journals. Journal of the Association for Information Science and Technology, 70(3):256–270, March 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Zigron:2019:HWY

[ZB19] Shimrit Zigron and Jenny Bronstein. “Help is where you find it”: The role of weak ties networks as sources of information and support in virtual health communities. Journal of the Association for Information Science and Technology, 70(2): 130–139, February 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Zuccala:2018:MAB

[ZBBW18] Alesia Zuccala, Mads Breum, Kasper Bruun, and Bernd T. Wunsch. Metric assessments of books as families of works. Journal of the Association for Information Science and Technology, 69(1):146–157, January 2018. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Zhang:2018:USC

[ZBDX18] Chenwei Zhang, Yi Bu, Ying Ding, and Jian Xu. Understanding scientific collaboration: Homophily, transitivity, and preferential attachment. Journal of the Association for Information Science and Technology, 69(1):72–86, January 2018. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Zhang:2014:ESG

[ZBKG14] Yan Zhang, Ramona Broussard, Weimao Ke, and Xuemei Gong. Evaluation of a scatter/gather interface for supporting distinct health information search tasks. Journal of the Association for Information Science and Technology, 65(5):1028– 1041, May 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Zhao:2015:IRR

[ZCvdH15] Yun Wei Zhao, Chi-Hung Chi, and Willem-Jan van den Heuvel. Imperfect referees: Reducing the impact of multiple biases in peer review. Journal of the Association for Information Science and Technology, 66(11):2340–2356, November 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Zahedi:2017:MRF

[ZCW17] Zohreh Zahedi, Rodrigo Costas, and Paul Wouters. Mendeley readership as a filtering tool to identify highly cited publications. Journal of the Association for Information Science and Technology, 68(10):2511–2521, October 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Zheng:2023:GCB

[ZCYN23] Xiang Zheng, Jiajing Chen, Erjia Yan, and Chaoqun Ni. Gender and country biases in Wikipedia citations to scholarly publications. Journal of the Association for Information Science and Technology, 74(2):219–233, February 2023. CO-DEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Zhang:2020:PHS

[ZCZ^{+}20] Jin Zhang, Ye Chen, Yuehua Zhao, Dietmar Wolfram, and Feicheng Ma. Public health and social media: a study of Zika virus-related posts on Yahoo! Answers. Journal of the Association for Information Science and Technology, 71(3): 282–299, March 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Zhao:2023:JSI

[ZDP^{+}23] Yuxiang (Chris) Zhao, Jia Tina Du, Natalie Pang, Jaya Raju, and Hui Yan. JASIST special issue on ICT4D and intersections with the information field. Journal of the Association for Information Science and Technology, 74(12): 1345–1349, December 2023. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Zhai:2018:MDI

[ZDW18] Yujia Zhai, Ying Ding, and Fang Wang. Measuring the diffusion of an innovation: a citation analysis. Journal of the Association for Information Science and Technology, 69(3): 368–379, March 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Zhai:2021:IAB

[ZDZ21] Yujia Zhai, Ying Ding, and Hezhao Zhang. Innovation adoption: Broadcasting versus virality. Journal of the Association for Information Science and Technology, 72(4):403–416, April 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Zemel:2017:TAR

[Zem17] Alan Zemel. Texts as actions: Requests in online chats between reference librarians and library patrons. Journal of the Association for Information Science and Technology, 68 (7):1687–1697, July 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Zhu:2019:EUM

[ZF19] Xiaohua Zhu and Mark Antony Freeman. An evaluation of U.S. municipal open data portals: A user interaction framework. Journal of the Association for Information Science and Technology, 70(1):27–37, January 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Zhang:2019:EMT

[ZFH^{+}19] Xiaojie Zhang, Yulin Fang, Wei He, Yixiang Zhang, and Xinmei Liu. Epistemic motivation, task reflexivity, and knowledge contribution behavior on team wikis: a cross-level moderation model. Journal of the Association for Information Science and Technology, 70(5):448–461, May 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Zhitomirsky-Geffet:2016:PID

[ZGB16] Maayan Zhitomirsky-Geffet and Yair Bratspiess. Professional information disclosure on social networks: the case of Facebook and LinkedIn in Israel. Journal of the Association for Information Science and Technology, 67(3):493–504, March 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Zhitomirsky-Geffet:2017:ACU

[ZGBIL17] Maayan Zhitomirsky-Geffet, Judit Bar-Ilan, and Mark Levene. Analysis of change in users’ assessment of search results over time. Journal of the Association for Information Science and Technology, 68(5):1137–1148, May 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Zhitomirsky-Geffet:2018:CRJ

[ZGBIL18] Maayan Zhitomirsky-Geffet, Judit Bar-Ilan, and Mark Levene. Categorical relevance judgment. Journal of the Association for Information Science and Technology, 69(9): 1084–1094, September 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Zuccala:2015:CWR

[ZGCB15] Alesia Zuccala, Raf Guns, Roberto Cornacchia, and Rens Bod. Can we rank scholarly book publishers? A bibliometric experiment with the field of history. Journal of the Association for Information Science and Technology, 66(7): 1333–1347, July 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Zhou:2022:SSB

[ZGE22] Hongyu Zhou, Raf Guns, and Tim C. E. Engels. Are social sciences becoming more interdisciplinary? Evidence from publications 1960–2014. Journal of the Association for Information Science and Technology, 73(9):1201–1221, September 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Zhitomirsky-Geffet:2017:TMO

[ZGEJ17] Maayan Zhitomirsky-Geffet, Eden S. Erez, and Bar-Ilan Judit. Toward multiviewpoint ontology construction by collaboration of non-experts and crowdsourcing: the case of the effect of diet on health. Journal of the Association for Information Science and Technology, 68(3):681–694, March 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Zheng:2022:UEM

[ZGL^{+}22] Han Zheng, Dion Hoe-Lian Goh, Edmund Wei Jian Lee, Chei Sian Lee, and Yin-Leng Theng. Understanding the effects of message cues on COVID-19 information sharing on Twitter. Journal of the Association for Information Science and Technology, 73(6):847–862, June 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Zhitomirsky-Geffet:2017:IPS

[ZGPB17] Maayan Zhitomirsky-Geffet, Gila Prebor, and Orna Bloch. Improving proverb search and retrieval with a generic multidimensional ontology. Journal of the Association for Information Science and Technology, 68(1):141–153, January 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Zimmer:2017:RSC

[ZH17] J. Christopher Zimmer and Raymond M. Henry. The role of social capital in selecting interpersonal information sources. Journal of the Association for Information Science and Technology, 68(1):5–21, January 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Zhang:2014:LIT

[Zha14a] Lei Zhang. Linking information through function. Journal of the Association for Information Science and Technology, 65(11):2293–2305, November 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Zhang:2014:BQA

[Zha14b] Yan Zhang. Beyond quality and accessibility: Source selection in consumer health information searching. Journal of the Association for Information Science and Technology, 65(5):911–927, May 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Zhang:2014:SSH

[Zha14c] Yan Zhang. Searching for specific health-related information in MedlinePlus: Behavioral patterns and user experience. Journal of the Association for Information Science and Technology, 65(1):53–68, January 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Zhang:2016:USU

[Zha16] Yan Zhang. Understanding the sustained use of online health communities from a self-determination perspective. Journal of the Association for Information Science and Technology, 67(12):2842–2857, December 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Zhang:2019:BRC

[Zha19] Jin Zhang. Book review: Coping with Illness Digitally. Stephen A. Rains (Ed.). Cambridge, MA: MIT Press, 2018. 240 pp. $35.00 (hardcover). (ISBN 978-0-262-03828-7). Journal of the Association for Information Science and Technology, 70(10):1167–1169, October 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Zhu:2018:FCB

[ZHGY18] Jun Zhu, Lixin Han, Zhinan Gou, and Xiaofeng Yuan. A fuzzy clustering-based denoising model for evaluating uncertainty in collaborative filtering recommender systems. Journal of the Association for Information Science and Technology, 69(9):1109–1121, September 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Zimmer:2017:BRD

[Zim17] Michael Zimmer. Book review:: Digital research confidential: the secrets of studying behavior online. Edited by Eszter Hargittai and Christian Sandvig. Cambridge, MA: MIT Press, 2015. 288 pp. $26.00 (Paperback) (ISBN: 978-0-262-52820-7). Journal of the Association for Information Science and Technology, 68(7):1802–1803, July 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Zitt:2015:BRD

[Zit15] Michel Zitt. Book reviews: Les Dérives de l’Évaluation de la Recherche: du bon usage de la bibliométrie (The Excesses of Research Evaluation: The Proper Use of Bibliometrics)

— by Yves Gingras. Journal of the Association for Information Science and Technology, 66(10):2171–2176, October 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Zhou:2023:UEW

[ZJHT23] Zhongyun Zhou, Xiao-Ling Jin, Carol Hsu, and Zhenya Tang. User empowerment and well-being with mHealth apps during pandemics: a mix-methods investigation in China. Journal of the Association for Information Science and Technology, 74(12):1401–1418, December 2023. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Zhang:2014:BVP

[ZK14] Yin Zhang and Sonali Kudva. E-books versus print books: Readers’ choices and preferences across contexts. Journal of the Association for Information Science and Technology, 65 (8):1695–1706, August 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Zhou:2016:NOC

[ZL16] Qiuju Zhou and Loet Leydesdorff. The normalization of occurrence and co-occurrence matrices in bibliometrics using Cosine similarities and Ochiai coefficients. Journal of the Association for Information Science and Technology, 67(11):

2805–2814, November 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Zhang:2015:PUD

[ZLCB15] Xiangmin Zhang, Jingjing Liu, Michael Cole, and Nicholas Belkin. Predicting users’ domain knowledge in information retrieval using multiple regression analysis of search behaviors. Journal of the Association for Information Science and Technology, 66(5):980–1000, May 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Zhang:2020:UAP

[ZLF20] Yanli Zhang, Xinmiao Li, and Weiguo Fan. User adoption of physician’s replies in an online health community: an empirical study. Journal of the Association for Information Science and Technology, 71(10):1179–1191, October 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Zhang:2023:LLG

[ZLY23] Li Zhang, Wei Lu, and Jinqing Yang. LAGOS-AND: a large gold standard dataset for scholarly author name disambiguation. Journal of the Association for Information Science and Technology, 74(2):168–185, February 2023. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Zhang:2014:EEC

[ZLZG14] Zhongfeng Zhang, Qiudan Li, Daniel Zeng, and Heng Gao. Extracting evolutionary communities in community question answering. Journal of the Association for Information Science and Technology, 65(6):1170–1186, June 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Zornic:2014:LEH

[ZMJ14] Nikola Zornic, Aleksandar Markovic, and Veljko Jeremic. Letter to the Editor: How the top 500 ARWU can provide a misleading rank. Journal of the Association for Information Science and Technology, 65(6):1303–1304, June 2014. CO-DEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Zhao:2020:FAA

[ZPL^{+}20] Yuxiang (Chris) Zhao, Xixian Peng, Zhouying Liu, Shijie Song, and Preben Hansen. Factors that affect asker’s pay intention in trilateral payment-based social Q&A platforms: From a benefit and cost perspective. Journal of the Association for Information Science and Technology, 71(5):516–528, May 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Zhang:2024:IPD

[ZPPW24] Dan Zhang, Loo G. Pee, Shan L. Pan, and Jingyuan Wang. Information practices in data analytics for supporting public health surveillance. Journal of the Association for Information Science and Technology, 75(1):79–93, January 2024. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Zhu:2023:PCU

[ZQC^{+}23] Yongjun Zhu, Lihong Quan, Pei-Ying Chen, Meen Chul Kim, and Chao Che. Predicting coauthorship using bibliographic network embedding. Journal of the Association for Information Science and Technology, 74(4):388–401, April 2023. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Zhang:2021:TDS

[ZRdR21] Yangjun Zhang, Pengjie Ren, and Maarten de Rijke. A taxonomy, data set, and benchmark for detecting and classifying malevolent dialogue responses. Journal of the Association for Information Science and Technology, 72(12):1477–1497, December 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Zhang:2016:DRI

[ZRG16] Lin Zhang, Ronald Rousseau, and Wolfgang Glänzel. Diversity of references as an indicator of the interdisciplinarity of journals: Taking similarity between subject fields into account. Journal of the Association for Information Science and Technology, 67(5):1257–1265, May 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Zhang:2014:AIS

[ZS14a] Pengyi Zhang and Dagobert Soergel. Advances in information science: Towards a comprehensive model of the cognitive process and mechanisms of individual sensemaking. Journal of the Association for Information Science and Technology, 65(9):1733–1756, September 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Zhao:2014:TAC

[ZS14b] Dangzhi Zhao and Andreas Strotmann. In-text author citation analysis: Feasibility, benefits, and limitations. Journal of the Association for Information Science and Technology, 65(11):2348–2358, November 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Zhao:2014:KBR

[ZS14c] Dangzhi Zhao and Andreas Strotmann. The knowledge base and research front of information science 2006–2010: an author cocitation and bibliographic coupling analysis. Journal of the Association for Information Science and Technology, 65(5):995–1006, May 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Zhao:2016:DUA

[ZS16] Dangzhi Zhao and Andreas Strotmann. Dimensions and uncertainties of author citation rankings: Lessons learned from frequency-weighted in-text citation counting. Journal of the Association for Information Science and Technology, 67(3): 671–682, March 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Zheng:2019:CER

[ZS19] Xin Zheng and Aixin Sun. Collecting event-related tweets from Twitter stream. Journal of the Association for Information Science and Technology, 70(2):176–186, February 2019. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Zhang:2020:CMS

[ZS20] Pengyi Zhang and Dagobert Soergel. Cognitive mechanisms in sensemaking: a qualitative user study. Journal of the Association for Information Science and Technology, 71(2): 158–171, February 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Zhang:2023:AJS

[ZSAW23] Yin Zhang, Il-Yeol Song, Theresa Anderson, and Dan Wu. About JASIST special issue on “Data Science in the iField”.

Journal of the Association for Information Science and Technology, 74(6):601–605, June 2023. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Zhao:2018:TFA

[ZSC18] Dangzhi Zhao, Andreas Strotmann, and Alicia Cappello. Intext function of author self-citations: Implications for research evaluation practice. Journal of the Association for Information Science and Technology, 69(7):949–952, July 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Zeng:2022:JSI

[ZSG^{+}22] Marcia Lei Zeng, Chris Alen Sula, Karen F. Gracy, Eero Hyvönen, and Vânia Mara Alves Lima. JASIST special issue on digital humanities (DH). Journal of the Association for Information Science and Technology, 73(2):143–147, February 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Zubiaga:2015:RTC

[ZSMF15] Arkaitz Zubiaga, Damiano Spina, Raquel Martínez, and Víctor Fresno. Real-time classification of Twitter trends. Journal of the Association for Information Science and Technology, 66(3):462–473, March 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Zhang:2015:QHI

[ZSX15] Yan Zhang, Yalin Sun, and Bo Xie. Quality of health information for consumers on the web: a systematic review of indicators, criteria, tools, and evaluation results. Journal of the Association for Information Science and Technology, 66 (10):2071–2084, October 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Zhang:2022:QHW

[ZT22] Yan Zhang and Ciaran B. Trace. The quality of health and wellness self-tracking data: a consumer perspective. Journal of the Association for Information Science and Technology, 73(6):879–891, June 2022. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Zou:2020:UDN

[ZTA20] Jie Zou, George Thoma, and Sameer Antani. Unified deep neural network for segmentation and labeling of multipanel biomedical figures. Journal of the Association for Information Science and Technology, 71(11):1327–1340, November 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Zhu:2015:MAI

[ZTLV15] Xiaodan Zhu, Peter Turney, Daniel Lemire, and André Vellino. Measuring academic influence: Not all citations are equal. Journal of the Association for Information Science and Technology, 66(2):408–427, February 2015. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Zubiaga:2018:LAP

[Zub18] Arkaitz Zubiaga. A longitudinal assessment of the persistence of Twitter datasets. Journal of the Association for Information Science and Technology, 69(8):974–984, August 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Zuccala:2014:MLA

[ZvSvB14] Alesia Zuccala, Maarten van Someren, and Maurits van Bellen. A machine-learning approach to coding book reviews as quality indicators: Toward a theory of megacitation. Journal of the Association for Information Science and Technology, 65(11):2248–2260, November 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Zimmer:2020:EII

[ZVW20] Michael Zimmer, Jessica Vitak, and Philip Wu. Editorial introduction: “Information privacy in the digital age”. Journal of the Association for Information Science and Technology, 71(9):997–1001, September 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Zhang:2023:WOR

[ZW23] Dong Zhang and Chong Wu. What online review features really matter? An explainable deep learning approach for hotel demand forecasting. Journal of the Association for Information Science and Technology, 74(9):1100–1117, September 2023. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Zhang:2023:DSC

[ZWH^{+}23] Yin Zhang, Dan Wu, Loni Hagen, Il-Yeol Song, Javed Mostafa, Sam Oh, Theresa Anderson, Chirag Shah, Bradley Wade Bishop, Frank Hopfgartner, Kai Eckert, Lisa Federer, and Jeffrey S. Saltz. Data science curriculum in the iField. Journal of the Association for Information Science and Technology, 74(6):641–662, June 2023. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Zhang:2023:HQT

[ZWTX23] Xing Zhang, Durong Wang, Yuyao Tang, and Quan Xiao. How question type influences knowledge withholding in social Q&A community. Journal of the Association for Information Science and Technology, 74(10):1170–1184, October 2023. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Zhang:2018:HBJ

[ZWW18] Peng Zhang, Peiling Wang, and Qiang Wu. How are the best JASIST papers cited? Journal of the Association for Information Science and Technology, 69(6):857–860, June 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Zhou:2018:EPN

[ZXLC18] Hao Zhou, Lu Xiao, Yongmei Liu, and Xiaohong Chen. The effect of prediscussion note-taking in hidden profile tasks. Journal of the Association for Information Science and Technology, 69(4):566–577, April 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Zhao:2018:UDD

[ZXPS18] Yuxiang (Chris) Zhao, Xiaojuan Xu, Xixian Peng, and Shijie Song. Understanding the determinants and dynamic process of user exodus in social networking sites: Evidence from Kaixin001. Journal of the Association for Information Science and Technology, 69(4):553–565, April 2018. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Zhang:2015:UCN

[ZY15] Mi Zhang and Christopher C. Yang. Using content and network analysis to understand the social support exchange patterns and user behaviors of an online smoking cessation intervention program. Journal of the Association for Information Science and Technology, 66(3):564–575, March 2015. CO-DEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Zhao:2018:DSM

[ZYL18] Mengnan Zhao, Erjia Yan, and Kai Li. Data set mentions and citations: a content analysis of full-text publications. Journal of the Association for Information Science and Technology, 69(1):32–46, January 2018. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Zhu:2017:UGB

[ZYS17] Yongjun Zhu, Erjia Yan, and Il-Yeol Song. The use of a graph-based system to improve bibliographic information retrieval: System design, implementation, and evaluation. Journal of the Association for Information Science and Technology, 68(2):480–490, February 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Zhang:2016:CKP

[ZYZ^{+}16] Juan Zhang, Qi Yu, Fashan Zheng, Chao Long, Zuxun Lu, and Zhiguang Duan. Comparing keywords plus of WOS and author keywords: a case study of patient adherence research. Journal of the Association for Information Science and Technology, 67(4):967–972, April 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Zhang:2020:POT

[ZZ20] Min Zhang and Yin Zhang. Professional organizations in twittersphere: An empirical study of U.S. library and information science professional organizations-related tweets. Journal of the Association for Information Science and Technology, 71(4):491–496, April 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Zhang:2021:EKE

[ZZ21a] Yingyi Zhang and Chengzhi Zhang. Enhancing keyphrase extraction from microblogs using human reading time. Journal of the Association for Information Science and Technology, 72(5):611–626, May 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Zuo:2021:UPF

[ZZ21b] Zhiya Zuo and Kang Zhao. Understanding and predicting future research impact at different career stages — a social network perspective. Journal of the Association for Information Science and Technology, 72(4):454–472, April 2021. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Zuo:2017:SEU

[ZZE17] Zhiya Zuo, Kang Zhao, and David Eichmann. The state and evolution of U.S. iSchools: From talent acquisitions to research outcome. Journal of the Association for Information Science and Technology, 68(5):1266–1277, May 2017. CO-DEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Zhao:2014:ACS

[ZZL^{+}14] Star X. Zhao, Paul L. Zhang, Jiang Li, Alice M. Tan, and Fred Y. Ye. Abstracting the core subnet of weighted networks based on link strengths. Journal of the Association for Information Science and Technology, 65(5):984–994, May 2014. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Zhang:2020:JMC

[ZZL20] Yingyi Zhang, Chengzhi Zhang, and Jing Li. Joint modeling of characters, words, and conversation contexts for microblog keyphrase extraction. Journal of the Association for Information Science and Technology, 71(5):553–567, May 2020. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Zhang:2016:SNA

[ZZLS16] Jin Zhang, Shanshan Zhai, Hongxia Liu, and Jennifer Ann Stevenson. Social network analysis on a topic-based navigation guidance system in a public health portal. Journal of the Association for Information Science and Technology, 67 (5):1068–1088, May 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Zhang:2016:OSD

[ZZSX16] Jin Zhang, Shanshan Zhai, Jennifer Ann Stevenson, and Lixin Xia. Optimization of the subject directory in a government agriculture department web portal. Journal of the Association for Information Science and Technology, 67(9): 2166–2180, September 2016. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).

Zhang:2023:BIC

[ZZY23] Yuhao Zhang, Guangchun Zheng, and Hui Yan. Bridging information and communication technology and older adults by social network: an action research in Sichuan, China. Journal of the Association for Information Science and Technology, 74(12):1437–1448, December 2023. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Zhang:2017:SEP

[ZZZL17] Yi Zhang, Guangquan Zhang, Donghua Zhu, and Jie Lu. Scientific evolutionary pathways: Identifying and visualizing relationships for scientific topics. Journal of the Association for Information Science and Technology, 68(8):1925–1939, August 2017. CODEN ???? ISSN 2330-1635 (print), 2330-1643 (electronic).